A Selected Bibliography of Publications by, and about, Max Planck

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 November 2022 Version 1.69

Title word cross-reference

12 × 5 [Pic25]. 14 × 22 [C.49]. $2.40 [Pic25]. 2.7^{∘} [Noe73]. $3.75 [C.49]. $6 [Hie68]. α [Pla06d, Pla06e, Pla07a, Why50]. β [Pla06d, Pla06e, Pla07a]. C_{2} [Has14, Men17, VD47]…. [Hof08a]. h [Ano47c, Bir19, Mil16].

-. [Sch65]. -Strahlen [Pla06d, Pla06e, Pla07a].

. [Sch65].

/Almost [KSB63]. /Aufsätze [Ano48e, Ano48f]. /Butterflies [Ano46c]. /Death [Ano46c]. /Dr. [Ano46c]. /Einigkeit [MSQM63]. /Experiment [Ano46b]. /Fast [KSB63]. /Ferromagnetism [HD48]. /Friedrich [MKD^{+}91]. /Fritzsch [MKD^{+}91]. /Goethe [Ano46c]. /Is [MSQM63]. /Ist [MSQM63]. /Lenz [Ano48i]. /Max [BT58, Cah99, DT58, vBL96]. /Max-Planck-Medaille [BT58, DT58]. /Nobel [Ano46c]. /Papers [Ano48e]. /Pasteur [Ano46c]. /President [Ano46c]. /The [Ano46c]. /Unity [MSQM63].

0-520-05710-4 [San87].

1 [Pla15a, Pla46a]. 1/B [Ano46b]. 1/Zerstörung [Ano46b]. 100th [HD01, Max58, vL58]. 10th [Pla57d]. 11 [You68]. 12 [Pla31e]. 12.14.1930/Albert [Ano69]. 130th [Gri87]. 14.12.1930/Albert [Ano69]. 15 [Pla14f, Pla16e, Pla23h]. 15.25 [San87]. 150 [Sta10]. 150th [Ano08f, Ano08a, Bec09, Jac08b]. 15s [McK27]. 16 [Pla94a]. 1646 [Pla46a]. 18 [Pla27b]. 1889/91 [Pla15c]. 1890 [Pla50a, Sti84]. 1893 [Pla94a]. 1894 [Pla94a]. 19 [MSQM63]. 1900 [Brü48b]. 1902 [Gae03]. 1906 [Pla06f]. 1908 [Pla09c]. 1909 [Pla09b, Pla10a]. 1910 [Pla10c, Pla10d, Pla10e, Pla11a]. 1911 [Pla13e, Str11a]. 1911/sous [LdB12]. 1913 [Pla13c, Pla13d, Pla14c, Pla14f, Pla14g, Pla16e, Pla23h]. 1914 [Jüt17, Pla14d]. 1918 [Pla18a]. 1920 [Jüt21, Pla20c]. 1920/Jakob [PS58, PS58]. 1923 [Ano24]. 1925 [PP79]. 1927 [Pla27a, Pla27b]. 1928 [Pla28c, Pla29e, Pla29f]. 1929 [Pla29c, Pla29b, Pla29m, Pla29k, Sch29]. 1930 [Pla31e]. 1931 [Hei31]. 1932 [Pla32b, Sim06]. 1935 [Pla35a, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. 1936 [Pla36c, Pla39c, Pla52d, Pla53d, Pla55e]. 1937 [Pla38h]. 1938 [Ano38a, Pla38a, Pla38g]. 1939 [Pla39b]. 1943 [Ano43b]. 1945 [FH00, Hop96]. 1947 [Ano08c, Bor63b]. 1947/Cottam [KAR89]. 1948 [Ano48j, Ano49a, DdCA48]. 1949 [Hop96]. 195-264 [Jüt18a]. 1958 [Hof96, KMP59, Max58, vL58]. 1959b [CS79b]. 1961 [Ano61]. 1961/Max [Ano61]. 1961/Preis [Ano61]. 1964 [Kas64]. 1967 [Pun67]. 1969 [Rut69]. 1970 [Haa70]. 1973 [KZR^{+}73]. 1974 [Haa75]. 1978/Karl [BWN78]. 1978/Karl-Scheel-Preis [BWN78]. 1978/Nominierung [BWN78]. 1981 [Ano81, BCS82]. 1981/Frühjahrstagungen [Ano81]. 1981/Haushaltsplan [Ano81]. 1997 [Ano97a, MH97].

2 [Ano48a, PP79, Pla20c, Pla29e, Pla29f]. 2.6.1920 [Pla20b]. 2008 [Bec09]. 2011 [Pla11c, Pla11d, Par12]. 20th [Sch96]. 21 [Pla11c, Pla11d, Pla32b, Pla36c]. 23 [Ano38a, Ano43b, Ano48j, Ano49a, Bec09, DdCA48, Pla38a]. 23rd [Max58]. 24.4.1958 [vL58]. 25 [Haa75, Pla37a]. 25-jährigen [Pla37i]. 25th [Ham72, Pla37i]. 26 [Pla18a]. 27 [Pla35a, Pla39c, Pla52d, Pla53d, Pla55e]. 28 [Pla27a, Pla29b, Pla29k]. 28.10.1919 [Pla20a]. 28th [Sch29]. 29 [Pla10c, Pla10d, Pla10e, Pla11a]. 29th [Pla28c]. 2nd [KGSO55].

3 [Ber14, Dri01, Pla14d]. 3-927579-08-4 [Kra97]. 30 [Pla06f]. 300th [Pla46a]. 31.1.1930 [Ano69]. 3hyphen [Dri01].

4 [MH97, Pla29c]. 4th [Pla91b].

5 [Pla06a]. 5.70 [Ano39]. 50 [Sch00b, Wei68]. 50th [Ano97a, Ano97b, Eck97a, MH97, RH97]. 540 [Dri01]. 540-41274-3 [Dri01]. 5th [KGSO55].

6 [Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. 60th [Ber15, Din59, Pla18a, PS58, Wes57]. 65th [GSU07].

7 [Gae03]. 70 [Ben30]. 70th [Pla38b].

8 [Zil33]. 8. [Reg34]. 8.75 [Sch49]. 80th [Ano38a, HRR10b, Pla38g, SR37, Som38a, Zen38, vL38].

9 [Pla09c]. 90th [Brü48a, HD48]. 978 [Ber14]. 978-3-8442-5258-3 [Ber14]. 9s [Cra27].

A. [Rut69]. A.I.C [Cra27]. a.o [BDR86]. A.R.C.Sc [Cra27]. Abänderung [Kan69]. Abb [Ano34, Ano39, Sch55, Wil37]. Abbildungen [Ano36, Jüt22]. abermals [Ano24, Her24, Jüt22]. Abhandlung [Pla00a, Wes40, Pla87c, Pla87d, Pla87e, Pla91b, Pla89j]. Abhandlungen [Ano02, Ano58, BBG^{+}58, Cla91, Ebe63, Max58, Pla23b, Pla28d, Pla58f, Pla95, Pla97, PHR07, Bor63a, Kir98a, Kir98b, Pla15a]. ableiten [Cla89]. Ableitung [Pla01g, Pla23b, Pla95, Pla97, PHR07]. Ablenkbarkeit [Pla06d, Pla06e, Pla07a]. Ablenkungsmessungen [Pla07c]. abnormal [Pla05a]. Above [Kir98b]. absolute [Hei07, Her77, Jac08b, Pla16f, Pla25g, Pla21a, Pla22a]. Absoluten [Her77, Jac08b, Pla25g]. absorbed [Pla00f, Pla02c]. absorbierte [Pla00f, Pla02c]. Absorption [Kir98a, Pla03h, Pla95a, Pla96a]. Absorptionsvermögen [Kir98a]. Abteilung [BDR86]. academic [Hei31, Hof08b, Lem08]. Academy [Ano48c, FH00, Pla20b, Pla20c, Pla23g, Pla48f, Sim06, vL58, Ano48c, Hof00, KK75, Kir79, LMOS15, PP79, Pla15a, Pla17a, Pla48f, Pla57d, Sim06]. access [Ber14]. Accommodation [Hui13]. accompanying [Brü38]. according [Pla19c]. accroissement [Duh88]. accuracy [Gré19]. Accurate [WSNO98]. achievement [Jor48]. Acht [Pla10a]. Achtzig [Hei38]. achtzigsten [Pla38a, Zen38]. acid [Kir98b]. Ackeret [PS58, PS58]. Ackermann [BBG^{+}58]. Across [Hei74]. Act [Büh98a, HRR10b, Dar00]. Action [Ano21, Ish15, Ish17, PB17, Pla08a, Pla08b, Pla09a, Pla12a, Pla14a, Pla14b, Pla25a, Pla43c, Ros08, Som11, Som12, dB48]. adaptation [HW07, Hac07]. addendum [Pla01b, Pla34c, Pla40c, Pla01c]. Addition [BT58, DT58]. Additions [BT58, DT58]. Address [DdCA48, Pla25c, Pla48c, Pla57d, Pla29b, Pla38a, Sch29]. addresses [JDB^{+}97, Kir79, SCP19]. addresses/Looking [JDB^{+}97]. Adenauer [Ano70, Ano70]. adjustment [Gré19]. ADLON [JDB^{+}97, JDB^{+}97]. ADLON/in [JDB^{+}97]. ADLON/kurzgefaßt/Auf [JDB^{+}97]. admits [Abb01]. Adolf [Alb93, HT99, Mul00, Pla08d, Pla47b, Sch00b, Sch04, HT99, O’F56, Pla47b]. Adresse [Pla29b, Pla38a, Sch29]. Adressen/Blick [JDB^{+}97]. Adrien [Ano52]. Advanced [Ano37]. Advancement [Pla31e, Pla37a]. advent [Pye85b]. Advisory [Pla33g]. Aether [Pla02b]. Affiliation [Anoxx]. after [BA01, Fad96, Kae48, Löh12, Sch03]. aftermath [Hen07]. against [Ano46c, Bad08, Jae79, Noe73, Hof96, Pla96b]. Ahnenerbe [Ano41b, Mos42]. Ahnenerbe-Stiftung-Verlag [Ano41b, Mos42]. Aiding [Ano47c]. air [Pla89a]. Akad [Pla20b, Pla23g]. Akademie [Ano48c, FH00, Hof00, Jüt21, KK75, Kir79, Koy52, PP79, Pla15a, Pla17a, Pla20c, Pla48f, Sch49, Sim06, Som49, vL58, Pla57d]. Akademie-Ansprache [Ano48c]. Akademie-Ansprachen [Koy52, Pla48f, Sch49, Som49]. Akademie-Verlag [Sch49, Som49]. Akademie-Vorträge [Sim06]. Akademische [Hei31]. akademischer [Hof08b, Lem08]. Akt [Büh98a, Dar00]. Aktion [Pla08a, Pla08b, Pla09a]. Alax [Hac07]. Albert [Adl03, Ano69, Bre08, Gut15, Her03, Hof08f, Koh02, Mic68, Ano69, Bre08, Gut15, Hof08f, KKRS96, Koh02, Prz86, Prz11, Sim06]. alchemist [Ano53b]. alchemist/International [Ano53b]. Alfred [Adl03, BCS82, Kut14a, Kut14a]. algebra [Bae01]. Allgemeine [Pla15f, Pla16c, Pla21c, Pla28b, Pla37e, Som11]. allgemeinen [Pla93c, Pla08a, Pla08b, Pla09a]. allgemeines [Pla89m, Pla91a]. Allgemeinheit [Pla89n]. Allgemeinverständliche [BDR86]. allgemeinverständlichen [Hei88a, Hei06]. Allgemeinwissens [Dri01]. Allow [Sei08]. Almost [KSB63]. already [Bre08]. alt [SH53, Hei38]. alt/25 [SH53]. alten [Pla46b]. alternative [Lon84, Lon03c]. Alwin [MdB63]. Am [Ano38a, AH93, Ano49a, Bec09, Dri01, EP06, Flü47, Haa75, Hei31, Jüt17, Jüt21, PP79, Pla94a, Pla06f, Pla09c, Pla10c, Pla10d, Pla10e, Pla11c, Pla11d, Pla11a, Pla14d, Pla14f, Pla16e, Pla20b, Pla20c, Pla20a, Pla23h, Pla27a, Pla27b, Pla28c, Pla29b, Pla29k, Pla29e, Pla29f, Pla31e, Pla32b, Pla35a, Pla35c, Pla35d, Pla36c, Pla37d, Pla38a, Pla39c, Pla45a, Pla45b, Pla52d, Pla53d, Pla55e, Sch29, Som48, Sta10, Vog61, WvLuAE18, vL58, Pla06a]. Ambr [Ben30, Jüt21, Jüt22]. Ambrosius [Ano24, Ano30b, Ano48i, Wei68]. amendment [Kan69]. American [WC91, BBG^{+}58, HMF98]. Amerika [Ano46b]. Amerika/Nobel [Ano46b]. Amerika/Nobel-Preis [Ano46b]. Amerikanischer [Ano46b]. Amplitude [SDF14]. Amplitude-driven [SDF14]. Analog [CG00, CG03, CC03]. Analyse/Peters [GBM^{+}49]. Analysis [PB17]. andere [Hei69, Pla90b]. anders [MAS63]. Anhang [Mei47, Sch00b]. Anlass [Max58]. anlässlich [HD01, Haa75, Pla13h]. Anlayisi [Pla15g]. Anmerkungen [Alb93, Wes40]. Anna [Pla30b, Pla30b]. Annalen [Gae03, Hof08a, Hof08c, Pye83, Pye85a, Pye08]. anniversary [Ano46c, Ano97a, Ano97b, Eck97a, Gri87, Ham72, HD48, MH97, Pla15a, Pla17a, Pla22b, Pla24a, Pla26a, Pla28a, Pla29b, Pla30c, Pla35e, Pla37b, Pla37i, Pla46a, Pla57d, RH97, Sch29, Wes57]. announced [Ano04]. announcement [Pla27b]. Annual [Pla20a]. anormale [Pla05a].

Anpassung [HW07, Hac07]. Anschauliche [BBG^{+}58]. Ansprache [Ano48c, Pla13a, Pla15a, Pla17a, Pla19a, Pla22b, Pla24a, Pla26a, Pla28a, Pla30c, Pla32b, Pla35a, Pla37b]. Ansprachen [Koy52, Pla18a, Pla48f, Sch49, Som49, WvLuAE18]. answer [Ano69]. anthology [Hen96]. anthropomorphizing [Weg10]. antithetisch [Pla89c]. Antoine [Pla89a, Pla89b, Pla89h, Pla89f, Pla89g, Pla98a, Pla89a, Pla89b]. Anton [Fis10b, Fis12, Fis10b, Fis12]. Antoon [Pla28c]. Antritt [Pla14f, Pla16e, Pla23h]. Antrittsrede [Pla25c, Pla57d, BP24, EP06, FP27, MP19, PP25, PP79, Pla29c, SP29, Sch61, vLP21]. Antrittsreden [Kir79, Pla37f, Pla38f, SCP19]. Antwort [Ano69]. Anwendung [KGSO55, Hol13, Pla89t]. Anwendungen [Kir98b]. anyone [Hof08a]. Apfel [Büh98b]. Apology [Koe01]. Appendix [Seg07a, Mei47]. apple [Büh98b]. Application [Som12, Bad09, KGSO55, Hol13, Som12]. Applications [Sch17, Kir98b, Ris96]. Applied [BT58, DT58]. April [Ano38a, Ano43b, Ano48j, Ano49a, Bec09, DdCA48, Jüt18a, Max58, Pla11c, Pla11d, Pla13c, Pla13d, Pla14c, Pla18a, Pla38a, Pla38g, Sta10, WvLuAE18, van48, vL58, Pla13c, Pla13d, Pla14c]. Arbeit [Ein07a, Her76]. Arbeiten [Ber15, vL18]. Arbeitet [Ano46b, Ano46c]. Arbeitet/Präsident [Ano46b, Ano46c]. Arbeitsgem [BPL^{+}54]. Arbeitsgemeinschaft [HD48]. Arbeitstagung [HD48]. arbitrary [Pla22i]. Archimedes [Pic08]. Archiv [Kra97]. archival [Uli97, Ull13]. archivalischen [Uli97, Ull13]. Archives [Ano14]. Armin [Dri01, Kra02a, Kra02b]. Arnold [Cra27, Eck97b, Jüt18a, Eck97b, Pla38b, Pla38c]. Aroldo [Pla89d]. Arrhenius [Pla92b, Cra84, Pla92b]. Art [Cla91]. artefacts [Qui17]. Arthur [Adl03]. ARTICLE [Pla40a, Pla42d]. Arzeliès [KZR^{+}73]. Arzimowitsch [KZR^{+}73]. ask [BA01]. Assemble [Tib13]. Assessing [Gré19]. assessment [Thi69]. Association [Pla29e, Pla29f]. assumption [Pla13h, Pla14g]. Assumptions [v.B18]. Astronautical [KGSO55]. Astronautische [KGSO55]. Astronomie [Ano48i]. Astronomie/Max [Ano48i]. Astronomisches [Ano46b]. astronomy [FF91]. asymmetric [Pla18b]. asymmetrischen [Pla18b]. Äther [Pla01a]. atmosphärischen [Pla89a, Pla89f]. atmospheric [Pla89a]. Atoll [Ano46b]. Atom [Ano23, Ano47c, HN70, MKD^{+}91, JDB^{+}97, Pet14, Pla16f]. Atombomben [Ano46b, Ano46c]. Atombomben/Nobel [Ano46b, Ano46c]. Atombomben/Nobel-Preis [Ano46b, Ano46c]. Atomenergie [GBM^{+}49]. Atomenergie/Klein [GBM^{+}49]. Atomfallen [JDB^{+}97]. Atomfallen/kurzgefaßt/Max [JDB^{+}97]. Atomforschung [Ano46b]. Atomforschung/Schiffssteuerung [Ano46b]. Atomforschung/Wieder [Ano46b]. Atomic [Cra84, FR13, KZR^{+}73, Pla23c, Hei58, Pla24c, Bro36]. Atomicity [Din46c]. Atomism [Kri86, Yos95a]. Atomismus [Yos95a]. Atomkontrolle [Ano46b]. Atomkontrolle/Ehrung [Ano46b]. Atomlehre [Hei58]. Atommodells [Pla24c]. Atomphysik [GBM^{+}49, MKD^{+}91]. Atomphysik/Ignatovich [MKD^{+}91]. Atomphysik/Max [GBM^{+}49]. atoms [Qui17]. Atomtheorie [Pla23c]. Atomzertrümmerung [Ano48i].

Attempt [Pla40b, Pla40c, Pla41c]. attitude [Löh12]. auch [JDB^{+}97, Wes57]. Audretsch [KAR89]. Auffindung [Pla43c]. Aufgabe [MKD^{+}91]. Aufl [Ebe63, Her24, Sac69, Sch55, Wil37]. Auflage [Ano13, Ano24, Ano30b, Ano36, Ano39, Ano41a, Cla91, Jüt11, Jüt18b, Jüt22, Sch65, Wei68]. Aufnahme [KK75, Kir79, Pla57d]. Aufsätze [Ano48e, Ano48f, Jos95, Pla22h]. Aufwind [JDB^{+}97]. Aufwind/Trägheitsfusion/Erstes [JDB^{+}97]. Aug [Jüt17]. Auge [Bre08]. August [BCS82, Jüt18a, Pla14d]. aula [DdCA48, Ano48j]. Ausdehnung [Pla89a, Pla89f]. ausgedehnt [Pla89g]. Ausgewählte [Bor63a]. ausländischer [Ano29b]. Auslieferung [Kra97]. auspices [LdB12]. Außenwelt [Hei31, Pla31e, Pla40a]. Aussenweltproblem [GBM^{+}49]. Ausstellung [Ano97a, Ano11, Bad08, Uns67]. Ausstellungseröffnung [Ano11]. Austrian [Ano11]. Auswahl [Hei88a, Hei06]. Auswirkung [Pla33g]. Auswirkungen [Hei49]. Auszug [Pla48d]. auteur [Pla08a]. author [Pla08a]. authority [Eck97a]. Authors [Ano48e, Ano48f]. Autobiographical [BDR86]. Autobiographie [Pla60a, Pla91a]. autobiographies [Her73a]. Autobiographisches [BDR86]. Autobiography [A.50, Dou49, Pla49g, Pla50b, Pla68, Pla71, Pla14, Rus50, C.49, Pla48c, Pla60a, Pla90c, Pla91a, Pla58g]. Autonomie [HW07]. autonomy [HW07]. Autoren [Ano48e, Ano48f]. Autorität [Eck97a]. aux [Som12]. avril [Ano48b]. avtobiografija [Pla58g]. award [Ano94, Fie75, Haa75, Ano46c]. Awarded [Ano29c]. Awarding [KGSO55]. Awards [Ano04]. ax [Ein06]. ayant [Duh88].

B [Ano46b, Ber60, Ano34, Ano39, Wil37]. back [BA01, Frö01]. Bad [Ano46b, Ano46c, BWN78, Ano46c]. Baden [Som48]. Bahnen [Pla13h, Pla14f, Pla16e, Pla23h]. balance [Pla13i, Pla35b]. Ballif [KZR^{+}73]. Ballistic [Hop96]. Ballistics [Ber14]. Ballroom [Ano11]. Band [BDR86, Cla91, Fra31, Pla37a, Pla15c]. Barth [Ano24, Ano30b, Ano48i, Ben30, Flü47, Jüt17, Jüt21, Jüt22, Wei68, Her24]. Barth-Verlag [Flü47]. Baruch [Ano46b]. based [Bad08, Qui17]. Basel [Sac69]. basic [Hei58, Pla06c, HKS^{+}68]. basics [JDB^{+}97]. bath [SNE01]. Battle [Som36, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c]. Baur [Ano48i]. Bauten [Cah99]. Bayrischer [Ano46b]. Bd [Ebe63, Hei43a, Sch00b, Thi71]. Be [BT58, DT58, Cha97, Cla89, HS97, Pla42f, Qui17]. beams [Her01]. beantworten [Ano69]. Beck [Sta10]. Beckel [GBM^{+}49]. Becker [Ano48i]. Becksmann [GBM^{+}49]. Bedeutung [Her58, Ish15, Pla04b, Pla06d, Pla06e, Pla07a, Pla10h, Pla13c, Pla13d, Pla14c, Pla31c, Som11]. been [Bre08]. befindlichen [Pla89x]. before [Pad83, Pla20b, Sch03, vL58]. Began [Kae48]. Beginn [Lem08]. beginning [Lem08]. beginnings [Kle61]. Begleittext [Brü38]. Begründers [Bad08]. Begründung [Ano41b, Din39, Kan70, Mos42, Pla12b, Pla26d, Thi71]. behind [Pic08]. beider [Hof08b]. Beim [Ano46b, Pla14f, Pla16e, Pla23h, Som38b]. being [Mei58, Pla32g, Pla32h, Pla32i, Pla32j, Pla33d, Pla49a, Pla49e, Pla49h, Pla49i, Pla49j, Pla57a, Pla57b, Pla57c, Pla64b, Wes58]. Beings [Pla19b]. Beirates [Pla33g]. Beiser [HKS^{+}68]. Beiträge [Sch00b, HKS^{+}68]. Beitragsvermeidung [JDB^{+}97]. Bekker [BBG^{+}58]. beliebige [Pla88e]. beliebiger [Pla22i]. belief [Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c, Hei35]. Beliefs [Ano31, Mes07]. Believe [MKD^{+}91]. belong [Har59]. Belyaev [MKD^{+}91]. Bemerkung [Kir98b, Pla23d, Pla06b, Pla15b, Pla15c, Pla16a]. Bemerkungen [Pla89o, Pla89r, Pla92a, Pla00a, Pla08a, Pla08b, Pla09a, Pla35b, Spr89]. Benndorf [SS53]. Benndorf/Max [SS53]. beobachtet [JDB^{+}97]. beobachtet/Laserlicht [JDB^{+}97]. Berechnung [Hol13, Pla98a]. bereits [Bre08]. Bergmann-Schäfer [WF63]. Bericht [Ano46b]. Berichterstatter [Ano36]. Berichtigung [BPL^{+}54, Ein07a]. Berkeley [San87]. Berlin [Ano02, Ano30b, Ano41b, Ano97a, Ber60, Ber14, BWN78, Dri01, FH00, Fra31, Hof96, Jüt18a, Kra97, Mos42, Pla94a, Pla13h, Pla14f, Pla16e, Pla23g, Pla23h, Pla28c, Pla36c, Pla08a, Sch49, Sch00b, Sch65, vL58, Ano29a, Ano36, Ano46d, Ano48c, Ano11, FH00, Hof96, Hof99, KK75, Kir79, LMOS15, Pla94a, Pla14d, Pla14f, Pla14g, Pla15a, Pla16e, Pla17a, Pla23g, Pla23h, Pla28c, Pla35e, Pla35c, Pla35d, Pla36c, Pla37d, Pla45a, Pla45b, Pla48f, Pla08a, Sch49, Som49, Koy52]. Berlin-Dahlem [Kra97, Ano36, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. Berliner [Ano46b, Jüt17, Jüt18a, KK75, Kir79, Pla13h, Pla14d, Pla15a, Pla17a, Sch00b]. Berlins [LMOS15]. Bernhard [HT99, Mul00, Sch00b, Sch04, HT99]. Bertele [GBM^{+}49]. Bertelsmann [Ano48i]. Bertelsmann-Verlag [Ano48i]. besonderenVorstellung [Cla91]. Besprecher [Ano41a]. Besprechung [Pla07c, HPE^{+}33, Hei35, Hei43a, Som33]. Bestandsverzeichnis [HD01]. Bestehens [Pla35e]. Bestimmung [BBG^{+}58]. Besuch [Alb93, Ano46a, Pla47b]. betrachtet [Pla99a]. between [Boy79, BFG81, HT99, HW07, Hof08b, Hop96, Kir98a, Koh02, Pla90b, Pla90e, Pla01a, Pla02b, Pla13i, Pla40b, Pla40c, Pla41c, Thi69, Wei89, vP06, Hof08f]. bewegende [Cla89]. bewegten [Pla07d]. bewegter [Pla07f, Pla08f]. Bewegung [Cla91]. Beweis [Pla94c, Pla95b]. Beyond [Cas09, SDF14, MKD^{+}91]. Beziehungen [Hof08b, Pla89c, Pla88b]. Bezold [Pla07e]. bezüglich [Pla89m]. Bibliographie [Cah99]. Bibliography [Ano77, Ole81, Low77]. Bibliometrische [Cah99]. Bibliothek [HS97]. Bieberbach [BP24]. Big [Hür22, HRR10b]. Bikini [Ano46b]. Bikini-Atoll [Ano46b]. Bild [Cah99, Sch49, Som49]. Bilddokumenten [Her73a]. Bilder [Ano48e, Ano48f]. Bilder/Autoren [Ano48e, Ano48f]. Bildern [Ano41a]. Bildnis [Pla48c]. Bildnisse [Ano53a]. binärer [Pla90b, Pla90e, Pla02e]. binary [Pla90b, Pla90e, Pla02e]. Biographical [BDR86]. Biographie [Sch65]. Biographisches [BDR86]. Biography [Nye06, Pla58a]. Biologisch [Cah99]. Biologisch-Medizinischen [Cah99]. biophysics [MSQM63]. Biophysik [MSQM63]. birth [Gri87, Nau16, Pla17a, Pla46a]. Birthday [Bec09, Kem38, SR37, Ano38a, Ano57, Ano08f, Ano08a, Ber15, Brü48a, Din59, GSU07, HRR10b, Jac08b, Max58, Pla18a, Pla38a, Pla38b, Pla38c, Pla38g, PS58, Som18, Som38a, Wei58, Wes57, Zen38, vL38, vL58]. birthday/The [Wes57]. bis [Ano39, Büh98b, Fis10b, Fis12, Kan70, KK75, PS58, Sch78c, Thi71]. bisherige [Jüt21, Pla20b, Pla20c, Pla22c, Pla58e]. Björkborn [BCS82]. Bl [MSQM63]. Bl. [MSQM63]. Black [Kuh78, Kuh87, Lon03a, Lon03b, Mul94, Pla12b, Pla12a]. Black-body [Kuh78, Kuh87, Lon03a, Lon03b]. blackbody [Gea08, Seg07a, War12]. Blaschke [BBG^{+}58]. Blick [Frö01, JDB^{+}97]. Bligh [Cra27]. Board [Pla33g]. Bodenstein [Pla25c]. Bodies [Pla16d, Pla19e, Pla22d, Pla31b, Pla32g, Pla49e, Pla57a, Pla64b, Pla80, Pla16f]. Body [MKD^{+}91, Kir98a, Kuh78, Kuh87, Lon03a, Lon03b, Pla12b, Pla12a]. Boethius [Thi71]. Böhme [Ano36]. Bohr [Adl03, Bro36, Pla19c, Ano23, HN70, MR82d, MR82e, Pla19c, Pla23c, Pla24c]. Bohrsche [Pla23c]. Bohrschen [Pla24c]. Bok [Wor63]. Boltzmann [BFG81, Kör61, Bad09, Boy79, Koc91, Kör61, Sch00a, Yos95a]. Boltzmannstraße [Kra97]. bomb [Cas09, Hop96]. bombs [Ano46c]. Book [A.50, Ano32, Ano14, Ber14, BDR86, Cah99, Cas86, Cra27, Dou49, Fal50, Fra25, Gea09, HB87, Hen60, Her03, Hib91, Hie68, Klo63, Koy52, Len49, McK27, Mic68, Mul00, Ole81, Par12, Pye88, RS49, Ros12, Rus50, San87, Sch04, Sta10, Swe89, Thi71, Tre01, You68, Hei35, Hei43a]. booklet [Pla33j]. books [Hop96]. Bosch [Her03, Koh02, Koh02]. Bose [JDB^{+}97, FMM96, JDB^{+}97]. bound [Sti91]. boundary [Pla30g, Pla31a, Pla31f, Pla33b, Pla33e]. bouquet [Her01]. Boyer [Sac84]. Bragg [Ole81]. Brain [Bre08]. Braun [Pla34c, Pla34a, Pla34b, Pla34c]. Braunschweig [Ebe63]. breakthroughs [Lig05a, Lig05b]. Breeding [HH63, MAS63]. Breitkopf [Pla93d]. brief [JDB^{+}97]. brief/Max [JDB^{+}97]. brief/student [JDB^{+}97]. Briefe [Klo63, Prz63]. Briefen [Eck97b]. Briefmarken [Ano53c]. Brieftagebuch [Tob08, HT99, Sch00b, Mul00, Sch04]. Briefwechsel [Kör61]. Brillouin [Pla88e, Pla88f, Pla89q]. Brockhaus [BPL^{+}54]. Brockhaus/Berichtigung [BPL^{+}54]. broke [BA01]. broken [Bro15]. Brookes [HKS^{+}68]. brosch [Ano39]. broschiert [Sch00b, Som49]. Brought [Ano46d]. Bruhat [BBG^{+}58]. Brüssel [Pla13e]. Brussels [Pla13e]. Bruxelles [LdB12]. Buchanan [Pla88c]. Buchbesprechung [Jac02, Jac08a]. Buchbesprechungen [BDR86]. Bücher [BBG^{+}58, BBG^{+}58]. bumpy [Bad15]. Butterflies [Ano46c]. buys [Ano14].

C [Adl03, BDR86, Pla89e, Pla93b, Pla93d, BDR86]. C. [Pla93b, Pla94b]. Cabinet [HRR10b]. calculate [Hol13]. Calculation [KZR^{+}73]. California [San87]. called [Din39]. Can [Cha97, Bra14b, Cla89, Pla42f, MKD^{+}91]. canonical [Pla08e]. can’t [Hof08a]. Carathéodory [SCP19]. Carbon [FSBR09]. career [Lem08]. Carl [HT99, Her03, Koh02, Mul00, Pla07b, Pla15c, Sch00b, Sch04, Sch55, HT99, Koh02]. Carnof [Pla88d]. Carnofsehen [Pla89t]. Carnot [Pla92a, Pla92a]. case [Lüs00, Pic25, Rub12, Weg10]. Cassidy [Hop96]. Cast [Ano29a]. Catastrophe [Tib13]. causal [Pla32a]. causality [Pla23g, Pla32c, Pla32d, Pla32e, Pla37c, Pla41a, Pla48a, Pla52a, Pla53a, Pla55a, Pla55b, Pla58b, Pla65a, Pla67a, Pla33c]. Cautious [Bru02]. cavity [Iro05, Pla07d]. cavity-mode [Iro05]. celebrate [Pla15a, Pla24a, Pla28a, Pla29k, Pla29l]. Celebration [Hof96, Hof99, Brü38, Pla17a, Pla22b, Wei58, Pla38g]. cemetery [Ham72]. Censorship [Sch22]. Centennial [Hof99, Hol99]. Central [Ano11]. Century [Ano47c, Bra09, Eck97b, Hof48, Lig05a, Lig05b, Yos95a, Sch96]. ceremonial [Pla35e]. Ceremony [Ano46d, Ano48d, Haa75, Pla14d, Pla27a]. Cermak [SH53]. Ch [BT58, DT58, Pla88b, Pla89h, Pla89f, Pla89g, Pla98a]. Ch. [Pla89a, Pla89b]. Chair [AH93]. Chairman [Pla22f]. change [Bre08]. Changeable [Hui13]. changed [MAS63]. channel [SNE01]. chapter [Pla39c, Pla52d, Pla53d, Pla55e]. Characteristics [Ano21]. Charge [Luc30, JDB^{+}97]. Charles [Ano48j]. Chatelier [Pla89p, Pla34a, Pla34b, Pla34c, Pla34a, Pla34b, Pla34c]. Chemical [Pla08a, Pla21a, Pla22a, Pla88a]. Chemie [Ber15, BBG^{+}58, Hei69, Pla88b, Ano48i]. Chemie/Becker [Ano48i]. Chemie/Spitzer [BBG^{+}58]. Chemiker [Ano46b, Ano46c]. Chemiker/Das [Ano46b, Ano46c]. Chemisch [Cah99]. Chemisch-Physikalisch-Technischen [Cah99]. chemische [Pla88a, Pla21a, Pla22a]. chemischen [Pla89p]. Chemischer [BBG^{+}58, Ano02]. chemist [Ano46c]. Chemistry [Cra84, Par12, Ber15, Hei69, KZR^{+}73]. Chemistry/Max [KZR^{+}73]. Chemistry/Max-Planck-Medaille [KZR^{+}73]. chest [Iva59b]. chimique [Pla08a]. choice [Pla70]. Christentums [Löh12]. Christianity [Löh12]. chromodynamics [Lüs00]. ciência [FV12]. cinematography [KGSO55]. city [Pla10a, Ano21]. Clara [Wes57]. Classic [Ped75]. Classical [Nee80, Boy83, BRS03, CG00, CG03, CC03, Cha97, Kra76, TP76]. Clausiu [Pla05b]. Clausius [Pla92a]. Clausius’sche [Pla92a]. Clausthaler [HD48, HD48]. clears [Bre08]. Clerk [Pla31c, Tho31]. Climate [Bre08]. clock [Ano11]. close [Kir98b]. closing [Kut14b]. cloth [Pic25]. Club [Bra14b]. cm [C.49]. Co [Ano13, Cra27, Jüt11, Jüt18b, McK27, Sch65]. Cobalt [VD47]. cognition [Pla10i]. coin [Ano70]. Colleagues [Hof08f]. Collected [Ano58, BDR86, KKRS96]. Collection [Fra25, Iva59a, Pla22h, Pla25d, Mic68, Pic25]. College [Ano21]. Collision [KZR^{+}73]. Collocation [Sch17]. Columbia [Pla10a, Pla09e, Pla10a, Pla15g, Pla12a, Pla15e]. combined [Bec09]. Coming [Hop96]. Commemorating [Ano08e, MH97, Her72]. Commemoration [Pla28c, JDB^{+}97, Tho31, Ano49a, Brü48a, HD48]. commemoration/WE [JDB^{+}97]. commemoration/WE-Heraeus-program [JDB^{+}97]. commencement [Pla14f, Pla16e, Pla23h]. Comment [BF12, Boy84, BFG81, CG03, CC03, Kra76, Sac84, TP76, Pla15b].

Comments [Alb93, Kir98b, Pla00a, Pla35b, Spr89, Wes40]. Commission [BT58, DT58]. communication [Pla97c, Pla97d, Pla97b, Pla98c, Pla99c, Pla15d, Pla15e, Pla17d, Pla17e, Pla29a, Pla29n, Pla31f, Pla33e, Pla41c]. Company [Pic25]. Competes [Bla88]. complete [Pla59a]. completion [MR00]. Compression [Pla89a, Pla89f]. comprimirt [Pla89g]. compromises [Hei07]. Computing [Luc30, KZR^{+}73]. concept [Pla32c, Pla32d, Pla32e, Pla37c, Pla41a, Pla48a, Pla52a, Pla53a, Pla55a, Pla55b, Pla58b, Pla65a, Pla67a]. concepts [Lon84, Lon03c]. Concerning [Nag82]. condensate [JDB^{+}97]. conductive [Pla05a]. conductors [Pla89y]. conférence [Pla08a]. confinement [JDB^{+}97]. conflict [Hof08f]. confusing [Ano53b]. congratulations [SR37]. Congress [KGSO55, Pla13e, Str11a]. Congressitus [Ano59]. connection [Pla13h, Pla14g]. conoscenza [Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a]. conscience [Hei88c]. consensus [Lev17]. consequences [Mun36]. conservation [Pla87a, Pla08c, Pla13b, Pla21b, Pla15b, Weg10]. Conservative [CM08]. Constant [Ano12, BF12, Bir19, Cha14, Gra69, Has14, Hol99, Inf61, KM12, Lon45, Luc30, Men17, Noe73, VD47, WB78, Why50, WSNO98, Ano47c, Cha97, FGH91, MR98, Mül08b, Peg77, Pla06b, Pla15c, Pla21a, Pla22a, Sti91]. constante [Pla89g]. constantem [Pla89g]. Constants [RB01, Gré19, Mey08, Qui17]. constitution [Pla87b, Pla87f, Pla88l, Pla09d]. Constructions [Set03]. Contemporary [Sch22, CH01, Pla13c, Pla13d, Pla14c, Thi69]. content [OT97, Pla10h]. Context [Hof01, Bro83, Sin08]. continuità [Ros08]. Continuities [Dar00]. Continuity [Koc91, Ros08]. contraction [Pla10f]. contradiction [Pla00b]. contro [Ros12]. control [KGSO55]. controversial [Wol08]. controversy [Thi69]. conversion [Gor91]. cooling [JDB^{+}97]. cooperation [Wol08]. cooperative [Wol08]. Copley [Ano29c]. core [Pla93a]. corps [Pla09c, War12]. corpuscular [Pla40b, Pla40c, Pla41c]. Corrected [BT58, DT58]. Correction [Din46b, Ein07a]. corrections [Par94]. Correspondence [Pet14, Hof08b, Kör61, Prz11]. cosmic [Bre08]. Cottam [KAR89]. couple [Bad09]. Courageous [GSU07]. Couragierte [GSU07]. Courant [BBG^{+}58]. Crazy [Arr06]. creation [Kan70]. creative [Shp58a]. Creativity [MKD^{+}91]. creators [Str11b]. Cremer [GBM^{+}49]. crisis [Pla29g, Yos95a]. critical [Pla09f, Sch78c]. criticism [Sch91, Vog61, Yos95c, Pla00d]. crossroads [ET07]. cuánticas [W^{+}87, WdC02]. Cuestiones [W^{+}87, WdC02]. Curies [Kae48]. current [Pla20b, Pla20c, Pla22c, Pla58e]. curriculum [Pla20b]. cycles [Pla05b].

D [BPL^{+}54, Hib91, Kra97, Pic25, Pla89k, Pla89l, Pla06f, Pla20b, Pla23g, Pla38g]. D-14195 [Kra97]. d’action [Pla12a, Som12]. Daeves [GBM^{+}49]. Dagmar [Cah99]. Dahlem [Kra97, Ano36, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. damped [Pla96c]. dampened [Pla97e]. Dampf [Pla89w]. Dämpfe [Pla89g, Pla89b, Pla89h, Pla98a]. Dampfes [Kir98b, Pla89g].

Dampfspannung [Pla88i, Pla88k, Pla88h, Pla98a]. Dämpfung [Pla97a]. damping [Pla97a]. Dankworte [Pla27a]. dans [Pla63a, Rub12]. daraus [Cla89]. dargebracht [Pla30f]. dargestellt [Ano39, Ano34]. dark [Kra12]. Darstellende [GBM^{+}49]. Darstellung [Sch78c]. Darwin [Ano48j, DdCA48]. Dasselbe [Pla89p, Pla89p]. date [Pic25]. David [Hop96]. d’azione [Ros08]. dazu [Jüt22, Pla89o]. De-anthropomorphizing [Weg10]. Dead [Ano47c, Ano47d]. Dear [KSB63]. Death [Ano47a, Ano97a, Ano97b, Eck97a, Ham72, MH97, Pla57d, RH97, Wes57, Ano46c]. debate [Löh12]. Debatte [Löh12]. Debye [Pla37f, Pla19c, Pla37f]. Debyeschen [Pla19c]. decades [Haa70]. December [Pla09c, Pla29e, Pla29f]. déchirée [Hei88c]. Deeper [Bla92]. defiant [Bra14b]. Definition [CS79b, Pla25b, Ros79, Ros59a, Ros59b]. deflection [Pla06d, Pla06e, Pla07a, Pla07c]. Deformable [Pla16d, Pla19e, Pla22d, Pla31b, Pla32g, Pla49e, Pla57a, Pla64b]. Deformationen [Pla88f, Pla89q]. Deformierbarer [Moo23, Pla16d, Pla19e, Pla22d, Pla31b]. degrees [Pla15d, Pla15e]. deinen [MKD^{+}91]. del [Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a, W^{+}87, WdC02]. Delivered [Pla09e, Pla12a, Pla15g, Pla15e]. Democratic [WC91]. demone [Ros12]. Denkens [JDB^{+}97]. Denker [BPL^{+}54, Har38, Har43, Har48, Har53a, Har53b, Har58, Har64a, Har64b, Hen60]. Denker/Morrison [BPL^{+}54]. Denkmal [Uns58]. Denkmal/Kongressitis/Internationale [Ano59]. density [Pla03e, Pla05a]. dependence [MR98]. depleted [Pla30g, Pla31a, Pla31f, Pla33e]. Derivation [Boy83, SNE01, Boy84, Pla01g, Pla23b, Pla95, Pla97, PHR07, Sac84]. derivations [Kra76, TP76]. Derived [Din46c, Foo16]. deriving [Iro05]. derniers [Pla60a, Pla91a]. Description [Bro36]. Design [Kut14a]. Design-Fehler [Kut14a]. Designer [Kut14b]. Designierten [BWN78]. Desk [Pla30b]. Despair [Büh98a, Dar00]. desselben [Kir98b]. Determination [Men17, Mil16, VD47]. Determinism [Pla38e, Pla47a, Pla48b, Pla52b, Pla53b, Pla55c, Pla58c, Pla65b, Pla67b]. Determinismo [Pla47a]. Determinismus [Pla38e, Pla48b, Pla52b, Pla53b, Pla55c, Pla58c, Pla65b, Pla67b]. Deutsche [Ano46b, BWN78, HW07, vL58]. Deutschen [Ano46b, Ano48a, Ano48c, Koy52, Pla18a, Pla36c, Pla39c, Pla48f, Pla52d, Pla53d, Pla55e, Sch49, Som49, WvLuAE18, Pla29e, Pla29f]. Deutscher [Ber60, Ebe63, Pla30f]. Deutschland [Ano46b, Eck97a, Sti84, Pla31c]. Deutschland/Amerikanischer [Ano46b]. Deutschland/Max [Ano46b]. Deutung [Pla16a]. developed [Pla03c]. Development [Ber14, Pla20d, Pla22g, Pla10a, Cra27, Fal59, Gri87, JDB^{+}97, Pla91a, Pla20b, Pla20c, Pla22c, Pla58e, Cla91]. Developments [HKS^{+}68]. Deventen [Pla88b]. Dezember [Pla29e, Pla29f, Pla09c]. Dialogue [EM32, Ein77]. Diamond [JDB^{+}97]. diary [HT99, Tob08]. diatomic [Hol13, Pla15c]. Dibble [KZR^{+}73]. Dichte [Pla89w, Pla03e, Pla05a]. dictatorship [Hof00].

Did [Har59]. died [Ano47a]. Dielectric [Gra69]. dielectrics [Pla03h]. dies [Ano47e]. dieses [Pla06a]. difference [Pla90b, Pla90e, Pla27e, Pla29a, Pla29n]. different [Pla80]. Differential [Sch17, GBM^{+}49]. Differential- [GBM^{+}49]. difficulties [MR82d, MR82e]. Diffusion [Pla92c, Pla35f]. Diktatur [Hof00]. dilemma [Gré19]. Dilemmas [Cas86, HB87, Hop96, Pye88, San87, Swe89, Hei86a, Hei88b, Hei00]. dilemmi [Hei88b]. dilute [Pla87b, Pla87f, Pla88a, Pla88i, Pla88k, Pla88j, Pla88l, Pla90b, Pla90e, Pla27e, Pla29n]. diluted [Pla29a]. dimensional [Bae01]. DIN [Som49]. Dinge [Hei69, MSQM63]. Dingler [Ano41b, Mos42]. Dipole [Fok14, Pla15h]. dipoles [Fok14, Pla15h]. Dirac [Ano52]. Direct [Mil16]. Dirk [Cah99, Kra97]. Disagreements [Dar01]. Discipline [WC91]. disciplines [Sti84]. Discontinuities [Dar00]. Discontinuity [BDH^{+}00, Kuh78, Kuh87]. Discovered [Ano47c]. discoveries [Bra09, Lig05a, Lig05b, Seg07b]. discovery [Ano46c, Hei58, Kae48, MR82a, Pla37i, Pla43c]. Discusses [Ano23]. Discussion [Bla78, Nag82, Pla00a, Pla09d]. discussions [LdB12, RS80]. disintegration [Fis07, Fis10a]. Diskussion [Pla09d]. Dispersion [Pla02d, Pla05a, Pla04a, Pla07g, Pla08g, Pla04a, Pla07g, Pla08g]. displacement [Pla06b]. Dissertation [OT97]. Dißmann [Kra03]. dissociation [Pla88j, Pla02e, Pla19c]. Dissoziation [Pla88j]. Dissoziationstheorie [Pla02e]. Dissoziationswärme [Pla19c]. Distribution [Pla00i, Pla01d, Pla69b, Pla78a, Ano48a, Boy79, BFG81, Kan70, NK06, Pla94c, Pla95b, Pla01a, Pla02b, Pla15h, Pla22i, Sch78d]. Disziplinen [Sti84]. Divided [Hof99]. DM [Ano89, Dri01, Ebe63, Sac69, Sch49, Sch00b, Sch65, Som49, Wei68, You68, Ber60, Sch55]. do [FV12]. Dobzhansky [Adl03]. doctor [Pla29k, Pla29l]. doctorate [Pla29b, Sch29]. Doctors [Hop96]. Doctrine [Ano37]. document [Her76]. documentary [HMF98]. Documents [Bla92, Her73a]. Doga [Pla15g]. Doktorjubiläum [Pla29b, Sch29]. Doktorjubiläums [Pla29k, Pla29l]. domaine [Rub12]. d’onde [Rub12]. Dörge [GBM^{+}49]. Dr [Ano02, Ano29b, Ano30b, Ano34, Ano36, Ano41b, Ano41a, Ano46b, Ano46c, Fra31, Her24, McK27, Mis33, Mos42, Pla19d, Pla15c, Sch55, Wil37]. Dr. [Ano23, Ano29a, Ano46c, Pla07b, Pla08d]. Drei [Ano69]. Dreistellige [GBM^{+}49]. Dresden [Pla29e, Pla29f, Pla29e, Pla29f]. Dritte [Ano36, Pla97b]. Dritten [HW07]. Dritter [Pla37a, Pla15c]. driven [Bro15, SDF14]. Drohen [Bre08]. Druck [Pla03e, Pla90a]. Drucken [Pla89g]. Drude [Pla06a, Pla06f, Pla06a, Pla06f]. Dtsch [Pla06f, Pla38g]. Duhem [Pla89p, Pla89i, Pla89r, Pla89s]. Durch [Ano46b, Ano49b, Ano49c, JDB^{+}97, Kan69, Pla95a, Pla96a, Pla96c, Pla97e, Yos95a]. durchstrahlten [Pla01g]. durée [Wal94]. Dürfet [BPL^{+}54]. during [Hac07, Kir79]. Dutch [van48]. Dutton [Pic25]. duty [Fis08b]. Dynamic [Pla14d]. Dynamics [Nee80, Pla7 , MR98, Pla06d, Pla06e, Pla07f, Pla07a, Pla08a, Pla08b, Pla08f, Pla09a, Pla17b, KZR^{+}73]. Dynamics/Lipson [KZR^{+}73]. Dynamik [Pla06d, Pla06e, Pla07f, Pla07a, Pla08f, Pla17b, Pla08a, Pla08b, Pla09a].

Dynamische [Jüt17, Pla14d].

E. [Klo63]. e.V [Ebe63]. each [Pla14i, Pla14j, Pla25f]. Early [Kan76]. Ears [Hui13]. Easter [Pla11c, Pla11d]. Eckener [Ano46b, Ano46c, Ano46c]. écrits [Pla60a, Pla91a]. Ed [Hie68, Pla15c, Ber14]. Eddington [Adl03]. edited [Mic68]. Edition [Ber14, Tre01, Cla91, McK27, Ber14]. Editor [Hof08a]. Editors [Ber60]. edits [Pye83, Pye85a, Pye08]. Eduard [Pla15f]. Edward [Cra27]. Effect [Har98, Pla00b, Pla13e]. Effecte [Pla89m]. effective [CG00, CG03, CC03]. effects [Pla15i]. Effektes [Pla00b]. Ehren [Som38b]. Ehrenmitglieds [Pla38g]. Ehrung [Ano46b]. Eigenschaften [Pla03g]. Eight [Pla09e, Pla15g, Pla98, Pla12a, Pla15d, Pla15e, Pla10a]. Eighth [Pla7 ]. Eightieth [Kem38, Pla38a]. Eighty [Hei38]. Einakter” [HRR10a, HRR10b]. einatomiger [Pla08e, Pla16f, Pla25h]. einem [Ano34, Ano36, Ano39, Ano41a, Pla00f, Pla02c, Pla04a, Pla07d, Pla15h, Pla39c, Pla52d, Pla53d, Pla55e, Sch55, Wil37]. Einfache [Pla88b]. Einfluss [Pla89s]. Einführung [Fra31, BBG^{+}58, Fra31, HKS^{+}68, Moo23, Pla16c, Pla16d, Pla19e, Pla21c, Pla22d, Pla22e, Pla30a, Pla28b, Pla31b, Pla37e, Pla15f]. Einführungsmathematik [Ano48i]. Einheit [Pla09b, Pla09c]. einige [Kir98b, Sch65, Hah57, Pla89r]. einiger [Pla88b]. Einigkeit [KSB63, MSQM63]. einleitende [Ano11]. Einstein [Adl03, Ano69, Hie68, Hof08f, JDB^{+}97, Klo63, Mic68, Pla09d, Prz63, Prz86, Sim06, Acz01, Acz03, Ano21, Ano29a, Ano69, Bla88, Bre08, BRS03, FMM96, Gut15, Hof08f, JDB^{+}97, KKRS96, Lev04, MR82d, MR82e, Mun36, Pla09d, Prz11, Pye85b, RS80, Sch00a, SDF14, Sim06, Sto13, Yos95a]. Einstein-Kondensat [JDB^{+}97]. Einsteins [Ano69]. Einwand [Pla92b]. Eismassen [Bre08]. Eispunkt [Kir98b]. Eitz [Pla93d, Pla93b]. Elaboration [Gol76]. electric [Pla95a]. electrical [Fok14, Pla96a, Pla96c, Pla97e, Pla97a]. Electricität [Pla90c]. electricity [Pla90d, Pla90c, Pla98b, Pla99a, Pla01e, Pla22e, Pla32h, Pla35g, Pla35h, Pla49h, Pla57b, Pla78b, Pla93c]. electrische [Pla97e]. electrischer [Pla96a, Pla97a]. electrodynamics [Boy83, Pla15i]. electrolysis [Pla92c]. electrolyte [Fal59, Pla31a, Pla90e]. Electrolyten [Pla90c]. electrolytes [Pla90d, Pla90c, Pla90b, Pla90e, Pla02e, Pla30g, Pla31f, Pla33b, Pla33e, Pla35f, Pla35g, Pla35h]. Electromagnetic [Ano21, Boy83, FMM96, Pla02d, Pla03h, Pla23b, Pla95, Pla97, PHR07]. Électromagnétisme [Pla39a]. Electron [Luc30]. electronic [Pla19d]. electrons [Pla06d, Pla06e, Pla07a, Pla13i]. Elektricität [Pla01e]. elektrische [Pla96c]. elektrischer [Fok14, Pla95a]. Elektrizität [Pla90d, Pla99a, Pla22e, GBM^{+}49, Pla98b, Pla78b]. Elektrizitätserregung [Pla35g, Pla35h]. elektro [Pla03h]. elektro-magnetischen [Pla03h]. Elektrodynamik [Pla15i]. Elektrolyse [Pla92c]. Elektrolytarbeiten [Fal59]. Elektrolyte [Pla31f, Pla33e, Pla90b, Pla02e, Pla30g]. Elektrolyten [Pla90d, Pla35f, Pla35g, Pla35h]. elektromagnetischen [Pla02d, Pla23b, Pla95, Pla97, PHR07]. Elektronen [Pla13i, Pla06d, Pla06e, Pla07a]. Elektronenbahnen [Pla19d]. Elektronenmikroskopie [HKS^{+}68]. Elektronenmikroskopie/Brookes [HKS^{+}68]. Elektronenoptik [GBM^{+}49]. Elektronenoptik/Hahn [GBM^{+}49]. Elektronenstrahloszillographen [GBM^{+}49]. Elektronenstrahloszillographen/Cremer [GBM^{+}49]. élément [Som12]. élémentaires [Pla12a]. Elementare [KZR^{+}73]. elementaren [Pla13e]. Elementarquanta [Pla01e, Pla78b]. elementary [Pla01e, Pla13e, Pla78b, Pla03c]. Elements [Ano23]. Elfenbeinernen [Jos95]. elliptically [Pla00f, Pla02c]. elliptisch [Pla00f, Pla02c]. Elster [Ber15, MKD^{+}91, Ber15]. emergence [Hof08d, Pla20b, Pla22c, Pla20c, Pla58e]. Emergency [Pla29e, Pla29f]. Eminent [Ano47d]. Emission [FSBR09, Kir98a, Pla95a, Pla96a, Pla15b]. emissions [Pla11f]. emitted [Pla00f, Pla02c]. emittierte [Pla00f, Pla02c]. Emmissionsvermögen [Kir98a]. Emperor [Pla17a]. Employed [Ano23]. endgültige [Sch78a]. endothermische [Pla89t]. energetics [Pla96b]. Energetik [Pla96b]. Energie [Fok14, Hol13, Pla01a, Pla02b, Pla22i, Pla87a, Pla00f, Pla02c, Pla08c, Pla13b, Pla21b, Pla15b, Pla88g, Pla11c, Pla11d]. Energien [LMOS15]. Energies [LMOS15]. Energieschwankungen [Pla23d, Pla23d]. Energieverteilung [Ano48a, Kan70, Pla00i, Pla01d, Pla15h, Pla78a, Thi71]. Energieverteilungsgesetz [Sch78d]. Energy [Pla00i, Pla01d, Pla69b, Pla78a, Ano48a, Fok14, Hol13, Kan70, Kra12, Pla87a, Pla00f, Pla01a, Pla02c, Pla02b, Pla08c, Pla13b, Pla15h, Pla21b, Pla22i, Pla23a, Pla15b, Sch78d, Sin08, Weg10, Pla11c, Pla11d, Pla23d]. Engelmann [Ano02]. Engineers [HKS^{+}68]. English [McK27]. Enrico [KGSO55, KGSO55]. Entanglement [Acz01, Acz03]. Entdeckung [Ano46b, Hei58, Pla37i]. enthalten [Pla89g]. entire [Ano14]. entitled [Duh88]. entrepreneurship [JDB^{+}97]. Entropie [Pla87c, Pla87d, Pla87e, Pla91b, Pla16f, Pla21a, Pla22a, Duh88, Ros59a, Ros59b, Pla00c, Pla04b, Pla25b]. Entropiedefinition [Pla25e]. Entropiekonstante [Pla15c]. Entropy [CS79b, Kle63, Pla03b, Ros79, Duh88, Pla87c, Pla87d, Pla87e, Pla91b, Pla04b, Pla05b, Pla15c, Pla16f, Pla21a, Pla22a, Pla25b, Pla25e, Ros59a, Ros59b, Pla00c]. Entstehung [Gea09, Hof08d, Jüt21, Pla20b, Pla22c, Sti84, Pla20c, Pla58e]. entwickelt [Pla89c]. Entwicklung [Cla91, JDB^{+}97, Jüt21, Pla91a, Pla20b, Pla20c, Pla22c, Pla58e, Fal59]. entwined [Her01]. environment [PS33]. Epilogue [EM32, Ein77]. episodes [Bra09]. Epistemic [Ber14]. Epsilon [Hop96]. Equation [Has14, Lew30, Pla69a, Pla00g, Pla08e, Pla78c, Ris96, vK97]. Equations [Sch17, MR82c, Pla06c, SNE01]. equilibria [EH08b]. equilibrium [CG00, CG03, CC03, Pla83, Pla88a, Pla08b, Pla80]. era [Hac07, Hop96]. Erckmann [BBG^{+}58]. Erde [GBM^{+}49]. Erde/Becksmann [GBM^{+}49]. Erdgeschichte [GBM^{+}49]. Erdgeschichte/Kropp [GBM^{+}49]. Erforschung [Fis10b, Fis12]. Erhaltung [Pla87a, Pla88g, Pla08c, Pla13b, Pla21b, Pla15b]. erhobenen [Pla92b]. Erinnerungen [Hah57, Pla35e, Pla46b, BFV^{+}48, Dri01, HR01, Jac02, Kra02a, Kra02b, Pla30b, Pla49k, RH01, Len49]. Erinnerungsschrift [Ano48c, Koy52, Pla48f, Sch49, Som49]. Erkenntnis [Ano32, Mis33, Pla10i, Pla13h, Pla14f, Pla16e, Pla23h, Pla33h, Pla43b, Pla44b, Pla45e, Som33, Zil33, Hei43a, Pla34e, Reg34, HPE^{+}33]. Erklärung [Pla89o]. Ernest [Ole81]. Ernst [Ano34, Ano36, Ano39, Ano41a, Ano47e, Ano02, Ano08d, Bor48, Bor63b, Hai11, Kan08, Sac69, Thi69, Wil37, van10, Bla92, FS06, Hui13, Thi69, Weg10, Yos95b, KSB63]. Eröffnet [Ano46b]. Eröffnet/88 [Ano46b]. Eröffnung [Ano11]. Eröffnungsrede [Pla22f]. erregt [Pla96c, Pla97e]. Erregung [Pla90d, Pla90c]. ERS [Sch00b]. ERS-Verlag [Sch00b]. ersetzen [Ano70]. Erste [Ano46b, Pla97c]. Erstes [JDB^{+}97]. Erstveröffentlichung [Her76]. erweitert [Sch65]. erweiterte [Ano39, Ano41a, Jüt11]. Erweiterung [Pla88d, Pla17b]. Erwiderrung [Pla13f]. Erwiderung [BP24, EP06, FP27, MP19, PP25, PP79, Pla29c, SCP19, SP29, Sch61, WvLuAE18, vLP21, Pla92b, Pla03f, Pla10i, Pla18a, Pla23e, Pla25c, Pla37f, Pla38f]. Erwiderungen [Kir79]. Erwin [Adl03, Bro36, KK75, Mic68, Pla28d, Prz86, Wes57, Ano11, MR87, Pla28d, Prz11, Wes57]. erzählt [FS06]. escritos [W^{+}87, WdC02]. especial [Pla03c]. Essay [Hop96]. Essays [Fra25, Pla22h, Pla25d, Sch96, Cla91, Pic25, Pla28d]. Essen [KGSO55]. establishment [Din39]. Esteemed [Ano43b]. etc [Ein07a]. etc. [Ein07a]. Ether [Ano21, Pla01a, Pla02b]. Euer [Tob08]. Eulogy [van10, Haa75]. Europa [JDB^{+}97]. Europa/Quasiteilchen [JDB^{+}97]. Europäischer [Ano48i]. Europe [JDB^{+}97]. Europe/quasiparticles [JDB^{+}97]. Evaluation [Has14]. Evaporation [Pla82]. Even [Tib13]. events [Pla33c]. Ever [Fra05]. Everything [Law15]. Evidence [Noe73, Pla94c, Pla95b]. Evolution [Kut14a, Cra27, Kut14a]. Ewald [Hos78]. Exact [Pla49d, Pla42d, Pla42a, Pla42e]. Exactly [Tib13]. exakten [Ano02, Ano29b, Pla42d, Pla42a, Pla42e]. excerpt [Pla48d]. excitation [Pla90d, Pla35g, Pla35h, SDF14]. Excitations [KAR89]. excited [Pla96c, Pla97e]. excitement [Pla90c]. Exhibition [Uns67, Ano11, Bad08, Ano97a]. Exiles [Fra05]. expandierende [Bre08]. expanding [Bre08]. Expansion [Pla89a]. expansions [Sti91]. Expelled [MP01]. Experiment [Ano46b, Tib13]. Experimental [Hof01, Her58, BPL^{+}54, War12]. expérimentale [War12]. Experimentalphysik [WF63]. experimentelle [Her58]. Experimenting [Sch03]. Experiments [Pla89o]. expert [Ano47e]. Explain [Ano23]. Explains [Ano21]. Exploring [BRS03]. Exposition [Ano21]. extension [Pla17b]. external [Pla31e]. extinction [Pla04a]. Extinktion [Pla04a]. Extraordinary [Gib19]. eye [Bre08].

F [Ano48i, BPL^{+}54, Hei31, Pla89j]. FA [KGSO55, KGSO55]. face [Hof08a]. Fachhochschulen [JDB^{+}97]. Fachhochschulen/Einstein [JDB^{+}97]. Facsimile [Ano48a]. Facts [Law15, Mes07]. Faculty [Pla09c]. Failure [Nee80]. Faith [Pla30i, Pla88b]. Faithfully [Tob08]. Faksimile [Ano48a]. fakultät [Pla09c]. Falk [MKD^{+}91]. Fall [Lüs00, Men73]. Fallacy [Ano21]. Familie [vP06]. family [vP06]. famosos [W^{+}87, WdC02]. famous [W^{+}87, WdC02]. far [CG00, CG03, CC03]. Faraday [Pla00b, Pla00b]. Faraday-Effektes [Pla00b]. Farben-Fernsehen [Ano46b]. Farkas [Pla89n]. Farm [Hop96]. Fast [KSB63]. fate [HS97]. Favor [Bla88]. Feast [Som38b]. Februar [Pla94a, Pla36c]. February [Pla94a, Pla36c]. Fehler [Kut14a]. Feier [Hof96, Jüt17, Jüt18a, Pla14d, Pla15a, Pla17a, Pla22b, Pla24a, Pla28a, Pla29k, Pla29l, Pla35e, Pla38g]. Feld [BPL^{+}54]. Felde [Pla01g]. Fermi [KGSO55, KGSO55]. Fernsehen [Ano46b]. Ferromagnetism [HD48]. Ferromagnetismus [HD48]. Ferromagnetismus/Clausthaler [HD48]. Fest [Brü38]. Festen [Ano48i]. Festessen [Som38b]. Festival [BDR86, HRR10b]. Fest” [HRR10a, HRR10b]. Festrede [BDR86]. Festsaal [Ano11]. Festschrift [Ber15, GSU07, Pla30f, Ber60, Ber15, KMP59, Pla30f]. Festsitzung [Pla35e]. Festvortrag [Ano11]. Few [MKD^{+}91]. Few-Body [MKD^{+}91]. Fiat [Ano48i]. Ficker [FP27]. field [Fok14, Haa70, Pla01g]. Fields [Ano21, Ber15]. Fifth [Pla99c]. fiftieth [Sch29]. Fifty [Kae48, SB51, Pla29b, Pla30f, Pla62]. fifty-years [Pla29b]. Figuren [Ano13, Fra31, Jüt18b]. film [Pla42c]. filmed [Bre08]. Filmportrait [Pla42c]. filozoficznych [Pla70]. final [Lon84, Sch78a]. final-year [Lon84]. Fine [Why50]. Fine-Structure [Why50]. firm [Ano46c]. First [Ano16, Ano29a, JDB^{+}97, Tib13, Her76, Pla36c, Str11a, Pla87c, Pla97c, Pla15d, Pla17d, Pla22f]. Fischer [FS06, Jac08a]. fisica [Ros08]. fisico [Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a]. físicos [W^{+}87, WdC02]. Fission [Sim12]. Fits [Tib13]. Fizicheskaja [Pla26e]. fizik [Pla29o]. fizycznego [Pla70]. flaw [Kut14a]. Flawed [Tib13, Iro05]. Flöttmann [Ano48i]. flüchtiger [Pla88i, Pla88k, Pla88h]. fluctuations [Pla23d]. fluid [Pla84]. Flüssigkeit [Pla88c, Pla89w]. Flüssigkeitsstrahlen [Pla84]. Focus [Hol99]. Fokker [vK97, Ris96, vK97]. Forderung [Cah99, Ebe63, Hei31, Pla31e, Pla37a]. foreword [Cra27, Pla34d]. Forgetting [Sim12]. Formation [WC91]. Formelsammlung [Ano48i]. formerly [Pla60b]. Formula [Ano23, Din46c, Din46b, Din46a, Ken46, Boy83, CG00, CG03, CC03, Rub12, Seg07a, War12]. formulation [MR82b, Pla14e]. formule [Rub12, War12]. Formulierung [Pla14e]. Forscher [Ein13]. Forschung [Ano46b, Ano46c, BPL^{+}54, GBM^{+}49, GBM^{+}49, Pla26b]. Forschung/Der [BPL^{+}54]. Forschung/Goethe [Ano46b, Ano46c]. Forschung/Goethe-Preis [Ano46b, Ano46c]. Forschungsunternehmen [Ano46b, Ano46c]. Fortbildung [BWN78]. fortunes [Hei00]. forward [Frö01]. foster [Bra14b]. Foundation [Hof01, MR82d, MR82e, Pla03f, Pla03c]. Founder [Mas12, Bad08, Pla14d]. Fourth [Pla98c]. fractional [JDB^{+}97]. Frage [Pla25h]. Fragen [Ano69]. Fragen/Max [Ano69]. Franck [Kas64]. Franck/Wilhelm [Kas64]. Frank [Ber60, C.49, Dou49]. Franklin [Pla27b, Pla27b]. Franklin-Institut [Pla27b]. Franz [Hib91, Gre90]. Französ [Pla11c, Pla11d]. Frau [Mei03].

Fraunhofer [Kir98a]. Fraunhoferschen [Kir98a]. Frederick [Pla17a, Pla24a, Pla28a, Pla37b]. Free [Ano37, Boy79, BFG81, Gri98, Pla13i, Pla22i, Pla23g, Pla36a, Pla36c, Pla37h, Pla39c, Pla45d, Pla48e, Pla52d, Pla53d, Pla55e, Pla90b]. freedom [Pla15d, Pla15e]. Freezing [VD47, Kir98b]. freie [Pla22i]. freien [Pla13i]. Freiheitsgraden [Pla15d, Pla15e]. French [Duh88, Hei88c, Lev04, Pla11c, Pla11d, Pla12a, Pla13g, Pla41b, Pla60a, Pla63a, Pla89, Pla91a, PD93, Pla08a, Ros59a, Ros59b, Rub12, Som12, War12]. frequency [KGSO55]. Frequenz [BBG^{+}58]. Freunden [Tob08]. Friedenspreis [MdB63]. Friedhof [Ham72]. Friedrich [Ebe63, Jüt17, MKD^{+}91, Pla14d, Pla14f, Pla16e, Pla23h, Pla30f, Pla37i, Sta10, PP79, Pla14d, Pla14f, Pla16e, Pla23h, Pla37i]. Friedrich-Wilhelms-Universität [Jüt17, Pla14d]. Friedrichs [Pla17a, Pla24a, Pla28a, Pla37b]. friends [Tob08]. Fritz [Par12]. Fritzsch [MKD^{+}91]. front [Pla20c]. Frontiers [Hei74, KZR^{+}73]. Frontiers/Arzimowitsch [KZR^{+}73]. frühjahr [Pla10a]. Frühjahrstagung [BWN78]. Frühjahrstagungen [Ano81, JDB^{+}97]. Fully [BBG^{+}58, Gra69]. fundamental [Gré19, Lüs00, MR82c, Qui17]. fundamentalen [Lüs00]. Fundamentals [KZR^{+}73]. funeral [Pla48c]. Fünfstellige [GBM^{+}49]. Fünfte [Her24, Pla99c]. fünfzig [Pla30f, Pla62]. fünfzigjährigen [Pla29b, Sch29]. Funktionen [Ano48i]. Funktionen/Schulz [Ano48i]. further [Fal59]. fusion [JDB^{+}97]. fusion/First [JDB^{+}97]. Future [Ano53b, Peg77].

G [Adl03, Ano34, Ano39, MP19, Pla89k, Pla89l, Pla89o, Wil37]. G. [MP19]. Galilei [Büh98b]. Galileo [Büh98b]. Gas [Gra69, Hol13, Pla83, Pla94c, Pla95b, Pla09f, Pla22i]. Gase [Pla89g, Pla89e, Pla89o, Pla15c, Pla08e, Pla15c, Pla25h, Pla89d]. Gases [Pla89g, Pla08e, Pla15c, Pla25h, Pla15c, BBG^{+}58, Hol13]. Gases/Wirk [BBG^{+}58]. Gasgemengen [Pla83]. Gasmolecülen [Pla95b]. Gasmolekülen [Pla22i, Pla94c]. Gastheorie [Pla09f, Pla13c, Pla13d, Pla14c]. Gastvortrag [Pla29h]. Gastvorträgen [Ano29b]. gauge [Lüs00]. Gaynor [C.49, Dou49]. geb [Ano24, Ano39, Ano41a, Dri01, Jüt11, Jüt18b, Jüt22, Mis33, Reg34, Sch55, Wil37, Zil33]. Gebiete [Pla23b, Pla95, Pla97, PHR07]. Gebieten [Ber15]. Gebrauch [Fra31, Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f]. gebrochenzahliger [JDB^{+}97]. Geburtsfestes [Pla17a]. Geburtstag [Ano38a, Ano08f, Bec09, Brü48a, HD48, Jac08b, Pla18a, Pun67, Sta10, WvLuAE18, vL38, vL58, Ano57, Ano08a, Ber15, Din59, GSU07, HRR10a, HRR10b, Pla38b, Pla38c, SR37, Som38a, Zen38]. Geburtstag/J [Pun67]. Geburtstage [Pla38a, Som18]. Geburtstages [Jüt18a, Max58, Pla38g]. Geburtstags [Pla46a]. Gedächtnis [Jüt17, Mei47, Pla94a, Pla14d, Som47]. Gedächtnisfeier [Som48]. Gedächtnisrede [Pla06f, Pla28c, SR37, Pla23f]. Gedächtnisreden [Kir79]. gedämpft [Pla96c, Pla97e]. Gedanken [Kro52, Uns67]. Gedenken [JDB^{+}97, MH97, Her72]. Gedenken/WE [JDB^{+}97]. Gedenken/WE-Heraeus-Programm [JDB^{+}97]. Gedenkfeier [Ano49a, Brü48a, HD48]. geehrter [KSB63]. Gegen [Ano46b, Ano46c, Jae79, Pla96b]. Gegensätze [Wol08]. Gegenwart [Pla89i]. gegenwärtige [Pla13c, Pla13d, Pla14c]. Gegner [Löh12]. Geh [Ano29b, Ano41a, Ano34, Ano36, Jüt17, Jüt18a, Jüt21, Jüt22, Reg34, Zil33]. Gehalten [Pla14f, Pla16e, Pla20b, Pla23h, Flü47, Hei31, Jüt17, Jüt21, PP79, Pla06f, Pla09c, Pla10a, Pla10c, Pla10d, Pla10e, Pla11c, Pla11d, Pla11a, Pla13e, Pla13h, Pla14d, Pla15a, Pla17a, Pla20c, Pla20a, Pla28c, Pla29e, Pla29f, Pla31e, Pla32b, Pla35a, Pla35e, Pla35c, Pla35d, Pla36c, Pla37d, Pla38h, Pla42e, Pla45a, Pla45b, WvLuAE18]. gehaltenen [Pla48c]. Gehört [Har59]. Geiger [Pla37f]. Geist [Pla33j]. Geisteswissenschaften [MKD^{+}91, Pla37a]. Geisteswissenschaften/Vischer [MKD^{+}91]. geistige [Hof96]. Geitel [Ber15, Ber15]. Gelehrten [Ano46b]. Gelehrten/Neuer [Ano46b]. Gelehrter [Ano29b, Ano46b, Hof08b]. Gelehrter/Farben [Ano46b]. Gelehrter/Farben-Fernsehen [Ano46b]. Geleit [Pla33j]. Geleitwort [Ano34, Ano36, Ano39, Ano41a, Sch55, Wil37, Pla34d]. Gemeinverständlich [Ano39, Ano34]. General [BT58, DT58, Pla16c, Pla21c, Pla28b, Pla33d, Pla7 , Pla49a, Pla15f, Pla93c, Pla08a, Pla08b, Pla09a, Pla37e, Pla91a]. generalized [NK06]. generally [Hei88a, Hei06]. Generation [Hür22]. Genesis [Her71, HN70, Pla20d, Sin08]. Genocide [Hop96]. Geologie [GBM^{+}49]. Geometrie [BBG^{+}58, GBM^{+}49]. Geometrie/Ackermann [BBG^{+}58]. Geometrie/von [GBM^{+}49]. Geometrische [GBM^{+}49]. Georg [Ano11, Kör61, Kör61, Pun67]. George [Ano11]. Gerlach [GBM^{+}49]. German [Alb93, Ano48a, Bad08, Brü48b, C.49, Dou49, EH08b, HB87, Hie68, Hop96, McK27, Pla13f, Pla29e, Pla29f, Pye88, San87, Swe89, Ull13, vP06, AH93, Ano16, Ano29c, Ano30a, Ano38a, Ano43b, Ano44, Ano46c, Ano47a, Ano47d, Ano48a, Ano48e, Ano48f, Ano48c, Ano48g, Ano49a, Ano49b, Ano49c, Ano53a, Ano53b, Ano53c, Ano57, Ano59, Ano69, Ano70, Ano97a, Ano97b, Ano08f, Ano08a, Ano11, Arr06, Bec09, Ber15, BFV^{+}48, BuDH07, BP24, BDR86, Bor63a, Bre08, Brü38, Brü48a, Brü48b, Büh98a, Büh98b, Cah85, Cah88, Cla91, Cla89, Din39, Din59, Dys70, EH08a, Eck97a, Eck97b, Ein06, Ein07a, Ein07b, Ein13, EP06, Fal59, FP27, Fie75, Fis00, FH00, FS06, Fis07, Fis08a, Fis08b, Fis10b, Fis10a, Fis12, Fok14, Frö01, GSU07]. German [Gri98, Gut15, Haa70, Haa75, Hac07, Hah57, Ham72, HD48, Har38, Har43, Har48, Har53a, Har53b, Har58, Har59, Har64a, Har64b, HD01, HH63, HMF98, Hei86a, Hei88a, Hei00, Hei06, Hei31, HPE^{+}33, Hei35, Hei38, Hei43a, Hei47b, Hei49, Hei58, Hei86b, Hei69, HT99, Hen07, Her72, Her73b, Her73a, Her77, HR01, Her76, Her58, Hof48, Hof96, HS97, Hof00, HW07, Hof08b, Hof08d, Hof08e, Hof08f, Hof09, HRR10b, Hof10, Hol13, Hop96, Ish15, Jac08b, Jae79, JDB^{+}97, Jor48, Jos95, Kan69, Kan70, Kir94, Kir98a, Kir98b, KSB63, KK75, Kir79, Koh02, Kör61, Kre67, Kro52, KGSO55, Kut14a, Kut14b, Lan57, LMOS15, Lem08, Löh12, Lüs00, Max58, MH97, Mei47, Mei58, Mei01, Mei03, Men73, MSQM63, MAS63, Mey08, MP19]. German [Mül08b, Mul94, OT97, PP25, PP79, Pla79, Pla80, Pla81, Pla82, Pla83, Pla84, Pla87b, Pla87f, Pla87a, Pla87c, Pla87d, Pla87e, Pla88a, Pla88i, Pla88k, Pla88j, Pla88l, Pla89a, Pla89b, Pla89y, Pla90d, Pla90c, Pla90a, Pla90b, Pla90e, Pla91a, Pla91b, Pla92a, Pla92b, Pla92c, Pla93c, Pla93a, Pla93b, Pla94c, Pla94a, Pla95a, Pla95b, Pla96a, Pla96c, Pla96b, Pla97e, Pla97c, Pla97d, Pla97b, Pla97a, Pla97f, Pla98c, Pla98b, Pla99b, Pla99c, Pla99a, Pla00a, Pla00f, Pla00c, Pla00h, Pla00d, Pla00i, Pla00b, Pla00g, Pla01e, Pla01d, Pla01f, Pla01b, Pla01c, Pla01g, Pla01a, Pla02c, Pla02d, Pla02a, Pla02e, Pla02b, Pla03h, Pla03f, Pla03a, Pla03e, Pla03g, Pla04b, Pla04a, Pla05a, Pla05c, Pla06b, Pla06a, Pla06d, Pla06e, Pla06f, Pla06c, Pla06g, Pla07f, Pla07a, Pla07c, Pla07g, Pla07d, Pla08a, Pla08b]. German [Pla08f, Pla08e, Pla08c, Pla08g, Pla09a, Pla09d, Pla09b, Pla09c, Pla09f, Pla10a, Pla10f, Pla10h, Pla10i, Pla10c, Pla10d, Pla10e, Pla10b, Pla10j, Pla11f, Pla11b, Pla11a, Pla11e, Pla12b, Pla12c, Pla12d, Pla13a, Pla13c, Pla13d, Pla13e, Pla13i, Pla13h, Pla13b, Pla13k, Pla13j, Pla14d, Pla14c, Pla14f, Pla14a, Pla14b, Pla14e, Pla14i, Pla14j, Pla15a, Pla15h, Pla15b, Pla15c, Pla15i, Pla15d, Pla15e, Pla16f, Pla16a, Pla16c, Pla16d, Pla16e, Pla16b, Pla17a, Pla17b, Pla17f, Pla17g, Pla17h, Pla17d, Pla17e, Pla17c, Pla18a, Pla18b, Pla19a, Pla19c, Pla19e, Pla19d, Pla19f, Pla19b, Pla20b, Pla20c, Pla20a, Pla21a, Pla21c, Pla21d, Pla21b, Pla21g, Pla21f, Pla22a, Pla22b, Pla22c, Pla22d, Pla22e, Pla30a, Pla22f, Pla22i, Pla22h, Pla22j, Pla23b, Pla23c, Pla23e, Pla23d, Pla23f, Pla23a, Pla23g, Pla23h, Pla23i, Pla24a]. German [Pla24b, Pla24c, Pla25c, Pla25h, Pla25b, Pla25a, Pla25g, Pla25e, Pla25f, Pla26a, Pla26d, Pla26b, Pla27a, Pla27b, Pla27e, Pla27f, Pla28a, Pla28d, Pla28b, Pla28c, Pla29c, Pla29b, Pla30b, Pla29a, Pla29k, Pla29l, Pla29d, Pla29g, Pla29e, Pla29f, Pla29n, Pla29h, Pla29i, Pla29j, Pla30c, Pla30g, Pla30f, Pla30h, Pla30i, Pla30d, Pla30e, Pla31b, Pla31a, Pla31f, Pla31c, Pla31e, Pla32b, Pla32a, Pla32d, Pla32e, Pla33b, Pla33j, Pla33a, Pla33e, Pla33c, Pla33g, Pla33f, Pla33h, Pla34d, Pla34a, Pla34b, Pla34c, Pla34e, Pla35a, Pla35b, Pla35e, Pla35c, Pla35d, Pla35f, Pla35g, Pla35h, Pla36a, Pla36c, Pla37b, Pla37e, Pla37f, Pla37i, Pla37a, Pla37c, Pla37d, Pla37h, Pla38a, Pla38b, Pla38c, Pla38e, Pla38f, Pla38g, Pla38h, Pla38d, Pla39c, Pla40a, Pla40b, Pla40c, Pla41a, Pla41c, Pla42c, Pla42d, Pla42a, Pla42e, Pla42f]. German [Pla43c, Pla43b, Pla44a, Pla44b, Pla45a, Pla45b, Pla45e, Pla45f, Pla45d, Pla46a, Pla46b, Pla46c, Pla47b, Pla48b, Pla48a, Pla48f, Pla48c, Pla48d, Pla48e, Pla49b, Pla49k, Pla50a, Pla52b, Pla52a, Pla52c, Pla52d, Pla53b, Pla53a, Pla53c, Pla53d, Pla54c, Pla55c, Pla55a, Pla55b, Pla55d, Pla55e, Pla57d, Pla58c, Pla58e, Pla58b, Pla58f, Pla58d, Pla58a, Pla62, PP64, Pla65b, Pla65a, Pla65c, Pla66, Pla67b, Pla67a, Pla67c, Pla78b, Pla78a, Pla78d, Pla78c, Pla82, Pla88b, Pla90c, Pla90b, Pla95, Pla97, PHR07, Pla08b, Pla10b, Pla15a, Pla15f, Pla15c, Pla15b, Pla15i, PS58, Prz63, RH97, RTKK75, RH01, Sch96, SCP19, Sch78c, Sch78a, Sch78b, Sch78d, Sch29, SP29, Sch61, Sch91, Sei08, Sim12, Sim06, Som11, Som18, Som33, Som36, SR37, Som38a, Som38b, Som47, Som48].

German [Spr89, Sti84, Thi69, Thi88, Tob08, Uli97, Ull96, Uns58, Uns67, Vog61, Wei89, Wes40, Wes57, Wes58, Wol08, WC91, Yos95a, Yos95d, Zac92, Zen38, Zim34, dB48, vK97, vBL96, vL18, vLP21, vL38, vL50, vL58]. German-Speaking [Cah88]. German/Latin [Her73b]. Germany [Eck97a, Hop96, JM17, Mac93, Pla31c, Pla31d, Set03, Sti84]. Ges [Pla94a, Pla06f, Pla38g]. Ges. [Pla11c, Pla11d]. Gesammelte [BDR86, Pla22h]. Gesamtwerkes [Ebe63]. gesättigter [Pla89b, Pla89h]. Geschichte [AH93, Bad08, Bre08, Cah99, Har59, Kra97, LMOS15, Pla43c, Sch00b, Ano48i, vBL96]. Geschwindigkeitsverteilung [Pla22i]. Geschwindigkeitsverteilungsgesetzes [Pla94c]. Geschwindigkeitsvertheilungsgesetzes [Pla95b]. Gesellschaft [Ano48a, Bec09, Brü48b, Cah99, Ebe63, Hei31, Her03, HW07, Koh02, Kra97, Pla20a, Pla28c, Pla31e, Pla35e, Pla36c, Pla37a, Pla39c, Pla50a, Pla52d, Pla53d, Pla55e, Som48, Sta10, WvLuAE18, vBL96, Gri98, Hof08e, Pla18a, Pla42e, Zac92]. Gesellschaften [Ebe63]. Gesetz [Pla88c, Pla89m, Pla01d, Pla78a]. Gesetze [Cla89, Pla89p, Pla89e, Pla13e]. Gesetzes [Ano48a, Pla00i, Pla12b]. Gesetzlichkeit [Pla26b]. Gesetzmässigkeit [Pla14d, Jüt17]. Gestatten [Sei08]. Gestirne [Gut15]. geworden [MAS63]. Giants [Ano47c]. Gibbs [Pla03c, Kör61]. Gibt [Pla49b]. Gießen [BWN78]. Gießen/Frühjahrstagung [BWN78]. Gift [MP01]. Gill [Cah99]. ging [Bre08]. Giris [Pla15g]. Gittereichtheorie [Lüs00]. given [DdCA48]. Glänzendsten [Gut15]. Glaube [Pla30i, Pla88b]. gleichförmig [Pla07d]. Gleichförmige [Pla10f]. Gleichgewicht [Pla83, Pla88a, Pla13i, Pla35b, Pla89v]. Gleichgewichte [EH08b, Pla08b]. Gleichgewichts [Pla89p]. Gleichgewichtszustände [Pla80]. Gleichung [vK97]. Glenys [Cah99]. Global [BF12, KM12]. Go [Fra05]. God [Kut14a]. God-less [Kut14a]. Goethe [BBG^{+}58, Som38b, Ano46c, Som38b]. Goethe-Preis [Ano46b, Ano46c]. Goethe-Saal [Som38b]. Going [Bro36, Pla32k, Pla33i, Pla77, Pla15j, Pla33h, Pla34e, Pla43b, Pla44b, Pla45e, Pla59a, Pla81]. golden [Pla29k, Pla29l]. goldenen [Pla29k, Pla29l]. Goldmacher [Ano53b]. Goldmacher/Internationales [Ano53b]. Gott [Kut14a]. Gott-lose [Kut14a]. Gottfried [Pla46a]. Göttingen [Flü47, Ham72, Pla13c, Pla13d, Pla14c, Ano48d, Ano48j, Ano49a, DdCA48, Pla13c, Pla13d, Pla14c]. Göttinger [Ham72]. Gouy [Pla89m]. GPS [Ano12]. Gräber [Ham72]. grandes [Rub12]. Gratulation [SR37]. Gravelius [Ano48i, GBM^{+}49]. graves [Ham72]. gravity [CH01, Pad83]. Gray [BBG^{+}58]. Great [Ano31, Brü38, Büh98b, Kae48, Lig05a, Lig05b, Pic08]. Greatest [Acz01, Ano21, Hür22]. Green [McK27]. Greenberg [Hib91]. Greenwood [KZR^{+}73]. Gregor [Fie75, Hun75, Fie75]. Grenzen [Pla42d, Pla42e]. Grenzschicht [Pla30g, Pla31f, Pla33e]. Griechische [BBG^{+}58]. Großbritannien [JDB^{+}97]. Grosse [BPL^{+}54]. großen [Brü38]. Grösste [Ano46b, Ano46c]. Grosswetterlagen [Ano48i]. Grosszahl [GBM^{+}49]. Grosszahl-Forschung [GBM^{+}49]. Grotrian [HRR10b]. Grotrians [HRR10a, HRR10b]. Group [Ano48g]. Grund [GBM^{+}49].

Grundbegriffe [Sch65]. Grundbeziehungen [GBM^{+}49]. Grundfragen [Hei58]. Grundgleichungen [Pla06c]. Grundlage [Bad08, Pla03f]. Grundlagen [JDB^{+}97]. Grundriß [Pla93c]. Grundzahl [Ano24]. Grützmacher [Ano61]. Gruyter [Sch65]. Guest [Pla29h]. Guggenheim [BBG^{+}58]. Günther [SH53]. Güntherschulze [SH53]. Guthenick [Pla23e, Pla23e].

H [Adl03, Ano02, Ano34, Ano39, Ano41a, Jac02, Klo63, Mic68, Ped75, Pic25, Sac69, Wil37]. H. [Pla78e, Prz11]. habe [Hof08e]. Haber [Par12]. Habilitationsakten [Sch00b]. Hahn [Ano46a, Ano46b, Ano49b, Ano49c, GBM^{+}49, Pla25c, Ano49b, Ano49c, Pla25c, Sim12]. halbes [Eck97b]. half [Eck97b]. Hall [Hop96, Som38b]. Haltung [Löh12]. Hamack [Som38b, Som38b]. Hamack-Hauses [Som38b]. Hamburg [Ano81, Uns67]. Handbook [BBG^{+}58]. Handbook/Guggenheim [BBG^{+}58]. Handbuch [Kay02]. Handelns [JDB^{+}97]. Hannover [Haa70]. Hans [Ber15, Din59, Dri01, Hen60, Kra02a, Kra02b, Thi71, Ber15, Din59, Pun67, SS53]. Hans-Georg [Pun67]. Hans-Joachim [Din59, Din59]. Hanser [Sch55]. Hanser-Verlag [Sch55]. Hardly [Hof08e]. Hardy [Adl03]. harmful [Ano53b]. harmonic [Sch78b]. harmonischen [Sch78b]. Harmonium [Pla93b]. harmony [Pla93b]. Harnack [Hei31, Pla31e, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla31e, Pla35c, Pla35d, Pla37d, Pla38g, Pla45a, Pla45b]. Harnack-Haus [Hei31, Pla31e, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla38g]. Härtel [Pla93d]. Hartmann [BPL^{+}54, Hen60]. harvest [Bra09]. hatte [OT97]. Häufigkeits [GBM^{+}49]. Häufigkeits-Analyse [GBM^{+}49]. Häufigkeits-Analyse/Peters [GBM^{+}49]. Hauke [Cah99, Kra03]. Hauptsatz [Mül08b, Pla79, Pla89j]. Hauptsatzes [Pla89n, Pla93a, Pla26d]. Hauptversammlung [Pla20a]. Haus [Ano97a, Hei31, Pla31e, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Ano97a, Pla38g]. Hauses [Som38b]. Haushaltsplan [Ano81]. Hawking [Pic08]. Heads [Ano30a]. Health [Wei89, Hop96]. Hearing [Hui13]. Heat [Pla06g, Pla10h, Pla13j, Pla14h, Pla19f, Pla21f, Pla23i, Pla32i, Pla49i, Pla57c, Pla59b, Pla66, Pla88a, Pla91b, Pla12b, Pla15a, Pla15c, Pla15i, Cla91, Cla89, Ein07b, Kir98a, Kir98b, Pla90d, Pla90c, Pla91a, Pla00c, Pla10j, Pla12c, Pla12d, Pla13i, Pla24b, Cla91]. heavy [JDB^{+}97]. Heft [Jüt18a]. Heftes [Pla33j]. Heidelberg [BWN78, Dri01, KGSO55, KGSO55]. Heidelberg/Der [KGSO55]. Heidelberg/Frühjahrstagung [BWN78]. Heidelberg/The [KGSO55]. Heidenheim [Som48]. Heilbron [Ano89, Hop96, KAR89, San87]. Heiliger [KSB63]. Heiligers [KSB63, WF63]. Heinrich [Cah99, Pla23f, Pla94a, Pla23f]. Heisenberg [Adl03, BDR86, BDR86, Cas09, Hop96, Tre01]. Held [Ano48j, Pla20b, BCS82, DdCA48, Hei31, PP79, Pla06f, Pla09c, Pla10a, Pla10c, Pla10d, Pla10e, Pla11c, Pla11d, Pla11a, Pla13h, Pla14d, Pla14f, Pla14g, Pla15a, Pla16e, Pla17a, Pla20c, Pla23h, Pla28c, Pla31e, Pla32b, Pla35a, Pla35e, Pla35c, Pla35d, Pla36c, Pla37d, Pla38h, Pla42e, Pla45a, Pla45b]. Helm [Kör61, Kör61]. Helmholtz [Ano02, KK75, Pla30b, Pla30b]. Hendrik [Prz86, Pla28c]. Hengstenberg [Rut69]. Henri [Adl03, Pla21d]. Henry [Ole81, Pla94b]. Henschel [Sch04]. Hentschel [Mul00, Sch00b]. Hentschel/Renate [Sch00b]. Heppenheim [Pla19d, Pla19d]. Heraeus [JDB^{+}97]. Herausgeg [HKS^{+}68, Ebe63]. Herausgegeben [Ano02, Jüt18a, Klo63]. Herleitung [Pla88b]. Hermann [Dri01, KK75, You68, Jac02, Kra02a, Kra02b, Pla94b]. Herr [KSB63]. Herrn [Pla92b, Pla29c]. Hertz [Pla94a, Pla94a]. Hettner [Pla16a]. Hettnerschen [Pla16a]. Heuser. [Ano41a]. Hi [MKD^{+}91]. Hi-Tech [MKD^{+}91]. Hi-Tech/Scribal [MKD^{+}91]. high [HS97, KGSO55]. high-frequency [KGSO55]. Higher [Bae01]. Higher-dimensional [Bae01]. Hilbert [BBG^{+}58, Lev04]. Himmel [Dri01]. Hintertreppe [Fis10b, Fis12]. hinweg [Wol08]. Hirn [Pla89o]. Hirnforschung [Bre08]. Hirth [Ano34, Ano36, Ano39, Ano41a, Wil37]. Hirzel [Ano89, Fra31, Mis33, Reg34, Zil33]. Histoire [Lev04]. Historian [Dar01]. Historians [Gea02]. Historic [Koe01]. Historical [AH93, Alb93, Sch78c, Sin08]. historical-critical [Sch78c]. historisch [Sch78c]. historisch-kritischer [Sch78c]. Historischen [AH93, Alb93]. History [Ber14, Cah88, Hop96, JM17, LMOS15, Nye06, WP85, AH93, Ano46c, Bad08, BN13, Bre08, Eck97b, ET07, Har59, HMF98, Kan76, Lev04, Pla43c, vBL96]. Hitler [Alb93, HMF98, Jae79, MP01, O’F56, Pla47b]. Hitler^{≫} [Alb93]. Hochfrequenzkinematographie [KGSO55]. Hochgeehrter [Ano43b]. hohem [HS97]. hohen [Pla89e]. Höhere [GBM^{+}49, HKS^{+}68]. Hohlraum [Pla07d]. homogenen [Pla04a]. homogeneous [Pla04a]. Hon. [Pla19d]. Hönger [BBG^{+}58]. Honor [Ano29a, Ano46d, Iva59a, Pla35e, Som38b]. honorary [Pla38g]. Honoree [Ano43b]. Honors [Ano48g]. Horst [Ano36]. hospital [Ano47a]. House [Hei31, Pla31e, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla38g, Som38b]. Hove [Haa75, Haa75]. Hrn [Pla00a, Pla00d, Pla13f, Pla19d, Sch29]. Hrsg [Dri01, Pla15c, Sch04]. Hrsgg [Sch00b]. Hubble [Bru02]. huge [Bre08]. Hugo [Ano41b, Mos42]. Human [Bro36, Mei58, Mei01, Wes58]. Humanist [O’F59]. humanities [Pla37a]. Humanökologie [MKD^{+}91]. Humorous [Ano21]. Hundred [Hoh99, Par12, Fad96]. hundredth [Ano57, Wei58]. Hunger [Bra14a]. hunt [Seg07a]. hydrogen [Pla19c]. Hygiene [Hop96]. Hypothese [Pla88j, Pla11f, Pla12c, Pla12d, Pla13e, Pla12a, Pla13g, Pla08a]. Hypotheses [Tib13]. Hypothesis [Cra84, Hol13, Nau16, Pla88j, Pla11f, Pla11b, Pla12a, Pla12c, Pla12d, Pla13c, Pla13d, Pla13e, Pla13g, Pla14c, Pla14e, Pla15d, Pla15e, Pla08a].

I/Solomon [Ano48i]. ice [Bre08]. ich [Hof08e]. idea [Seg07c]. Ideas [Ein54a, Ein82, Pla33a, Pla33g, Pla33f]. Ideen [Pla33g, Pla33f]. if [Tib13]. Ignatovich [MKD^{+}91]. ihr [BPL^{+}54]. ihre [Fal59, Yos95a, vBL96]. ihren [OT97]. ihrer [Hof08b, Pla06d, Pla06e, Pla07a, vBL96]. ihres [Bad08]. II [BT58, DT58, Ebe63, Kir98a, Kir98b, Kir79, Pla97d, Pla15e, Pla29n, Pla31f, Gra69, Hei43a, Kir98a, Kir98b, Kir79, Pla17a, Pla24a, Pla28a, Pla32g, Pla37b, Pla49e, Pla57a, Pla64b, Sim12]. III [BT58, DT58, Ebe63, Kir98a, Kir98b, Pla33e, Kir98a, Kir98b, Pla32h, Pla49h, Pla57b, WF63]. illus [Hie68]. im [Ano13, Ano34, Ano36, Ano39, Ano41a, Ano48a, Brü38, Cah99, Fok14, Hei31, Hei35, Her03, HW07, Hof08b, Hof08f, JDB^{+}97, Jüt18b, Koh02, Lem08, Pla00i, Pla01d, Pla01g, Pla10a, Pla19d, Pla26b, Pla31e, Pla33c, Pla35c, Pla35d, Pla37d, Pla38g, Pla38h, Pla42c, Pla45a, Pla45b, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c, Pla78a, Sch00b, Sch55, Som36, Som38b, Uli97, Ull13, Wil37, Zim34]. image [Her73a]. images [Ano48e, Ano48f, Pla63a]. images/Authors [Ano48e]. images/Authors/Papers [Ano48f]. Immortals [Ano47c]. Impact [AP00, Hei49, Pla33a, Pla33g]. Importance [Som11, Pla04b, Pla06d, Pla06e, Pla07a, Pla31c]. improved [Thi88]. Improvement [Pla69a, Pla00g, Pla78c]. Inaugural [EP06, Kir79, Pla25c, Pla37f, Pla38f, Pla57d, BP24, FP27, MP19, PP25, PP79, Pla29c, SCP19, SP29, Sch61, vLP21]. Incandescent [FSBR09]. inches [Pic25]. Inconsistency [Boy79, BFG81]. increase [Pla87c, Pla87d, Pla87e, Pla91b, Pla05b]. increasing [Duh88]. indefinitely [Bra14b]. independent [MSQM63]. Indeterminism [Pla48b, Pla52b, Pla53b, Pla55c, Pla58c, Pla65b, PS33, Pla38e, Pla47a, Pla67b, Sch91]. indeterminismo [Pla47a]. Indeterminismus [Pla48b, Pla52b, Pla53b, Pla55c, Pla58c, Pla65b, Pla38e, Pla67b]. Indeterminismuskritik [Sch91]. indeterminizm [PS33]. inertial [JDB^{+}97]. inferred [Cla89]. influence [Pla31d]. information [Pla91a]. Inhalt [Pla10h, OT97]. Initiations [Pla41b, Pla89, PD93]. Innsbruck [KGSO55, KGSO55]. Innsbruck/2 [KGSO55]. Innsbruck [?]2ndKruger:1955:VMP. Inönu [HKS^{+}68]. Inquiry [Gib19]. ins [Bre08]. instead [Ano46c]. Institut [AH93, Pla27b]. Institute [AH93, Ano11, BBG^{+}58, Cah99, Par12, Pla27b, vBL96, WC91, Cah99]. Instituten [Cah99]. Institutional [Cah85]. institutions [vBL96]. instruction [Pla15f]. Instrumentation [HKS^{+}68]. Instruments [Ber14]. insulators [Pla02d]. Integralrechnung [GBM^{+}49]. Integralrechnung/Reutter [GBM^{+}49]. Intellectual [Ano47c]. intelligence [HMF98]. Intelligent [Kut14b]. Intelligente [Kut14b]. intended [Ano70]. Intensitätsparameter [Pla35b]. intensity [Pla35b]. interference [Pla62]. Interferenz [Pla62]. International [Ano53b, Ano59, BT58, DT58, KGSO55]. Internationale [KGSO55, Ano59]. Internationales [Ano53b, KGSO55]. Internet [JDB^{+}97, JDB^{+}97]. Internet-Adressen [JDB^{+}97]. Internet-Adressen/Blick [JDB^{+}97]. interpolation [Boy83]. interpretation [Pla16a]. Intersection [Par12]. Interviews [Ano31]. Introduction [Pla32f, Pla32g, Pla32h, Pla32i, Pla32j, Pla33d, Pla49a, Pla49e, Pla49h, Pla49i, Pla49j, Pla57a, Pla57b, Pla57c, Pla64b, Hie68, KAR89, Pla16c, Pla16d, Pla19e, Pla21c, Pla22d, Pla22e, Pla30a, Pla28b, Pla31b, Pla37e, Pla41b, Pla89, PD93, Pla15f, Pla15g, Fal50]. introductory [Ano11]. Invariance [BF12, KM12, WB78]. Inventory [HD01, Ull96]. Inverted [Ano53b]. Invertierte [Ano53b]. investigation [Pla09f]. Involving [Luc30]. Ion [Pla00f, Pla02c]. Ionen [JDB^{+}97]. Ionized [BBG^{+}58, Gra69]. ions [JDB^{+}97]. irradiated [Pla01g]. irreverent [Bra14b]. Irreversibility [Nee80]. Irreversible [Pla99b, Pla97c, Pla97d, Pla97b, Pla98c, Pla99c, Pla00h, Pla01f, Pla01b, Pla01c, Pla05b, Pla78d, Pla99b]. irreversibler [Sch65]. Isaac [Ano46d]. ISBN [Ber14, Dri01, Kra97, San87]. Ishiwara [PB17]. Isotherme [Pla89d]. isotropen [Pla02d, Pla03h]. isotroper [Pla80]. isotropic [Pla80, Pla02d, Pla03h]. issledovanij [Pla26e]. Issue [AP00, Ano08e, Ano48k]. Issues [PS33, Pla45f, W^{+}87, WdC02]. Italian [Hei88b, Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a, Ros08]. IV [BT58, DT58, Pla98c, Pla32j, Pla49j]. ivory [Jos95]. Iwand [Din59, Din59]. IX [Mis33, Zil33]. Izbrannye [Ano75].

J [Ano41a, Ano53c, Ano89, Ben30, Flü47, Her24, Hie68, Jüt17, Mic68, Pla88c, Pla88d, Pla89n, Pla03c, Pun67]. J. [Ano53c, San87]. jahr [Ano46c]. Jahre [AH93, Ano46b, Ano61, Brü38, HD01, Hei38, Pla37a, Pla62, Rut69, SH53, Wes57, PS58]. Jahre/Josef [Rut69]. Jahre/Max [Rut69]. Jahre/Max-Planck-Medaille [Rut69]. Jahre/Nobelpreise [Ano61]. Jahre/Paul [SH53]. Jahre/Winfried [Rut69]. Jahre/Zur [Wes57]. JahreMax [MdB63]. JahreMax-Planck-Medaille [MdB63]. Jahren [Pla30f]. Jahres [Pla06a]. Jahrestag [Pla37b, Pla26a, Pla30c]. Jahrestages [Pla15a, Pla17a, Pla24a, Pla28a, Pla22b]. Jahrg [Jüt18a]. Jahrhundert [Eck97b]. Jahrhundertwende [Yos95a, Hof48]. jährigen [Pla35e, Pla37i]. Jahrzehnte [Haa70]. Jakob [PS58, PS58]. James [Bro36, Kas64, Pla31c, Tho31]. Januar [Pla32b]. January [Pla32b]. Japanischen [Ano46b]. Jay [Hop96]. Jeans [Gor91]. Jedność [Pla70]. Jehova [Kut14b]. Jehovah [Kut14b]. JET [JDB^{+}97, JDB^{+}97]. JET/JET [JDB^{+}97, JDB^{+}97]. jets [Pla84]. jetzt [JDB^{+}97]. Jews [BA01]. Joachim [Din59, Din59]. Joh [Ano24, Jüt21, Jüt22]. Johann [Ano30b, Wei68]. John [GSU07, Hop96, GSU07]. Johnsen [Pla23e]. Jones [Pic25]. Joos [Ano48i]. Josef [Rut69]. Josiah [Kör61, Kör61]. Joule [Pla89o]. Journey [FR13, Lev17]. Jubilar [Ano43b]. Jubiläum [Pla37i]. jubilee [HD01, Pla29k, Pla29l]. Jubileums [HD01]. Juli [PP79, Pla06a, Pla29c, Pla46a, Pla15a]. Julius [Ber15, Jüt18a, Ber15]. July [PP79, Pla06a, Pla29c, Pla46a, Pla15a]. Jun [PB17]. June [Pla20c, Pla28c, Pla29b, Pla29k, Pla35a, Sch29]. Juni [Flü47, Jüt21, Pla20c, Pla28c, Pla29b, Pla29k, Pla35a, Sch29]. Jürgen [Ano11, GSU07, GSU07]. justice [MSQM63]. justification [Pla12b, Pla26d].

Kafka [Hib91, Gre90]. Kaiser [Cah99, Hei31, Her03, Koh02, Pla20a, Pla31e, Pla37a, vBL96, Koh02, Pla20a, Pla37a, Pla42e]. Kaiser-Wilhelm- [Cah99, vBL96]. Kaiser-Wilhelm-/Max-Planck-Gesellschaft [Cah99, vBL96]. Kaiser-Wilhelm-Gesellschaft [Pla42e]. Kaisers [Pla17a]. Kalorimetrische [BBG^{+}58]. Kampf [Hei35, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Som36, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c]. Kangro [Ped75, Thi71]. kann [Pla42f]. kanonische [Pla08e]. Kapitel [Sch65]. Karl [Ano47e, Ano08d, Bor48, Bor63b, Hai11, Kan08, van10]. Karl-Scheel-Preis [BWN78]. Karlskoga [BCS82]. Karlsruhe [Hei69]. Karsten [HT99, Mul00, Sch00b, Sch04, HT99]. kart [Sac69]. Kartina [Pla29o]. Kaufmann [Pla06d, Pla06e, Pla07a, Pla07c]. Kaufmannschen [Pla06d, Pla06e, Pla07a, Pla07c]. Kaum [Hof08e]. Kausalbegriff [Pla32d, Pla32e, Pla37c, Pla41a, Pla48a, Pla52a, Pla53a, Pla55a, Pla55b, Pla58b, Pla65a, Pla67a]. Kausalgesetz [Pla23g]. Kausalität [Pla33c]. Kay [MKD^{+}91]. keeping [JDB^{+}97]. keiner [OT97]. Kern [Pla93a]. Kernphysik [BDR86]. Kernphysikalisches [Ano53b]. Kettenreaktion [GBM^{+}49]. Key [Tib13]. Kiel [Qui13, Uns58, Uns67]. Killing [Hop96]. Kinematographie [BBG^{+}58]. Kinetic [Pla15c, Pla09f, Pla13c, Pla13d, Pla14c]. Kinetics [Lew30]. Kinetik [Pla89x]. Kinetische [Pla15c, Pla13c, Pla13d, Pla14c]. kinetischen [Pla89o, Pla09f]. King [Pla17a, Pla24a, Pla28a, Pla37b]. Kirchhoff [Sch78c, Aga67, Sch03, Sch78c]. Klassiker [Ano02]. Klaus [Mul00, Sch00b, Sch04]. Klein [GBM^{+}49, Hie68, Mic68]. kleine [Bad08]. Kleiner [Ano11]. kleinsten [Fis10b, Fis12, Pla14a, Pla14b, Pla25a]. Klenke [Cah99]. Klimawandel [Bre08]. Knipping [Pla37i, Pla37i]. Knorr [Ano34, Ano36, Ano39, Ano41a, Wil37]. Know [Law15]. knowing [HS97]. Knowledge [Ber14, Pla13h, Pla14f, Pla14g, Pla16e, Pla23h, Som33, Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a, HPE^{+}33, Hei43a]. Kockel [Ber60]. Kohl [Her03]. Kohlensäure [Pla89i]. Kollegen [Hof08f]. Komitee [Ano29b]. Kompendium [BBG^{+}58]. Kondensat [JDB^{+}97]. Kongreß [KGSO55, Pla13e]. Kongressitis [Ano59]. König [Pla17a, Pla24a, Pla28a, Pla37b]. königl [Pla20b]. Königlich [Jüt21, Pla20c]. Königsberg [Pla10c, Pla10d, Pla10e, Pla11a, Pla10c, Pla10d, Pla10e, Pla11a]. Konstante [Pla06b, Pla21a, Pla22a]. Konstanten [Mey08]. Konstitution [Pla87b, Pla09d]. Kontraktion [Pla10f]. Kontrovers [Wol08]. Kontroverse [Thi69]. kooperativ [Wol08]. Korper [Kir98a, Ano48i, Pla80, Pla19e, Pla22d, Pla31b, Pla16f, Pla16d, Moo23, Pla89x]. Korpuskularmechanik [Pla40c, Pla41c, Wes40, Pla40b]. Korrespondenz [Hof08b]. kosmische [Bre08]. Kraft [Cla89, Pla38f, Pla38f]. Kreisels [Pla18b]. Kretzschmar [HKS^{+}68, Sac69, You68]. Krise [Yos95a]. Krisis [Pla29g]. Kristie [Hop96]. Kritik [Vog61, Yos95c, Pla00d]. kritische [Pla09f]. kritischer [Sch78c]. Kropp [GBM^{+}49]. Krystallisation [Pla89k, Pla89l]. Kuantum [Pla15g]. Kuhn [Yos95c]. Kunde [Ano46b]. Kundgebung [Pla29e, Pla29f]. Kurd [Pla07d]. Kurzes [BBG^{+}58]. kurzgefaßt [JDB^{+}97]. Kurzzeitphotographie [KGSO55].

L [Ano89, Pla89o, San87]. Laboratorium [Ano53b].

Laboratorium/Zukunftsmusik [Ano53b]. laboratory [Ano53b]. laboratory/Future [Ano53b]. Ladenpreis [Som49]. Ladung [JDB^{+}97]. Landau [HKS^{+}68]. Lande [Pla19d, Pla19d]. lange [Pla03g]. Langevin [Ano46a]. Langevin/Besuch [Ano46a]. Laplace [KGSO55]. Laplace-Transformation [KGSO55]. largest [Ano46c]. laser [JDB^{+}97]. Laserkühlung [JDB^{+}97]. Laserlicht [JDB^{+}97]. lasse [Cla89]. last [Pla60a, Pla91a]. Later [Eck08]. Latin [Her73b]. lattice [Lüs00]. Laudatio [Haa75]. Laudationes [BDR86]. Laudations [BDR86]. lauds [Ano29a]. Laue [Ano48i, Pla29m, Pla37i, Pla39b, Pla48c, Ano49b, Ano49c, Pla37i, Pla48c, RS80, vLP21]. Laues [vLP21]. Laufbahn [Lem08]. Laughing [HRR10b]. Laurikainen [MKD^{+}91]. Lautdokuments [Her76]. Law [Aga67, Mül08b, Pla93a, Pla01d, Pla26d, Pla69b, Pla78a, v.B18, Ano48a, FMM96, Gar76, Iro05, Kan70, Kan76, KSB63, Par94, Pla79, Pla81, Pla94c, Pla95b, Pla06b, Pla12b, Pla12a, Pla23b, Pla34a, Pla34b, Pla34c, Pla38d, Pla95, Pla97, PHR07, Sch03, Sch78a, Sch78d, Yos95d, FSBR09, Foo16, Gra69, Kan69, Pla23g]. Lawrence [Ole81]. laws [Cla89, Pic08, Pla01g, Pla13e, Pla26b, Pla26e]. layer [Pla30g, Pla31a, Pla31f, Pla33b, Pla33e]. “großes [HRR10a, HRR10b]. “Lachkabinett” [HRR10a, HRR10b]. “physikalischer [HRR10a, HRR10b]. leap [Fis10b, Fis12]. least [Pla14a, Pla14b, Pla25a]. Leben [Ano89, Ano97a, EI08, Hei88a, Hei06, KAR89, Mül08b]. Lebenslauf [Pla20b, PS58]. Lebenswerkes [Hei49]. Lebenszeit [JDB^{+}97]. Lecons [Pla13g, Pla08a]. Lecture [Ano23, Pla11c, Pla11d, Pla23g, Pla29e, Pla29f, Pla7 , Ano11, Fis08a, Hei69, Pla09c, Pla13c, Pla13d, Pla14c, Pla29h, Pla31e, Pla35c, Pla35d, Pla36c, Pla37d, Pla39c, Pla45a, Pla45b, Pla52d, Pla53d, Pla55e, Haa70, Hei31, Pla10c, Pla10d, Pla10e, Pla11a, Pla20b, Pla20c, Pla29g, Pla35e, Pla38h, Pla42e, vL58]. Lectures [Fra25, HR01, Pla09e, Pla10a, Pla22h, Pla25d, Pla98, Pla12a, Pla15d, Som33, Gri98, Max58, Pic25, Pla15g, Pla33h, Pla43b, Pla44b, Pla45e, Pla58f, Pla90b, Pla15f, Pla15e, Sim06, Kir94, MKD^{+}91, Pla97f, Pla05c, Pla06g, Pla11e, Pla13g, Pla13k, Pla13j, Pla17c, Pla19f, Pla21g, Pla21f, Pla22j, Pla23i, Pla27f, Pla30h, Pla49k, Pla54c, PP64, Pla66, Pla08a, Pla10b, Pla15i, RH01, Ein06]. Lehrbuch [BBG^{+}58, WF63, HKS^{+}68]. Lehrer [Hof08b]. Lehrstuhl [AH93]. Leibniz [Pla46a, EP06, LMOS15, Pla15a, Pla22b, Pla26a, Pla30c, Pla46a]. Leibnizischen [Pla22b, Pla26a, Pla30c]. Leibnizschen [LMOS15, Pla15a]. Leibniztag [EP06]. Leiden [Pla09c]. Leinen [Ano34, Ano36]. Leipzig [Ano02, Ano13, Ano24, Ano30b, Ben30, Flü47, Fra31, Hei31, Her24, Jüt11, Jüt17, Jüt18b, Jüt21, Jüt22, Mis33, Pla93d, Pla94a, Pla09b, Pla39c, Pla52d, Pla53d, Pla55e, Reg34, Wei68, Zil33, Pla94a, Pla09b, Pla39c, Pla52d, Pla53d, Pla55e]. Leistung [Jor48]. Leitern [Pla89y]. Lekcii [Pla00e]. length [MR98]. Lense [GBM^{+}49]. Lenz [Ano48i]. Leo [Fra05]. Léon [Haa75, CS79a, Haa75]. Leopold [Mul00, Sch00b, HT99, Mul00, Sch04]. less [Kut14a]. letter [Tob08, HT99]. Letters [Prz86, Prz11, Eck97b, Prz63, Hie68, Mic68]. letzten [Haa70]. Levensbericht [van48]. liberated [Bra14b]. Library [Ano11, Ber14, C.49, Hie68, HS97]. Licht [HKS^{+}68, Kir98a]. Licht-und [HKS^{+}68]. Lichte [Pla26b]. Lichtes [Pla04a, Gae03, Pla02a]. Lichtquanten [Pla27b]. Lichts [Pla20a, Pla19b]. Liebenow [Pla07b]. Liebig [Ano53c, Ano53c]. Life [Ano97a, Hei88a, Hei06, Hop96, dACB58, Hei49, JDB^{+}97, Mül08a, Mül08b, PS58, van48]. Lifton [Hop96]. Light [FSBR09, Gam85a, Pla19b, Pla20a, Pla31g, Pla32j, Pla37g, Pla49j, Pla96, Pla15h, JDB^{+}97, Kir98a, Pla02a, Pla04a, Pla26b, Pla27c, Pla59a, Pla26e]. light-quanta [Pla27c]. L’image [Pla63a]. Limits [Pla49d, BRS03, Pla42d, Pla42e]. linear [Pla01g]. linearen [Pla01g]. lines [Kir98a, Pla15b]. Linien [Kir98a]. linked [Pla19d]. links [Abb01]. Lipson [KZR^{+}73]. Lise [Ano46b, Ano49b, Ano49c, Büh98b, Ano49b, Ano49c, Büh98b]. List [Mei47]. Listen [Tib13]. Literatur [HD01, Kra03]. Literature [HD01]. lithographische [Pla93d]. Liveing [Pla89k, Pla89l]. Lives [Ano23]. Local [BF12, KM12, Pla39c, Pla52d, Pla53d, Pla55e]. localization [MR98]. Logarithm [Ano48i]. Logarithmen [GBM^{+}49]. Logarithmen/Seeliger [GBM^{+}49]. loi [Pla12a]. London [Cra27, McK27, Ano46d]. long [Pla03g, Rub12]. Longmans [McK27]. longue [Wal94]. longueurs [Rub12]. Look [Bla92, Frö01]. Looking [JDB^{+}97]. Looks [O’F59]. l’orchestre [Pla94b]. Lorentz [Klo63, Pla27a, Pla28c, Pla78e, Prz63, Pla10f, Pla27a, Pla28c, Prz63, Prz86, Prz11, Hie68, Mic68]. Lorentz-Kontraktion [Pla10f]. Lorentz-Medaille [Pla27a]. Lorenz [Sta10]. löscht [Bre08]. lose [Kut14a]. lost [Bre08]. Lösung [Pla89k, Pla89l, Pla89x, Pla03e]. Lösungen [Pla88i, Pla88k, Pla88j, Pla88h, Pla90b, Pla90e, Pla29n, Pla87b, Pla87f, Pla88a, Pla88l, Pla27e, Pla89s]. Lösungs [Pla89r]. Lösungs- [Pla89r]. Lösungstheorie [Pla03f]. lotta [Ros12]. Low [Lon45]. Lowood [Ole81, Ole81]. Ltd [McK27]. Ludendorff [Pla23e, Pla23e]. Ludwig [Ano47e, Ano02, Bor48, Bor63b, Hai11, Kör61, van10, Ano08d, Kan08, Kör61, MdB63]. Luft [Pla89a, Pla89f, Pla89i]. lunga [Ros12].

M [Ano24, Ano48a, Ano75, Cra27, Ein06, Gae03, Hei31, Klo63, MP19, Pla88b, Pla88e, Pla88f, Pla88g, Pla88h, Pla89q, Pla22b, Pla22f, Pla26a, Pla29b, Pla35a, Pla37i, Pla38a, Pla38g, SCP19, Som33, WvLuAE18, van48, vLP21, Ano34, Ano39, Sch55, Wil37, Ano34, Ano41b, Mis33]. M. [Ano48a, Duh88, Gam85a, Mos42, MP19, Pla22b, Pla26a, Pla29b, Pla35a, Pla37i, Pla38a, Pla38g, SCP19, Shp58a, Shp58b, Wil37, van48, vLP21, LdB12]. même [Pla08a]. Mach [SR37, Yos95b, Bla88, Bla92, Hui13, Pla10i, Pla92, SR37, Thi69, Weg10]. Machschen [Pla10i]. Macht [Koh02, Her03]. Macke [Ber60]. Macrakis [Hop96]. made [Cli87]. magnetischen [Pla03h]. magnetism [Pla22e, Pla32h, Pla49h, Pla57b, Pla93c]. Magnetismus [GBM^{+}49, Pla22e]. Magnetismus/Brüche [GBM^{+}49]. magneto [Pla00b]. magneto-optical [Pla00b]. Magneton [Cha14]. magnetooptischen [Pla00b]. Magnus [Ano97a, Ano97a]. Magnus-Haus [Ano97a, Ano97a]. Mai [Pla27a, Pla27b, Pla38h]. Maier [Kas64]. Maier/Max [Kas64]. Maier/Max-Planck-Medaille [Kas64]. Mainzer [KAR89]. Maksa [Iva59b, Lau58, Pla00e]. Man [HB87, Hop96, Mei51, Pye88, San87, Swe89, Hei86a, Hei00, Cas86]. Management [Pla33g]. Managing [MKD^{+}91]. Mankind [Ano31, Dys70]. March [Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. Märchen [Jos95]. Maren [Kra03]. Mariotte’schen [Pla89e]. Mark [Ano36, Hop96, Ano30b]. markedly [Pla03e]. markedly-variable [Pla03e]. marks [JDB^{+}97]. marks/inertial [JDB^{+}97]. Martin [Hie68, Mic68, Ano61]. Marx [Sei08, Sei08]. März [Pla35c, Pla35d, Pla37d, Pla45a, Pla45b]. más [W^{+}87, WdC02]. masses [Bre08]. masters [Pla33j]. material [Pla33j]. materials [Pla88k]. Materie [Pla01e, Pla78b, JDB^{+}97, Pla01a, Pla02b, Pla44a]. Mathematical [BT58, BBG^{+}58, DT58, Kir94, KGSO55, Pla99a]. Mathematicians [Acz01, Acz03]. mathematics [Ber15, FF91]. Mathematik [Ber15, Ano48i, GBM^{+}49, HKS^{+}68]. Mathematik/Daeves [GBM^{+}49]. Mathematik/Dörge [GBM^{+}49]. Mathematik/Gravelius [Ano48i]. Mathematik/Landau [HKS^{+}68]. mathematisch [Pla93d]. mathematisch-reine [Pla93d]. mathematische [Kir94]. Mathematischen [Ano48i, Pla99a]. matrix [MR82b]. Matteo [Ber14]. Matter [Pla44a, Wei07, JDB^{+}97, Pla88i, Pla01e, Pla01a, Pla02b, Pla78b]. Maurice [Ano52]. Max [A.50, Adl03, Alb93, Ano02, Ano13, Ano24, Ano29b, Ano30b, Ano32, Ano34, Ano36, Ano37, Ano38b, Ano39, Ano41a, Ano46a, Ano46b, Ano46c, Ano47a, Ano47d, Ano47e, Ano48h, Ano48i, Ano49a, Ano49b, Ano49c, Ano53a, Ano53b, Ano53c, Ano58, Ano61, Ano69, Ano70, Ano97b, Ano08f, Ano08d, Ano11, Bad08, Bec09, Ben30, Ber14, BFV^{+}48, Bre08, Bro36, Brü48a, Büh98a, Cah99, Cra27, Dou49, Duh88, Dys70, EH08b, Ebe63, Eck97a, EP06, Fal59, Fal50, Fie75, Fis07, Fis10b, Fis10a, Fis12, Flü47, Fra31, Fra25, GBM^{+}49, Gri98, Gut15, Haa70, Haa75, Hah57, Ham72, HD48, Har59, HD01, HB87, Hei88a, Hei06, Hei31, HPE^{+}33, Hei35, Hei43a, Hei43b, Hei49, Hei69, HT99, Her03, Her72, Her77, Her24, Hib91, Hof96, HS97]. Max [Hof00, Hof08e, Hof08f, Hof09, HRR10a, HRR10b, Hop96, Jac08a, Jac08b, JDB^{+}97, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21, Jüt22, Kan69, Kan08, KAR89, KSB63, Koh02, Kör61, Kra97, Kra03, Kro52, KGSO55, Lem08, Len49, McK27, Mei47, MSQM63, Mic68, Mis33, Mul00, Nee80, OT97, Par12, Prz86, Pye88, RH97, Reg34, RS49, Ros12, Rus50, San87, Sch00b, Sch65, Sch04, Sch78d, Sch29, Sch61, Sch91, Sch55, Som33, Som36, Som38b, Som48, Spr89, Sta10, Swe89, SS53, Thi69, Tob08, Tre01, Ull96, Uns67, Vog61, Wei68, Wes57, WF63, Wil37, Wol08, Wor63, Yos95b, Zil33, vBL96, vL38, Abb01, And48, Ano29c, Ano31, Ano38a, Ano41b, Ano43a, Ano43b, Ano44, Ano46c, Ano47b, Ano47c, Ano48b, Ano48d, Ano48e, Ano48f]. Max [Ano48c, Ano48g, Ano48j, Ano48k, Ano49a, Ano49b, Ano49c, Ano52, Ano53a, Ano53b, Ano53c, Ano57, Ano58, Ano59, Ano69, Ano70, Ano75, Ano77, Ano81, Ano89, Ano94, Ano97a, Ano97b, Ano02, Ano04, Ano08f, Ano08a, Ano08b, Ano08c, Ano08e, Ano14, Anoxx, Bad08, Bad15, Bec09, Ber60, BFV^{+}48, BA01, BWN78, Bor48, Bor57, Bor63b, Bra14a, dACB58, dACB60, Bre08, Bro83, Brü48a, Brü48b, BPL^{+}54, Büh98a, C.49, CM08, Cas86, CS79b, Dah08, DdCA48, Del79, Din39, Din59, Din61, Dri01, Dys70, EH08a, EH08b, Ebe63, Eck97a, Eck97b, Eck08, Ein06, Ein13, Ein54b, Ein53, Ein91, EP06, Fad96, Fal59, Fie75, Fis00, Fis07, Fis08a, Fis10b, Fis10a, Fis12, Fli58, FV12, Fra48, Frö01, Gea08, Gla15, Gol76, Gre90, Gri98, Gri87]. Max [GMM47, GMM48, Gut15, Haa70, Haa75, Hac07, Hah57, Hai11, Ham72, HD48, Har98, Har38, Har43, Har44, Har48, Har53a, Har53b, Har58, Har59, Har64a, Har64b, HKS^{+}68, HD01, HH63, Hei86a, Hei88b, Hei88a, Hei00, Hei06, Hei07, HPE^{+}33, Hei35, Hei38, Hei43a, Hei47a, Hei47b, Hei49, Hei86b, Hei69, HT99, Her72, Her73b, Her73a, Her77, HR01, Her76, Her01, HS97, Hof99, Hof00, Hof08b, Hof08c, Hof08d, Hof08e, Hof08a, Hof08f, Hof09, HRR10b, Hof10, Hoh99, Hos78, Hui13, Hun75, Iva59a, Jac02, Jac08b, Jae79, JDB^{+}97, Jor48, Kan69, Kem38, Kho87, KSB63, Kle61, KMP59, Koe01, Koh02, Kör61, Kra02a, Kra02b, Kra00, Kre67, Kro52, KGSO55, Kut14b, Lan57, Law15, Lem08, Löh12, Low77, Lüs00, Max58, MH97, Mas12, Mei47]. Max [Mei51, Mei58, Mei01, MSQM63, MKD^{+}91, MAS63, Mey08, Moo23, Mos42, Mül08a, Mül08b, Mul94, Nag82, Nau16, O’F56, O’F59, OT97, PP79, Pic25, Pla18a, Pla19a, Pla29m, Pla29k, Pla29l, Pla39b, Pla42c, Pla48f, Pla48c, Pla57d, PB90, PS58, Prz11, Pye83, Pye85a, Pye08, Qui13, RH97, RB01, Ros59a, Ros59b, Ros79, Ros08, Rut29, Sac69, Sch49, Sch48, Sch78a, Sch78b, Sch78d, Sch91, SH53, Sin08, Som18, Som38a, Som38b, Som47, Som49, Spr89, Thi69, Thi88, Tib13, Tob08, Uli97, Ull96, Ull13, Uns58, Uns67, Vog61, WH88, Weg10, Wei07, Wei09, Wei58, Wes57, Wes58, Wol08, Zac92, Zaj83, Zen38, dB48, van10, vL38, vL47, vL48, vL50, vL58, Gea09, Hen60, Koy52, Ole81, Sta10, You68]. Max-Planck-Ausstellung [Uns67]. Max-Planck-Denkmal [Ano59, Uns58]. Max-Planck-Denkmal/Kongressitis/Internationale [Ano59]. Max-Planck-Feier [Hof96]. Max-Planck-Festschrift [Ber60, KMP59]. Max-Planck-Gesellschaft [Bec09, Cah99, Ebe63, Kra97, Sta10, vBL96, Zac92]. Max-Planck-Medaille [BT58, DT58, Fie75, Haa75, Kas64, KZR^{+}73, KGSO55, Pun67, Rut69, Ano52, Ano81, BWN78, Frö01, Hoh99, Hos78, Hun75, Lüs00, SH53]. Max-Planck-Vortrag [Dys70, Haa70, Hei69]. Maxwell [Pla31c, Bru02, FMM96, Pla94c, Pla95b, Pla98b, Pla99a, Pla03a, Pla31c, Pla31d, Tho31]. Maxwell’sche [Pla03a, Pla98b, Pla99a]. Maxwell’schen [Pla95b, Pla94c]. May [Pla27a, Pla27b, Pla38h]. Mayer [MdB63]. Mayer/Alwin [MdB63]. me [HS97]. mean [Fok14]. Meaning [Dar01, Pla49d, Ish15, Ish17, PB17, Pla10h, Pla42a, Pla42e]. Means [Luc30, Nye06]. measured [Cha97]. Measurement [Hol99, WSNO98, Gré19]. Measurements [Kan70, Hof48, Pla06d, Pla06e, Pla07a, Pla07c, Pla16a]. Mechanical [Pla10g, Pla15a, Pla15c, Cla91, Kir98b, Pla04b, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a]. Mechanics [Hie68, Mic68, Pla03c, Pla16c, Pla16d, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla32g, Pla49e, Pla57a, Pla64b, Pla15f, Prz63, CG00, CG03, CC03, Cha97, ET07, FGH91, MR82a, MR82c, MR82b, MR87, MR00, Pla06c, Pla28d, Pla33d, Pla37e, Pla40b, Pla40c, Pla41c, Pla49a, Prz86, Prz11, Wes40, HKS^{+}68]. Mechanics/Beiser [HKS^{+}68]. Mechanik [Moo23, Pla16c, Pla16d, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f, Pla06c, Pla37e]. Mechanische [Pla15a, Pla15c, Cla91, Kir98b, Pla04b]. mechanischen [Kir98b, Pla79, Pla89j, Pla89n, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a]. Medaille [Ano49b, Ano49c, Ano61, BT58, DT58, Fie75, Haa75, Kas64, KZR^{+}73, KGSO55, MdB63, Pla27a, Pun67, Rut69, Ano52, Ano81, BuDH07, BWN78, Frö01, Hoh99, Hos78, Hun75, Lüs00, SH53]. Medal [Ano29a, Ano29c, Ano49b, Ano49c, BuDH07, Fie75, Frö01, Haa75, KGSO55, Lüs00, Pla27a]. media [Pla04a]. Medical [Hop96]. Medicine [Ano46c, Hop96]. Medien [Pla05a]. Medium [Pla04a]. Medizin [Ano46b, Ano46c]. Medizin/Pasteur [Ano46b, Ano46c]. Medizinischen [Cah99]. Meer [Bre08]. Meeting [Pla20a, Dys70, Hof08e, Pla94a, Pla06f, Pla11c, Pla11d, Pla29c, Pla30c, Pla32b, Pla33g, Pla35a, Pla35e]. Meetings [JDB^{+}97]. meets [CH01]. mehreren [Pla15d, Pla15e]. Mein [Alb93, Pla47b]. meinen [OT97]. meiner [Ein07a]. meistert [Pla33j]. Meitner [Ano46b, Ano49b, Ano49c, Ano49b, Ano49c, Büh98b]. melting [Pla82]. Mem.R.S. [Ano38b, Ano48h]. member [Pla38g]. memoir [Duh88]. mémoire [Duh88]. Memorial [Ano48d, Ano48c, Ano48j, Pla48f, DdCA48, Kir79, SR37, Pla06f, Pla23f, Som48]. Memoriam [Ein54b, Ein53, Mei47, Pla46a, GMM48]. Memories [RH01, Hah57, Pla35e, Pla46b, Pla49k, BFV^{+}48, Pla30b]. memory [Pla94a, Pla14d, Som47]. Men [Cli87]. menneske [Har44]. Mensch [BPL^{+}54, Har38, Har43, Har48, Har53a, Har53b, Har58, Har64a, Har64b, Hen60, Mei58, Mei01, Wes58]. Menschheit [Dys70]. mental [Hen07]. mentality [Hen07]. Mérite [Wes57, Wes57]. merklich [Pla03e]. message [Pla19d]. Messtechnik [BBG^{+}58]. Messtechnik/Hönger [BBG^{+}58]. Messungen [Kan70, Pla06d, Pla06e, Pla07a, Thi71]. metafísica [FV12]. Metal [Pla03a]. Metalle [Pla03g]. metallic [Pla89y]. metallischen [Pla89y]. Metalloptik [Pla03a]. Metallurgy [Ber14]. metals [Pla03g]. metaphysical [FV12]. Method [Luc30, Iro05]. Methoden [BBG^{+}58]. methodology [Kon02]. Methods [KZR^{+}73, Sch17, Ris96, BBG^{+}58]. Michael [Hop96]. Mie [GBM^{+}49]. Millikan [SR37, Hol99, SR37]. Mind [Hür22]. minds [Pic08]. mir [HS97]. mira [Pla29o]. mirror [Hof08b, Lem08, Uli97, Ull13]. Mischung [Pla89u]. Mischungen [Kir98b]. Misprints [BT58, DT58]. místicos [W^{+}87, WdC02]. Misunderstood [Ano16]. Mitarbeit [Kra03]. Mitgliedsbeiträge [Ano81]. Mitt [Pla97d, Pla98c, Pla15d, Pla15e, Pla31f, Pla33e]. Mitt. [Pla29n]. Mitteilung [Pla97c, Pla97b, Pla99c, Pla19d, Pla27b, Pla41c, Pla17d, Pla17e]. Mittel [BBG^{+}58]. mittlere [Fok14]. mixtures [Kir98b, Pla83]. Mk [Jüt22, Ano02, Ano13, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21]. mode [Iro05]. Model [HN70, Tib13, Pad83, Pla19c, Pla24c, SNE01]. Modell [Pla19c]. Modern [Gib19, Pla31g, Pla37g, Pla96, Pla15g, Pla15h, Bra09, Din39, Hof08d, Hof10, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a, PS33, Pla59a, Pla63a, Pla29o, Seg07b, Sti84, Zaj83, HKS^{+}68]. moderne [Hof10]. Modernen [GBM^{+}49, Ano41b, Din39, Gea09, Hof08d, Mos42, Sti84]. modifications [MR82b]. modified [Pla14e]. moléculaires [Som12]. Molecular [Her01, Pla87b, Pla87f, Pla88l, SNE01, Som11, Som12]. Molecularconstitution [Pla87f, Pla88l]. Moleculargeschwindigkeiten [Pla88c]. Moleculartheorie [Pla89u]. Molecülen [Pla88c]. molecules [Pla94c, Pla95b, Pla15d, Pla15e, Pla22i]. Molekeln [Pla15d, Pla15e]. molekulare [Pla87b]. Molekularphysik [Som11]. Möller [Pun67]. moments [Büh98b]. monatomic [Pla08e, Pla25h]. monde [Pla63a]. mondo [Pla42b, Pla43a, Pla49c, Pla54a, Pla54b, Pla64a, Pla93a]. Monika [Hop96]. monument [Ano59, Qui13, Uns58]. monument/Congressitus/International [Ano59]. mood [Pla93b]. Moral [Ano31]. Moritz [Hof08b, EI08, Hof08b]. morning [Ano47a]. Morrison [BPL^{+}54]. Mosengeil [Pla07d]. Mössbauer [Pun67]. Mössbauer/Max [Pun67]. Mössbauer/Max-Planck-Medaille [Pun67]. Most [Bir19, Gut15, W^{+}87, WdC02]. motion [Cla91]. motive [Cla89]. Mott [MKD^{+}91]. moving [Pla07f, Pla07d, Pla08f]. MPI [Ano11]. Mr [KSB63, Pla29c, Sch29, Pla92b]. Mr. [Duh88, Pla00a, Pla00d, Pla03b, Pla13f, Pla29c]. Müller [MP19]. Multiple [Fra05]. München [Ano34, Ano36, Ano39, Ano41a, BWN78, Sac69, Sch55, Wil37]. München-Basel [Sac69]. München/Frühjahrstagung [BWN78]. Münchener [BDR86]. Münchner [Tob08]. mundo [W^{+}87, WdC02]. Munich [BDR86, You68, TEW02, Tob08]. Münster [Ano81]. Münze [Ano70]. Murphy [Bro36, Bro36]. Music [Ano53b]. Must [Tib13]. Musterbeispiele [Ano48i]. Mutter [BBG^{+}58]. My [Alb93, Ein07a, OT97, Pla58a, Pla47b]. Mystery [Acz01, FR13]. mystical [W^{+}87, WdC02].

N [Cra27, Hop96, Hof96]. Nach [Ano46b, Ano46a, Fis00, Frö01, Her77, Jac08b, Löh12, Pla89o, Pla93b, Pla19c, Pla39c, Pla52d, Pla53d, Pla55e]. Nachruf [Hei47b, SR37, vL50]. Nachtrag [Pla01b, Pla34c, Pla40c, Pla01c, Pla07c]. Nachweis [Ano48e, Ano48f]. Nägel [Pla38f, Pla38f]. nahe [Kir98b]. Name [Sei08]. Namens [Sei08]. Namens-Überraschung [Sei08]. Nanotube [FSBR09]. Natanson [Pla89o]. National [Her03, Hop96, Wei89, Hen96, Koh02, RW94, Wal94]. National-sozialismus [Her03]. Nationalsozialismus [Koh02]. Natur [Cla91, Gae03, Kut14a, MKD^{+}91, Pla02a, Pla24b]. Natur- [MKD^{+}91]. Natural [Pla49f, Fis00, FV12, Pla93b, Pla33c, PS33, Pla38d, WH88, Pla40a].

Naturanschauung [Pla10c, Pla10d, Pla10e, Pla11a, Pla10b]. Nature [Gol76, Inf61, Pla10g, Pla44a, Pla15g, RB01, Cla91, Kut14a, Pla02a, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a, Pla24b, Pla36a, Pla36c, Pla37h, Pla39c, Pla45d, Pla48e, Pla52d, Pla53d, Pla55e, Pla90b, Bro36]. Naturgeschehen [Pla33c]. Naturgesetzlichkeit [Pla38d]. natürlichen [Fis00]. natürlicher [Pla93b]. Naturwissenschaft [AH93, HD48, Pla38h, Pla48d, AH93, BBG^{+}58, Pla40a, Tre01]. Naturwissenschaft/Erckmann [BBG^{+}58]. Naturwissenschaften [Sch00b, Jüt18a]. naturwissenschaftlichen [Pla09c]. Nauchnaja [Pla58g]. Nauheim [BWN78]. nauki [PS33]. Nazi [Hop96, Abb01, Hac07, Mac93, MP01]. Nazis [Hop96]. Nazism [Hop96, Hop96, Wei89]. near [Ano46c]. nebst [WvLuAE18]. Necessary [v.B18]. Need [Law15]. neglected [Hof08e]. Nernst [Pla89x, Pla13g, Pla08a, Men73, Pla10h, Pla12c, Pla12d, Pla13g, Pla08a, SNE01]. Nernstschen [Pla10h]. Nernstsches [Pla12c, Pla12d]. net [McK27]. Neue [Ano46b, Pla98a, Pla11b, Pla25b, Ano70, Pla13h, Pla14f, Pla16e, Pla23h]. Neuen [GBM^{+}49, Ano29b, Ano30b, Ano53c, Ben30, Pla29d, Pla29g, Pla29e, Pla29f, Pla29h, Pla29i, Pla29j, Pla30d, Pla30e]. Neuer [Ano46b]. neuere [Pla96b, Pla12c, Pla12d]. neueren [Hol13, Pla91a, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a]. neuerer [Pla26b]. neues [Pla93b]. Neufeld [Hop96]. Neutrons [MKD^{+}91]. Neutrons/Belyaev [MKD^{+}91]. newer [Hol13]. Newton [Ano46d, Büh98b]. Newtons [Büh98b]. nicht [Pla42f]. nichtleitenden [Pla05a]. Nichtleitern [Pla02d, Pla03h]. Nickel [VD47]. Nicolò [Ber14]. Niederfrequenz [BBG^{+}58]. Niels [Adl03, Bro36]. No [Ano21, Ano31, Pic25]. Nobel [BCS82, Jüt21, Pla20b, Ano46c, Ano47c, Cra84, Fri81, JDB^{+}97, Nag82, Pla20b, Pla20c]. Nobel-Preis [Ano46b, Ano46c]. Nobel-Vortrag [Jüt21, Pla20b]. Nobelpreis [JDB^{+}97]. Nobelpreise [Ano61]. Nobelvortrag [Pla20c]. noir [Pla12a, War12]. Nomenclature [BT58, DT58]. Nomenclature/Additions [BT58, DT58]. Nominations [KK75]. Nominierung [BWN78]. Non [Ano77, Ole81, Low77, Pla05a, Som12]. non-conductive [Pla05a]. Non-Technical [Ano77, Ole81, Low77]. nonconformity [Hac07]. none [OT97]. Nonequilibrium [Hoh99]. Nonkonformität [Hac07]. nonperiodic [Som12]. Normal [Pla00i, Pla01d, Pla69b, Pla78a, Ano48a, Pla04a, Pla05a]. Normale [Pla05a]. normaler [Pla04a]. Normalspectrum [Ano48a, Pla01d]. Normalspektrum [Pla00i, Pla78a]. North [Adl03]. Norwegian [Har44]. Note [Har98, Has14, LSW95, Rut29, Pla92c, Pla97a]. Noted [Ano47c]. Notes [Bla78]. Notgemeinschaft [Pla29e, Pla29f]. notions [Pla32a]. Notiz [Pla92c, Pla97a]. Nov [Pla31e]. Nova [Ber14]. November [Hei31, Pla06f, Pla31e, Pla39c, Pla52d, Pla53d, Pla55e, Som48]. novembre [LdB12]. novyh [Pla26e]. Nr [Ano02]. NS [Hac07]. NS-Zeit [Hac07]. Nuclear [Ano53b, BPL^{+}54, Hop96, Sim12, Ano46c, BDR86, Kae48]. Nuclear-Fission [Sim12]. Nucleus [FR13]. Numerisches [Ano48i].

O [Ano46a, Pla47a]. Öberhambach [Pla19d, Pla19d]. obituary [Hei47b, SR37, vL50]. objection [Pla92b]. oblivion [Bre08]. obra [dACB58]. obrazu [Pla70]. Observability [AP00]. Observations [Pla23a]. observed [JDB^{+}97]. observed/Laser [JDB^{+}97]. occasion [HD01, Max58]. October [Ano97a, MH97, Pla14f, Pla16e, Pla23h, van48]. octobre [Ano48b, LdB12]. odd [Bad09]. Oettingen [Pla89c]. öffentl [Pla15a]. Offentliche [Pla29c]. öffentlichen [Pla22b, Pla29e, Pla29f, Pla30c, Pla32b, Pla35a]. Offered [Luc30, Pla30f]. Offers [Koe01]. Office [Pla13h, Pla14g]. Ogg [McK27]. Oktober [Ano97a, BWN78, MH97, Pla29m, Pla39b, Pla14f, Pla16e, Pla23h]. Old [tH67, Hei38, Pla46b]. One [Ano47c, Gib19, Hoh99, Wei58, HRR10b, Par12]. one-act [HRR10b]. open [Ber14]. opening [Ano11, Pla22f]. Operation [Hop96]. Operations [BT58, DT58]. Opinions [Ein54a, Ein82]. Oppenheimer [Pun67]. Oppenheimer/Unruhe [Pun67]. opponent [Löh12]. opposite [Wol08]. optical [Pla00b, Pla03g, Pla04a, KZR^{+}73]. optics [Pla03a]. Optik [Ein06]. optisch [Pla04a]. optischen [Pla03g]. orbits [Pla19d]. Order [Mer38, Fis00, Pla49b]. Ordnung [Fis00]. Organic [KZR^{+}73]. organizer [Kho87, Uli97, Ull13]. Origin [Pla22g, Pla10a, Sch00a, MSQM63, Sti84, Pla33a, Pla33g]. Original [Ped75, Kan72b, Kan72a]. Originated [Ano47c, Ano47d, Ano47a]. Ortsgruppe [Pla39c, Pla52d, Pla53d, Pla55e]. oscillations [Pla96c]. oscillator [Sch78b]. oscillators [Pla13i]. osmotic [Pla90a, Pla03e]. osmotischen [Pla90a, Pla03e]. Österreichischen [Ano11]. Ostertagung [Pla11c, Pla11d]. Ostwald [Kör61]. Ostwalds [Kör61, Ano02]. Oszillatoren [Pla13i]. Oszillators [Sch78b]. Other [A.50, Dou49, Pla49g, Pla50b, Pla68, Pla71, Pla14, Rus50, C.49, Hei69, Pla14i, Pla14j, Pla25f, Pla90b]. Ott [Pla30f]. Otto [Ano49b, Ano49c, Ano49b, Ano49c, Sim12]. our [Uns67]. Outline [Pla93c]. overall [Som11]. Overproduction [Pla33f]. Overview [Pla82].

P [Jac08a, Pic25, Pla89p, Pla89r, Pla89s, Pla37i, Ano48a, Ber60]. P. [Pla37i]. Paalzow [Pla08d]. Paarläufe [FS06]. pages [C.49, Hei31, Pic25]. Pain [Ano21]. pair [FS06]. Pamjati [Lau58]. Papapetrou [Ber60]. Paper [Pla20a, You68, Ein07a, Hof08a, Pla91b, Wes40, Pla13e]. Papers [Ano48e, Ano48f, CS79a, Hof08c, Jos95, Ped75, Pla49g, Pla50b, Pla68, Pla71, Pla14, WC91, C.49, Kan72b, KKRS96, Kan72a, A.50, Dou49, Rus50]. para [FV12]. Paradox [Ano31]. Parametern [Lüs00]. parameters [Lüs00, Pla35b]. Paris [KGSO55, Pla94b, Pla11c, Pla11d, KGSO55, Pla11c, Pla11d]. Paris/FA [KGSO55, KGSO55]. Parker [Pla88d]. Part [Cah99, MR87, Kör61]. Parthey [Cah99]. Partial [Sch17]. particle [Wes40]. particles [Boy79, BFG81, Fis10b, Fis12, Pla20a]. Paschen [Kan69, Kan69, PP79]. Paschen-Wiensche [Kan69]. Paschen. [PP25]. Päsler [Sch65]. Pasteur [Ano46b, Ano46c, Ano46c]. path [Shp58a]. paths [Pla13h, Pla14f, Pla14g, Pla16e, Pla23h, HPE^{+}33, Hei43a, Som33]. patriotism [vP06]. Patriotismus [vP06]. Patterns [Hoh99]. Paul [Hop96, Hos78, Pla06a, SH53, Ano46a, Ano52, Pla06a, Pla06f]. Pauli [BT58, DT58, MKD^{+}91]. Pauli/Falk [MKD^{+}91]. pedagogy [BN13]. Peierls [MdB63]. Peierls/Friedenspreis [MdB63]. Pellat [Pla89t]. Pennemunde [Hop96]. People [Tib13]. Perceptions [Pla15g]. periodic [Pla23d]. périodiques [Som12]. periodischer [Pla23d]. Permanente [Pla88f, Pla89q]. person [Har38, Har43, Har44, Har48, Har53a, Har53b, Har58, Har64a, Har64b]. personal [Hah57, Pla35e, Pla46b]. personality [Ano97a]. persönliche [Hah57, Pla35e, Pla46b]. Persönlichkeit [Ano97a]. Perspective [Sch22, Fri81]. Perspectives [Bla92, ET07]. Peters [Ano48i, GBM^{+}49]. Petra [Cah99, Kra03]. Pflichtgefühl [Fis08b]. Phantasies [Ano31]. phase [Pla16b]. Phasenraumes [Pla16b]. phenomena [Som12]. phénomènes [Som12]. Phenomenon [Tib13]. phil [Pla07b]. Philadelphia [Pla27b, Pla27b]. Philosoph [HKS^{+}68, Kre67, Sac69, You68]. Philosoph/Strubecker [HKS^{+}68]. Philosopher [Kre67]. philosophers [Acz03]. Philosophical [C.49, Hie68, MKD^{+}91, Pla36c, Pla39c, Pla52d, Pla53d, Pla55e, Hei58, Kro52, Pla70, Vog61, BDR86]. Philosophie [Har59, Her77, KAR89, Lan57]. Philosophischen [Pla39c, Pla52d, Pla53d, Pla55e, Hei58, Kro52, Pla36c, Vog61]. Philosophy [Gol76, Pla36b, CH01, ET07, Har59, Her77, Lan57, Mun36, Pla59a, Pla63b, Sce16]. Philosphers [Acz01]. Photoelectric [Mil16]. Photograph [Qui13]. Photographie [BBG^{+}58]. Photographie/Photographie [BBG^{+}58]. photography [KGSO55]. photon [Par94]. Photons [Gra69, Pla27b]. Phys [MSQM63]. Phys. [MSQM63]. Physicae [Her73b]. Physical [Ano48a, BT58, Brü48b, DT58, HW07, Pla06f, Pla11c, Pla11d, Pla18a, Pla35e, Pla38g, Pla50a, Pla60b, Pla93b, Pye83, Pye85a, Pye08, Som48, Wor63, v.B18, Gré19, HPE^{+}33, Hei43a, HRR10b, Max58, Pla09b, Pla09c, Pla10i, Pla13h, Pla14f, Pla14g, Pla16e, Pla16b, Pla23h, Pla27b, Pla27c, Pla29g, Pla42b, Pla43a, Pla43c, Pla49c, Pla54a, Pla54b, Pla64a, Pla70, Pla93a, Ros08, Som33, Hof08e, Pla94a, Pla26b, Pla28c, Pla58f, Pla26e]. Physicist [Ano47c, Ano47d, Anoxx, Ham72, JM17, Ano47a, Fis07, Fis10a, Hof09, HRR10b, Wei09]. Physicists [Hür22, KK75, Kir79, RTKK75, TEW02, Brü38, Cli87, Dys70, Hen07, Seg07b, W^{+}87, WdC02, dR85, HW07, KK75, Kir79, RTKK75, HKS^{+}68]. Physics [Acz01, Ano21, Ano53b, Ano11, BT58, BBG^{+}58, Cah85, Cah88, CH01, Cra84, DT58, Fra25, Fri81, Gam85b, Hen96, Her73b, JM17, Mes07, MKD^{+}91, Par12, Ped75, Pla09e, Pla10a, Pla30a, Pla22h, Pla25d, Pla30f, Pla31g, Pla32g, Pla32i, Pla33d, Pla36b, Pla37g, Pla49a, Pla49e, Pla49i, Pla49j, Pla57a, Pla64b, Pla96, Pla98, Pla12a, Pla15d, Pla15h, TEW02, Tib13, Bad09, BN13, Bae01, Ber15, BDR86, Bra09, Büh98b, BRS03, Cas09, Cli87, Din39, Eck97a, Eck97b, ET07, FF91, Haa70, Hei35, Her58, Hof08d, Hof10, Kae48, Kan72b, Kir94, Lon84, Lon03c, Mul94, Pic25, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a, Pla15g, Pla29d, Pla29g, Pla29e, Pla29f, Pla29h, Pla29i, Pla29j, Pla30d, Pla30e, Pla31c, Pla31d, Pla32a, Pla32f, Pla32d, Pla32e, Pla32h, Pla37c, Pla41b]. physics [Pla41a, Pla48a, Pla49h, Pla52a, Pla53a, Pla55a, Pla55b, Pla57b, Pla58b, Pla59a, Pla60b, Pla63a, Pla63b, Pla65a, Pla67a, Kan72a, Pla89, PD93, Pla15e, Pla29o, Set03, Som11, Spr89, Sti84, Str11b, WP85, Zaj83, Zim34, BPL^{+}54, BBG^{+}58, HKS^{+}68, Hei69, JDB^{+}97, KZR^{+}73, Pla32j, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla52c, Pla53c, Pla55d, Pla57c, Pla58d, Pla65c, Pla67c, Som36, Fal50, Ped75]. Physics/Blaschke [BBG^{+}58]. Physics/Inönu [HKS^{+}68]. Physics/Werner [BPL^{+}54]. Physik [Ano02, Ano29b, Ano34, Ano11, Ber15, Büh98b, Eck97a, Fra31, Gae03, Haa70, Hei35, KAR89, Pla94a, Pla06f, Pla10a, Pla10c, Pla10d, Pla10e, Pla10b, Pla11c, Pla11d, Pla11a, Pla29g, Pla29e, Pla29f, Pla29h, Pla30f, Pla31c, Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c, Som36, Spr89, Sti84, Pla38g, Ano30b, Ano36, Ano39, Ano41b, Ano41a, Ano48i, Ben30, Din39, GBM^{+}49, HKS^{+}68, Hei69, Her58, Hof08d, Hof08a, Hof10, JDB^{+}97, Kir94, MKD^{+}91, Mos42, Pla30a, Pla29d, Pla29i, Pla29j, Pla30d, Pla30e, Pla32d, Pla32e, Pla37c, Pla41a, Pla48a, Pla52a, Pla53a, Pla55a, Pla55b, Pla58b, Pla65a, Pla67a, Pye83, Pye85a, Pye08, Sch55, Wil37, Zim34, Gea09]. Physik- [Spr89]. Physik-Nobelpreis [JDB^{+}97]. Physik/Landau [HKS^{+}68]. Physik/Mie [GBM^{+}49]. Physik/Peters [Ano48i]. Physik/Planck [Ano48i]. Physik/Waerden [HKS^{+}68]. Physik/Westphal [Ano48i]. Physikalisch [Cah99]. Physikalische [Brü48b, Ebe63, HW07, Hof08e, Pla50a, Max58, Pla16b, Pla27b, Pla29g, Ano58, BBG^{+}58, Pla22h, Pla26b, Pla58f]. Physikalischen [Ano48a, BBG^{+}58, Pla18a, Pla28c, Pla35e, Som48, WvLuAE18, Ano32, HPE^{+}33, Hei43a, Mis33, Pla09b, Pla09c, Pla10i, Pla13h, Pla14f, Pla16e, Pla23h, Pla33h, Pla34e, Pla43c, Pla43b, Pla44b, Pla45e, Reg34, Som33, Zil33]. Physikalischer [Ebe63]. Physiker [Brü38, Fis07, Fis10a, Hof09, HRR10a, HRR10b, Jac08a, KK75, Kir79, Ano48i, Ano02, HW07, KK75, Kir79, MKD^{+}91, RTKK75]. Physiker-Gräber [Ham72]. Physiker/Kay [MKD^{+}91]. Physiker [?]ReihenentwicklungenAnonymous:1948:RWG. Physikern [KK75, Kir79]. Physikertagung [BWN78, Dys70, Haa70, Hei69]. Physikertagung/Fortbildung/Frühjahrstagung [BWN78]. Physikgeschichte [Eck97b]. physique [Pla41b, Pla63a, Pla89, PD93]. picture [Pla09c, Pla70]. pictures [Pla09b]. Pinpointing [Ano12]. Pionen [Lüs00]. pions [Lüs00]. pism [Pla70]. Planck [Abb01, Adl03, Alb93, Ano02, Ano13, Ano24, Ano29b, Ano30b, Ano34, Ano36, Ano38a, Ano38b, Ano39, Ano41b, Ano41a, Ano43b, Ano44, Ano46a, Ano46b, Ano46c, Ano48e, Ano48f, Ano48c, Ano48h, Ano48i, Ano48j, Ano49a, Ano49b, Ano49c, Ano52, Ano53b, Ano57, Ano59, Ano61, Ano70, Ano81, Ano89, Ano97a, Ano02, Ano04, Ano08a, Ano08c, Ano11, Ano14, Bad08, Bec09, Ben30, BF12, Ber60, Ber14, BA01, BuDH07, BT58, BWN78, Bor63b, Boy84, BFG81, Bre08, Bro36, Brü48b, BPL^{+}54, BBG^{+}58, Büh98a, C.49, Cah99, CG03, Cas86, CC03, Cra27, DT58, Din39, Din59, Dou49, Dri01, Dys70, EH08a, Ebe63, Eck97b, Ein06, Ein07a, Ein13, Fal59, Fie75, Fis00, Fis07, Fis08a, Fis10b, Fis10a, Fis12, Flü47, Fra31, Frö01, Gae03]. Planck [Gea09, Gri98, GMM48, Gut15, Haa70, Haa75, Ham72, HD48, Har38, Har43, Har48, Har53a, Har53b, Har58, Har59, Har64a, Har64b, HKS^{+}68, HD01, HH63, HB87, Hei88a, Hei06, Hei31, HPE^{+}33, Hei35, Hei38, Hei43a, Hei43b, Hei47b, Hei86b, Hei69, Hen60, HT99, Her03, Her72, Her73b, Her73a, Her77, HR01, Her76, Her24, Hie68, Hof96, Hof00, Hof08b, Hof08d, Hof08e, Hof08f, Hof10, Hoh99, Hop96, Hos78, Hun75, Jac02, Jac08a, Jae79, JDB^{+}97, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21, Jüt22, Kas64, KAR89, KSB63, Klo63, KMP59, Koe01, Koh02, Kör61, Koy52, Kra97, Kra02a, Kra02b, KZR^{+}73, Kre67, KGSO55, Kut14b, Lan57, Lem08, Löh12, Lüs00, Max58, MH97, McK27, Mei47, Mei58, Mei01, MSQM63, MKD^{+}91, MAS63, Mey08, Mic68]. Planck [Mis33, MdB63, Moo23, Mos42, Mül08b, Mul00, Nee80, OT97, Ole81, Par12, Pic25, PS58, Prz63, Prz86, Pun67, Pye88, Reg34, Ros12, Rut69, Sac69, San87, Sch49, Sch00b, Sch65, Sch04, Sch78c, Sch78a, Sch78b, Sch78d, Sch29, Sch55, SH53, Som18, Som36, SR37, Som38a, Som38b, Som47, Som48, Som49, Sta10, Swe89, TP76, Thi69, Thi88, Tob08, Tre01, Uli97, Ull13, Uns58, Uns67, Wei68, Wes57, Wes58, WF63, Wil37, Wol08, Yos95a, You68, Zac92, Zen38, Zil33, dB48, van48, vBL96, vL58, A.50, Aga67, And48, Ano16, Ano21, Ano23, Ano29a, Ano29c, Ano30a, Ano31, Ano32, Ano37, Ano38a, Ano43a, Ano43b, Ano44, Ano46c, Ano46d, Ano47a, Ano47b, Ano47c, Ano47d, Ano47e, Ano48a, Ano48b, Ano48d]. Planck [Ano48e, Ano48f, Ano48c, Ano48g, Ano48k, Ano49a, Ano49b, Ano49c, Ano53a, Ano53b, Ano53c, Ano57, Ano58, Ano59, Ano69, Ano70, Ano75, Ano77, Ano94, Ano97a, Ano97b, Ano08f, Ano08a, Ano08b, Ano08d, Ano08e, Ano12, Anoxx, Bad08, Bad09, Bad15, Bae01, Bec09, BFV^{+}48, BuDH07, BP24, Bir19, Bla78, Bla88, Bor48, Bor57, Boy79, Boy83, Bra14a, Bra14b, dACB58, dACB60, Bre08, Bro83, Bro15, Brü48a, Brü48b, Bru02, Büh98a, BRS03, Cah99, CH01, CG00, CM08, Cha14, Cha97, CS79b, Dah08, Dar01, DdCA48, Del79, Din46c, Din46b, Din46a, Din39, Din59, Din61, Duh88, Dys70, EH08a, EH08b, Eck97a, Eck97b, Eck08, Ein06, Ein07b, Ein13, Ein54b, Ein53, Ein91, EP06, Fad96, Fal59, Fal50, FSBR09, FP27, FGH91, Fis00, Fis07, Fis08a]. Planck [Fis10b, Fis10a, Fis12, Fli58, Foo16, FV12, FMM96, Fra48, Gam85a, Gar76, Gea02, Gea08, Gla15, Gol76, Gor91, GBM^{+}49, Gra69, Gre90, Gri98, Gri87, GMM47, Gut15, Hac07, Hah57, Hai11, Ham72, HD48, Har98, Har38, Har43, Har44, Har48, Har53a, Har53b, Har58, Har59, Har64a, Har64b, Has14, HD01, HH63, Hei86a, Hei88b, Hei88a, Hei88c, Hei00, Hei06, Hei07, Hei31, Hei38, Hei47a, Hei47b, Hei49, Hei58, Hei86b, Hei69, HT99, Her72, Her73b, Her73a, Her77, HR01, Her76, Her01, Her58, HN70, Hof96, HS97, Hof99, Hof00, Hof01, Hof08b, Hof08c, Hof08d, Hof08e, Hof08a, Hof08f, Hof09, HRR10b, Hof10, Hol13, Hol99, Hui13, HTD78, Inf61, Iro05, Iva59a, Jac08b, Jae79, JDB^{+}97, Jor48, Kan69, Kan70, Kan72b, Kan76]. Planck [Kan08, Kem38, Ken46, KM12, Kho87, KSB63, Kle61, Kle63, Kle66, Kle85, Koc91, Koh02, Kon02, Kör61, Kra76, Kra97, Kra00, Kre67, Kri86, Kro52, Kuh84, Kut14b, Lan57, Law15, Lem08, Len49, Lev04, LSW95, Lew30, Löh12, Lon45, Lon03a, Lon03b, Low77, Luc30, Max58, MH97, Mas12, MR82d, MR82e, Mei47, Mei51, Mei58, Mei01, Men17, MSQM63, Mes07, MAS63, Mey08, Mil16, MP19, Mül08a, Mül08b, Mul94, Mun36, NK06, Nag82, Nau16, Noe73, O’F56, O’F59, OT97, Pad83, Par94, PP25, PP79, Peg77, Pet14, Pla13f, Pla18a, Pla19a, Pla22b, Pla22f, Pla23e, Pla24a, Pla25c, Pla26a, Pla29c, Pla29b, Pla29k, Pla29l, Pla30c, Pla35a, Pla37f, Pla38a, Pla38f, Pla38g, Pla42c, Pla48f, Pla48c, Pla57d, Kan72a, Pla92, Dar00, PS58, Prz11, Pye83]. Planck [Pye85a, Pye08, Qui13, RH97, Ris96, RB01, Ros59a, Ros59b, Ros79, RS49, RS80, Ros08, Rub12, Rus50, Rut29, Sac84, Sch48, Sch00a, Sch03, SCP19, Sch78c, Sch78a, Sch78b, Sch78d, SP29, Sch61, Sch91, SNE01, Seg07a, Seg07c, Sei08, SDF14, Shp58a, Sin08, Som11, Som18, Som33, SR37, Som38a, Som38b, Som47, Spr89, SS53, Thi69, Thi88, Tib13, Tob08, Uli97, Ull96, Ull13, Uns58, Uns67, VD47, Vog61, WH88, War12, WvLuAE18, WB78, Weg10, Wei07, Wei09, Wei58, Wes40, Wes57, Wes58, Why50, WSNO98, Wol08, Wor63, Yos95a, Yos95b, Yos95c, Yos95d, Zaj83, Zen38, dB48, v.B18, vK97, van10, vL18, vLP21, vL38, vL47, vL48, vL50, vL58, Hib91, Kra03, Fra25, Ped75]. Planck-Gleichung [vK97]. Planck-Medaille [Ano49b, Ano49c, Ano61, BuDH07]. Planck-Plastik [KSB63, WF63]. Planck-Plastik/Bergmann-Schäfer [WF63]. Planck-Plastik/Einigkeit [KSB63]. Planck-Scale [Mes07]. Planck-Statue [Ano53b]. Planck-Statue/Staatsmänner [Ano53b]. Planck-type [Pet14]. Planck/Arbeitsgemeinschaft [HD48]. Planck/Der [Wes57]. Planck/Nachweis [Ano48e, Ano48f]. Planck/O [Ano46a]. Planck/Physicist [Ham72]. Planck/Physiker [Ham72]. Planck/Physiker-Gräber [Ham72]. Planck/Proof [Ano48e, Ano48f]. Planck/The [Wes57]. Planck/Todesurteil [Ano46b, Ano46c]. Planckia [Dri01]. Planckian [Sch22]. Plancks [Ano69, Hac07, HRR10a, HRR10b, Jor48, Lem08, Löh12, Mei47, Pla18a, Sch00b, SR37, Vog61, WvLuAE18, vP06, Hei49, Kro52, Spr89, Wes40, vL18, vP06, MR98, Sti91]. Planck’s-constant [MR98, Sti91]. Plancksche [Ein07a, Ein07b, Hei58, Som11]. Planckschen [Her58, Hol13, Kan70, Mül08b, Thi71, Yos95d]. Plane [MKD^{+}91]. Plank [Ano75]. Planka [Iva59b, Lau58, Pla00e, Shp58b]. Plant [HH63, MAS63]. Plasmaphysik [KZR^{+}73]. Plasmaphysik/Smith [KZR^{+}73]. plastic [KSB63]. plastic/Unity [KSB63]. Plastik/Bergmann [WF63]. Plastik/Einigkeit [KSB63]. Platonic [Her77]. Platonische [Her77]. play [HRR10b]. pledge [BA01]. Poincaré [Adl03, Pla21d, Kri86, Lev04, Pla21d]. Point [Ano23, KZR^{+}73, Boy83, Kir98b, Kra12]. Points [VD47]. Poisson [SNE01]. Polemics [Pla92]. Polish [PS33, Pla70]. political [Hac07]. Politics [Hop96, Sim12, Hof08f, Wei89]. Politik [Hof08f]. politischen [Hac07]. Poll [Ano69]. Poll/Three [Ano69]. Polyederverband [Pla19d]. polyhedra [Pla19d]. polytechnics [JDB^{+}97]. polytechnics/Einstein [JDB^{+}97]. popular [BDR86, Pla42f, Pla42f]. popularizer [Kho87]. Portrait [Fis08a, Pla48c]. Portraits [LMOS15, Ano53a]. Porträts [LMOS15]. portrayal [Pla42c]. Portuguese [dACB58, FV12]. Position [BF12, KM12, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a]. Positionierung [Hac07]. Positioning [BF12, KM12, Hac07]. positivism [Vog61, Hei31, Pla31e]. Positivismus [Hei31, Pla31e, Vog61]. Post [JDB^{+}97]. potential [Pla90b, Pla90e, Pla27e, Pla29a, Pla29n]. Potentialdifferenz [Pla90b, Pla90e, Pla27e, Pla29n]. Pour [Wes57, Duh88, War12, Wes57]. Power [Hop96, Cla89, Koh02, Pla23d]. Powers [Hop96]. Pp [Cra27, McK27, San87, Ber14, Hie68, You68]. Praeceptor [Her73b]. Praktische [GBM^{+}49]. Praktischen [Ano48i]. Präsident [Ano46b, Ano46c]. Präsidenten [BWN78, Her03, Koh02]. Präsidenten/3 [BWN78]. precision [Gré19]. predecessor [Abb01]. Preface [Pla33j]. prehistory [Yos95d, Kan70]. Preis [Ano02, Ano30b, Ano34, Ano36, Ano39, Ano41b, Ano41a, Ano46b, Ano46c, Ano61, Ben30, BWN78, Ebe63, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21, Jüt22, Mis33, Mos42, Reg34, Sch65, Sch55, Wei68, Wil37, Zil33]. Preisgeb [Ano13]. Prejudice [Cah88]. Preludes [Kra12]. Present [Pla20d]. presentation [Dys70, Ano49b, Ano49c]. presentations [Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b]. presented [Pla13e, Pla20a, Pla27b, Pla29e, Pla29f, Pla33g]. President [Ano11, Ano46c]. Presidents [Koh02]. Press [San87]. pressure [Pla88i, Pla88k, Pla90a, Pla03e]. preuß [Pla23g]. Preußische [FH00]. Preußischen [Sim06]. prevention [JDB^{+}97]. Price [C.49, Cra27, McK27, Pic25]. Pride [Cah88]. primary [Hen96]. Primordial [Noe73]. Princip [Pla87c, Pla87d, Pla87e, Pla88g, Pla91b, Pla15b, Pla92a]. principe [Duh88]. Principle [HTD78, LSW95, Pla7 , Duh88, Gor91, Pet14, Pla87a, Pla87c, Pla87d, Pla87e, Pla91b, Pla08a, Pla08b, Pla08c, Pla09a, Pla13b, Pla14a, Pla14b, Pla21b, Pla25a, Pla15b, Bla78, Pla92a, Pla06c]. Principles [BT58, DT58, Pla03c, Set03]. Princips [Pla89t]. Prinzip [Pla87a, Pla06c, Pla08a, Pla08b, Pla08c, Pla09a, Pla13b, Pla14a, Pla14b, Pla21b, Pla25a, Pla34a, Pla34b, Pla34c]. Prize [Ano46c, JDB^{+}97, Ano47c, Fri81, Nag82]. Prizes [Cra84]. Probable [Bir19]. Probing [FSBR09]. Probleme [GBM^{+}49]. Problèmes [BBG^{+}58]. Problèmes/Roth [BBG^{+}58]. problems [Pla46c, Set03]. Proceedings [Ano48a, Sim06]. processes [Pla97c, Pla97d, Pla97b, Pla98c, Pla99b, Pla99c, Pla00h, Pla01f, Pla01b, Pla01c, Pla78d, KZR^{+}73]. Processes/Arzeliès [KZR^{+}73]. Proctor [Hop96]. products [Pla05a]. Prof [Ano29b, Ano30b, Ano41b, Ano46b, Mis33, Mos42]. Prof. [Ano16, Ano38b, Ano47a, Ano47e, Ano48h]. Professor [Ano02, Ano29c, Cra27, Fra31, Ano29a, Ano37, Ano47a, Ano47d, Ano47e, Rut29]. professore [Ros12]. Profile [Cah99]. program [JDB^{+}97]. Programm [JDB^{+}97]. Progress [Ano31]. Project [Sim12]. Prologue [EM32, Ein77]. Promoting [Bra14b]. Promotions [Sch00b]. Promotions- [Sch00b]. Proof [Ano48e, v.B18, Ano48f]. proofs [Sch03]. properties [Pla03g]. prosperity [Bra14b]. Proved [Acz01, Acz03]. Prozesse [Sch65]. prüft [Thi88]. Prussian [Pla23g, Sim06, FH00]. Przibram [Hie68, Klo63, Mic68]. przyrodnicze [PS33]. Pseudo [Pla46c]. Pseudo-problems [Pla46c]. Psychology [Hop96]. PTR [Hof48, Hof48]. public [Pla15a, Pla22b, Pla29e, Pla29f, Pla32b, Pla35a, Pla29c]. publication [Her76]. Publications [Mei47]. Pulfrich [Pla15c]. Pure [BT58, DT58, Lem08, Pla10h]. Puschl [Pla89e]. put [Shp58b]. Pütter [Jüt18a].

QED [Par94]. Quanta [Gam85a, Kle63, Kle66, Kle85, LdB12, Pla01e, Pla11f, Pla13g, Pla23a, Pla27c, Pla78b, Pla08a, SDF14, Sin08]. Quantelung [Pla18b, Pla25h]. Quanten [Pla12c, Pla12d]. Quanten-Hypothese [Pla12c, Pla12d]. Quantenchromodynamik [Lüs00]. Quantenemission [Pla11f]. Quantenfeldtheorie [Haa70]. Quantenhypothese [Hol13, Pla13c, Pla13d, Pla14c, Pla15d, Pla15e, Kan70, Pla14e, Thi71]. Quantensprung [Fis10b, Fis12]. Quantenstatistik [Pla23d, Pla24c]. quantentheoretischen [Pla16a]. Quantentheorie [Bad08, Hei86b, Jüt21, Klo63, Pla20b, Pla20c, HD01, Pla17b, Pla21d, Pla22c, Pla58e]. Quantenwelt [Arr06]. Quantenwirkungen [Pla15i]. Quantisierung [Sch78b]. Quantitätsparameter [Pla35b]. quantités [Pla12a]. Quantities [BT58, DT58]. Quantities/Addition [BT58, DT58]. quantity [Pla35b]. quantization [Fad96, Pla18b, Pla25h, Sch78b, Seg07c]. quanto [Ros08]. Quantum [Ano31, Ano47c, Ano47e, Arr06, BDH^{+}00, Dar01, Eck08, ET07, Gam85b, Gea02, HKS^{+}68, Her71, Hof01, Hür22, KZR^{+}73, Lew30, Mas12, Nee80, Ped75, Pla20c, Pla20d, Pla22g, Pla10a, Pla15g, SB51, Tib13, tH67, Ano47a, Ano47d, Bad08, BN13, Bad15, Boy79, BFG81, CH01, Cas09, Cli87, Cra27, FGH91, Fis10b, Fis12, Haa70, HD01, Hei86b, Her01, Her58, Hol13, Ish15, Ish17, Kan70, Kan72b, Kle61, Kuh78, Kuh87, Lüs00, MR82a, MR82c, MR82d, MR82e, MR00, MR98, Nau16, Pad83, PB17, Pla12a, Pla12c, Pla12d, Pla13c, Pla13d, Pla13e, Pla13g, Pla14c, Pla14e, Pla15i, Pla15d, Pla15e, Pla16a, Pla17b, Pla20b, Pla21d, Pla22c, Pla23d, Pla24c, Pla43c, Pla58e, Kan72a, Pla08a, Ros08, SDF14, Som11, Som12, Sto13, Str11b, dB48, W^{+}87, WdC02]. Quantumtheorie [Her58]. quarks [Seg07b]. quasiparticles [JDB^{+}97]. Quasiteilchen [JDB^{+}97]. Quelle [Alb93]. Quellen [Bec09, Sta10, Lem08]. Quelleninventar [Cah99, Kra97, Ull96]. Quest [Hop96, Sto13]. question [Pla25h]. questions [Ano69, Hei58]. questions/Max [Ano69]. quiet [JDB^{+}97]. quiet/Spring [JDB^{+}97].

R [MdB63, Pic25, Pun67, Sch17]. Race [Hop96, Wei89]. Racial [Hop96]. Radar [Ano46b]. Radar/Zwei [Ano46b]. radiant [Pla00c, Pla13i]. Radiation [Aga67, Din46c, Din46b, Din46a, Ein07a, Has14, HN70, Ken46, Men17, Noe73, Pla06g, Pla13j, Pla14h, Pla19f, Pla21f, Pla23i, Pla59b, Pla66, Pla88a, Pla91b, Pla12b, Pla15i, v.B18, Bad15, Boy79, Boy83, BFG81, Ein07b, Fok14, FMM96, Gea08, Kan70, Kan76, Kon02, Lon03a, Lon03b, Par94, Pla96c, Pla97e, Pla97c, Pla97d, Pla97b, Pla98c, Pla99b, Pla99c, Pla00h, Pla01f, Pla01b, Pla01c, Pla07d, Pla10j, Pla11b, Pla12b, Pla12a, Pla13e, Pla23b, Pla24b, Pla78d, Pla95, Pla97, PHR07, Sch03, Sch78c, Sch78a, Seg07a, War12, Yos95d, Hof48, Ein06, Pla09d]. Radiological [dR85]. Radiozünder [Ano46b]. Radium [Kae48, Ano46c]. Radiumfund [Ano46b, Ano46c]. raised [Pla92b]. rally [Pla29e, Pla29f]. Ranked [Ano47c, Ano47e]. rapports [LdB12]. Rat [Ano29b, Ano41a]. Ratio [Pla14i, Pla14j, Pla25f]. rational [Pla03c, Pla49b]. Raum [KAR89]. Raumakustik [GBM^{+}49]. Raumakustik/Simon [GBM^{+}49]. ray [Pla62]. rayonnement [LdB12, Pla12a, Rub12, War12]. rays [Pla06d, Pla06e, Pla07a, Seg07b]. Reaction [Pla08a, Pla08b, Pla09a]. Reactionen [Pla89t]. reactions [Gar76]. Readings [Hib91, Gre90]. Ready [Fra05]. Reaktion [Pla09a]. Reaktionswärmen [BBG^{+}58]. Reaktionswärmen/Kurzes [BBG^{+}58]. real [Hei31, Pla31e, Pla40a]. reale [Hei31, Pla31e, Pla40a]. Realism [AP00, Sch00a, FV12]. realismo [FV12]. Realität [Pla27b, MKD^{+}91]. Reality [AP00, RS80, Pla27b, Pla27c]. realm [Kae48]. reasoning [Lon84, Lon03c]. reatises [Kir98b]. Receives [Ano29a]. receiving [KK75]. reception [MR82c, Pla57d]. Recheninstitut [Ano46b]. Recheninstitut/Zwei [Ano46b]. Rechnen [Ano48i]. Rechnen/Baur [Ano48i]. Recht [KSB63, MSQM63]. Recollections [HR01]. Recommendations [BT58, DT58, BT58, DT58]. Recommendations/Additions [BT58, DT58]. record [JDB^{+}97]. Recording [Brü38, Kir79]. Rector [Pla13h, Pla14f, Pla14g, Pla16e, Pla23h]. Recueil [BBG^{+}58]. Red [Inf61]. Rede [Jüt17, Pla94a, Pla13h, Pla14d, SR37, Pla14f, Pla16e, Pla23h]. Reden [Ano32, Dri01, HR01, Jac02, Jos95, Kra02a, Kra02b, Mis33, Pla22h, Pla23e, Pla33h, Pla43b, Pla44b, Pla45e, RH01, Som33, Zil33]. reference [Pla03c]. Regelungsmathematik [KGSO55]. Regia [BBG^{+}58]. Regia/Courant [BBG^{+}58]. Regime [MP01]. regularity [Pla14d]. Reich [Hop96, HW07]. Reihenentwicklungen [Ano48i]. reine [Pla93d, Pla10h]. reiner [Lem08]. Reinhard [Sac69]. Reinhardt [You68]. Reise [Ano46a]. Rektor [Ano11]. Rektorats [Pla14f, Pla16e, Pla23h, Pla13h]. Relation [Foo16, Pla98a, Hof00]. relations [Hof08b]. relationship [Kir98a]. relative [Pla25g]. Relativen [Pla25g]. relativistic [Boy79, BFG81, KZR^{+}73]. Relativität [Pla06c]. relativité [Lev04]. Relativity [Pla7 , Pye83, Pye85a, Pye85b, Pye08, Ano21, Gol76, Lev04, Pla06c]. Relativity’s [Ano21]. relevância [FV12]. Religion [Din59, HD48, Löh12, Pla49f, Pla38h, Pla48d, Tre01]. Religious [Anoxx]. Reluctant [Bra14a, Kra00]. Remark [Kut14b]. Remarks [Pla92a, Pla06b, Pla08a, Pla08b, Pla09a, Pla15c, Pla16a, Pla23d]. remembrance [Mei47]. Renate [Mul00, Sch00b, Sch04]. Renn [Ano11]. Renneberg [Hop96]. replace [Ano70]. Reply [Boy84, CG03, PP25, Sch61, TP76, Ano69, EP06, MP19, Mun36, Pla03f, Pla10i, Pla29c, SCP19, vLP21, BP24, FP27, PP79, Pla13f, Pla18a, Pla23e, Pla25c, Pla37f, Pla38f, SP29]. report [van48]. representation [Sch78c]. Reprint [SP29, Pla37i]. Republic [Hof00, WC91]. Republik [Hof00]. Research [BN13, Ber14, HH63, Hop96, MAS63, Ano46c, Ano94, Bre08, Mac93, Pla26b, Pla26e]. Research/What [MAS63]. Researcher [Ein13]. researchers [Bra14b]. resistance [HMF98, vP06]. resonance [Pla95a, Pla96a, Pla96c, Pla97e]. Resonanz [Pla96c, Pla97e, Pla95a, Pla96a]. resonator [Pla01g]. Resonators [Pla01g]. response [Ano69, Pla92b]. responses [Kir79]. Responsibility [Gri98]. Rethinking [Uns67]. RETRACTED [Pla40a, Pla42d]. Reuber [BBG^{+}58]. réunion [LdB12]. Reutter [GBM^{+}49]. reversibler [Pla89m]. Review [A.50, Ano32, Ano48i, Ber14, Bro36, Cah88, Cah99, Cas86, Cra27, Dou49, Fal50, Fra25, Gea09, HB87, Hei35, Hei43a, Hen60, Her03, Hib91, Hie68, Klo63, Koy52, Len49, McK27, Mic68, Moo23, Mul00, Ole81, Par12, Ped75, Pye88, RS49, Ros12, San87, Sch04, Sta10, Swe89, Thi71, Tre01, Wor63, You68, Hop96, Pla07c, Ein06, Som33, Som36]. reviewed [Thi88]. Reviews [BDR86, Rus50, HPE^{+}33]. Revisited [LSW95]. Revisiting [BDH^{+}00, Kuh84]. Revolution [Cah85, Kae48, BRS03]. Revolutionär [Bad08, Fis08b]. revolutionaries [Bru02]. Revolutionary [Bad08, Bra14a, CM08, Kra00, Seg07c, Wei09, Fis08b]. Revolutionize [Tib13]. Revolutions [Koc91]. Rezension [Ein06, Kra02a, Kra02b, Kra03, Sch00b, Som36]. Rezensionen [Ano02, Ano13, Ano24, Ano29b, Ano30b, Ano34, Ano36, Ano39, Ano41b, Ano41a, Ano48i, Ano89, Ben30, BPL^{+}54, Dri01, Flü47, Fra31, Gae03, GBM^{+}49, HKS^{+}68, Her24, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21, Jüt22, KAR89, KZR^{+}73, MKD^{+}91, Mis33, Mos42, Pla88b, Pla88c, Pla88d, Pla88e, Pla88f, Pla88g, Pla88h, Pla89d, Pla89h, Pla89f, Pla89g, Pla89p, Pla89e, Pla89i, Pla89j, Pla89m, Pla89k, Pla89l, Pla89n, Pla89q, Pla89o, Pla89c, Pla89t, Pla89r, Pla89s, Pla89w, Pla89v, Pla89u, Pla89x, Pla93d, Pla94b, Pla98a, Pla03c, Reg34, Sac69, Sch49, Sch65, Sch55, Som49, Wei68, Wil37, Zil33, Kra97]. Riecke [Pla15f]. riesige [Bre08]. rise [JDB^{+}97, MR82d, MR82e, MR87, Men73, Zaj83]. rise/Inertial [JDB^{+}97]. road [Bad15]. Robert [Hop96]. Rocket [Hop96]. role [Gri87, Haa70]. Rolle [Haa70]. romoting [Hei31]. Röntgenstrahlen [Pla62]. Roos [Dri01, Jac02]. Rosenfeld [CS79a]. Ross [Kra02a, Kra02b]. rotating [Fok14, Pla15h]. Rotation [Pla10f]. rotational [Hol13, Pla17f, Pla17g, Pla17h, Pla17d, Pla17e]. Rotationsspektrums [Pla17d, Pla17e, Pla17f, Pla17g, Pla17h]. rotatorischen [Hol13]. Roth [BBG^{+}58]. rotierender [Fok14, Pla15h]. Royal [Pla20b, Ano29c, Pla20c]. rubbish [Hof08a]. Rubens [Pla16a, Pla16a, Pla23f]. Rubinowicz [Rut69]. Rudolf [Pla94a]. Ruhm [MKD^{+}91]. Rundblicke [Pla22h]. Rundfrage [Ano69]. Rundfrage/Drei [Ano69]. Rundfunkröhren [BBG^{+}58]. Rundfunkröhren/Mutter [BBG^{+}58]. Runge [HT99, Mul00, Sch00b, Sch04, HT99]. Runges [Sch00b]. runs [FS06]. Russel [Kut14a]. Russian [Ano75, Pla26e, Pla29o, Pla58g]. Russische [Ano46b]. Russischer [Ano46b]. Rutherford [Ole81, SR37, Rut29, SR37]. rutschen [Bre08]. Rziha [BPL^{+}54].

S [Ano13, Ano34, Ano41b, Ano41a, Ano48a, Ano89, Dri01, Ebe63, Fra31, Jüt17, Jüt18a, Jüt18b, Mis33, Mos42, Pla93d, Reg34, Sac69, Sch55, Wil37, Zil33, Hof96, Hof99, Flü47, Fra31]. S. [Klo63]. Saal [Som38b]. safety [Gré19]. salts [Pla88j]. Salzburg [Dys70]. Salze [Pla88j]. same [Pla08a]. Sanden [GBM^{+}49]. Sättigungsgesetzes [Pla81]. saturated [Pla89b]. saturation [Pla81]. Saturday [Pla38g]. Satz [Kir98b, Pla17b]. Sätze [Pla00d]. Satzes [Pla88d]. Saves [Tib13]. say [Hof08a]. Says [Ano16]. Sbornik [Iva59b]. Scale [Mes07, Bae01, CH01]. scaling [MR98]. Schädliche [Ano53b]. Schäfer [WF63]. Schallplatte [Brü38]. Scheel [BWN78]. Scheinlichkeitsrechnung [BBG^{+}58]. Scheinlichkeitsrechnung/Gray [BBG^{+}58]. Scheinprobleme [Flü47, Pla46c]. Scherzer [GBM^{+}49]. Schicksal [HS97]. Schiffssteuerung [Ano46b]. Schlick [EI08, Hof08b, Hof08b]. Schlusswort [Kut14b]. Schmelzen [Pla82]. Schmetterlinge [Ano46b, Ano46c]. Schmidt [Pla30f, SCP19]. Schmidt-Ott [Pla30f]. Scholar [Ano16]. scholars [Hof08b]. school [Ano44]. Schreiber [KSB63]. Schreibtisch [Pla30b]. Schriften [BDR86, Hei88a, Hei06, BBG^{+}58]. Schröder [Tre01]. Schrödinger [Adl03, Ano11, Bro36, Hie68, Klo63, Pla29c, Prz86, KK75, MR87, Pla28d, Prz63, Prz11, SP29, Wes57]. Schrödinger-Symposiums [Ano11]. Schroedinger [Mic68]. Schule [Ano44]. Schulz [Ano48i]. Schumann [Rut69]. schwarzen [Pla12b]. Schwed [Pla20b]. Schweden [Ano46a]. Schweden/nach [Ano46a]. Schwedischen [Jüt21, Pla20c]. Schwefelsäure [Kir98b]. Schweiz [Ano46a]. Schweiz/nach [Ano46a]. Schwere [Pla89s]. schweren [JDB^{+}97]. schwingenden [Pla00f, Pla02c]. Schwingungen [Pla96c, Pla97e, Pla23d, Pla97a]. Schwingungsgesetze [Pla01g]. Science [AH93, Ano47c, Ano48g, Bro36, Gib19, HD48, Hei88a, Hei06, LMOS15, Mer38, MKD^{+}91, Nye06, Pla32k, Pla33i, Pla49d, Pla49f, Pla77, Pla15j, RW94, Sch96, Tre01, Wal94, WC91, Ano53b, Ano59, Brü38, FS06, FV12, GSU07, Gri87, Hei86a, Hei00, Hof08f, Hop96, Kho87, Koh02, Lev17, Lig05a, Lig05b, Mei03, Men73, MSQM63, Pic08, Pla30f, Pla33h, PS33, Pla34e, Pla38h, Pla40a, Pla42d, Pla42e, Pla42f, Pla43b, Pla44b, Pla45e, Pla46c, Pla48d, Pla59a, Pla81, Sce16, Ull13, AH93, BCS82, Bro36, Hop96, Pla09c, Pla29e, Pla29f, Pla30i, Pla37a, Pla88b, San87, HB87, Pye88, Swe89]. science/The [Ano53b]. Sciences [Ano48c, FH00, Hei31, Pla20b, Pla23g, Pla48f, Sim06, vL58, Pla42a, PP79, Pla20c, Pla31e]. scientia [Ber14]. Scientific [AP00, Hop96, Koc91, Nye06, Pla48c, Pla49g, Pla50b, Pla68, Pla71, Pla14, Sch00a, WC91, Ano69, BRS03, C.49, Her76, Jor48, Kör61, Mac93, Pla33a, Pla33g, Pla33f, Pla58a, RS80, Sti84, Uli97, Pla45f, Pla60a, Pla90c, Pla91a, Pla58g, A.50, Dou49, Rus50]. scientifique [Pla60a, Pla91a]. Scientist [Ano29a]. Scientists [Acz01, Ano30a, Ano31, Ano46d, MP01, MKD^{+}91, Acz03, Sce16, Str11b, Ano53c]. Scribal [MKD^{+}91]. sea [Bre08]. search [Fis00, Her77, Jac08b]. Sechsstellige [Ano48i]. sechzigstem [WvLuAE18]. sechzigsten [Som18]. Second [JM17, Cla91, Pla79, Pla29n, Löh12, Mül08b, Pla93a, Pla97d, Pla15e, Pla17e, Pla26d, Pla29a, Pla31f, Pla41c]. Secret [Hop96]. secretary [Hof00, Pla13f, Pla29c]. Secular [Noe73]. Seeliger [GBM^{+}49]. sehen [Pla88d]. Sehr [KSB63, Pla88j, Pla89e]. sein [Pla42f]. Seine [Vog61, Bre08, Kan69, Pla17b, Som11, Uns67]. seinem [Pla94a]. Seinen [Ano46b, Ano48c, Koy52, MH97, Pla48f, Sch49, Som49, Tob08]. seiner [Ano46a, Pla31c]. seines [Jüt18a, Max58, Pla29k, Pla29l, Pla46a]. seit [MAS63]. Seite [Pla99a]. Seiten [Ano02, Ano24, Ano30b, Ano36, Ano39, Ano89, Ben30, Hei31, Jüt21, Jüt22, Kra97, Sch49, Sch00b, Sch65, Som49, Wei68, Zil33]. Sekretär [Hof00]. Sekretars [Pla29c, Pla13f]. Sektion [Cah99]. selbst [Cla89]. selbständige [MSQM63]. Selbstbiographie [Pla48c, Ano48i, MKD^{+}91, Pla58a, PB90, Pla90c, RS49]. Selbstbiographie/Joos [Ano48i]. Selbstbiographie/Laurikainen [MKD^{+}91]. Selbstdarstellung [Pla42c, Pla82]. Selbstunterricht [Fra31, Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f]. Selbstzeugnissen [Her73a]. Selected [CS79a, Ano75, Bor63a]. selection [Hei88a, Hei06]. selective [Pla03h]. selektiven [Pla03h]. self [Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla42c, Pla15f, Pla58a]. self-biography [Pla58a]. self-instruction [Pla15f]. self-portrayal [Pla42c]. sense [Fis08b, Pye83, Pye85a, Pye08]. sentence [Ano46c]. separation [Hof96]. September [Haa75, Pla10c, Pla10d, Pla10e, Pla11a]. Series [BDR86]. serious [KSB63]. service [DdCA48, Ano48j]. session [Pla15a, Pla22b]. sets [JDB^{+}97]. setzt [JDB^{+}97]. Seven [PHR07, Pla23b, Pla95, Pla97, Sce16]. seventh [McK27]. seventieth [Pla38c]. several [Pla15d, Pla15e]. Shaped [Gib19]. Shift [Inf61]. Shining [Gut15]. Shook [Gam85b]. short [KGSO55]. short-time [KGSO55]. Shows [Ano23]. SI [Qui17]. sich [Cla89, JDB^{+}97]. side [Pla99a]. Sieben [PHR07, Pla23b, Pla95, Pla97]. siebzigsten [Pla38c]. significance [FV12, Her58, Pla13c, Pla13d, Pla14c]. signification [Pla42d]. Simon [GBM^{+}49]. Simplified [Pla01g]. Simson [Wes57, Wes57]. since [MAS63]. Single [FSBR09, Hür22, Pla16f]. single-atom [Pla16f]. Sinn [Pla42a, Pla42d, Pla42e]. Sir [Adl03, Rut29, Ano48j]. Sitzg [Pla06f]. Sitzung [Hof08e, Pla94a, Pla15a, Pla22b, Pla29c, Pla30c, Pla32b, Pla33g, Pla35a]. Sitzungsberichte [Sim06]. sixtieth [Som18]. Sixty [FR13]. Skurrile [Arr06]. slip [Bre08]. Small [Ano11, Bad08]. smallest [Fis10b, Fis12]. Smith [KZR^{+}73]. so-called [Din39]. Social [Mer38, Bro83]. Socialism [Hop96, Hen96, Hop96, Koh02, RW94, Wal94]. société [Pla08a]. Society [Ano29c, Gib19, Hei31, Koe01, Par12, Pla94a, Pla28c, Pla31e, Pla36c, Pla37a, Som48, BCS82, Abb01, Ano48a, Ano14, Bec09, BA01, Brü48b, Gri98, HD48, HW07, Hof08e, Koh02, Pla06f, Pla11c, Pla11d, Pla18a, Pla20a, Pla35e, Pla38g, Pla39c, Pla42e, Pla50a, Pla52d, Pla53d, Pla55e, Pla08a, vBL96]. Society/Clausthaler [HD48]. Socratic [EM32, Ein77]. sogenannten [Ano41b, Din39, Mos42]. Sohn [Ebe63]. soll [Ano70]. Solomon [Ano48i]. solution [Pla03f, Pla03e, Ris96]. solutions [Pla87b, Pla87f, Pla88a, Pla88i, Pla88k, Pla88j, Pla88l, Pla90b, Pla90e, Pla27e, Pla29a, Pla29n]. Solvay [Pla13e, LdB12, Pla13e, Str11a]. Solvay-Kongreß [Pla13e]. som [Har44]. Some [Gar76, Kir98b, Hah57]. Sommerfeld [Eck97b, Pla38b, Pla38c, Eck97b, MR82d, MR82e, Pla38b, Pla38c]. Sonderdruck [Pla37i, SP29]. Sonnabend [Pla38g]. Sonne [Ano53b]. Sonne/Max [Ano53b]. Soul [Bra14a]. Source [Ull96, Alb93, Her76]. Sources [Ber14, Bec09, Hen96, Lem08, HKS^{+}68]. sous [LdB12]. sovremennoj [Pla29o]. Sowie [Pla15f, Fra31, Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b]. sozialismus [Her03]. Space [Ano21, Ano31, Pla16b]. Spanish [Pla47a, W^{+}87, WdC02]. Spannung [Kir98b, Pla98a]. Spannungen [Pla98a]. Speaking [Cah88]. Special [Ano08e, Ano21, Ano48k, Gol76]. specific [Ein07b]. specifics [Cla91]. specifische [Pla88e]. Spectra [Ano23]. Spectral [Ano23, Kan70, Pla00g, Pla15b, Pla16a, Pla78c]. Spectroscopie [Kay02]. Spectrum [Pla00i, Pla01d, Pla69a, Pla69b, Pla78a, Boy79, Boy83, Boy84, BFG81, Kra76, Lon03a, Pla17f, Pla17g, Pla17h, Pla17d, Pla17e, Sac84, TP76, Ano48a]. speculations [Kra12]. Speech [Pla13h, Pla14g, BP24, EP06, FP27, MP19, PP25, PP79, Pla06f, Pla14f, Pla16e, Pla22f, Pla23f, Pla23h, Pla28c, Pla29c, SP29, Sch61, SR37, vLP21, Pla94a, Pla13a, Pla14d, Pla15a, Pla17a, Pla19a, Pla22b, Pla24a, Pla26a, Pla28a, Pla30c, Pla32b, Pla35a, Pla37b, SR37]. Speeches [BDR86, HR01, Jos95, RH01, Som33, Ano48c, Kir79, Pla18a, Pla23e, Pla37f, Pla38f, Pla48f, Pla33h, Pla43b, Pla44b, Pla45e]. spektralen [Kan70, Thi71]. Spektralgleichung [Pla00g, Pla78c]. Spektrallinien [Pla15b]. Spektralmessung [Pla16a]. spezifischen [Ein07b]. Spiegel [Hof08b, Lem08, Uli97, Ull13]. Spirit [Gib19, Pla33j]. spiritual [Hof96]. Spitzer [BBG^{+}58]. Spline [Sch17]. Splitting [Ano47c]. Spokesman [HB87, Hop96, Pye88, San87, Swe89, Hei86a]. Spookiest [Acz01, Acz03]. spring [Pla10a, JDB^{+}97]. Springer [Dri01, Jüt18a, Klo63, Ano14]. Springer-Verlag [Klo63]. srodowiska [PS33]. Staatsmänner [Ano53b]. stabiles [Pla35b]. stable [Pla35b]. staircase [Fis10b, Fis12]. stamps [Ano53c]. Stand [Ano16]. Standpunkt [Pla88e, Pla88f, Pla89q]. Stärke [KSB63, MSQM63]. Stärke/Sehr [KSB63]. Stärke/Vom [MSQM63]. Starkstromtechnik [BPL^{+}54]. Starkstromtechnik/Veröffentlichungen [BPL^{+}54]. Stars [Gut15]. State [Pla20d, Pla08e]. Statement [Ano16]. statements [Pla00d]. states [Pla80, Sti91, JDB^{+}97]. States/UK [JDB^{+}97]. Statesmen [Ano53b]. stationär [Pla01g]. stationären [Pla07d]. stationary [Pla01g, Pla07d]. statistica [Ros12]. Statistical [CS79b, Dah08, Ros79, CG00, CG03, CC03, Pla03c, Pla14d, Pla17b, Pla25b, Pla25e, Ros59a, Ros59b]. statistics [Bad09, FMM96, Pla23d, Pla23a, Pla24c]. statistique [Ros59a, Ros59b]. statistische [Jüt17, Pla14d, Pla25b, Pla25e]. statistischen [Pla17b]. Statt [Ano46b, Ano46c]. statue [Ano53b].

Statue/Staatsmänner [Ano53b]. statue/Statesmen [Ano53b]. Stella [Dri01]. Stellung [Pla10c, Pla10d, Pla10e, Pla10b, Pla11a]. Sterne [BPL^{+}54]. Sternstunden [Büh98b]. Stifters [Jüt17, Pla14d]. Stiftung [Ano41b, Mos42]. still [Kae48]. Stillhalten [JDB^{+}97]. Stillhalten/Frühjahrstagungen [JDB^{+}97]. Stimme [Brü38]. Stimmung [Pla93b]. Stochastic [Dah08]. Stockholm [Ano46a, Jüt21, Pla20b, Pla20c, Pla20c]. Stockholm/nach [Ano46a]. Stoff [Pla33j]. Stoffe [Pla88i, Pla88k, Pla88h]. Stoffen [Pla89u]. Story [Acz01, Gam85b, MP01, Acz03]. Strahlen [Pla06d, Pla06e, Pla07a]. strahlender [Pla00c, Pla13i]. Strahlung [Ein07a, Ein07b, Pla96c, Pla97e, Pla09d, Pla12b]. Strahlungs [Pla07d]. Strahlungsfeld [Fok14]. Strahlungsgesetz [Kan69, Sch78a]. Strahlungsgesetze [Pla23b, Pla95, Pla97, PHR07]. Strahlungsgesetzes [Kan70, Thi71, Yos95d]. Strahlungshypothese [Pla11b]. Strahlungsmessungen [Hof48]. Strahlungstheorie [PHR07, Pla23b, Pla95, Pla97]. Strahlungsvorgäne [Pla99b]. Strahlungsvorgänge [Pla97c, Pla97d, Pla97b, Pla98c, Pla99c, Pla01b, Pla01c, Pla00h, Pla01f, Pla78d]. Streitfragen [Pla45f]. strength [Kir98a, KSB63, MSQM63]. strength/Dear [KSB63]. strength/The [MSQM63]. Strictly [Sch22]. Strubecker [HKS^{+}68]. Structure [Why50, BRS03, Pla16b]. struggle [Hei35]. Struktur [Pla16b]. student [JDB^{+}97, Pla09c]. studentenkorps [Pla09c]. Students [Ano21]. Studien [vBL96]. Studienzeit [Tob08]. Studierender [JDB^{+}97]. Studies [Hui13, WC91, vBL96]. study [Fis10b, Fis12, Tob08]. Stuttgart [Ano89, AH93]. sua [dACB58]. subliming [Pla82]. Sublimiren [Pla82]. Subsequent [Pla07c]. Success [Law15]. Suche [Fis00, Her77, Jac08b]. Sucher [MKD^{+}91]. sulfuric [Kir98b]. Summer [WC91]. Sun [Ano53b]. Sun/Max [Ano53b]. suo [Ros12]. Superhero [Tib13]. Superlattice [KAR89]. Superposition [Pla23d]. supposed [Pla00b]. Surface [KAR89]. surmounting [Yos95a]. surprise [Sei08]. Survey [Fra25, Pla22h, Pla25d, Pla60b, Pla93b, Wor63, Pic25]. Surviving [Mac93, Hop96]. svete [Pla26e]. Swabian [Sto13]. Swastika [Hop96, Mac93]. Sweden [BCS82]. Swedish [Pla20b, Pla20c]. świata [Pla70]. Swinburne [Pla03b]. Swiss [KKRS96]. Symbols [BT58, DT58]. Symbols/Additions [BT58, DT58]. Symposium [Ano11, KGSO55, BCS82]. Symposiums [Ano11]. Synthese [Pla40b, Pla40c, Pla41c, Wes40]. synthesis [Pla40b, Pla40c, Pla41c, Wes40]. System [BF12, KM12, Pla93b, Pla15h, Sti84]. Systeme [Pla07f, Pla08f]. Systems [Dah08, Sti84, Pla07f, Pla08f, MKD^{+}91]. Systems/Elster [MKD^{+}91]. Szilard [Fra05].

T [Yos95c]. T. [Yos95c]. Table [Som38b]. Tafel [Pla93d]. Tafeln [Ano48i, GBM^{+}49]. Tafeln/Bertele [GBM^{+}49]. Tafeln/Gravelius [GBM^{+}49]. Takes [Ano23, Ano23]. tale [Jos95, Sce16]. Talk [Som38b, Pla08a]. Talks [Ano21]. Tartaglia [Ber14]. Tausend [Ano39]. Teacher [Her73b, Hof08b, Kho87]. teachers [OT97]. teaching [Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b]. Tech/Scribal [MKD^{+}91]. Technical [Ano77, Ole81, Low77]. Technik [AH93, Sch00b]. Technischen [Cah99]. Technology [Hop96, RW94, AH93, BCS82]. Teil [Kör61]. Teilchen [Fis10b, Fis12]. Tells [Ano21, FS06]. Temperament [Bro36]. Temperatur [Pla00c, Pla04b]. Temperaturänderung [Pla89g]. Temperaturänderungen [Pla89g]. Temperature [Lon45, CG00, CG03, CC03, Pla00c, Pla04b, Pla11c, Pla11d, Pla11c, Pla11d]. Temperaturen [Kir98b, Pla80, Pla98a, Pla89e]. temperatures [Kir98b, Pla80]. tension [Kir98b]. tenth [Wes57]. tenue [LdB12]. Teorisine [Pla15g]. termodinamike [Pla00e]. Test [Ano13, BF12, KM12, FGH91]. Teuerungszuschläge [Jüt22]. Text [Jüt18b]. textbooks [BN13]. Texts [Hib91, Gre90]. thanks [Pla27a]. Their [AP00, Ano16, Ole81, Fal59, Hof08b, Hof08a, Pla06d, Pla06e, Pla07a, Seg07b, vBL96]. them [Pic08]. themselves [Cla89]. Theodosius [Adl03]. theorem [Pla12c, Pla12d, Pla13g, Pla17b, Pla08a, Pla05b, Pla10h]. théorème [Pla13g, Pla08a]. Theoretical [Fal50, JM17, Lon84, Lon03c, Pla09e, Pla10a, Pla30a, Pla32g, Pla32i, Pla32j, Pla33d, Pla49a, Pla49e, Pla49i, Pla49j, Pla57a, Pla57c, Pla64b, Pla98, Pla12a, Pla15d, Din39, Eck97a, Pla15g, Pla31c, Pla31d, Pla32f, Pla32h, Pla49h, Pla57b, Pla15e, Set03, Pla30f]. Theoretiker [Lem08]. Theoretische [Fra31, Pla10a, Pla30a, Pla31c, MKD^{+}91, Pla30f]. Theoretischen [HKS^{+}68, Ano41b, Din39, Eck97a, Fra31, Mos42]. Theorie [Ano24, Ano48a, Ein06, Ein07a, Ein07b, Fra31, Her24, Jüt22, Pla81, Pla84, Pla89o, Pla89y, Pla92c, Pla97a, Pla98b, Pla99a, Pla00i, Pla02d, Pla03h, Pla06g, Pla07g, Pla08g, Pla10i, Pla10j, Pla13j, Pla17f, Pla17g, Pla17h, Pla17d, Pla17e, Pla19f, Pla21f, Pla22e, Pla23i, Pla35f, Pla35g, Pla35h, Pla66, Pla15c, Pla15i, Sch78c, LdB12, Lev04, Som12, Pla03a, Pla07d, Wei68]. Theorieauffassung [Spr89]. Theorien [Kan70, Pla12c, Pla12d, Pla14i, Pla14j, Pla25f, Thi71]. theories [CH01, Kan70, Pla12c, Pla12d, Pla14i, Pla14j, Pla25f]. theorist [Lem08]. Theory [Acz01, Ano21, Ano23, Ano47c, Bro36, Eck08, Ein07a, Gam85b, Gol76, Gra69, Her71, HN70, Hof01, Inf61, Mas12, MKD^{+}91, Nee80, Pla00i, Pla06g, Pla10g, Pla13j, Pla14h, Pla17g, Pla17h, Pla19f, Pla20c, Pla20d, Pla21f, Pla22g, Pla23c, Pla23i, Pla32h, Pla32i, Pla32j, Pla49h, Pla49i, Pla49j, Pla57b, Pla57c, Pla59b, Pla60b, Pla66, Pla69b, Pla88a, Pla91b, Pla93b, Pla93c, Pla10a, Pla12b, Pla15a, Pla15g, Pla15c, Pla15i, SB51, Tib13, tH67, Acz03, Ano47a, Ano47d, Ano47e, Ano48a, Bad08, Bad15, Boy79, BFG81, Cla91, Cla89, Cli87, Ein06, Ein07b, FMM96, Haa70, HD01, Hei58, Hei86b, Her01, Her58, Kan70, Kir98b, Kle61, Kon02, Kuh78, Kuh87, Lev17, Lon03b, Lüs00, MR82d, MR82e, Pet14, Pla81, Pla84, Pla89y, Pla91a, Pla92c, Pla97a, Pla98b, Pla99a, Pla02d]. theory [Pla03h, Pla03f, Pla03a, Pla07g, Pla08g, Pla09f, Pla10i, Pla10j, Pla13c, Pla13d, Pla14c, Pla17b, Pla17f, Pla17d, Pla17e, Pla20b, Pla21d, Pla22c, Pla22e, Pla23b, Pla35f, Pla35g, Pla35h, Pla58e, Pla95, Pla97, PHR07, Sch03, Sch78c, Som12, Spr89, HKS^{+}68, Lev04, Pla07d, Cra27, Wor63]. Theory/Herausgeg [HKS^{+}68]. there [Pla49b]. thermal [Boy79, BFG81, Ein06, Pla13e, Sch78c]. Thermochemie [Pla93c]. thermochemistry [Pla93c]. Thermodymamics [Dah08]. thermodynamic [EH08b, Pla83, Pla12c, Pla12d, Pla08b]. thermodynamical [vL18]. Thermodynamics [Kle66, Kle85, Pla03d, Pla21e, Pla26c, Pla27d, Pla45c, Pla69c, Pla90a, EH08a, Mül08a, Mül08b, Par94, Pla79, Pla97f, Pla00a, Pla00b, Pla02e, Pla05c, Pla10h, Pla11e, Pla13g, Pla13k, Pla17c, Pla21g, Pla22j, Pla27f, Pla30h, Pla54c, PP64, Pla08a, Pla10b, BBG^{+}58, Mül08b, Pla93a, Pla26d, McK27, Pla03c]. Thermodynamics/Bruhat [BBG^{+}58]. Thermodynamik [Ano02, Mül08b, Pla00a, Pla02e, Sch65, Ano13, EH08a, Jüt11, Jüt18b, Pla97f, Pla00b, Pla05c, Pla10h, Pla11e, Pla13k, Pla17c, Pla21g, Pla22j, Pla26d, Pla27f, Pla30h, Pla54c, PP64, Pla10b, Sch65, Pla88e, Pla88f, Pla89q]. thermodynamique [Pla13g, Pla08a]. Thermodynamische [EH08b, vL18, Pla83, Pla89v, Pla12c, Pla12d, Pla08b]. thermodynamischen [Pla89c]. thermodynamischer [Pla88b]. Thermoelectricität [Pla89y]. thermoelectriciy [Pla89y]. thesis [OT97]. Thilo [BBG^{+}58]. things [Hei69, MSQM63]. thinker [Har38, Har43, Har44, Har48, Har53a, Har53b, Har58, Har64a, Har64b]. thinned [Pla33b]. Third [McK27, Pla87e, HW07, Pla97b, Pla33e, Pla37a]. Thirty [Gam85b]. Thomas [Hop96]. Thought [MKD^{+}91]. thoughts [Kro52]. threaten [Bre08]. Three [Ano69]. Tilley [KAR89]. Time [Ano31, WB78, BCS82, KGSO55, Lev17, Pad83, Tob08, Uns67]. times [Pla46b]. Tischrede [Som38b]. titled [Pla60b]. titre [Duh88]. Tobies [Sch00b, Sch04, Mul00]. Todes [Wes57]. Todestag [Ano97a, Ano97b, Eck97a, Pla57d, RH97]. Todestagam [MH97]. Todestages [Ham72]. Todesurteil [Ano46b, Ano46c]. Tölz [Ano46b, Ano46c, Ano46c]. Tölz/Dr [Ano46b, Ano46c]. Tonsystem [Pla93d]. Too [Hür22, KSB63]. top [Pla18b]. topic [Pla23b, Pla95, Pla97, PHR07]. torn [Hei88c]. Tot [Ano46b]. Tot/Erste [Ano46b]. tower [Jos95]. tradition [Uli97, Ull13]. Trägheitsfusion [JDB^{+}97]. training [JDB^{+}97]. Trans [Sch22]. Trans-Planckian [Sch22]. Transcripts [Hop96]. Transfer [Lon45]. transform [KGSO55]. Transformation [KGSO55, Pla94b]. Transforms [KZR^{+}73]. Transforms/Greenwood [KZR^{+}73]. Transgressive [Gre90, Hib91]. Translated [Hie68, C.49, Dou49, McK27, Mic68, Pic25]. Transport [HKS^{+}68]. Transurane [Ano46b]. traps [JDB^{+}97]. traps/in [JDB^{+}97]. Traueransprache [Pla48c]. treasury [FF91]. Treatise [Pla03d, Pla21e, Pla26c, Pla27d, Pla45c, Pla69c, Pla90a, McK27]. treatises [Max58, Pla23b, Pla58f, Pla95, Pla97, PHR07, Kir98a, Pla15a]. treatment [Pla87c, Pla87e]. Trennung [Hof96]. Treulichst [Tob08]. Tribute [Ein91]. Trigonometr [Ano48i]. trudy [Ano75]. True [Bla78, MP01, Lev04].

Truman [Ano46b, Ano46c, Ano46c]. truth [RS80]. Turkish [Pla15g]. Turm [Jos95]. turn [Hof48, Yos95a]. turns [JDB^{+}97]. Tvorcheskij [Shp58b]. twentieth [Lig05a, Lig05b]. twentieth-century [Lig05a, Lig05b]. Two [Tib13, Hof08b, Pla90b, Pla90e, Pla00d, Gut15]. type [Cla91, Pet14].

u [KAR89, MKD^{+}91, Pla93d, Pla48c]. u.a. [BDR86]. U.I.P [BT58, DT58]. überarb [Sch55]. Übergangsschicht [Pla89w]. Überlieferung [Uli97, Ull13]. Übernahme [Pla13h]. Überproduktion [Pla33f]. Überraschung [Sei08]. Überreichung [Ano49b, Ano49c]. Überwindung [Yos95a]. Ueber [Pla83, Pla87f, Pla87c, Pla87d, Pla87e, Pla88k, Pla88l, Pla88h, Pla89m, Pla89l, Pla89s, Pla90d, Pla90c, Pla90e, Pla91b, Pla95b, Pla97e, Pla00h, Pla00g, Pla01e, Pla01d, Pla01f, Pla02c, Pla02a, Pla02b]. Uhr [Ano11]. UK [JDB^{+}97]. Ullmann [Kra97, Cah99]. Ulrike [Her03]. Ultracold [MKD^{+}91]. Ultraviolet [Tib13]. umgearb [Her24]. umgearbeitete [Ano24, Jüt22]. Umsturz [Ano34, Ano36, Ano39, Ano41a, Sch55, Wil37, Zim34]. Umwandlungen [Pla89m]. unbounded [Kae48]. uncertainty [Cas09, Hop96]. Uncovered [Hür22]. Uncovering [Wei07]. undergraduates [Lon84]. Underlying [v.B18]. understand [Fad96]. understandable [Hei88a, Hei06]. understanding [OT97]. Unification [Hop96, Wei89, Zac92]. Uniform [Pla10f]. uniformly [Pla07d]. Union [BT58, DT58]. unique [FGH91]. United [JDB^{+}97]. Units [BT58, DT58, WH88]. Units/Misprints [BT58, DT58]. Unity [KSB63, Pla09b, Pla09c, Pla70, MSQM63]. universal [Ish15, Ish17, Mey08, PB17]. Universe [Pad83, Pla31g, Pla37g, Pla96, Pla15h, Tib13, Pla59a, Bre08, Dys70]. Universelle [Ish15]. universellen [Mey08]. Universitaet [Pla14f, Pla16e, Pla23h]. Universität [AH93, Ano02, Ano46b, Ano11, Fra31, Jüt17, Pla09c, Pla13h, Pla14d]. Universität/Russische [Ano46b]. Universitätslehrern [OT97]. universities [Bra14b]. University [AH93, Ano48j, Ano11, Pla13h, Pla14f, Pla16e, Pla23h, San87, OT97, Pla10a, DdCA48, Pla09e, Pla09c, Pla14d, Pla14g, Pla15g, Pla12a, Pla15e]. Universum [Bre08]. Unlikely [Acz01, Acz03]. Unobservability [AP00]. Unobservable [AP00]. Unravelling [FR13]. Unruhe [Pun67]. unsere [Uns67]. unsigned [Sch29]. unternehmerischen [JDB^{+}97]. Untersuchung [Pla09f]. unveränderte [Ano30b]. unwilling [Seg07c]. Upheaval [Zim34]. Upright [Cas86, HB87, Hop96, Pye88, San87, Swe89, Hei86a, Hei00]. Urans [GBM^{+}49]. Urans/Gerlach [GBM^{+}49]. Ursprung [MSQM63, Pla33g]. Urteil [Thi69]. USA [Ano46b, JDB^{+}97, MAS63, Ano46a, MAS63]. USA./Prof [Ano46b]. USA/Großbritannien [JDB^{+}97]. Use [KGSO55, Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f].

V [Ano46b, Fra31, Ano53c, Iva59b, Pla37i, Pla26e, vLP21, Pla32i, Pla49i, Pla57c].

V-1 [Ano46b]. V-1/B [Ano46b]. V-1/Zerstörung [Ano46b]. vacuum [Kra12]. Vahlen [Pla38f]. Valerie [Hib91]. valiant [Sto13]. Valleriani [Ber14]. Value [Bir19, Alb93, HS97]. vapor [Kir98b, Pla88i, Pla88k]. vapors [Pla89b]. variable [Pla03e, Pla05a]. variabler [Pla03e, Pla05a]. Variation [Noe73, Peg77]. variations [Pla23a]. VDI [Pla33g, Pla33g]. VDPG [Ano61]. ve [Pla15g]. VEB [Ber60]. Veit [Ano13, Jüt11, Jüt18b]. velocity [Pla94c, Pla95b, Pla22i]. veränderte [Pla14e]. Veranstaltung [Ano29b]. Verantwortung [Gri98]. Verband [Ebe63]. verbessert [Thi88]. Verbesserung [Pla00g, Pla78c]. Verdampfen [Pla82]. verdünnten [Pla88a, Pla88i, Pla88k, Pla88j, Pla88h, Pla90b, Pla90e]. verdünnter [Pla87b, Pla87f, Pla88l, Pla27e, Pla29n, Pla30g, Pla31f, Pla33e]. Verdünnungswärme [Pla89r]. Vereinfachte [Pla01g]. verfilmt [Bre08]. Verflüssigung [Pla89i]. Vergessen [Bre08]. Verhalten [Pla89e]. Verhältnis [Hof00, Kir98a, Pla14i, Pla14j, Pla25f]. Verhältnisbestimmung [Tre01]. Verhandlungen [Ano48a]. verheiligert [MSQM63]. véridique [Lev04]. Vérification [Rub12, War12, War12, Rub12]. Verlag [Ano02, Ano13, Ano24, Ano30b, Ano34, Ano36, Ano39, Ano41b, Ano41a, Ano48i, Ano89, Ben30, Ber60, BBG^{+}58, Ebe63, Flü47, Fra31, Her24, Jüt11, Jüt17, Jüt18a, Jüt18b, Jüt21, Jüt22, Klo63, Mis33, Mos42, Reg34, Sac69, Sch49, Sch00b, Sch55, Som49, Wei68, Wil37, Zil33]. Verlage [Ano46b]. Verlagsgesellschaft [Hei31]. Verlegt [Ano46b]. Verleihung [Fie75, Haa75, Pla27a, KGSO55]. verloren [Bre08]. Vermehrung [Pla87c, Pla87d, Pla87e, Pla91b]. vermeintlicher [Pla00b]. vernünftige [Pla49b]. Veröffentlichungen [BPL^{+}54, Kra97]. Veröfftentlichunger [Mei47]. versäumt [Hof08e]. Verschiebungsgesetzes [Pla06b]. verschiedenen [Pla80, Pla89g]. verschiedener [Pla98a]. Verschwenden [BPL^{+}54]. Verschwenden/Rziha [BPL^{+}54]. Verständnis [OT97]. Versuch [Tre01, Pla40b, Pla40c, Pla41c]. versus [Gré19, Pet14]. Verteilung [Pla01a, Pla02b]. Vertheilung [Pla88c]. verwandelt [JDB^{+}97]. Verwirrende [Ano53b]. very [Pla88j]. Verzeichnis [Mei47]. Verzweiflung [Dar00]. Verzweiflung. [Büh98a]. VI [Fra31, Ano31]. Via [BBG^{+}58]. vibrating [Pla00f, Pla02c]. vibration [Pla01g]. vibrations [Pla97e, Pla97a, Pla23d]. vida [dACB58]. Vienna [Ano11, MR87, Pla29h]. Vierst. [Ano48i]. Vierstellige [GBM^{+}49]. Vierte [Ano13, Ano39]. view [HD48, Lon84, Lon03c, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a, Pla29g, Pla29i, Pla29o]. Vieweg [Ebe63]. Views [WC91, Som36]. VIII [Ano13, Pic25]. Violi [Pla89d]. Vischer [MKD^{+}91]. vision [Bro15]. visit [Alb93, Pla47b]. vitae [Pla20b]. vitenskapsmann [Har44]. Vögler [Her03, Koh02, Koh02]. Voice [Brü38]. Vol [KKRS96]. volatile [Pla88i, Pla88k]. völlig [Sch55]. Volume [BDR86, Cla91, Pla32g, Pla32i, Pla32j, Pla49e, Pla49i, Pla49j, Pla57a, Pla57c, Pla64b, Pla32h, Pla33d, Pla37a, Pla49a, Pla49h, Pla57b, Pla15c, Tho31]. Volumes [Pla89b]. Volumina [Pla89b, Pla89h]. Vom [MSQM63, Mül08b, Pla30b, Ano69, Ebe63, Jos95, Pla88e, Pla88f, Pla89q, Pla89w, Pla15a, Pla18a, Pla29c, Brü38, Pla25g, Pla36c, Pla37h, Pla39c, Pla45d, Pla48e, Pla52d, Pla53d, Pla55e, Pla90b]. vor [Ano46a, Jüt21, Pla20b, Pla20c, vL58]. Vorgänge [Ano02]. vorgelegt [Pla27b]. Vorgeschichte [Yos95d, Kan70, Thi71]. Vorgetragen [Pla33g]. Vorlesung [Fis08a]. Vorlesungen [Pla10a, Ano13, Ano24, Ein06, GBM^{+}49, Her24, Jüt11, Jüt18b, Jüt22, Kir94, Pla97f, Pla05c, Pla06g, Pla11e, Pla13k, Pla13j, Pla17c, Pla19f, Pla21g, Pla21f, Pla22j, Pla23i, Pla27f, Pla30h, Pla54c, PP64, Pla66, Pla10b, Pla15i, Sch65]. vorn [Frö01]. Vorsitzenden [Pla22f]. Vorstandes [Pla33g]. Vortag [Flü47]. Vortrag [Dys70, Haa70, Hei31, Hei69, Jüt21, Pla10c, Pla10d, Pla10e, Pla11c, Pla11d, Pla11a, Pla13c, Pla13d, Pla13e, Pla14c, Pla20b, Pla20a, Pla23g, Pla29g, Pla29e, Pla29f, Pla31e, Pla35e, Pla35c, Pla35d, Pla36c, Pla37d, Pla38h, Pla39c, Pla42e, Pla45a, Pla45b, Pla52d, Pla53d, Pla55e, vL58, Pla09c, Ano29b]. Vorträge [Ano58, Dri01, Ebe63, HR01, Jac02, Kra02a, Kra02b, Max58, Sim06, Som33, BBG^{+}58, Mis33, Pla33h, Pla43b, Pla44b, Pla45e, Pla49k, Pla58f, Pla90b, RH01, Zil33, Ano32]. Vorträge/Goethe [BBG^{+}58]. Vorträgen [Fra31, Gri98, Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f].

W [Ber60, BT58, DT58, MKD^{+}91, Pla89x, Pla00d, Pla03c, Pla37i]. W. [Pla37i]. Waals [Pla89w, Pla89v, Pla89u]. Waddington [Adl03]. Waerden [HKS^{+}68]. Wahlvorschläge [KK75]. Wahr [BBG^{+}58]. während [Hac07]. Wald [Pla89j]. Walden [Ano48i]. Walker [Hop96]. Wallace [Kut14a, Kut14a]. Walter [HRR10a, HRR10b, Sch65, HRR10b]. Walther [MdB63, Men73]. War [Ano16, Hop96, Sim12, BA01, Bro15, Löh12]. wäre [HS97]. Wärme [Cla91, Cla89, Fra31, Kir98a, Pla88e, Pla90d, Pla90c, Ein07b, Pla00c, Pla13i]. Wärmeinhalt [Pla89g]. Wärmelehre [Cla89]. Wärmemenge [Pla89g]. Wärmestrahlung [Pla13e, Sch78c, Ano24, Her24, Jüt22, Pla06g, Pla10j, Pla13j, Pla19f, Pla21f, Pla23i, Pla24b, Pla66, Pla15i, Wei68, Ein06]. Wärmetheorem [Pla12c, Pla12d]. Wärmetheorems [Pla10h]. Warmetheorie [Cla91, Cla91, Kir98b, Pla89j, Pla15c, Pla79, Pla91a, Pla93a, Pla15a, Pla89n]. Warum [Ano48i, Pla42f]. Was [Ano16, Ano47c, Bre08, MAS63, MSQM63]. Washingtoner [Ano46b]. Wasser [Kir98b]. Wasserdampfes [Kir98b]. Wasserstoffs [Pla19c]. water [Kir98b]. Wave [Hie68, Mic68, Prz63, MR87, Pla28d, Pla40b, Pla40c, Pla41c, Prz86, Prz11, Wes40]. wavelengths [Rub12]. waves [Pla95a, Pla96a, Pla03g]. way [Hei07]. ways [Wol08]. WE-Heraeus-program [JDB^{+}97]. WE-Heraeus-Programm [JDB^{+}97]. Wege [HPE^{+}33, Hei43a, Mis33, Pla33h, Pla34e, Pla43b, Pla44b, Pla45e, Reg34, Som33, Zil33, Ano32]. Weindling [Hop96]. Weird [Str11b]. weissen [Pla02a, Gae03]. Weiterbildung [JDB^{+}97]. weitere [Fal59]. Weizsäcker [MdB63]. welche [Cla89, Pla96c, Pla97e]. well [Pla16c, Pla19e, Pla21c, Pla22d, Pla28b, Pla31b, Pla15f]. Wellen [Pla95a, Pla96a, Wes40, Pla03g]. Wellen- [Wes40]. Wellenmechanik [Pla28d, Pla40b, Pla40c, Pla41c, Prz63, Klo63]. Welt [Bre08, Fis07, Fis10a, Jac08a]. Weltall [Dys70]. Weltanschauung [Pla35c, Pla35d, Pla37d, Pla45a, Pla45b, HD48, Pla29g, Pla52c, Pla53c, Pla55d, Pla58d, Pla65c, Pla67c, Hei35, Som36]. Weltbild [Ano29b, Ano30b, Ano34, Ano36, Ano39, Ano41a, Ben30, Pla29h, Pla29i, Pla29j, Pla30d, Pla30e, Sch55, Wil37, Zim34]. Weltbildes [Pla09b, Pla09c]. Weltgeschichte [Ano46b, Ano46c]. Weltgeschichte/Schmetterlinge [Ano46b, Ano46c]. Weltkrieg [Löh12]. Weltordnung [Pla49b]. Weltrekordmarken [JDB^{+}97]. Weltrekordmarken/Trägheitsfusion [JDB^{+}97]. Wentzel [Fie75, Hun75]. werden [Pla97e]. werden. [Pla96c]. Werk [Ano97a, EI08]. Werke [BDR86]. Werkstatt [Pla30b]. Werner [Adl03, BPL^{+}54, Hop96, Tre01, BDR86]. Wert [Alb93]. Werte [HS97]. Wesen [Pla19b, Pla20a, Pla36a, Pla36c, Pla37h, Pla39c, Pla44a, Pla45d, Pla48e, Pla52d, Pla53d, Pla55e, Pla90b]. Wesendonck [Pla00a]. Westphal [Ano48i]. Wetterdienst [Ano46b]. Wetterdienst/Radiozünder [Ano46b]. Where [Pla32k, Pla33i, Pla77, Pla81, Pla15j, Pla33h, Pla34e, Pla43b, Pla44b, Pla45e, Pla59a, Bro36]. which [Cla89, Pla96c]. white [Pla02a]. Whitehead [Adl03]. who [Ano47a, Ano47d, Ano47e, Cli87]. wichtiger [Pla88b]. wider [Bad08, Hof96]. Widerspruch [Pla00b]. Widerstand [vP06]. Widerstreit [Hof08f]. wie [Bre08]. Wieder [Ano46b]. Wiederkehr [Ham72, Pla46a, Wes57]. Wien [Ano29b, Ano11, Kan69, Klo63, Sch78d, Wol08, Pla00d, Pla00g, Pla06b, Pla29h, Pla69a, Pla78c, Sch78d, Wol08]. Wiensche [Kan69]. Wiens’chen [Pla00g, Pla06b, Pla78c]. Wild [LMOS15]. wilde [LMOS15]. Wilfried [Tre01]. Wilhelm [Cah99, Hei31, Her03, Kas64, Koh02, Kör61, Pla14d, Pla14f, Pla16e, Pla20a, Pla23h, Pla31e, Pla37a, Pla42e, Pla46a, Wol08, vBL96, Kör61, Pla07e, Wol08]. Wilhelm-Gesellschaft [Cah99, Hei31, Her03, Koh02, Pla20a, Pla31e, Pla37a]. Wilhelm-Universitaet [Pla14f, Pla16e, Pla23h]. Wilhelm/Max [vBL96]. Wilhelms [Jüt17, Pla14d]. Will [Ano37, Bad08, Tib13, Gri98, Pla23g, Pla36a, Pla36c, Pla37h, Pla39c, Pla45d, Pla48e, Pla52d, Pla53d, Pla55e, Pla90b]. Willard [Kör61]. Willen [Bad08]. Willensfreiheit [Pla23g, Pla36c, Pla39c, Pla52d, Pla53d, Pla55e, Pla90b, Gri98, Pla36a, Pla37h, Pla45d, Pla48e]. William [Ole81, Ole81]. Williams [Pic25]. Willy [Sch78d]. Winfried [Rut69]. Winkler [Ano11]. Winner [Ano47c]. wird [Pla89g]. Wirk [BBG^{+}58]. Wirken [Vog61]. Wirkung [EI08, Pla14a, Pla14b, Pla25a]. Wirkungsquanten [Pla13e]. Wirkungsquantum [Som11, Mül08b, dB48]. Wirkungsquantums [Ish15, Pla43c]. Wiss [Pla20b, Pla23g, Pla33g]. wissen [HS97]. Wissenschaft [Ano53b, Brü38, GSU07, Hei88a, Hei06, Her03, Hof08f, KAR89, Koh02, MSQM63, Pla29e, Pla29f, Pla30f, Pla42e, Pla42f, Ano59, Ano89, FS06, Flü47, Mei03, Pla30i, Pla42d, Pla46c, Pla88b].

Wissenschaft/Die [Ano53b]. Wissenschaften [Ano02, Ano29b, Ano48c, Ber60, Cah99, Ebe63, FH00, Hei31, Jüt21, Koy52, PP79, Pla20c, Pla37a, Pla48f, Sch49, Sim06, Som49, vL58, Pla31e, Pla42a, Cah99]. Wissenschaftler [Ano46b, Ano53c]. Wissenschaftliche [MKD^{+}91, Pla48c, Ano69, Her76, Jor48, PB90, Ano48i, Pla45f, Pla90c, RS49]. Wissenschaftlichen [Kör61, Pla33f, Pla58a]. wissenschaftlicher [Pla33g, Sti84]. Wissenschaftsakademie [LMOS15]. Wissenschaftsorganisator [Uli97, Ull13]. Without [Ano21, Din46c, Luc30]. Wohnsitz [Ano46b]. Wohnsitz/Lise [Ano46b]. Wolfskehl [Pla13c, Pla13d, Pla14c, Pla13c, Pla13d, Pla14c]. Wolfskehl-Vortrag [Pla13c, Pla13d, Pla14c]. Woman [Mei03]. wonders [Pla38d]. words [Ano11, Pla27a]. Work [Eck08, Kle66, Mei51, Nee80, Ano97a, Hei49, Her76, Kle85, Vog61, vL18]. Works [Ano21, Ano58, BDR86, Ano46c, Ano75, Bor63a, dACB58, Fal59, Pla59a, Ber15]. Workshop [HD48, Hei69, Pla30b, EP06, Haa70]. World [Arr06, Cah88, Hür22, Löh12, Men73, Som36, Wei07, WC91, Ano46c, Bre08, FF91, Fis07, Fis10a, HD48, Hei31, JDB^{+}97, Pla09b, Pla09c, Pla29g, Pla29i, Pla31e, Pla40a, Pla42b, Pla43a, Pla49c, Pla49b, Pla54a, Pla54b, Pla63a, Pla64a, Pla70, Pla93a, Pla29o, W^{+}87, WdC02, Zim34, Sim12]. worldview [Pla29h, Pla29j, Pla30d, Pla30e]. Worte [Ano11]. would [HS97]. wplyw [PS33]. Writings [Ano77, KKRS96, Ano48c, BDR86, Hei88a, Hei06, Low77, Pla48f, Pla60a, Pla70, Pla91a, W^{+}87, WdC02, Ole81, Ole81]. wspólczesnej [PS33]. Wunder [Pla38d]. Wurde [MSQM63]. Württemberg [Som48]. Württemberg-Baden [Som48]. Wybór [Pla70].

X [Jüt18b, Pla62, Seg07b]. X-ray [Pla62]. X-rays [Seg07b]. XI [Ebe63, Thi71]. XII [Ebe63, Jüt22]. xiii [San87]. xiv [McK27]. XV [Ebe63]. xx [Hie68].

Yale [Ano23]. Year [FR13, Lon84, Mul94, Pla06a]. Years [AH93, Gam85b, Hoh99, Par12, SB51, Ano46c, Fad96, Hei38, Kae48, KKRS96, Pla29b, Pla30f, Pla37a, Pla62]. York [C.49, Hie68, Pic25, Pla10a, Pla10a]. young [Pye85b]. youth [Bra14b].

Zagadnienia [PS33]. Zahl [MKD^{+}91]. zakonomernost [Pla26e]. Zehnte [Sch55]. zehnten [Wes57]. Zeilinger [Fis10b, Fis12]. Zeit [Uns67, Hac07, KAR89]. Zeit/Heilbron [KAR89]. Zeiten [Pla46b]. zeitgenössisches [Thi69]. Zeitmarken [GBM^{+}49]. Zentralbibliothek [Ano11]. Zerfallen [Fis07, Fis10a, Jac08a]. zero [Boy83, Kra12]. zero-point [Boy83, Kra12]. Zerstörung [Ano46b]. Zimmer [Ano34, Ano36, Ano39, Ano41a, Sch55, Wil37]. Züchtungsforschung [MAS63, HH63]. Züchtungsforschung/Was [MAS63]. zueinander [Pla14i, Pla14j, Pla25f]. zuerst [Pla36c]. Zukunftsmusik [Ano53b]. Zum [Ano38a, Ano43b, Ano57, Eck97a, Jac08b, MH97, Pla19e, Pla22d, Pla31b, Pla15f, Sta10, Alb93, Ano97a, Bec09, Ber15, Brü48a, Din59, Fis10b, Fis12, Fra31, GSU07, HD48, Her72, JDB^{+}97, Jüt17, Mei47, Mül08b, Pla89e, Pla08a, Pla08b, Pla09a, Pla14d, Pla16c, Pla21c, Pla25g, Pla26a, Pla28b, Pla29b, Pla29m, Pla30c, Pla37b, Pla38a, Pla38b, Pla38c, Pla39b, Pun67, Sch29, Som18, Som38a, Som47, Zen38, Ano97b, Ano08f, Pla33j, Pla37i, Pla57d, RH97, Vog61, vL38]. zum- [Pla38a]. Zur [Ano46b, Ano48i, BBG^{+}58, Hac07, Pla37a, Wes57, Ano02, Ano29b, Ano32, Ano46c, Brü38, Cah99, Ebe63, HKS^{+}68, Hei31, HPE^{+}33, Hei43a, Hol13, Jüt18a, Kan70, KK75, Klo63, Kra97, Löh12, Mis33, Pla89w, Pla91a, Pla92c, Pla97a, Pla10c, Pla10d, Pla10e, Pla10b, Pla11a, Pla15a, Pla16a, Pla17a, Pla22b, Pla23d, Pla24a, Pla28a, Pla28d, Pla29k, Pla29l, Pla31e, Pla33h, Pla34e, Pla35e, Pla43b, Pla44b, Pla45e, Pla46a, Pla57d, Prz63, Reg34, Sch00b, Som33, Spr89, Thi71, Uns67, Zil33, Ano48a, Fie75, Ham72, Pla84, Pla89x, Pla89y, Pla00i, Pla02d, Pla02e, Pla03h, Pla07f, Pla07g, Pla08f, Pla08g, Pla09f, Pla10i, Pla10j, Pla11f, Pla17f, Pla17g, Pla17h, Pla17d, Pla17e, Pla18b, Pla24c, Pla25h, Pla35f, Pla35g, Pla35h, Pla43c, Sch91, Sti84, Wes40, Yos95d]. Zurich [MR87]. zurück [Frö01]. Zusammenarbeit [Wol08]. zusammengestellt [Bec09, Sta10]. Zusammenhang [Kir98a]. Zustandsänderung [Pla88e]. Zustandsgleichung [Pla08e]. Zwei [Ano46b, Pla89u, Pla90b, Pla90e, Gut15]. zweiatomigen [Hol13]. zweiatomiger [Pla15c]. zweier [Pla00d]. Zweifel [HKS^{+}68]. Zweite [Ano30b, Pla89j, Pla41c]. Zweiten [Löh12, Mül08b, Pla79, Pla89j, Pla89n, Pla93a, Pla26d]. zwischen [HT99, Her03, HW07, Kir98a, Koh02, Mul00, Pla90b, Pla90e, Pla98a, Pla01a, Pla02b, Pla13i, Pla40b, Pla40c, Pla41c, Sch00b, Sch04, Thi69, vP06, Hof08f]. Zyklotrons [Ano46b]. Zyklotrons/Um [Ano46b].

References

A:1950:BRB

[A.50] R. L. A. Book review: Scientific Autobiography and Other Papers by Max Planck. Philosophy of Science, 17(4):361, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

            (electronic). URL http://www.jstor.org/stable/185944.

Abbott:2001:MPS

[Abb01] Alison Abbott. Max Planck Society admits to its predecessor’s Nazi links. Nature, 411(6839):726, June 14, 2001. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Aczel:2001:EGM

[Acz01] Amir D. Aczel. Entanglement: The Greatest Mystery in Physics: The Unlikely Story of How Scientists, Mathematicians, and Philosphers Proved Einstein’s Spookiest Theory. Four Walls Eight Windows, New York, NY, USA, 2001. ISBN 1-56858-232-3. xviii + 284 pp. LCCN QC174.12 .A29 2001. URL http://www.loc.gov/

            catdir/enhancements/fy0831/2002069338-d.html.

Aczel:2003:EUS

[Acz03] Amir D. Aczel. Entanglement: the unlikely story of how scientists, mathematicians, and philosophers proved Einstein’s spookiest theory. Plume, New York, NY, USA, 2003. ISBN 0-452-28457-0 (paperback). xviii + 284 pp. LCCN QC174.12 .A29 2003.

Adler:2003:HPM

[Adl03] Mortimer J. Adler, editor. Henri Poincaré, Max Planck, Alfred North Whitehead, Albert Einstein, Sir Arthur Eddington, Niels Bohr, G. H. Hardy, Werner Heisenberg, Erwin Schrödinger, Theodosius Dobzhansky, C. H. Waddington, volume 56 of Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, 2003. ISBN 0-85229-531-6. xvi + 749 pp. LCCN AC1 .G72 1990.

Agassi:1967:KPR

[Aga67] Joseph Agassi. The Kirchhoff–Planck radiation law. Science, 156(3771):30–37, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/156/3771/30.full.pdf;   http:/
/www.sciencemag.org/content/156/3771/30.short.

Albrecht:1993:NTG

[AH93] Helmuth Albrecht and Armin Hermann, editors. Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. (German) [Science and technology in history: 25 Years of the Chair of History of Science and Technology at the Historical Institute of the University of Stuttgart]. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. 401 pp. LCCN Q183.4.G293 U556 1993.

Albrecht:1993:FMP

[Alb93] Helmuth Albrecht. Max Planck: Mein Besuch bei Adolf Hitler— Anmerkungen zum Wert einer historischen Quelle. (German) [Max Planck: My visit to Adolf Hitler — Comments on the value of a historical source]. In Albrecht and Hermann [AH93], pages 41–64. ISBN 3-928186-15-9. LCCN Q183.4.G293 U556 1993.

Andrade:1948:MP

[And48] E. N. Da C. Andrade. Max Planck. Nature, 161(4086):284, February 21, 1948. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v161/n4086/pdf/161284a0.pdf.

Anonymous:1902:RBK

[Ano02] Anonymous. [Rezensionen:] Ostwalds Klassiker der exakten Wissenschaften Nr. 124, Abhandlungen zur Thermodynamik chemischer Vorgänge. von H. Helmholtz. Herausgegeben von Dr. Max Planck, Professor der Physik an der Universität Berlin. Verlag von Engelmann in Leipzig. 83 Seiten. Preis 1,40 Mk. Zeitschrift für Elektrochemie, 8(41):799, 1902. CODEN ZEELAI. ISSN 0005-9021.

Anonymous:1913:RBT

[Ano13] Anonymous. [Rezensionen:] Vorlesungen über Thermodynamik. von Max Planck. Vierte Auflage. VIII und 288 S. mit 5 Figuren im Test. Verlag von Veit & Co., Leipzig. 1913. Preisgeb. 7.50 Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 19(13):549–550, 1913. CODEN ZEAPAA. ISSN 0005-9021.

Anonymous:1916:GSS

[Ano16] Anonymous. German scholars’ stand: Their first war statement was misunderstood, says Prof. Planck. New York Times, ??(??):13, June 6, 1916. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/97914442/.

Anonymous:1921:EEF

[Ano21] Anonymous. Einstein explains fallacy of ether: Space without electromagnetic fields has no characteristics, he tells students. Humorous at City College. talks about “the greatest pain of physics” in exposition of his Relativity theory. “Special Relativity’s” action. How Planck’s theory works. New York Times, ??(??):12, April 21, 1921. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/98436746/.

Anonymous:1923:DAN

[Ano23] Anonymous. Discusses atom from new point: Dr. Bohr, in Yale lecture takes up the theory of spectral lives. Takes up Planck formula. He shows how it is employed to explain the spectra of the elements. New York Times, ??(??):17, November 14, 1923.

CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/103174188/.

Anonymous:1924:RBT

[Ano24] Anonymous. [Rezensionen:] Vorlesungen über die Theorie der Wärmestrahlung. Von Max Planck. 5., abermals umgearbeitete Auflage. 221 Seiten. Grundzahl 7, geb. 9, M. Verlag von Joh. Ambrosius Barth, Leipzig 1923. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 30(7–8):189B–189, 1924. CO-DEN ZEAPAA. ISSN 0005-9021.

Anonymous:1929:DPL

[Ano29a] Anonymous. Dr. Planck lauds Einstein: Professor receives first medal cast in Berlin in honor of scientist. New York Times, ??(??):15, June 30, 1929. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/104972983/. Report of 29 June

1930 of first award of Planck Medal, to Albert Einstein.

Anonymous:1929:RKV

[Ano29b]  Anonymous.  [Rezensionen:] Komitee zur Veranstaltung

von Gastvorträgen ausländischer Gelehrter der exakten Wissenschaften, Wien Geh. Rat Prof. Dr. Max Planck: Das Weltbild der neuen Physik. Vortrag. Angewandte Chemie, 42(24):672, 1929. CODEN ANCEAD. ISSN 1521-3757.

Anonymous:1929:RSC

[Ano29c] Anonymous. The Royal Society Copley Medal awarded to German professor [Max Planck]. Manchester Guardian, ??(??):10, December 2, 1929. ISSN 0307-756X. URL http://pqasb.

pqarchiver.com/guardian/doc/477987079.html.

Anonymous:1930:PHG

[Ano30a] Anonymous. Planck heads German scientists. New York Times, ??(??):3, July 19, 1930. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/98958782/. Report of Max

Planck’s election on 18 July 1930 as President of the Kaiser Wilhelm Society for Scientific Research.

Anonymous:1930:RWN

[Ano30b] Anonymous. [Rezensionen:] Das Weltbild der neuen Physik. Von Prof. Dr. Max Planck, Berlin. Zweite, unveränderte Auflage.

Leipzig 1929. Verlag von Johann Ambrosius Barth. 52 Seiten. Preis 2,70 Mark. Archiv der Pharmazie (Weinheim, Germany), 268(1):80, 1930. CODEN ARPMAS. ISSN 0365-6233 (print), 1521-4184 (electronic).

Anonymous:1931:IGS

[Ano31] Anonymous. Interviews with great scientists VI: Max Planck the paradox of the quantum space and time no moral progress of mankind beliefs and “phantasies”. The Observer (London), ??(??):17, January 25, 1931. URL http://pqasb.pqarchiver.

com/guardian/doc/481317093.html.

Anonymous:1932:BRB

[Ano32] Anonymous. Book review: Wege zur physikalischen Erkenntnis: Reden und Vorträge, by Max Planck. Journal of Physical Chemistry, 37(8):1083, January 1, 1932. CODEN JPCHAX. ISSN 0092-7325 (print), 1943-300X (electronic). URL http://pubs.

acs.org/doi/pdf/10.1021/j150350a013.

Anonymous:1934:REZ

[Ano34] Anonymous. [Rezensionen:] Ernst Zimmer, Umsturz im Weltbild der Physik, gemeinverständlich dargestellt. Mit einem Geleitwort von Dr. Max Planck. 262 S. m. 58 Abb. München 1934, Verlag Knorr und Hirth, G. m. b. H. Preis geh. 4,50 M., Leinen 5,70 M. Zeitschrift für Angewandte Mathematik und Mechanik, 14(3):190, 1934. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Anonymous:1936:RUI

[Ano36] Anonymous. [Rezensionen:] Umsturz im Weltbild der Physik. Von Ernst Zimmer. Mit einem Geleitwort von Dr. Max Planck und 59 Abbildungen. Dritte Auflage. Verlag Knorr und Hirth, München 1936. 270 Seiten. Preis geh. 4,50 Mark, Leinen 5,70 Mark. Berichterstatter: Horst Böhme, Berlin-Dahlem. Archiv der Pharmazie (Weinheim, Germany), 274(8):503, 1936. CO-DEN ARPMAS. ISSN 0365-6233 (print), 1521-4184 (electronic).

Anonymous:1937:FWD

[Ano37] Anonymous. Free will doctrine advanced by Professor Max Planck. The Science News-Letter, 31(830):158, March 6, 1937. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (elec-

tronic). URL http://www.jstor.org/stable/view/10.2307/
3914163.

Anonymous:1938:MPG

[Ano38a] Anonymous. Max Planck. Zum 80. Geburtstag. Am 23 April 1938. (German) [Max Planck’s 80th birthday, 23 April 1938]. Annalen der Physik (Berlin), 424(1–2):1, 1938. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Anonymous:1938:PMP

[Ano38b] Anonymous. Prof. Max Planck, For. Mem.R.S. Nature, 141 (3573):720, April 23, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v141/n3573/pdf/141720a0.pdf.

Anonymous:1939:RUI

[Ano39] Anonymous. [Rezensionen:] Umsturz im Weltbild der Physik. Gemeinverständlich dargestellt von Ernst Zimmer. Mit einem Geleitwort von Max Planck. Vierte, erweiterte Auflage, 13. bis 16. Tausend. 290 Seiten mit 61 Abb. Verlag Knorr und Hirth G. m. b. H., München 1938. Preis brosch. RM 4.50, geb. RM 5.70. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 45(5):426, 1939. CODEN ZEAPAA. ISSN 0005-9021.

Anonymous:1941:RUI

[Ano41a] Anonymous. [Rezensionen:] Umsturz im Weltbild der Physik. Von Ernst Zimmer. Mit einem Geleitwort von Geh. Rat Dr. Max Planck. 4., erweiterte Auflage, 1938. Verlag Knorr & Hirth, München. 290 S. mit 61 Bildern. Preis geh. 4,50 RM., geb. 5.70 RM. Besprecher: H. J. Heuser. Archiv der Pharmazie (Weinheim, Germany), 279(3):105, 1941. CODEN ARPMAS. ISSN 0365-6233 (print), 1521-4184 (electronic).

Anonymous:1941:RPD

[Ano41b] Anonymous. [Rezensionen:] Prof. Dr. Hugo Dingler, Max Planck und die Begründung der sogenannten modernen theoretischen Physik. 32 S. Berlin 1939, Ahnenerbe-Stiftung-Verlag. Preis 0,80 M. Zeitschrift für Angewandte Mathematik und Mechanik, 21(1): 61, 1941. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Anonymous:1943:MP

[Ano43a] Anonymous. Max Planck. Das Reich, ??(??):1, April 11, 1943.

Anonymous:1943:MPA

[Ano43b] Anonymous. Max Planck. Zum 23. April 1943. Hochgeehrter Jubilar!. (German) [Max Planck. 23 April 1943. Esteemed honoree!]. Annalen der Physik (Berlin), 434(5):421, 1943. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Anonymous:1944:MPS

[Ano44] Anonymous. Max Planck über die Schule. (German) [Max Planck about the school]. Physikalische Blätter, 1(7):84–85, 1944. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1946:PLB

[Ano46a] Anonymous. Paul Langevin/Besuch bei Max Planck/O. Hahn vor seiner Reise nach Stockholm/nach der Schweiz/nach Schweden/nach USA. Physikalische Blätter, 2(9):239–241, 1946. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1946:RBTa

[Ano46b] Anonymous. Radiumfund Bei Bad Tölz/Dr. Eckener Arbeitet/Präsident Truman Zur Forschung/Goethe-Preis für Max Planck/Todesurteil Gegen Chemiker/Das Grösste Forschungsunternehmen Der Weltgeschichte/Schmetterlinge Statt Atombomben/Nobel-Preis für Medizin/Pasteur 50 Jahre Tot/Erste Kunde Von V-1/B. Baruch Über Die Forschung in Deutschland/Max Planck Verlegt Seinen Wohnsitz/Lise Meitner an der Washingtoner Universität/Russische Entdeckung in der Atomforschung/Wieder Astronomisches Recheninstitut/Zwei Berliner Verlage Wieder Eröffnet/88 Deutsche Wissenschaftler Nach Amerika/Nobel-Preis für Deutschen Gelehrten/Neuer Bayrischer Wetterdienst/Radiozünder Gegen V-1/Zerstörung der Japanischen Zyklotrons/Um Die Atomkontrolle/Ehrung Russischer Gelehrter/Farben-Fernsehen in USA./Prof. Hahn Wieder in Deutschland/Amerikanischer Bericht Zur Deutschen Atomforschung/Schiffssteuerung Durch Radar/Zwei Neue Transurane?/ Experiment Beim Bikini-Atoll. Physikalische Blätter, 2(1):17–22, 1946. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1946:RBTb

[Ano46c] Anonymous. Radiumfund bei Bad Tölz/Dr. Eckener arbeitet/ Präsident Truman zur Forschung/Goethe-Preis für Max Planck/ Todesurteil gegen Chemiker/das grösste Forschungsunternehmen der Weltgeschichte/Schmetterlinge statt Atombomben/Nobel-Preis für Medizin/Pasteur 50 jahr. (German) [Radium discovery near Bad Tölz /Dr. Eckener works /President Truman on research /Goethe Award for Max Planck /Death sentence against chemist /The largest research firm in world history /Butterflies instead of nuclear bombs /Nobel Prize for Medicine /Pasteur 50 years anniversary]. Physik Journal, 2(1):17–22, 1946. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1946:SHI

[Ano46d] Anonymous. Scientists honor Isaac Newton in London; Planck brought from Berlin for ceremony. New York Times, ??(??): 12, July 16, 1946. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107778779/.

Anonymous:1947:DPP

[Ano47a] Anonymous. Death of Prof. Planck. Professor Max Planck, the German physicist who originated the quantum theory, died at 3.30 this morning in hospital. Nottingham Evening Post, ??(??):1,

October 4, 1947. URL http://www.britishnewspaperarchive.
co.uk/viewer/bl/0000321/19471004/009/0001.

Anonymous:1947:MP

[Ano47b] Anonymous. Max Planck. Die Naturwissenschaften, 34(1):1, January 1947. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF00633309.

Anonymous:1947:MPD

[Ano47c] Anonymous. Max Planck dead; noted physicist, 89: Winner of Nobel Prize in 1919, originated quantum theory, aiding in splitting atom; discovered constant h; one of the intellectual giants of century, he was ranked with science immortals. New York Times, ??(??):68, October 5, 1947. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/108076771;   http://search.proquest.com/hnpnewyorktimes/
docview/108080837.

Anonymous:1947:PEP

[Ano47d] Anonymous. Planck, eminent physicist, dead. originated quantum theory. Professor Max Planck, the German physicist who originated the quantum theory. Gloucestershire Echo, ??(??):1,

October 4, 1947. URL http://www.britishnewspaperarchive.
co.uk/viewer/bl/0000320/19471004/014/0001.

Anonymous:1947:PPQ

[Ano47e] Anonymous. Prof. Planck, ‘quantum theory’ expert, dies. Professor Max Karl Ernst Ludwig Planck, who ranked with . . .. Hull Daily Mail, ??(??):4, October 4, 1947. URL http:/

/www.britishnewspaperarchive.co.uk/viewer/bl/0000324/
19471004/066/0004.

Anonymous:1948:FVD

[Ano48a] Anonymous. Faksimile aus den Verhandlungen der Deutschen Physikalischen Gesellschaft 2 (1900) S. 237: Zur Theorie des Gesetzes der Energieverteilung im Normalspectrum; von M. Planck. (German) [Facsimile of the Proceedings of the German Physical Society 2 (1900) p. 237: On the theory of the law of energy distribution in the normal spectrum; by M. Planck]. Physikalische Blätter, 4(4):146–151, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1002/phbl.19480040404/abstract. Special issue on

Max Planck.

Anonymous:1948:MPA

[Ano48b] Anonymous. Max Planck (23 avril 1858–4 octobre 1947). Revue de métaphysique et de morale, 53(1):85–86, January 1948. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://

www.jstor.org/stable/view/10.2307/40899306.

Anonymous:1948:MPS

[Ano48c] Anonymous. Max Planck in seinen Akademie-Ansprache: Erinnerungsschrift der Deutschen Akademie der Wissenschaften zu Berlin. (German) [Max Planck in his Academy speeches: Memorial writings of the German Academy of Sciences in Berlin]. Akademie-Verlag, Berlin, Germany, 1948. 204 pp. LCCN ????

Anonymous:1948:MPM

[Ano48d] Anonymous. Max Planck memorial ceremony at Göttingen. Nature, 161(4098):751–752, May 15, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v161/n4098/pdf/161751a0.pdf.

Anonymous:1948:MPNa

[Ano48e] Anonymous. Max Planck/Nachweis der Bilder/Autoren:/ Aufsätze. (German) [Max Planck/proof of images/authors:/ papers]. Physikalische Blätter, 4(4):179–180, April 1948. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.

19480040408/abstract. Special issue on Max Planck.

Anonymous:1948:MPNb

[Ano48f] Anonymous. Max Planck/Nachweis der Bilder/Autoren:/ Aufsätze. (German) [Max Planck/proof of images/authors/ papers]. Physikalische Blätter, 4(4):179–180, 1948. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1948:NGS

[Ano48g] Anonymous. New German science group honors Max Planck. The Science News-Letter, 54(4):60–61, July 24, 1948. CODEN SNLEAI. ISSN 0096-4018 (print), 2326-1285 (electronic). URL

http://www.jstor.org/stable/view/10.2307/3925056.

Anonymous:1948:PMP

[Ano48h] Anonymous. Prof. Max Planck, For. Mem.R.S. Nature, 161 (4080):47–48, January 10, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v161/n4080/pdf/161047c0.pdf.

Anonymous:1948:RWG

[Ano48i] Anonymous. [Rezensionen:] Walden: Geschichte Der Chemie/ Becker: Geschichte Der Astronomie/Max, Laue, Geschichte Der Physik/Westphal: Physik/Planck, Ambrosius Barth, Wissenschaftliche Selbstbiographie/Joos: Fiat Review, Physik der Festen Körper I/Solomon, Flöttmann, Bertelsmann-Verlag: Warum Atomzertrümmerung?/Lenz: Einführungsmathematik für Physiker/Reihenentwicklungen in der Mathematischen Physik/ Peters: Sechsstellige Tafeln Der Trigonometr. Funktionen/ Schulz: Formelsammlung Zur Praktischen Mathematik/Gravelius: Vierst. Tafeln F. Logarithm. Und Numerisches Rechnen/Baur: Musterbeispiele Europäischer Grosswetterlagen. Physikalische Blätter, 4(2):77–82, 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1948:SCD

[Ano48j] Anonymous. Sir Charles Darwin at the Max Planck Memorial Service held on 23 April 1948 in the Aula of the University of Göttingen. Notes and Records of the Royal Society of London, 6 (1):69–70, December 1948. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/view/10.2307/531259.

Anonymous:1948:SIM

[Ano48k] Anonymous. [Special issue on Max Planck]. Physikalische Blätter, 4(4):134–180, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.v4.4/issuetoc.

Anonymous:1949:GMP

[Ano49a] Anonymous. Gedenkfeier für Max Planck am 23. April 1948 in Göttingen. (German) [Commemoration for Max Planck on 23 April 1948 in Göttingen]. Angewandte Chemie, 61(4):151, 1949. CODEN ANCEAD. ISSN 1521-3757.

Anonymous:1949:UPM

[Ano49b] Anonymous. Überreichung der Planck-Medaille an Otto Hahn und Lise Meitner durch Max von Laue. (German) [Presentation of the Planck Medal to Otto Hahn and Lise Meitner by Max von Laue]. Physikalische Blätter, 5(10):471–473, 1949. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1949:UPMb

[Ano49c] Anonymous. Überreichung der Planck-Medaille an Otto Hahn und Lise Meitner durch Max von Laue. (German) [Presentation of the Planck Medal to Otto Hahn and Lise Meitner by Max von Laue]. Physik Journal, 5(10):471–473, October 1949. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Anonymous:1952:MPM

[Ano52] Anonymous. Max-Planck-Medaille an Paul Adrien Maurice Dirac. Physikalische Blätter, 8(11):512–513, 1952. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1953:BMP

[Ano53a] Anonymous. Bildnisse von Max Planck. (German) [Portraits of Max Planck]. Physikalische Blätter, 9(4):190–191, 1953. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1953:SVW

[Ano53b] Anonymous. Die Schädliche und Verwirrende Wissenschaft/ Die Invertierte Sonne/Max Planck-Statue/Staatsmänner und Goldmacher/Internationales Kernphysikalisches Laboratorium/ Zukunftsmusik. (German) [The harmful and confusing science/the inverted sun/Max Planck statue/statesmen and alchemist/international nuclear physics laboratory/future music]. Physikalische Blätter, 9(9):429–431, 1953. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1953:WNB

[Ano53c] Anonymous. Wissenschaftler auf neuen Briefmarken: J. v. Liebig; Max Planck. (German) [Scientists on new stamps: J. v Liebig;. Max Planck]. Nachrichten aus Chemie und Technik, 1(17–18): 134, 1953. CODEN NCHTAD. ISSN 1868-0054.

Anonymous:1957:MPG

[Ano57] Anonymous. Max Planck Zum 100. Geburtstag. (German) [Max Planck’s hundredth birthday]. Annalen der Physik (Berlin), 456 (4–5):185, 1957. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Anonymous:1958:CWM

[Ano58] Anonymous. Collected works of Max Planck: Physikalische Abhandlungen und Vorträge von Max Planck. Nature, 182(4639): 820–821, September 27, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v182/n4639/pdf/182820a0.pdf.

Anonymous:1959:MPD

[Ano59] Anonymous. Max-Planck-Denkmal/Kongressitis/Internationale Wissenschaft. (German) [Max Planck monument/congressitus/ international science]. Physikalische Blätter, 15(2):94, 1959. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1961:MGJ

[Ano61] Anonymous. Martin Grützmacher 60 Jahre/Nobelpreise 1961/ Max Planck-Medaille 1961/Preis VDPG 1961. Physikalische Blätter, 17(11):525–526, 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1969:PEB

[Ano69] Anonymous. Planck und Einstein beantworten eine wissenschaftliche Rundfrage/Drei Fragen/Max Plancks Antwort vom 14.12.1930/Albert Einsteins Antwort vom 31.1.1930. (German) [Planck and Einstein to answer a scientific poll/three questions/ Max Planck’s response of 12.14.1930/Albert Einstein’s reply of 01.31.1930]. Physikalische Blätter, 25(12):558, 1969. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1970:NMA

[Ano70] Anonymous. Neue Münze: Adenauer soll Max Planck ersetzen. (German) [New coin: Adenauer intended to replace Max Planck]. Physikalische Blätter, 26(3):120, 1970. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1975:PMI

[Ano75] Anonymous. Plank M. — Izbrannye trudy. (Russian) [Max Planck — selected works]. Nauka, Moscow, USSR, 1975. 788

pp. URL http://librarun.org/book/10029/1.

Anonymous:1977:BMP

[Ano77] Anonymous. Bibliography of Max Planck’s non-technical writings. Newsletter on Science, Technology, & Human Values, 20 (??):10, June 1977. ISSN 2328-2436. URL http://www.jstor.

org/stable/view/10.2307/751580.

Anonymous:1981:MPM

[Ano81] Anonymous. Max-Planck-Medaille 1981/Haushaltsplan und Mitgliedsbeiträge 1981/Frühjahrstagungen Münster und Hamburg 1981. Physikalische Blätter, 37(1):23–24, 1981. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:1989:RMP

[Ano89] Anonymous. [Rezensionen:] Max Planck — ein Leben für die Wissenschaft. Von J. L. Heilbron, S. Hirzel Verlag Stuttgart 1988, 474 Seiten, DM 68,-. Physik in unserer Zeit, 20(4):127, 1989. CODEN PHUZAH. ISSN 1521-3943.

Anonymous:1994:MPR

[Ano94] Anonymous. The Max Planck research award. Journal of Aerosol Science, 25(1):209, 1994. CODEN JALSB7. ISSN 0021-8502 (print), 1879-1964 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021850294901929.

Anonymous:1997:MPL

[Ano97a] Anonymous. Max Planck: Leben–Werk–Persönlichkeit. Ausstellung zum 50. Todestag. Magnus-Haus, Berlin, Oktober 1997. (German) [Max Planck: Life—work—personality. Exhibition on the 50th anniversary of his death. Magnus-Haus, Berlin, October 1997]. Web site, October 1997. URL http://www.max-planck.

mpg.de/frameset.html.

Anonymous:1997:TMP

[Ano97b] Anonymous. Zum 50. Todestag von Max Planck. (German) [On the 50th anniversary of the death of Max Planck]. Physikalische Blätter, 53(10):964, 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Anonymous:2002:MEL

[Ano02] Anonymous. Max Ernst Ludwig Planck: Physiker. Report, Berlin-Brandenburgische Akademie der Wissenschaften Akademiebibliothek, Berlin, Germany, 2002. 32 pp.

Anonymous:2004:MPA

[Ano04] Anonymous. Max Planck Awards announced. Materials Today, 7(2):55, February 2004. CODEN MTOUAN. ISSN 1369-7021 (print), 1873-4103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1369702104000884.

Anonymous:2008:MPG

[Ano08a] Anonymous. Max Planck 150. Geburtstag. (German) [Max Planck’s 150th birthday]. Physik in unserer Zeit, 39(2):53, 2008. CODEN PHUZAH. ISSN 1521-3943.

Anonymous:2008:MPa

[Ano08b] Anonymous. Max Planck (1858–1947). Annalen der Physik (Berlin), 17 [i.e. 520](2–3):65–67, 2008. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Anonymous:2008:MPb

[Ano08c] Anonymous. Max Planck 1858–1947. Physik Journal, 7(3):26, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1104313/issue.html.

Anonymous:2008:PMK

[Ano08d] Anonymous. Planck, Max Karl Ernst Ludwig. In ????, editor, Complete Dictionary of Scientific Biography, page ?? Charles Scribner’s Sons, Detroit, MI, USA, 2008. ISBN 0-684-31559-9 (e-book). LCCN Q141 .C66 2008. URL http://www.

encyclopedia.com/topic/Max_Planck.aspx.

Anonymous:2008:SIC

[Ano08e] Anonymous. Special issue: Commemorating Max Planck. Annalen der Physik (Berlin), 17 [i.e. 520](1):3, 2008. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Anonymous:2008:GMP

[Ano08f] Anonymous. Zum 150. Geburtstag von Max Planck. (German) [On the 150th birthday of Max Planck]. Physik in unserer Zeit, 39(3):115, 2008. CODEN PHUZAH. ISSN 1521-3943.

Anonymous:2011:ESA

[Ano11] Anonymous. Erwin Schrödinger 1887–1961: eine Ausstellung der Österreichischen Zentralbibliothek für Physik; Ausstellungseröffnung und Eröffnung des Schrödinger-Symposiums; 13.1.2011, 19.30 Uhr, Kleiner Festsaal der Universität Wien; einleitende Worte: Rektor Georg Winkler; Festvortrag: Jürgen Renn, MPI Berlin. (German) [Erwin Schrödinger 1887–1961: an exhibition of the Austrian Central Library for Physics, and the opening of the exhibition opening Schrödinger Symposium, 01.13.2011, 19.30 clock, Small Ballroom of the University of Vienna; introductory words: President George Winkler lecture: Jürgen Renn, Max Planck Institute in Berlin]. Report, Internationales Erwin-Schrödinger-Institut für Mathematische Physik, Vienna, Austria, January 13, 2011.

Anonymous:2012:PPC

[Ano12] Anonymous. Pinpointing Planck’s constant with GPS. Web

site, March 15, 2012. URL http://physics.aps.org/synopsis-
for/10.1103/PhysRevLett.108.110801. See [KM12].

Anonymous:2014:MPS

[Ano14] Anonymous. The Max Planck Society buys the entire Springer Book Archives. Web press release., February 10, 2014. URL http:

//www.springer.com/about+springer/media/pressreleases?
SGWID=0-11002-6-1454444-0.

Anonymous:20xx:RAP

[Anoxx] Anonymous. The religious affiliation of physicist Max Planck.

Web site, 20xx. URL http://www.adherents.com/people/pp/
Max_Planck.html.

Agazzi:2000:RUO

[AP00] Evandro Agazzi and Massimo Pauri. The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism, volume 215 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346. vii +

377 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-015-9391-5.

ArroyoCamejo:2006:SQG

[Arr06] Silvia Arroyo Camejo. Skurrile Quantenwelt. (German) [Crazy Quantum World]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-29720-0. xiii + 246 pp. LCCN QC174.12 .A77 2006. EUR

29.95, SFR 51.00. URL http://www.springer.com/physics/
quantum+physics/book/978-3-540-29720-8.

Bettenworth:2001:MPS

[BA01] Vera Bettenworth and Alison Abbott. Max Planck Society broke pledge to ask Jews back after war. Nature, 409(6819):443, January 25, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Badino:2008:MPR

[Bad08] Massimiliano Badino. Max Planck Revolutionär wider Willen; eine kleine Geschichte der Quantentheorie und ihres Begründers; [auf Grundlage der Ausstellung “Max Planck — Revolutionär wider Willen” (26.4.–5.10.2008)]. (German) [Max Planck: Revolutionary against Will; a small history of quantum theory and its founder; [based on the exhibition “Max Planck — Revolutionary against Will” (26.4.–5.10.2008)], volume [20]08,1 of Spektrum der Wissenschaft: […], Biografie. Spektrum-der-Wiss.-Verl.-Ges., Heidelberg, Germany, 2008. ISBN 3-938639-89-X. 98 pp. LCCN

???? URL http://d-nb.info/98924024X/04.

Badino:2009:OCB

[Bad09] Massimiliano Badino. The odd couple: Boltzmann, Planck and the application of statistics to physics (1900–1913). Annalen der Physik (Berlin), 18 [i.e. 521](2–3):81–101, ???? 2009. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See also [Gea02].

Badino:2015:BRM

[Bad15] Massimiliano Badino. The bumpy road: Max Planck from radiation theory to the quantum (1896–1906). SpringerBriefs in history of science and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-20030-5, 3-319-20031-3 (e-book). ISSN 2211-4564. xii +

112 pp. LCCN QC173.98. URL http://www.springer.com/us/
book/9783319200309.

Baez:2001:HDA

[Bae01] John C. Baez. Higher-dimensional algebra and Planck scale physics. In Callender and Huggett [CH01], chapter 8, pages 177–198. ISBN 0-521-66280-X (hardcover), 0-521-66445-4 (paperback). LCCN QC178 .P48 2001. URL http://www.loc.

gov/catdir/description/cam021/00050775.html;   http:/
/www.loc.gov/catdir/toc/cam021/00050775.html.

Buchwald:1958:BBP

[BBG^{+}58] E. Buchwald, F. L. Bauer, H. Gericke, H. Witting, P. Grassmann, W. Döring, G. Dickel, J. Euler, L. Oster, A. Ebinger, and D. Lilienthal. Bücher Bücher: Planck: Physikalische Abhandlungen und Vorträge/Goethe: Schriften Zur Naturwissenschaft/ Erckmann: Via Regia/Courant Und Hilbert: Methods Of Mathematical Physics/Blaschke: Griechische Und Anschauliche Geometrie/Ackermann, Verlag: Einführung In Die Wahr Scheinlichkeitsrechnung/Gray: American Institute Of Physics Handbook/ Guggenheim: Thermodynamics/Bruhat: Recueil de Problèmes/ Roth und Bekker: Kalorimetrische Methoden Zur Bestimmung Chemischer Reaktionswärmen/Kurzes Lehrbuch Der Physikalischen Chemie/Spitzer: Physics Of Fully Ionized Gases/Wirk und Thilo: Niederfrequenz Und Mittel Frequenz Messtechnik/Hönger und Reuber: Rundfunkröhren/Mutter: Kompendium Der Photographie/Photographie Und Kinematographie. Physik Journal, 14(5):230–237, May 1958. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Bernhard:1982:STS

[BCS82] Carl Gustaf Bernhard, Elisabeth T. Crawford, and Per Sörbom, editors. Science, technology, and society in the time of Alfred Nobel: Nobel symposium 52, held at Björkborn, Karlskoga, Sweden, 17–22 August 1981. Published for the Nobel Foundation by Pergamon Press, Oxford, UK, 1982. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.).

Buttner:2000:RQD

[BDH^{+}00] Jochen Büttner, Olivier Darrigol, Dieter Hoffmann, Jürgen Renn, and Matthias Schemmel. Revisiting the Quantum Discontinuity, volume 150 of Preprint (Max-Planck-Institut für Wissenschaftsgeschichte). Max-Planck Institut für Wissenschaftsgeschichte, Berlin, Germany, 2000. ISSN 0948-9444. 68 pp. LCCN

Q124.6 .P74 no.150. URL http://www.mpiwg-berlin.mpg.de/
Preprints/P150.PDF.

Blum:1986:WHG

[BDR86] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 4. Allgemeinverständliche Schriften. Biographisches und Kernphysik: Autobiographisches, Laudationes, Buchbesprechungen, Kernphysik, Münchener Festrede u.a. (German) [Werner Heisenberg: Collected Works: Series C, Volume 4: Philosophical and popular writings. Biographical and nuclear physics. Autobiographical, Laudations, Book Reviews, Nuclear Physics, Munich Festival Speeches, a.o.]. Piper, München, West Germany, 1986. ISBN 3-492-02928-0. x + 505 pp. LCCN ????

Beck:2009:MPM

[Bec09] Lorenz Friedrich Beck, editor. Max Planck und die Max-Planck-Gesellschaft: zum 150. Geburtstag am 23. April 2008 aus den Quellen zusammengestellt. (German) [Max Planck and the Max Planck Society: the 150th Birthday on 23 April 2008 from combined sources], volume 20 of Veröffentlichungen aus dem Archiv der Max-Planck-Gesellschaft. Archiv der Max-Planck-Gesellschaft, Berlin, Germany, 2009. ISBN 3-927579-24-6 (hardcover). 356 pp. LCCN QC16.P6 M39 2008. URL http://

catalogue.bnf.fr/ark:/12148/cb42484404j.

Bennewitz:1930:RWN

[Ben30] ?. Bennewitz. [Rezensionen:] Das Weltbild der neuen Physik. Von Max Planck. 52 Seiten. Verlag J. Ambr. Barth, Leipzig 1929.

Preis RM. 2,70. Angewandte Chemie, 43(8):182, 1930. CODEN ANCEAD. ISSN 1521-3757.

Bergwitz:1915:AGP

[Ber15] Karl Bergwitz, editor. Arbeiten aus den Gebieten der Physik, Mathematik, Chemie: Festschrift Julius Elster und Hans Geitel zum 60. Geburtstag. (German) [Works from the fields of physics, mathematics, chemistry: Festschrift for Julius Elster and Hans Geitel’s 60th birthday]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1915. xi + 719 pp. LCCN QD455 .F47

1915. URL http://catalog.hathitrust.org/api/volumes/
oclc/47733990.html.

Bergmann:1960:MPF

[Ber60] Peter G. Bergmann. Max-Planck-Festschrift: B. Kockel, W. Macke, A. Papapetrou and W. Frank, Editors, 1958 (VEB Deutscher Verlag der Wissenschaften, Berlin, 1959. 413 p. {DM} 54). Nuclear Physics, 15(??):527, February/March 1960. CO-DEN NUPHA7. ISSN 0029-5582 (print), 1873-1546 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0029558260903382.

Bernardoni:2014:BRM

[Ber14] Andrea Bernardoni. Book review: Matteo Valleriani (ed.), Metallurgy, Ballistics and Epistemic Instruments. The Nova scientia of Nicolò Tartaglia. A New Edition (Sources 6, Edition open access). Berlin: Max Planck Research Library for the History and Development of Knowledge, 2013. 56+304 pp., ISBN 978-3-8442-5258-3. Nuncius, 29(1):251–253, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02901013.

Berengut:2012:CGP

[BF12] J. C. Berengut and V. V. Flambaum. Comment on “Global Positioning System Test of the Local Position Invariance of Planck’s Constant”. Physical Review Letters, 109(6):068901:1, August 2012. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://journals.aps.org/prl/
abstract/10.1103/PhysRevLett.109.068901. See [KM12].

Braden:1981:CIB

[BFG81] C. H. Braden, R. F. Fox, and H. A. Gersch. Comment on “Inconsistency between the Boltzmann distribution for relativistic free particles and the Planck spectrum for thermal radiation in quantum theory”. Physical Review D (Particles and Fields), 23(6):1455–1457, March 1981. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://journals.aps.org/prd/abstract/10.
1103/PhysRevD.23.1455. See [Boy79].

Bertholet:1948:EMP

[BFV^{+}48] Alfred Bertholet, Heinrich Ficker, Agnes Von Zahn-Harnack, Wilhelm Westphal, Felix Ehrenhaft, Axel Von Harnack, J. R. Partington, E. Grasdorf, and E. Lamla. Erinnerungen an Max Planck. (German) [Memories of Max Planck]. Physikalische Blätter, 4(4): 161–174, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19480040406/abstract. Special issue

on Max Planck.

Birge:1919:MPV

[Bir19] Raymond T. Birge. The most probable value of the Planck constant h. Physical Review, 14(4):361–368, October 1919. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://journals.aps.org/pr/abstract/10.

1103/PhysRev.14.361. From the abstract: “The mean value from seven distinct methods is (6.5543 ± 0.0025) × 10−27 erg sec”. A modern accurate value is h = 6.62606891(58) × 10^{−34}Js [WSNO98].

Blackmore:1978:DNP

[Bla78] John T. Blackmore. Discussion notes : Is Planck’s ‘Principle’ true? British Journal for the Philosophy of Science, 29(4):347– 349, December 1978. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.
org/content/29/4/347.full.pdf+html;   http://www.jstor.
org/stable/687097.

Blackmore:1988:MCP

[Bla88] John Blackmore. Mach competes with Planck for Einstein’s favor. Historia Scientiarum = International journal of the History of Science Society of Japan, 35(??):45–89, 1988. CO-DEN HISCDU. ISSN 0285-4821. URL https://ci.nii.

ac.jp/naid/110006435088;   https://ci.nii.ac.jp/naid/
110009837382.

Blackmore:1992:EMD

[Bla92] John Blackmore. Ernst Mach — A Deeper Look: Documents and New Perspectives, volume 143 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. xix + 463 pp.

LCCN C5.53; QC6.9. URL https://link.springer.com/book/
10.1007/978-94-011-2771-4.

Badino:2013:RPH

[BN13] Massimiliano Badino and Jaume Navarro. Research and pedagogy: a history of quantum physics through its textbooks, volume 2 of Max Planck research library for the history and development of knowledge. Studies. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5871-X. 302 pp. LCCN QC173.98. URL http://

www.edition-open-access.de/studies/2/.

Born:1948:MKE

[Bor48] M. Born. Max Karl Ernst Ludwig Planck. 1858–1947. Obituary Notices of Fellows of the Royal Society, 6(17):161–188, November 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118

(electronic). URL http://jstor.org/doi/10.2307/768916;
http://www.jstor.org/stable/pdfplus/768916.pdf.

Born:1957:MP

[Bor57] Max Born. Max Planck 1858–1947. In Hermann Heimpel, editor, Die Grossen Deutschen: deutsche Biographie, page ?? Propylaen-Verlag bei Ullstein, Berlin, Germany, 1957.

Born:1963:AAGb

[Bor63a] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:MKE

[Bor63b] Max Born. Max Karl Ernst Ludwig Planck 1858–1947. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 88, pages 626–646. LCCN ???? Reprinted from Obit. Notices Roy. Soc. 6, 161–181 (1948).

Boyer:1979:IBB

[Boy79] Timothy H. Boyer. Inconsistency between the Boltzmann distribution for relativistic free particles and the Planck spectrum for thermal radiation in quantum theory. Physical Review D (Particles and Fields), 19(12):3635–3640, June 1979. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500

(CD-ROM). URL http://journals.aps.org/prd/abstract/
10.1103/PhysRevD.19.3635. See comment [BFG81].

Boyer:1983:DPR

[Boy83] Timothy H. Boyer. Derivation of the Planck radiation spectrum as an interpolation formula in classical electrodynamics with classical electromagnetic zero-point radiation. Physical Review D (Particles and Fields), 27(12):2906–2911, June 1983. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500

(CD-ROM). URL http://journals.aps.org/prd/abstract/

10.1103/PhysRevD.27.2906. See comment [Sac84] and reply [Boy84].

Boyer:1984:RCB

[Boy84] Timothy H. Boyer. Reply to “Comment on Boyer’s derivation of the Planck spectrum”. Physical Review D (Particles and Fields), 29(10):2418–2419, May 1984. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://journals.aps.org/prd/abstract/10.
1103/PhysRevD.29.2418. See [Boy83, Sac84].

Bieberbach:1924:ABE

[BP24] Ludwig G. E. M. Bieberbach and Max Planck. Antrittsrede von Bieberbach. Erwiderung von Planck. (German) [Inaugural speech by Bieberbach. Reply by Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–3, ???? 1924. CODEN SPWPAI. ISSN 0371-2435.

Bruche:1954:RHM

[BPL^{+}54] E. Brüche, G. Pfotzer, W. Lohmann, W. Bader, and E. Moritz. [Rezensionen:] Hartmann: Max Planck als Mensch und Denker/ Morrison und Feld: Experimental Nuclear Physics/Werner: Die Sterne Dürfet ihr Verschwenden/Rziha: Starkstromtechnik/Veröffentlichungen D. Arbeitsgem. F. Forschung/Der Grosse Brockhaus/Berichtigung. Physikalische Blätter, 10(4):187–191, 1954. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Braben:2014:MPR

[Bra14a] D. W. Braben. Max Planck: A reluctant revolutionary with a hunger of the soul. In Promoting the Planck Club: how defiant youth, irreverent researchers and liberated universities can foster prosperity indefinitely [Bra14b], pages 38–49. ISBN 1-118-54635-0, 1-306-47279-2, 1-118-54642-3, 1-118-54638-5 (ePub), 1-118-54636-9 (Adobe PDF). LCCN Q141. URL http://proquest.

tech.safaribooksonline.de/9781118546383.

Braben:2014:PPC

[Bra14b] D. W. Braben. Promoting the Planck Club: how defiant youth, irreverent researchers and liberated universities can foster prosperity indefinitely. Wiley, New York, NY, USA, 2014. ISBN 1-118-54635-0, 1-306-47279-2, 1-118-54642-3, 1-118-54638-5 (ePub), 1-118-54636-9 (Adobe PDF). LCCN Q141. URL http://

proquest.tech.safaribooksonline.de/9781118546383.

Breuer:2008:KVE

[Bre08] Reinhard Breuer. Das kosmische Vergessen: das expandierende Universum löscht seine Geschichte. Was ging bereits verloren?; Hirnforschung: wie das Auge die Welt verfilmt; Klimawandel: Drohen riesige Eismassen ins Meer zu rutschen?; Geschichte: Max Planck und Albert Einstein. (German) [The cosmic oblivion: the expanding Universe clears its history. What has already been lost?; Brain research: how the eye filmed the world; Climate change: threaten huge masses of ice to slip into the sea?; History: Max Planck and Albert Einstein], volume Mai 2008 of Spektrum der Wissenschaft. Spektrum der Wissenschaft Verlagsgesellschaft, Heidelberg, Germany, 2008. 106 pp. LCCN ????

Brown:1936:RBS

[Bro36] G. Burniston Brown. Review: Where Is Science Going? by Max Planck, James Murphy; Atomic Theory and the Description of Nature by Niels Bohr; Science and the Human Temperament by Erwin Schrödinger, James Murphy. Philosophy, 11(43):366–367, July 1936. ISSN 0031-8191 (print), 1469-817X (electronic). URL

http://www.jstor.org/stable/3746205.

Broda:1983:MPS

[Bro83] E. Broda. Max Planck in the social context. Report IC/83/200, International Center for Theoretical Physics, Trieste, Italy,

September 1983. 31 pp. URL http://www-lib.kek.jp/cgi-
bin/img_index?8506283.

Brown:2015:PDV

[Bro15] Brandon R. Brown. Planck: driven by vision, broken by war. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-021947-5 (hardcover), 0-19-021948-3 (e-book). xviii + 258 pp. LCCN QC16.P6 B76 2015.

Buttner:2003:ELC

[BRS03] Jochen Büttner, Jürgen Renn, and Matthias Schemmel. Exploring the limits of classical physics: Planck, Einstein, and the structure of a scientific revolution. Studies in History and Philosophy of Modern Physics, 34(1):37–59, March 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980200031X.

Bruche:1938:GFP

[Brü38] E. Brüche. Vom großen Fest der Physiker im Jahre 1938. Begleittext zur Schallplatte ,,Stimme der Wissenschaft”. (German) [From the great celebration of physicists in 1938. recording accompanying “voice of science”]. Unknown, 1938. Listed in [Hof10, page 214], but publisher and year are uncertain.

Bruche:1948:GGM

[Brü48a] E. Brüche. Gedenkfeier zum 90. Geburtstag Max Planck. (German) [Commemoration of the 90th birthday of Max Planck]. Physikalische Blätter, 4(3):125–126, March 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19480040309.

Bruche:1948:MPP

[Brü48b] Ernst Brüche. Max Planck und die Physikalische Gesellschaft 1889–1900. (German) [Max Planck and the [German] Physical Society 1889–1900]. Physikalische Blätter, 4(4):152–160, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1002/phbl.19480040405/abstract. Special issue on

Max Planck.

Brush:2002:CRM

[Bru02] Stephen G. Brush. Cautious revolutionaries: Maxwell, Planck, Hubble. American Journal of Physics, 70(2):119–127, February 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Boer:1958:RIU

[BT58] J. De Boer and F. Trendelenburg. Recommendations of the International Union of Pure and Applied Physics Commission for Symbols, Units and Nomenclature/Additions to Ch. I: General Principles and Recommendations/Additions to Ch. II: Symbols for Physical Quantities/Addition to Ch. III: Mathematical Operations and Symbols/Additions to Ch. IV: Units/Misprints to Be Corrected in U.I.P. 6 (1955)/Max-Planck-Medaille an W. Pauli. Physikalische Blätter, 14(6):259–262, 1958. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Beyler:2007:PMG

[BuDH07] Richard Beyler and Michael Eckert und Dieter Hoffmann. Die Planck-Medaille. (German) [The Planck Medal]. In Hoffmann and Walker [HW07], pages 217–236. ISBN 3-527-40585-2 (hardcover). LCCN QC9.G3 P59 2007.

Buhrke:1998:LAV

[Büh98a] Thomas Bührke. “Ein Akt der Verzweiflung.” Max Planck (1858– 1947). (German) [“An act of despair”. Max Planck (1858–1947)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 106–126. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Boeters:1978:MPM

[BWN78] K. E. Boeters, H. Welker, and H. Nelkowski. Max-Planck-Medaille 1978/Karl-Scheel-Preis 1978/Nominierung des Designierten Präsidenten/3. 7. Oktober 1978 in Berlin: 42. Deutsche Physikertagung/Fortbildung/Frühjahrstagung Heidelberg/Frühjahrstagung Gießen/Frühjahrstagung München/Frühjahrstagung Bad Nauheim. Physikalische Blätter, 34(3):144–147, 1978. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

C:1949:MPS

[C.49] T. C. Max Planck’s scientific autobiography and other papers: translated from the German by Frank Gaynor. 192 pages, 14 × 22 cm. New York, The Philosophical Library, 1949. Price, $3.75. Journal of The Franklin Institute, 248(6):575–, 1949. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0016003249902975.

Cahan:1985:IRG

[Cah85] David Cahan. The institutional revolution in German physics, 1865–1914. Historical Studies in the Physical Sciences, 15(2): 1–65, ???? 1985. CODEN HSPSAS. ISSN 0073-2672. URL

http://www.jstor.org/stable/27757549.

Cahan:1988:RPP

[Cah88] David Cahan. Review: Pride and prejudice in the history of physics: The German-speaking world, 1740–1945. Historical Studies in the Physical and Biological Sciences, 19(1):173– 191, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print),

1533-8355 (electronic). URL http://www.jstor.org/stable/
27757620.

Cahan:1999:BRB

[Cah99] David Cahan. Book review: Institute im Bild. Part 1: Bauten der Kaiser Wilhelm-Gesellschaft zur Forderung der Wissenschaften by Glenys Gill; Dagmar Klenke; Bibliographie zur Geschichte der Kaiser-Wilhelm-/Max-Planck-Gesellschaft zur Forderung der Wissenschaften (1911–1994) by Petra Hauke; Bibliometrische Profile von Instituten der Kaiser Wilhelm-Gesellschaft zur Forderung der Wissenschaften (1923–1943): Institute der Chemisch-Physikalisch-Technischen und der Biologisch-Medizinischen Sektion by Heinrich Parthey; Quelleninventar Max Planck by Dirk Ullmann. Isis, 90(2):387–388, June 1999. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237109.

Cassidy:1986:BRM

[Cas86] David C. Cassidy. Book review: Max Planck, The Dilemmas of an Upright Man. Science, 233(4766):896, August 22, 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/view/
10.2307/1697233;   http://www.sciencemag.org/content/
233/4766/896.1.full.pdf.

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Cavalleri:2003:CAP

[CC03] G. Cavalleri and E. Cesaroni. Comment on “Analog of Planck’s formula and effective temperature in classical statistical mechanics far from equilibrium”. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 68(2):028101:1–028101:2, August 2003. CO-DEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://journals.aps.org/pre/abstract/10.
1103/PhysRevE.68.028101. See [CG00].

Carati:2000:APF

[CG00] Andrea Carati and Luigi Galgani. Analog of Planck’s formula and effective temperature in classical statistical mechanics far from equilibrium. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 61(5):4791–4794, May 2000. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376

(electronic). URL http://journals.aps.org/pre/abstract/

10.1103/PhysRevE.61.4791. See comment [CC03] and reply [CG03].

Carati:2003:RCA

[CG03] Andrea Carati and Luigi Galgani. Reply to “Comment on ‘Analog of Planck’s formula and effective temperature in classical statistical mechanics far from equilibrium’ ”. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 68(2):028102:1–028102:2, August 2003.

CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://journals.aps.org/pre/abstract/10.
1103/PhysRevE.68.028102. See [CC03, CG00].

Callender:2001:PMP

[CH01] Craig Callender and Nick Huggett, editors. Physics meets philosophy at the Planck scale: contemporary theories in quantum gravity. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-66280-X (hardcover), 0-521-66445-4 (paperback). x + 365 pp. LCCN QC178 .P48 2001. URL http://www.loc.

gov/catdir/description/cam021/00050775.html;   http:/
/www.loc.gov/catdir/toc/cam021/00050775.html.

Chalmers:1914:MPC

[Cha14] S. D. Chalmers. The magneton and Planck’s constant. Nature, 92(2312):687, February 19, 1914. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v92/n2312/pdf/092687a0.pdf.

Chang:1997:CPC

[Cha97] Hasok Chang. Can Planck’s constant be measured with classical mechanics? International Studies in the Philosophy of Science, 11(3):223–243, 1997. CODEN ???? ISSN 0269-8595 (print),

1469-9281 (electronic). URL http://www.tandfonline.com//
doi/abs/10.1080/02698599708573568.

Clausius:1891:MWA

[Cla91] R. Clausius. Die mechanische Wärmetheorie, 2. Auflage der Abhandlungen “Über die mechanische Warmetheorie”. 3 Band: Entwicklung der besonderenVorstellung von der Natur der Wärme als einer Art der Bewegung. (German) [The mechanical theory of heat, second edition of essays “On The mechanical theory of heat”. Volume 3: Development of the specifics on the nature of heat as a type of motion]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1891. ???? pp. Three volumes edited by Max Planck and Carl Pulfrich.

Clausius:1989:BKW

[Cla89] R. Clausius. Über die bewegende Kraft der Wärme und die Gesetze, welche sich daraus für die Wärmelehre selbst ableiten lasse. (German) [On the motive power of heat and laws which can themselves be inferred for the theory of heat], volume 99 of Ostwalds Klassiker. W. Engelmann, Leipzig, Germany, 1989. ???? pp. Edited by Max Planck.

Cline:1987:MWM

[Cli87] Barbara Lovett Cline. Men who made a new physics: physicists and the quantum theory. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-11027-3 (paperback). xii + 274 pp. LCCN QC15 .C4 1987. US$11.95. URL http://www.loc.

gov/catdir/description/uchi051/87010786.htm;   http:/
/www.loc.gov/catdir/enhancements/fy0608/87010786-t.htm;
 http://www.loc.gov/catdir/enhancements/fy0609/87010786-
b.htm.

Cardona:2008:MPC

[CM08] M. Cardona and W. Marx. Max Planck — a conservative revolutionary. Percorsi, 24(??):39–54, ???? 2008. URL http://www.

tddft.org/TDDFT2008/lectures/MC5.pdf.

Cranston:1927:BRB

[Cra27] John A. Cranston. [Book Review:] The evolution and development of the quantum theory. By N. M. Bligh, A.R.C.Sc., A.I.C. with a foreword by Professor Max Planck. Pp. 112. London: Edward Arnold & Co., 1926. Price 9s. Journal of the Society of Chemical Industry, London, 46(18):405, 1927. CODEN JSCIAN. ISSN 1934-9971.

Crawford:1984:AAH

[Cra84] Elisabeth Crawford. Arrhenius, the atomic hypothesis, and the 1908 Nobel prizes in physics and chemistry. Isis, 75(3):503–522, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232940.

Cohen:1979:SPL

[CS79a] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. xxxiv + 929 pp. LCCN Q174 .B67 vol. 21

QC7. URL https://link.springer.com/book/10.1007/978-
94-009-9349-5.

Cohen:1979:MPS

[CS79b] Robert S. Cohen and John J. Stachel. Max Planck and the statistical definition of entropy [1959b]. In Cohen and Stachel [CS79a], chapter 17, pages 235–246. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN

Q174 .B67 vol. 21 QC7. URL https://link.springer.com/
chapter/10.1007/978-94-009-9349-5_17.

Braga:1958:MPS

[dACB58] Carlos de Azevedo Coutinho Braga. Max Planck: sua vida e sua obra. (Portuguese) [Max Planck: his life and his works]. Centro de Estudos Humanísticos, Porto, Portugal, 1958. 35 pp.

Braga:1960:MP

[dACB60] Carlos de Azevedo Coutinho Braga. Max Planck. ????, Coimbra, Portugal, 1960. 18 pp.

Dahmen:2008:MPS

[Dah08] Silvio R. Dahmen. Max Planck on statistical thermodymamics and stochastic systems. In Hoffmann [Hof08c], pages 335–350. ISBN 3-527-40819-3 (hardcover). LCCN QC484 .M39 2008. URL

http://d-nb.info/987214098/04. All of the references in this

book to volumes of Beiblätter zu den Annalen der Physik are too large by 10: change XXII and XXIII to XII and XIII, respectively.

Planck:2000:CDP

[Dar00] Olivier Darrigol. Continuities and discontinuities in Planck’s Akt der Verzweiflung [act of despair]. Annalen der Physik (1900), 9 (??):951–960, ???? 2000. ISSN 1521-3889.

Darrigol:2001:HDM

[Dar01] Olivier Darrigol. The historians’ disagreements over the meaning of Planck’s quantum. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):219–239, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

deBroglie:1948:MPW

[dB48] Louis de Broglie. Max Planck und das Wirkungsquantum. (German) [Max Planck and the quantum of action]. Physikalische Blätter, 4(4):138–142, April 1948. CODEN PHBLAG.

ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/phbl.19480040402/

abstract. Special issue on Max Planck.

Darwin:1948:AGS

[DdCA48] Sir Charles Darwin and Professor E. N. da C. Andrade. Address given by Sir Charles Darwin at the Max Planck memorial service held on 23 April 1948 in the aula of the University of Göttingen. Notes and Records of the Royal Society of London, 6(1):69–70, December 1948. CODEN NOREAY. ISSN 0035-9149 (print),

1743-0178 (electronic). URL http://www.jstor.org/stable/
view/10.2307/531259.

DelRegato:1979:MP

[Del79] Juan A. Del Regato. Max Planck. International Journal of Radiation Oncology*Biology*Physics, 5(11–12):2097–2111, 1979. CO-DEN IOBPD3. ISSN 0360-3016 (print), 1879-355X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0360301679909659.

Dingler:1939:MPB

[Din39] Hugo Dingler. Max Planck und die Begründung der sogenannten modernen theoretischen Physik. (German) [Max Planck and the establishment of so-called modern theoretical physics]. Ahnenerbe-Stiftung Verlag, Berlin, Germany, 1939. 32 pp. LCCN ????

Dingle:1946:PRFc

[Din46a] Herbert Dingle. Planck’s radiation formula. Nature, 157(3996): 737, June 1, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v157/n3996/pdf/157737b0.pdf.

Dingle:1946:PRFb

[Din46b] Herbert Dingle. Planck’s radiation formula: a correction. Nature, 157(3991):556, April 17, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v157/n3991/pdf/157556b0.pdf.

See [Din46c].

Dingle:1946:PRFa

[Din46c] Herbert Dingle. Planck’s radiation formula derived without atomicity. Nature, 157(3990):515, April 20, 1946. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v157/

n3990/pdf/157515b0.pdf. See correction [Din46b].

Dinkler:1959:MPR

[Din59] Erich Dinkler. Max Planck und die Religion: Hans-Joachim Iwand zum 60. Geburtstag. (German) [Max Planck and religion: Hans-Joachim Iwand’s 60th birthday]. Zeitschrift für Theologie und Kirche, 56(2):201–223, ???? 1959. ISSN 0044-3549 (print),

1868-7377 (electronic). URL http://www.jstor.org/stable/
view/10.2307/23584125.

Dinkler:1961:PM

[Din61] Erich Dinkler. Planck, Max. In Die Religion in Geschichte und Gegenwart: Handwörterbuch für Theologie und Religionswissenschaft. (German) [Religion in history and today], volume 5, pages 404–405. Mohr, Tübingen, West Germany, third edition, 1961. LCCN BL31 .R42.

Douglas:1949:BRB

[Dou49] R. G. Douglas. Book review: Scientific Autobiography and Other Papers, by Max Planck, translated from the German by Frank Gaynor. Popular Astronomy: a review of astronomy and allied sciences, 57(??):525, ???? 1949. URL http://adsabs.harvard.

edu/abs/1949PA.....57..525D.

delRegato:1985:RP

[dR85] Juan A. del Regato. Radiological physicists. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-469-9. 188 pp. LCCN QC774.A2 D44 1985.

Driess:2001:RSP

[Dri01] Matthias Drieß. [Rezensionen:] Stella Planckia am Himmel des Allgemeinwissens: Max Planck: Vorträge, Reden, Erinnerungen. Hrsg. von Hans Roos und Armin Hermann. Springer, Berlin, Heidelberg, 2001. 224 S., geb., 69,90 DM. ISBN 3hyphen;540-41274-3. Nachrichten aus der Chemie, 49(11):1326, 2001. CODEN NACHFB. ISSN 1868-0054.

DeBoer:1958:RIU

[DT58] J. De Boer and F. Trendelenburg. Recommendations of the International Union of Pure and Applied Physics Commission for Symbols, Units and Nomenclature/Additions to Ch. I: General Principles and Recommendations/Additions to Ch. II: Symbols for Physical Quantities/Addition to Ch. III: Mathematical Operations and Symbols/Additions to Ch. IV: Units/Misprints to Be Corrected in U.I.P. 6 (1955)/Max-Planck-Medaille an W. Pauli. Physikalische Blätter, 14(6):259–262, 1958. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Duhem:1888:MMM

[Duh88] P. Duhem. Sur un mémoire de M. Max Planck ayant pour titre: sur le principe de l’accroissement de l’entropie . (French) [On a memoir of Mr. Max Planck entitled “On the principle of increasing entropy”]. Journal de Physique Th’eorique et Appliqu’ee, 7(1): 124–127, 1888. CODEN JPTAAT. ISSN 0368-3893.

Dyson:1970:MWM

[Dys70] Freeman J. Dyson. Menschheit und Weltall: Max-Planck-Vortrag von der Physikertagung in Salzburg. (German) [Mankind and the Universe: Max Planck presentation of the physicists meeting in Salzburg]. Physikalische Blätter, 26(1):7–14, January 1970. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19700260104/abstract.

Ebert:1963:MPP

[Ebe63] R. Ebert. Max Planck: Physikalische Abhandlungen und Vorträge, herausgeg. vom Verband Deutscher Physikalischer Gesellschaften und der Max-Planck-Gesellschaft zur Förderung der Wissenschaften e.V. Verlag Friedrich Vieweg & Sohn, Braunschweig 1958. 1. Aufl., Bd. I: XV, 776 S., Bd. II: XI, 716 S., Bd. III: XII, 426 S., Preis des Gesamtwerkes DM 150.–. Angewandte Chemie, 75(1):112, 1963. CODEN ANCEAD. ISSN 1521-3757.

Eckert:1997:ATP

[Eck97a] Michael Eckert. Die Autorität der theoretischen Physik in Deutschland. Zum 50. Todestag von Max Planck. (German) [The authority of theoretical physics in Germany. On the 50th anniversary of the death of Max Planck]. Physik in unserer Zeit, 28(6): 246–250, 1997. CODEN PHUZAH. ISSN 1521-3943.

Eckert:1997:MPA

[Eck97b] Michael Eckert. Max Planck und Arnold Sommerfeld: Ein halbes Jahrhundert Physikgeschichte in Briefen. (German) [Max Planck and Arnold Sommerfeld: A half century of physics history in letters]. Physikalische Blätter, 53(10):1014–1016, 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Eckert:2008:MPL

[Eck08] Michael Eckert. Max Planck’s later work on quantum theory. In Hoffmann [Hof08c], pages 643–652. ISBN 3-527-40819-3 (hardcover). LCCN QC484 .M39 2008. URL http://d-nb.info/ 987214098/04. All of the references in this book to volumes of Beiblätter zu den Annalen der Physik are too large by 10: change XXII and XXIII to XII and XIII, respectively.

Ebeling:2008:MPT

[EH08a] Werner Ebeling and Dieter Hoffmann. Max Planck und die Thermodynamik. (German) [Max Planck and thermodynamics]. In Über thermodynamische Gleichgewichte. (German) [On thermodynamic equilibrium] [Pla08b], pages vii–xiv. ISBN 3-8171-3299-9 (paperback). LCCN QC318.T47 P53 2008. Includes Planck’s 1880 habilitation thesis [Pla80], edited and introduced by Werner Ebeling and Dieter Hoffmann.

Eberling:2008:TGM

[EH08b] W. Eberling and D. Hoffmann. Über Thermodynamische Gleichgewichte von Max Planck. (German) [On the thermodynamic equilibria of Max Planck]. Verlag Harri Deutsch, Berlin, Germany, 2008.

Engler:2008:MSL

[EI08] Fynn Ole Engler and Mathias Iven, editors. Moritz Schlick: Leben, Werk und Wirkung, volume 1 of Schlickiana. Parerga, Berlin, Germany, 2008. ISBN 3-937262-82-2. LCCN B3329.S4874 M674 2008.

Einstein:1906:ORM

[Ein06] Albert Einstein. Optik: [Rezension von] M[ax] Planck: Vorlesungen über die Theorie der Wärmestrahlung. (German) [Review of Max Planck: Lectures on the theory of thermal radiation]. Beiblätter zu den Annalen der Physik, 30(15):764–766, ???? 1906. CODEN BEAPAI. ISSN 0375-846X. URL http://hdl.handle.

net/2027/uva.x000269580?urlappend=%3Bseq=772.

Einstein:1907:BMA

[Ein07a] A. Einstein. Berichtigung zu meiner Arbeit: Die Plancksche Theorie der Strahlung etc.. (German) [Correction to my paper: “The Planck Theory of Radiation etc.”]. Annalen der Physik (Berlin), 327(4):800, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See [Ein07b].

Einstein:1907:PTS

[Ein07b] Albert Einstein. Die Plancksche Theorie der Strahlung und die Theorie der spezifischen Wärme. (German) [Planck’s theory of radiation and the theory of specific heat]. Annalen der Physik (Berlin), 327(1):180–190, 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). See correction [Ein07a].

Einstein:1913:MPF

[Ein13] Albert Einstein. Max Planck als Forscher. (German) [Max Planck as researcher]. Die Naturwissenschaften, 1(45):1077– 1079, November 7, 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01493259. Reprinted in [KKRS96,

561–563].

Einstein:1982:MPM

[Ein53] Albert Einstein. Max Planck in memoriam. In Ideas and Opinions [Ein82], pages 78–79. ISBN 0-517-88440-2. LCCN AC35 .E526 1982x.

Einstein:1954:IO

[Ein54a] Albert Einstein. Ideas and Opinions. Crown Publishers, New York, NY, USA, 1954. vi + 377 pp. LCCN AC35 .E526 1954.

Einstein:1954:MPM

[Ein54b] Albert Einstein. Max Planck in memoriam. In Ideas and Opinions [Ein54a], pages 78–79. LCCN AC35 .E526 1954.

Einstein:1977:WSG

[Ein77] Albert Einstein. Prologue and epilogue: a Socratic dialogue. In Max Planck, editor, Where is Science Going?, page 221. AMS Press, New York, NY, USA, 1977. ISBN 0-404-14696-1. LCCN Q175 .P57 1977.

Einstein:1982:IO

[Ein82] Albert Einstein. Ideas and Opinions. Three Rivers Press, New York, NY, USA, 1982. ISBN 0-517-88440-2. 377 pp. LCCN AC35 .E526 1982x.

Einstein:1991:TMP

[Ein91] Albert Einstein. Tribute to Max Planck. In Ferris and Fadiman [FF91], page 590. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Einstein:1932:PES

[EM32] Albert Einstein and James Murphy. Prologue and epilogue: a Socratic dialogue. In Max Planck, editor, Where is Science Going?, page ?? W. W. Norton & Co., New York, NY, USA, 1932. LCCN Q175 .P57 1932. Reprinted in [Ein77].

Einstein:2006:AEM

[EP06] A. Einstein and M. Planck. Antrittsrede und Erwiderung von Max Planck am Leibniztag. (German) [inaugural speech and reply from Max Planck at the Leibniz Workshop]. In Albert Einstein: Akademie-Vorträge: Sitzungsberichte der Preußischen Akademie der Wissenschaften, 1914–1932. (German) [Albert Einstein: Academy lectures: Proceedings of the Prussian Academy of Sciences, 1914–1932] [Sim06], pages 1–7. ISBN 3-527-40609-3 (hardcover), 3-527-60895-8 (e-book), 3-527-62244-6 (e-book). LCCN QC71 .E36 2006.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http://www.loc.gov/

catdir/enhancements/fy0824/2006934045-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0824/2006934045-d.
html;   http://www.loc.gov/catdir/toc/fy0708/2006934045.
html.

Faddeev:1996:HWU

[Fad96] L. D. Faddeev. How we understand quantization a hundred years after Max Planck. Physikalische Blätter, 52(7–8):689–690, 1996. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Falkoff:1950:BRB

[Fal50] David L. Falkoff. Book review: Introduction to Theoretical Physics, by Max Planck. American Journal of Physics, 18(8): 527–528, 1950. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/18/8/10.1119/1.1932655.

Falkenhagen:1959:EMP

[Fal59] H. Falkenhagen. Die Elektrolytarbeiten von Max Planck und ihre weitere Entwicklung. (German) [The electrolyte works by Max Planck and their further development]. In Max-Planck-Festschrift 1958 [KMP59], pages 11–?? LCCN QC3 .M45.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Fischbach:1991:NTQ

[FGH91] Ephraim Fischbach, Geoffrey L. Greene, and Richard J. Hughes. New test of quantum mechanics: Is Planck’s constant unique? Physical Review Letters, 66(3):256–259, January 1991. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://journals.aps.org/prl/abstract/10.
1103/PhysRevLett.66.256.

Fischer:2000:PAW

[FH00] Wolfram Fischer and Rainer Hohlfeld. Die Preußische Akademie der Wissenschaften zu Berlin, 1914–1945. (German) [The Prussian Academy of Sciences in Berlin 1914–1945], volume 8 of Forschungsberichte / Interdisziplinäre Arbeitsgruppen. Akademie-Verlag, Berlin, Germany, 2000. ISBN 3-05-003327-4 (paperback). ISSN 0949-7285. xi + 594 pp. LCCN AS182.B38 P748 2000.

Fierz:1975:VMP

[Fie75] Markus Fierz. Zur Verleihung der Max-Planck-Medaille an Gregor Wentzel. (German) [On the award of the Max Planck Medal of Gregor Wentzel]. Physikalische Blätter, 31(8):368–369, 1975. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Fischer:2000:MPS

[Fis00] Ernst Peter Fischer. Max Planck und die Suche nach der natürlichen Ordnung. (German) [Max Planck and the search for the natural order]. Max-Planck-Forschung : das Wissenschaftsmagazin der Max-Planck-Gesellschaft, 1(??):36–42, ???? 2000.

ISSN 0341-7727. URL http://www.mpg.de/mpforschung.

Fischer:2007:PMP

[Fis07] Ernst Peter Fischer. Der Physiker: Max Planck und das Zerfallen der Welt. (German) [The physicist: Max Planck and the disintegration of the world]. Siedler, München, Germany, second edition, 2007. ISBN 3-88680-837-8. 351 pp. LCCN QC16.P6.

URL http://d-nb.info/984826645/04.

Fischer:2008:MPP

[Fis08a] Ernst Peter Fischer. Max Planck ein Portrait ; Vorlesung. (German) [Max Planck a Portrait: lecture]. Uni-Auditorium Wissenschaftsgeschichte. Komplett-Media, Grünwald, Germany, 2008. ISBN 3-8312-6225-X, 3-8312-9225-6, 3-8312-9522-0. LCCN ???? 1 CD-ROM (76 minutes).

Fischer:2008:RPG

[Fis08b] Ernst Peter Fischer. Revolutionär aus Pflichtgefühl. (German) [Revolutionary from a sense of duty]. Wiener Zeitung, ??(??):

??, April 18, 2008. URL https://www.mpiwg-berlin.mpg.de/
Presse-PDF/Planck_WienerZeitung_18.04..pdf.

Fischer:2010:PMP

[Fis10a] Ernst Peter Fischer. Der Physiker: Max Planck und das Zerfallen der Welt. (German) [The physicist: Max Planck and the disintegration of the world]. Pantheon, München, Germany, 2010. ISBN 3-570-55116-4. 351 pp. LCCN ????

Fischer:2010:HQE

[Fis10b] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Flint:1958:MP

[Fli58] H. T. Flint. Max Planck (1858–1947). Nature, 181(4616):1098– 1099, April 19, 1958. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v181/n4616/pdf/1811098a0.pdf. Written for the cen-

tennial of Planck’s birth.

Flugge:1947:RSW

[Flü47] S. Flügge. [Rezensionen:] Scheinprobleme der Wissenschaft von Max Planck. Vortag gehalten in Göttingen am 17. Juni 1946. J. A. Barth-Verlag, Leipzig 1947. 1,50 RM, 31 S. Angewandte Chemie, 59(5–6):183, 1947. CODEN ANCEAD. ISSN 1521-3757.

Franca:1996:MET

[FMM96] H. M. França, A. Maia, Jr., and C. P. Malta. Maxwell electromagnetic theory, Planck’s radiation law, and Bose–Einstein statistics. Foundations of Physics, 26(8):1055–1068, August 1996. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF02061403.

Fokker:1914:MER

[Fok14] A. D. Fokker. Die mittlere Energie rotierender elektrischer Dipole im Strahlungsfeld. (German) [The mean energy of rotating electrical dipoles in a radiation field]. Annalen der Physik (Berlin), 348(5):810–820, ???? 1914. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/andp.19143480507/
abstract;   https://en.wikipedia.org/wiki/Fokker%E2%80%

93Planck_equation. This paper introduces the famous equation now known as the “Fokker–Planck equation”; see also [Pla17b].

Foote:1916:NRD

[Foo16] Paul D. Foote. A new relation derived from Planck’s Law. Physical Review, 7(2):224, February 1916. CODEN PHRVAO.

ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

journals.aps.org/pr/abstract/10.1103/PhysRev.7.224.

Ficker:1927:AFE

[FP27] ?. Ficker and Max Planck. Antrittsrede von Ficker. Erwiderung von Planck. (German) [Inaugural speech by Ficker. Reply of Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):3–4, ???? 1927. CODEN SPWPAI. ISSN 0371-2435.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Franklin:1925:BRB

[Fra25] Wm. S. Franklin. Book review: A Survey of Physics: A Collection of Lectures and Essays by Max Planck. Science, 62(1611): 438–441, November 13, 1925. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/view/10.2307/1649209;   http://www.sciencemag.
org/content/62/1611/438.2.full.pdf.

Frank:1931:RDM

[Fra31] Ph. Frank. [Rezensionen:] Dr. Max Planck, Professor der theoretischen Physik an der Universität Berlin. Einführung in die Theorie der Wärme zum Gebrauch bei Vorträgen, sowie zum Selbstunterricht. Mit 7 Figuren. Einführung in die Theoretische Physik. V. Band. Verlag S. Hirzel, Leipzig 1930. VI 251 S. Zeitschrift für Angewandte Mathematik und Mechanik, 11(3):249, 1931. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Franck:1948:MP

[Fra48] James Franck. Max Planck: 1858–1947. Science, 107(2786): 534–537, May 21, 1948. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/view/10.2307/1675567;   http://www.sciencemag.
org/content/107/2786/534.full.pdf.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2005PhP....
.7..204F;   http://link.springer.com/article/10.1007/
s00016-004-0236-9;   http://www.springerlink.com/content/
1422-6960/.

Friedman:1981:NPP

[Fri81] Robert Marc Friedman. Nobel Physics Prize in perspective. Nature, 292(5826):793–798, August 27, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Frohlich:2001:MPM

[Frö01] Jürg Fröhlich. Max-Planck-Medaille: Ein Blick zurück, ein Blick nach vorn. (German) [Max Planck Medal: A look back, a look forward]. Physikalische Blätter, 57(7–8):53–58, 2001. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-1. LCCN ???? URL http://

deposit.ddb.de/cgi-bin/dokserv?id=2758637\%26prov=M\%
26dok\_var=1\%26dok\_ext=htm. 4 CDs.

Fan:2009:PPL

[FSBR09] Yuwei Fan, S. B. Singer, Raymond Bergstrom, and B. C. Regan. Probing Planck’s Law with incandescent light emission from a single carbon nanotube. Physical Review Letters, 102(18):187402:1– 187402:4, May 2009. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://journals.aps.

org/prl/abstract/10.1103/PhysRevLett.102.187402.

Fragozo:2012:MPR

[FV12] Fernando Fragozo and Antonio A. P. Videira. Max Planck e a relevância metafísica do realismo para a ciência natural. (Portuguese) [Max Planck and the metaphysical significance of realism to natural science]. Revista portuguesa de filosofia, 69(1–2): 51–72, ???? 2012. ISSN 0870-5283. URL http://www.jstor.

org/stable/view/10.2307/41955623.

Gaede:1903:RMP

[Gae03] ?. Gaede. [Rezensionen:] M. Planck: Über die Natur des weißen Lichtes. Annalen der Physik (4.), 7, 390. 1902. Zeitschrift für Psychologie und Physiologie der Sinnesorgane, 32(??):149–151, ????

1903. ISSN 0233-2302. URL http://vlp.mpiwg-berlin.mpg.
de/references?id=lit33191.

Gamow:1985:MPL

[Gam85a] George Gamow. M. Planck and light quanta. In Thirty Years That Shook Physics: the Story of Quantum Theory [Gam85b], chapter 1, pages 6–28. ISBN 0-486-24895-X (paperback). LCCN QC174.12 .G35 1985. US$4.95. URL http://www.loc.gov/

catdir/description/dover032/85006797.html.

Gamow:1985:TYS

[Gam85b] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover Publications, Inc., New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12 .G35 1985. US$4.95. URL http://www.loc.

gov/catdir/description/dover032/85006797.html.

Garber:1976:SRP

[Gar76] Elizabeth Garber. Some reactions to Planck’s law, 1900–1914. Studies in History and Philosophy of Science Part A, 7(2):89– 126, ???? 1976. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0039368176900133.

Gortler:1949:RLV

[GBM^{+}49] Görtler, Bilharz, Moritz, Dehlinger, Brüche, and Zenneck. [Rezensionen:] Lense: Vorlesungen über Höhere Mathematik/ Dörge: Differential- und Integralrechnung/Reutter: Darstellende Geometrie/von Sanden: Praktische Mathematik/Daeves und Beckel: Grosszahl-Forschung und Häufigkeits-Analyse/ Peters: Dreistellige Tafeln/Gravelius: Vierstellige Tafeln/ Bertele: Fünfstellige Logarithmen/Seeliger: Die Grundbeziehungen der Neuen Physik/Mie: Elektrizität und Magnetismus/Brüche und Scherzer: Geometrische Elektronenoptik/Hahn: Die Kettenreaktion des Urans/Gerlach: Probleme der Atomenergie/ Klein: Elektronenstrahloszillographen/Cremer: Raumakustik/ Simon, Grund und Forschung: Zeitmarken der Erde/Becksmann: Geologie als Erdgeschichte/Kropp: Das Aussenweltproblem der Modernen Atomphysik/Max Planck. Physikalische Blätter, 5(3): 145–151, 1949. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Gearhart:2002:PQH

[Gea02] Clayton A. Gearhart. Planck, the quantum, and the historians. Physics in Perspective (PIP), 4(2):170–215, May 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00016-002-8363-7. See also [Bad09].

Gearhart:2008:MPB

[Gea08] Clayton A. Gearhart. Max Planck and blackbody radiation. In Hoffmann [Hof08c], pages 395–418. ISBN 3-527-40819-3 (hardcover). LCCN QC484 .M39 2008. URL http://d-nb.info/ 987214098/04. All of the references in this book to volumes of Beiblätter zu den Annalen der Physik are too large by 10: change XXII and XXIII to XII and XIII, respectively.

Gearhart:2009:BRB

[Gea09] Clayton Gearhart. Book review: Max Planck: Die Entstehung der modernen Physik. Isis, 100(3):672–673, September 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/649181.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Glashow:2015:MP

[Gla15] Sheldon Lee Glashow. Max Planck. 74-minute video, directed

by Michele Calvano., 2015. URL http://www.amazon.com/Max-
Planck-Sheldon-Lee-Glashow/dp/B00ZGN16NQ.

Gruneisen:1947:MP

[GMM47] E. Grüneisen, F. Möglich, and A. Meiner. Max Planck. Annalen der Physik (Berlin), 436(7–8):317, 1947. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Gruneisen:1948:MPM

[GMM48] Eduard Grüneisen, Friedrich Möglich, and Arthur Meiner. Max Planck in memoriam. Annalen der Physik (Berlin), 438(1):1, 1948. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Goldberg:1976:MPP

[Gol76] Stanley Goldberg. Max Planck’s philosophy of nature and his elaboration of the Special Theory of Relativity. Historical Studies in the Physical Sciences, 7(??):125–160, ???? 1976. CODEN HSPSAS. ISBN 0-691-08169-7. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757355.

Gorham:1991:PPJ

[Gor91] Geoffrey Gorham. Planck’s principle and Jeans’s conversion. Studies in History and Philosophy of Science Part A, 22(3):471– 497, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print),

1879-2510 (electronic). URL http://www.sciencedirect.com/
science/article/pii/003936819190005D.

Grandy:1969:TPF

[Gra69] W. T. Grandy, Jr. Theory of photons in a fully ionized gas. II. Planck’s Law and the dielectric constant. Physical Review, 177(1):371–375, January 1969. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://journals.

aps.org/pr/abstract/10.1103/PhysRev.177.371.

Greenberg:1990:TRT

[Gre90] Valerie D. Greenberg. Transgressive readings: the texts of Franz Kafka and Max Planck. University of Michigan Press, Ann Arbor, MI, USA, 1990. ISBN 0-472-10158-7. viii + 224 pp. LCCN P302 .G69 1990.

Gregis:2019:AAM

[Gré19] Fabien Grégis. Assessing accuracy in measurement: the dilemma of safety versus precision in the adjustment of the fundamental physical constants. Studies in History and Philosophy of Science Part A, 74(??):42–55, April 2019. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/S0039368117302972.

Grigoryan:1987:MPH

[Gri87] A. T. Grigor′yan. Max Planck and his role in the development of science (on the 130th anniversary of his birth). In Studies in the history of physics and mechanics, 1987 (Russian), pages 152–159,

245. “Nauka”, Moscow, USSR, 1987.

Griffel:1998:GVV

[Gri98] Anton Griffel. Gesellschaft und Verantwortung: Zu Vorträgen von Max Planck über Willensfreiheit. (German) [Society and responsibility: lectures of Max Planck on free will]. ARSP: Archiv f”ur Rechts- und Sozialphilosophie / Archives for Philosophy of Law and Social Philosophy, 84(4):517–528, ???? 1998. CODEN

???? ISSN 0001-2343. URL http://www.jstor.org/stable/
view/10.2307/23679928.

Gibas:2007:CWF

[GSU07] Monika Gibas, Rüdiger Stutz, and Justus H. Ulbricht, editors. Couragierte Wissenschaft: eine Festschrift für Jürgen John zum 65. Geburtstag. (German) [Courageous science: a festschrift for Jürgen John on his 65th birthday]. Glaux-Verlag, Jena, 2007. ISBN 3-940265-09-8. 415 pp. LCCN DD281.6 .C68 2007. URL

http://d-nb.info/986332275/04.

Gutfreund:2015:ZGG

[Gut15] Hanoch Gutfreund. Zwei der Glänzendsten Gestirne: Max Planck und Albert Einstein. (German) [Two of the most shining stars: Max Planck and Albert Einstein]. In Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy] [LMOS15], pages 310– 343. ISBN 3-11-037598-2 (hardcover). LCCN AS182.B38 B477 2015.

Haag:1970:RQP

[Haa70] Rudolf Haag. Die Rolle der Quantenfeldtheorie in der Physik der letzten Jahrzehnte: Max-Planck-Vortrag von der Physikertagung 1970 in Hannover. (German) [The role of quantum field theory in physics in recent decades: Max Planck Lecture at the 1970 Physics Workshop in Hannover]. Physikalische Blätter, 26(12): 529–535, 1970. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Haag:1975:LLV

[Haa75] Rudolf Haag. Laudatio für Léon van Hove anläßlich der Verleihung der Max-Planck-Medaille am 25. September 1974. (German) [Eulogy for Léon van Hove at the award ceremony of the Max Planck Medal on 25 September 1974]. Physikalische Blätter, 31(2):82–84, 1975. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hachtmann:2007:ANP

[Hac07] Rüdiger Hachtmann. Anpassung und Nonkonformität. Zur politischen Positionierung Alax Plancks während der NS-Zeit. (German) [Adaptation and nonconformity. On the political positioning of Max Planck during the Nazi era]. In Gibas et al. [GSU07], pages 25–43. ISBN 3-940265-09-8. LCCN DD281.6 .C68 2007.

URL http://d-nb.info/986332275/04.

Hahn:1957:EPE

[Hah57] Otto Hahn. Einige persönliche Erinnerungen an Max Planck. (German) [Some personal memories of Max Planck]. Mitteilungen aus der Max-Planck-Gesellschaft zur Förderung der Wissenschaften, ??(??):242–246, 1957.

Haidar:2011:MKE

[Hai11] Riad Haidar. Max Karl Ernst Ludwig Planck. Photoniques, 52 (??):28–31, March/April 2011. CODEN ???? ISSN 1629-4475 (print), 2269-8418 (electronic). URL http://www.photoniques.

com/articles/photon/abs/2011/02/photon201152p28/photon201152p28.
html.

Hamla:1972:WTM

[Ham72] Ernst Hamla. Zur 25. Wiederkehr des Todestages von Max Planck/Physiker-Gräber auf dem Göttinger Friedhof. (German) [On the 25th anniversary of the death of Max Planck/physicist graves in the cemetery in göttingen]. Physikalische Blätter, 28 (10):443–446, 1972. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hartmann:1938:MPM

[Har38] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker]. Verlag Karl Siegismund, Berlin, Germany, 1938. 189 + 1 pp. LCCN Q171 .P7. URL http:

//catalog.hathitrust.org/api/volumes/oclc/992978.html.

Hartmann:1943:MPM

[Har43] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker]. Siegismund, Berlin, Germany, 1943. 200 pp. LCCN MLCS 87/03423 (Q).

Hartmann:1944:MPS

[Har44] Hans Hartmann. Max Planck som menneske og vitenskapsmann. (Norwegian) [Max Planck as a person and thinker]. Stenersen, Oslo, Norway, 1944. 158 pp. LCCN ????

Hartmann:1948:MPM

[Har48] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker]. S. Hirzel, Leipzig, Germany, second edition, 1948. 126. pp. LCCN ????

Hartmann:1953:MPMa

[Har53a] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker], volume 1 of Plinius-Bücher. Ott, Basel, Switzerland and Düsseldorf, West Germany, third edition, 1953. 255 pp. LCCN ????

Hartmann:1953:MPMb

[Har53b] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker], volume 490 of Ullstein-Bücher. Ullstein, Frankfurt am Main, West Germany, 1953. 173 pp. LCCN ????

Hartmann:1958:MPM

[Har58] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker], volume 1 of Plinius-Bücher. Ott, Basel, Switzerland and Düsseldorf, West Germany, 1958. 255 pp. LCCN ????

Hartmann:1959:GMP

[Har59] Hans Hartmann. Gehört Max Planck in die Geschichte der Philosophie?. (German) [Did Max Planck belong in the history of philosophy?]. Zeitschrift für philosophische Forschung, 13(1): 118–128, January/March 1959. ISSN 0044-3301. URL http://

www.jstor.org/stable/view/10.2307/20481034.

Hartmann:1964:MPMa

[Har64a] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker], volume 490 of Ullstein. Lang, Frankfurt, West Germany, 1964. 173 pp. LCCN ????

Hartmann:1964:MPMb

[Har64b] Hans Hartmann. Max Planck als Mensch und Denker. (German) [Max Planck as a person and thinker], volume 490 of Ullstein Buch. Ullstein, Berlin, Germany, 1964. 174 pp. LCCN ????

Harris:1998:NMP

[Har98] Randy Allen Harris. A note on the Max Planck effect. Rhetoric Society Quarterly, 28(4):85–89, Autumn 1998. CODEN ???? ISSN 0277-3945 (print), 1930-322X (electronic). URL http://

www.jstor.org/stable/view/10.2307/3886350.

Haskins:1914:NEC

[Has14] Charles N. Haskins. Note on the evaluation of the constant C_{2} in Planck’s radiation equation. Physical Review, 3(6):476– 478, June 1914. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://journals.aps.org/pr/
abstract/10.1103/PhysRev.3.476.

Heilbron:1987:BRB

[HB87] J. L. Heilbron and Lawrence Badash. Book review: The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. American Journal of Physics, 55(8):765–766, August 1987. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/55/8/10.1119/1.15019.

Harlem:1948:GGM

[HD48] ?. Harlem and ?. Dahme. Gedenkfeier zum 90. Geburtstag von Max Planck/Arbeitsgemeinschaft Ferromagnetismus/Clausthaler Arbeitstagung über Naturwissenschaft — Religion Weltanschauung. (German) [Commemoration of the 90th anniversary of Max Planck /Ferromagnetism Society/Clausthaler workshop on science — religion — world view]. Physikalische Blätter, 4(3): 125–126, March 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19480040309/abstract.

Hauke:2001:LMP

[HD01] Petra Hauke and Maren Dissmann. Literatur über Max Planck: Bestandsverzeichnis: anlässlich des Jubileums 100 Jahre Quantentheorie. (German) [Literature on Max Planck: Inventory: On the occasion of the 100th jubilee of quantum theory], volume 14 of Veröffentlichungen aus dem Archiv zur Geschichte der Max-Planck-Gesellschaft. Max-Planck-Gesellschaft, Berlin, Germany, 2001. ISBN 3-927579-14-9. 99 pp. LCCN Q111 .V47 Bd.14.

Heisenberg:1931:PRA

[Hei31] Werner Heisenberg. Positivismus und reale Außenwelt. Von M. Planck. Vortrag, gehalten am 12. November 1930 im Harnack-Haus der Kaiser Wilhelm-Gesellschaft zur Förderung der Wissenschaften. 35 Seiten. Akademische Verlagsgesellschaft, Leipzig 1931. (German) [Positivism and the real world outside. On Max Planck. Lecture, held on 12 November 1930 at the Harnack House the Kaiser Wilhelm Society for the F romoting the Sciences. 35 pages. academic Verlagsgesellschaft, Leipzig, 1931]. Angewandte Chemie, 44(10):189, 1931. CODEN ANCEAD. ISSN 1521-3757.

Heisenberg:1935:BMP

[Hei35] Werner Heisenberg. Besprechung: Max Planck, Die Physik im Kampf um die Weltanschauung. (German) [Book review: Max Planck, The physics in the struggle of belief]. Die Naturwissenschaften, 23(20):321, May 17, 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.

springer.com/article/10.1007/BF01494105.

Heisenberg:1938:MPA

[Hei38] Werner Heisenberg. Max Planck. Achtzig Jahre alt. (German) [Max Planck. Eighty years old]. Kölnische Zeitung, 201(??):1–2, April 23, 1938.

Heisenberg:1943:BMP

[Hei43a] Werner Heisenberg. Besprechung: Max Planck: Wege zur physikalischen Erkenntnis. Bd. II. (German) [Book review: Max Planck, Paths to physical knowledge]. Leipziger Neueste Nachrichten, Ausg. A, ??(??):??, June 25, 1943.

Heisenberg:1943:MP

[Hei43b] Werner Heisenberg. [Max Planck]. Völkischer Beobachter, ??(??): ??, March 1943. Incompletely cited in [Cas09, page 339, 450].

Heisenberg:1947:MP

[Hei47a] Werner Heisenberg. Max Planck. Göttinger Universitäts-Zeitung, 2(10):6, ???? 1947. ISSN 0170-138X.

Heisenberg:1947:MPN

[Hei47b] Werner Heisenberg. Max Planck [Nachruf]. (German) [Max Planck [obituary]]. Göttinger Universitäts-Zeitung, 2(21):1–2, ???? 1947. ISSN 0170-138X.

Heisenberg:1949:ALM

[Hei49] Werner Heisenberg. Die Auswirkungen des Lebenswerkes Max Plancks. (German) [The impact of the life work of Max Planck]. Angewandte Chemie, 61(4):115–117, 1949. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1958:PEPa

[Hei58] Werner Heisenberg. Die Plancksche Entdeckung und die philosophischen Grundfragen der Atomlehre. (German) [Planck’s discovery and the philosophical basic questions of atomic theory]. Die Naturwissenschaften, 45(10):227–234, January 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/wupx14qt11u1631q.

English translation in [Hei74].

Heitler:1969:PCA

[Hei69] W. Heitler. Physik, Chemie und andere Dinge: Max-Planck-Vortrag von der Physikertagung Karlsruhe. (German) [Physics, chemistry and other things: Max Planck’s lecture at the Physics Workshop in Karlsruhe]. Physikalische Blätter, 25(2):49–57, 1969. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Heisenberg:1974:AF

[Hei74] Werner Heisenberg. Across the Frontiers. Harper & Row, New York, NY, USA, 1974. ISBN 0-06-011823-7. ???? pp. LCCN ????

URL http://alsos.wlu.edu/information.aspx?id=941.

Heilbron:1986:DUM

[Hei86a] J. L. Heilbron. The dilemmas of an upright man: Max Planck as spokesman for German science. University of California Press, Berkeley, CA, USA and Los Angeles, CA, USA, 1986. ISBN 0-520-05710-4. xiii + 238 + 16 pp. LCCN QC16.P6 H45 1986.

Heisenberg:1986:MPQ

[Hei86b] Werner Heisenberg. Max Planck und die Quantentheorie (1943). (German) [Max Planck and quantum theory (1943)]. In Blum et al. [BDR86], pages 65–67. ISBN 3-492-02928-0. LCCN ????

Heilbron:1988:MPL

[Hei88a] J. L. Heilbron. Max Planck: Ein Leben für die Wissenschaft 1858– 1947. Mit einer Auswahl der allgemeinverständlichen Schriften von Max Planck. (German) [Max Planck: A Life in Science 1858– 1947. With a selection of generally understandable writings of Max Planck]. S. Hirzel Verlag, Stuttgart, Germany, 1988. ISBN 3-7776-0392-9. 474 pp. LCCN QC16.P5 H455 1988.

Heilbron:1988:DMP

[Hei88b] John L. Heilbron. I dilemmi di Max Planck. (Italian) [The dilemmas of Max Planck]. Saggi Scientifici. [Essays in Science]. Editore Boringhieri, Turin, Italy, 1988. ISBN 88-339-0427-X. 222 pp. Portavoce della scienza tedesca. [Spokesman for German science], Translated from the English by Riccardo Valla, Translation edited by Ivana Gambaro.

Heilbron:1988:PCD

[Hei88c] John L. Heilbron. Planck: 1858–1947: une conscience déchirée. (French) [Planck: 1858–1947: a torn conscience]. Un savant, une époque. Belin, Paris, France, 1988. ISBN 2-7011-1103-X. 255 pp. LCCN ????

Heilbron:2000:DUM

[Hei00] J. L. Heilbron. The dilemmas of an upright man: Max Planck and the fortunes of German science. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00439-6. xiii + 254 pp. LCCN QC16.P6 H45 2000.

Heilbron:2006:MPL

[Hei06] J. L. Heilbron. Max Planck: Ein Leben für die Wissenschaft 1858– 1947. Mit einer Auswahl der allgemeinverständlichen Schriften von Max Planck. (German) [Max Planck: A Life in Science 1858– 1947. With a selection of generally understandable writings of Max Planck]. S. Hirzel Verlag, Stuttgart, Germany, second edition, 2006. ISBN 3-7776-1438-6. xiii + 647 pp. LCCN ????

Heilbron:2007:MPC

[Hei07] J. L. Heilbron. Max Planck’s compromises on the way to and from the absolute. In Evans and Thorndike [ET07], pages 21–37. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. LCCN QC174.12 .Q346 2007. URL http://www.loc.gov/

catdir/enhancements/fy0824/2006934045-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0824/2006934045-d.
html;   http://www.loc.gov/catdir/toc/fy0708/2006934045.
html.

Henley:1960:BRB

[Hen60] Ernest M. Henley. Book review: Max Planck als Mensch und Denker by Hans Hartmann. Isis, 51(2):249, June 1960. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/226883.

Hentschel:1996:PNS

[Hen96] Klaus Hentschel, editor. Physics and National Socialism: an anthology of primary sources, volume 18 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-5312-0 (Basel), 0-8176-5312-0 (Boston). ci + 406 + civ pp. LCCN Q127.G3 P48 1996.

Hentschel:2007:MAM

[Hen07] Klaus Hentschel. The mental aftermath: the mentality of German physicists 1945–1949. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920566-3 (hardcover). 205 pp. LCCN QC9.G3 H46 2007. URL http://www.loc.gov/

catdir/toc/fy0805/2007280005.html.

Herrmann:1924:RBT

[Her24] ?. Herrmann. [Rezensionen:] Vorlesungen über die Theorie der Wärmestrahlung. Von Dr. Max Planck. Fünfte, abermals umgearb. Aufl. Leipzig 1923. Verlag J. A. Barth. Angewandte Chemie, 37(18):260, 1924. CODEN ANCEAD. ISSN 1521-3757.

Hertz:1958:BPQ

[Her58] Gustav Hertz. Die Bedeutung der Planckschen Quantumtheorie für die experimentelle Physik. (German) [The significance of Planck’s quantum theory for experimental physics]. Die Naturwissenschaften, 45(17):401–405, September 1, 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hermann:1971:GQT

[Her71] Armin Hermann. The Genesis of Quantum Theory (1899–1913). MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-08047-8. ix + 165 pp. LCCN QC174 .H4613 1971.

Hermann:1972:MPG

[Her72] Armin Hermann. Max Planck zum Gedenken. (German) [Commemorating Max Planck]. Physikalische Blätter, 28(10):437–442, 1972. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hermann:1973:MPS

[Her73a] Armin Hermann. Max Planck in Selbstzeugnissen und Bilddokumenten. (German) [Max Planck in autobiographies and image documents], volume 198 of Rororo Bildmonographien; Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, West Germany, 1973. ISBN 3-499-50198-8. 148 pp. LCCN QC16.P6 H47.

Hermann:1973:MPP

[Her73b] Armin Hermann. Max Planck Praeceptor Physicae. (German/ Latin) [Max Planck physics teacher]. Physikalische Blätter, 29 (11):483–487, 1973. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Herneck:1976:MPB

[Her76] F. Herneck. Max Planck über wissenschaftliche Arbeit Erstveröffentlichung eines Lautdokuments. (German) [Max Planck scientific work: first publication of a source document]. Die Naturwissenschaften, 63(11):530, November 1976. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF00596853.

Hermann:1977:SNA

[Her77] Armin Hermann. Die Suche nach dem Absoluten: Max Planck und die Platonische Philosophie. (German) [The search for the absolute: Max Planck and Platonic philosophy]. ????, ????, 1977. ???? pp. LCCN ????

Herschbach:2001:MBE

[Her01] Dudley Herschbach. Molecular beams entwined with quantum theory: A bouquet for Max Planck. Annalen der Physik (Berlin), 10 [i.e. 513](1–2):163–176, 2001. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL 2-W/abstract;

http://onlinelibrary.wiley.com/doi/10.1002/1521-3889(
200102)10:1/2%3C163::AID-ANDP163%3E3.0.CO.

Hering:2003:BRB

[Her03] Rainer Hering. Book review: Die Präsidenten der Kaiser Wilhelm-Gesellschaft im National-sozialismus: Max Planck, Carl Bosch und Albert Vögler zwischen Wissenschaft und Macht, by Ulrike Kohl. German Studies Review, 26(3):640–642, October 2003. CODEN GSREFO. ISSN 0149-7952 (print), 2164-

8646 (electronic). URL http://www.jstor.org/stable/view/
10.2307/1432776.

Haupt:1963:MPZ

[HH63] P. Haupt and H. Hartmann. Max Planck und die Züchtungsforschung. (German) [Max Planck and plant breeding research]. Physikalische Blätter, 19(7):328–329, 1963. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hibberd:1991:BRB

[Hib91] John Hibberd. Book review: Transgressive Readings: The Texts of Franz Kafka and Max Planck, by Valerie D. Greenberg. Colloquia Germanica, 24(3):256–257, ???? 1991. ISSN 0010-1338. URL

http://www.jstor.org/stable/view/10.2307/23980399.

Hiebert:1968:BRB

[Hie68] Erwin N. Hiebert. Book review: Letters on Wave Mechanics: Schrödinger, Planck, Einstein, Lorentz. K. Przibram, Ed. Translated from the German with an introduction by Martin J. Klein. Philosophical Library, New York, 1967. xx + 75 pp., illus. $6. Science, 160(3828):664–666, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/160/3828/664.full.pdf;   http:
//www.sciencemag.org/content/160/3828/664.short.

Hartmann:1968:RKB

[HKS^{+}68] Hans Hartmann, K. Kirchgäßner, U. E. Schröder, W. Kofink, A. Hermann, F. Stöckmann, G. Lehner, and H. Ebert. [Rezensionen:] Kretzschmar: Max Planck als Philosoph/Strubecker: Einführung in die höhere Mathematik/Landau: Lehrbuch der Theoretischen Physik/Landau: Lehrbuch der Theoretischen Physik/Waerden: Sources of Quantum Mechanics/Beiser: Modern Physics/Inönu, Zweifel: Developments in Transport Theory/Herausgeg: Beiträge zur Licht-und Elektronenmikroskopie/ Brookes: Basic Instrumentation for Engineers and Physicists. Physikalische Blätter, 24(5):235–239, 1968. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Heideking:1998:AIG

[HMF98] Jürgen Heideking, Christof Mauch, and Marc Frey, editors. American intelligence and the German resistance to Hitler: a documentary history. Widerstand, dissent and resistance in the Third Reich. Westview Press, Boulder, CO, USA, 1998. ISBN 0-8133-3636-8 (paperback). xxii + 457 pp. LCCN D810.S7 A58 1998.

Hirosige:1970:GBA

[HN70] Tetu Hirosige and Sigeko Nisio. The genesis of the Bohr atom model and Planck’s theory of radiation. Japanese Studies in the History of Science, 9(??):35–47, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Hoffmann:1948:SPJ

[Hof48] Fritz Hoffmann. Strahlungsmessungen der PTR um die Jahrhundertwende. (German) [Radiation measurements of the PTR at the turn of the century]. Physikalische Blätter, 4(4):143–145, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/
doi/10.1002/phbl.19480040403/abstract. Special issue on

Max Planck.

Hoffmann:1996:WGT

[Hof96] Dieter Hoffmann. Wider die geistige Trennung. Die Max-Planck-Feier(n) in Berlin 1958. (German) [Against spiritual separation. The Max Planck celebration(s) in Berlin in 1958]. Deutschland Archiv, 29(??):525–534, ???? 1996. ISSN 0012-1428.

Hoffmann:1999:DCM

[Hof99] Dieter Hoffmann. The divided centennial: The 1958 Max Planck celebration(s) in Berlin. Osiris (Series 2), 14(??):138–149, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (elec-

tronic). URL http://www.jstor.org/stable/301965.

Hoffmann:2000:VAR

[Hof00] Dieter Hoffmann. Das Verhältnis der Akademie zu Republik und Diktatur. Max Planck als Sekretär. (German) [The relation of the Academy to the Republic and the dictatorship. Max Planck as secretary]. In Die Preußische Akademie der Wissenschaften zu Berlin, 1914–1945. (German) [The Prussian Academy of Sciences in Berlin 1914–1945] [FH00], pages 53–85. ISBN 3-05-003327-4 (paperback). ISSN 0949-7285. LCCN AS182.B38 P748 2000.

Hoffmann:2001:ECP

[Hof01] Dieter Hoffmann. On the experimental context of Planck’s foundation of quantum theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(3–4):240– 259, October 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hoffmann:2008:YCS

[Hof08a] Dieter Hoffmann. “… you can’t say to anyone to their face: your paper is rubbish.” Max Planck as Editor of the Annalen der Physik. Annalen der Physik (Berlin), 17 [i.e. 520](5):273–301, 2008. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Hoffmann:2008:MPAa

[Hof08b] Dieter Hoffmann. Max Planck als akademischer Lehrer von Moritz Schlick und die Beziehungen beider Gelehrter im Spiegel ihrer Korrespondenz. (German) [Max Planck as an academic teacher of Moritz Schlick and the relations between the two scholars in the mirror of their correspondence]. In Engler and Iven [EI08], pages 31–58. ISBN 3-937262-82-2. LCCN B3329.S4874 M674 2008.

Hoffmann:2008:MPAb

[Hof08c] Dieter Hoffmann, editor. Max Planck: Annalen Papers. Wiley-VCH, Weinheim, Germany, 2008. ISBN 3-527-40819-3 (hardcover). xv + 857 pp. LCCN QC484 .M39 2008. URL http:/ /d-nb.info/987214098/04. All of the references in this book to volumes of Beiblätter zu den Annalen der Physik are too large by 10: change XXII and XXIII to XII and XIII, respectively.

Hoffmann:2008:MPE

[Hof08d] Dieter Hoffmann. Max Planck: die Entstehung der modernen Physik. (German) [Max Planck: the emergence of modern physics], volume 2442 of Beck’sche Reihe. Wissen. C. H. Beck, München, Germany, 2008. ISBN 3-406-56242-6 (paperback). 128 pp. LCCN QC16.P6 H64 2008.

Hoffmann:2008:MPK

[Hof08e] Dieter Hoffmann. Max Planck: Kaum eine Sitzung die ich versäumt habe. Max Planck und die Physikalische Gesellschaft. (German) [Max Planck: Hardly a meeting that I had neglected. Max Planck and the Physical Society]. Physik Journal, 7(3):27–33, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http:

//www.pro-physik.de/details/articlePdf/1104315/issue.
html;   http://www.pro-physik.de/details/news/1121445/
Kaum_eine_Sitzung_die_ich_versaeumt_habe.html.

Hoffmann:2008:ZWP

[Hof08f] Dieter Hoffmann. Zwischen Wissenschaft und Politik: Max Planck und Albert Einstein:. Kollegen im Widerstreit. (German) [Between science and politics: Max Planck and Albert Einstein: Colleagues in the conflict]. Spektrum der Wissenschaft (German translation of Scientific American), 5(??): 32–39, May 2008. CODEN SPEKDI. ISSN 0170-2971. URL

http://www.spektrum.de/magazin/max-planck-und-albert-
einstein-kollegen-im-widerstreit/947196.

Hoffmann:2009:PMP

[Hof09] Dieter Hoffmann. Der Physiker Max Planck. (German) [The physicist Max Planck]. In Beck [Bec09], pages 61–86. ISBN 3-927579-24-6 (hardcover). LCCN QC16.P6 M39 2008. URL http:

//catalogue.bnf.fr/ark:/12148/cb42484404j.

Hoffmann:2010:MPM

[Hof10] Dieter Hoffmann, editor. Max Planck und die moderne Physik. (German) [Max Planck and modern physics]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-87844-0 (hardcover). xii + 235 pp. LCCN

QC16.P6 M42 1997. URL http://link.springer.com/book/
10.1007/978-3-540-87845-2.

Hohenberg:1999:MPM

[Hoh99] Pierre Hohenberg. Max-Planck-Medaille: One hundred years of nonequilibrium patterns. Physikalische Blätter, 55(7–8):53–56, 1999. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Holm:1913:ANP

[Hol13] E. Holm. Anwendung der neueren Planckschen Quantenhypothese zur Berechnung der rotatorischen Energie des zweiatomigen Gases. (German) [Application of the newer Planck’s quantum hypothesis to calculate the rotational energy of the diatomic gas]. Annalen der Physik (Berlin), 347(16):1311–1320, ???? 1913. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19133471618/abstract.

Holton:1999:CFM

[Hol99] Gerald Holton. Centennial focus: Millikan’s measurement of Planck’s constant. Physical Review Focus, 3(??):23, April 22, 1999. CODEN PRFHAQ. ISSN 1539-0748. URL http://

link.aps.org/doi/10.1103/PhysRevFocus.3.23;   http://
physics.aps.org/story/v3/st23. From the article: “Millikan

[[Mil16]] determined h to have the value 6.57×10^{−27} erg-sec to ‘a precision of about 0.5 per cent,’ a value far better than had been obtained in any previous attempt.” A modern accurate value is h = 6.62606891(58) × 10^{−34}Js [WSNO98].

Hopper:1996:ERR

[Hop96] Teresa Hopper. Essay review: Recent books on Nazism and science: Uncertainty: The Life and Science of Werner Heisenberg by David Cassidy; The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science by John Heilbron; The Nazi Doctors: Medical Killing and the Psychology of Genocide by Robert Jay Lifton; Surviving the Swastika: Scientific Research in Nazi Germany by Kristie Macrakis; The Rocket and the Reich: Pennemunde and the Coming of the Ballistic Era by Michael Neufeld; Operation Epsilon: The Farm Hall Transcripts; Heisenberg’s War: The Secret History of the German Bomb by Thomas Powers; Racial Hygiene: Medicine under the Nazis by Robert N. Proctor; Science, Technology and National Socialism by Monika Renneberg and Mark Walker; German National Socialism and the Quest for Nuclear Power, 1939–1949 by Mark Walker; Health, Race and German Politics between National Unification and Nazism, 1870–1945 by Paul Weindling. Historical Studies in the Physical and Biological Sciences, 27(1):163– 176, ???? 1996. CODEN HSPSEW. ISSN 0890-9997 (print),

1533-8355 (electronic). URL http://www.jstor.org/stable/
27757772.

Hosemann:1978:MPM

[Hos78] Rolf Hosemann. Max-Planck-Medaille 1978 für Paul Ewald. Physikalische Blätter, 34(12):715–718, 1978. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Heisenberg:1933:BMP

[HPE^{+}33] W. Heisenberg, J. Picht, P. P. Ewald, K. Fajans, H. Cassel, et al. Besprechung: Max Planck, Wege zur physikalischen Erkenntnis. (German) [Reviews: Max Planck, Paths to physical knowledge]. Die Naturwissenschaften, 21(33):608–611, ???? 1933. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/v458v1565v5716jg.

Hermann:2001:MPV

[HR01] Armin Hermann and Hans Roos. Max Planck: Vorträge, Reden, Erinnerungen. (German) [Max Planck: Lectures, Speeches, Recollections]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41274-3 (hardcover). viii + 224 pp. LCCN QC3 .P65 2001. URL http:

//link.springer.com/book/10.1007/978-3-642-56594-6;
http://www.gbv.de/dms/bowker/toc/9783540412748.pdf;
http://www.gbv.de/dms/fazrez/FNUW200109261084460.pdf;
 http://www.zentralblattmath.org/zmath/en/search/?an=
0978.01057.

Hoffmann:2010:BLL

[HRR10a] Dieter Hoffmann, Hole Rößler, and Gerald Reuther. “Lachkabinett” und “großes Fest” der Physiker. Walter Grotrians “physikalischer Einakter” zu Max Plancks 80. Geburtstag. Berichte zur Wissenschaftsgeschichte, 33(1):30–53, March 2010. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Hoffmann:2010:LLF

[HRR10b] Dieter Hoffmann, Hole Rößler, and Gerald Reuther. “Lachkabinett” und “großes Fest” der Physiker. Walter Grotrians “physikalischer Einakter” zu Max Plancks 80. Geburtstag. (German) [“Laughing cabinet” and “big festival” of the physicist. Walter Grotrian’s “physical one-act play” for Max Planck’s 80th birthday]. Berichte zur Wissenschaftsgeschichte, 33(1):30–53, March 2010. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Hoffmann:1997:WWM

[HS97] Dieter Hoffmann and Thomas Stange. Das zu wissen wäre mir von hohem Werte über das Schicksal der Bibliothek von Max Planck. (German) [That would be me knowing the high value of the fate of the library of Max Planck]. Physikalische Blätter, 53(10):1019–1021, 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hentschel:1999:BZM

[HT99] Klaus Hentschel and Renate Tobies, editors. Brieftagebuch zwischen Max Planck, Carl Runge, Bernhard Karsten und Adolf Leopold. (German) [Letter diary between Max Planck, Carl Runge, Bernhard Karsten and Adolf Leopold], volume 24 of BBGNT. Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik [BBGNT. Berlin Contributions to the History of the Sciences and Technology]. ERS-Verlag, Berlin, Germany, 1999. ISBN 3-928577-34-4. 273 pp. Foreword, introduction and annotation by Klaus Hentschel and Renate Tobies, With an appendix containing the promotion and habilitation files of Planck and Runge.

Hull:1978:PP

[HTD78] David L. Hull, Peter D. Tessner, and Arthur M. Diamond. Planck’s principle. Science, 202(4369):717–723, 1978. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/202/4369/717.
abstract.

Hui:2013:CEE

[Hui13] Alexandra Hui. Changeable ears: Ernst Mach’s and Max Planck’s studies of accommodation in hearing. Osiris (Series 2), 28(1): 119–145, January 2013. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/10.1086/671366.

Hund:1975:MPM

[Hun75] Friedrich Hund. Max-Planck-Medaille für Gregor Wentzel. Physikalische Blätter, 31(12):543, 1975. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hoffmann:2007:PZA

[HW07] Dieter Hoffmann and Mark Walker, editors. Physiker zwischen Autonomie und Anpassung: Die Deutsche Physikalische Gesellschaft im Dritten Reich. (German) [Physicists between autonomy and adaptation: The German Physical Society in the Third Reich]. Wiley-VCH, Weinheim, Germany, 2007. ISBN 3-527-40585-2 (hardcover). xii + 675 pp. LCCN QC9.G3 P59 2007.

Infeld:1961:TRS

[Inf61] L. Infeld. Theory of the red shift and the nature of Planck’s constant. Nature, 191(4794):1184, September 16, 1961. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Irons:2005:WCM

[Iro05] F. E. Irons. Why the cavity-mode method for deriving Planck’s law is flawed. Canadian Journal of Physics = Journal canadien de physique, 83(6):617–628, 2005. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Ishiwara:1915:UBW

[Ish15] Jun Ishiwara. Universelle Bedeutung des Wirkungsquantums. (German) [The universal meaning of the quantum of action]. Proceedings of Tokyo Mathematico-Physical Society, 8(??):106–116, ???? 1915. See also English translation [Ish17] and commentary [PB17].

Ishiwara:2017:UMQ

[Ish17] Jun Ishiwara. The universal meaning of the quantum of action. European Physical Journal H, 42(4–5):523–536, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/

epjh/e2017-80041-1. See analysis [PB17].

Ivanenko:1959:CHM

[Iva59a] D. Ivanenko. Collection in the honor of Max Planck. Physics-Uspekhi, 2(2):337–342, February 1959. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://ufn.ru/

en/articles/1959/2/g/.

Ivanenko:1959:SVC

[Iva59b] D. Ivanenko. Sbornik v chest’ Maksa Planka. Uspekhi Fizicheskikh Nauk, 67(3):543–548, March 1959. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/

ru/articles/1959/3/d/.

Jacobi:2002:BBP

[Jac02] Manfred Jacobi. Buchbesprechung: Max Planck. Vorträge Reden Erinnerungen by H. Roos, A. Hermann. Physik in unserer Zeit, 33(2):90, 2002. CODEN PHUZAH. ISSN 1521-3943.

Jacobi:2008:BBP

[Jac08a] Manfred Jacobi. Buchbesprechung: Der Physiker. Max Planck und das Zerfallen der Welt. Von E. P. Fischer. Physik in unserer Zeit, 39(2):100, 2008. CODEN PHUZAH. ISSN 1521-3943.

Jacobi:2008:SNA

[Jac08b] Manfred Jacobi. Die Suche nach dem Absoluten. Zum 150. Geburtstag von Max Planck. (German) [The search for the absolute. On the 150th birthday of Max Planck]. Physik in unserer Zeit, 39(2):64–70, 2008. CODEN PHUZAH. ISSN 1521-3943.

Jaeger:1979:MPG

[Jae79] Robert Jaeger. Max Planck gegen Hitler. (German) [Max Planck against Hitler]. Physikalische Blätter, 35(10):484, 1979. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Jorda:1997:PNE

[JDB^{+}97] S. Jorda, E. Dreisigacker, Gunnar Berg, R. Scharf, I. Peschel, R. Bock, I. Hofmann, J. Redfearn, S. Jorda, and R. Scharf. Physik-Nobelpreis für die Entwicklung von Laserkühlung und Atomfallen/kurzgefaßt/Max Planck zum Gedenken/WE-Heraeus-Programm Weiterbildung in den Grundlagen unternehmerischen Denkens und Handelns jetzt auch für Fachhochschulen/Einstein im ADLON/kurzgefaßt/Auf Lebenszeit Studierender? Oder: Beitragsvermeidung durch Stillhalten/Frühjahrstagungen: Internet-Adressen/Blick in die USA/Großbritannien: Diamond und JET/ JET setzt Weltrekordmarken/Trägheitsfusion mit schweren Ionen im Aufwind/Trägheitsfusion/Erstes Bose–Einstein-Kondensat in Europa/Quasiteilchen mit gebrochenzahliger Ladung beobachtet/ Laserlicht verwandelt sich in Materie. (German) [Nobel Prize in Physics for the development of laser cooling and atom traps/in brief/Max Planck commemoration/we-heraeus-program training in the basics of entrepreneurship now for polytechnics/Einstein in ADLON/in brief/student’s life? or: Post prevention by keeping quiet/Spring Meetings: Internet addresses/looking at the United States/UK: Diamond and JET/JET sets world record marks/inertial confinement fusion with heavy ions on the rise/ inertial confinement fusion/first Bose–Einstein condensate in Europe/quasiparticles with the fractional charge observed/laser light turns into matter]. Physikalische Blätter, 53(11):1078–1088, 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Jungnickel:2017:SPH

[JM17] Christa Jungnickel and Russell McCormmach. The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-49564-X (hardcover), 3-319-49565-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxi + 460 pp. LCCN QC9.G3 J86 2017.

Jordan:1948:MPW

[Jor48] Pascual Jordan. Max Plancks wissenschaftliche Leistung. (German) [Max Planck’s scientific achievement]. Physikalische Blätter, 4(4):134–137, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19480040401/abstract. Spe-

cial issue on Max Planck.

Jost:1995:MET

[Jos95] Res Jost. Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers], volume 34 of Lecture notes in physics. New series M, monographs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. 286 pp. LCCN QC16.J67

J67 1995. URL http://www.springerlink.com/openurl.
asp?genre=issue\%26issn=0940-7677\%26volume=34. Edited

posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Juttner:1911:RBT

[Jüt11] F. Jüttner. [Rezensionen:] Vorlesungen über Thermodynamik. Von Max Planck. 3. erweiterte Auflage. Verlag von Veit & Co., Leipzig 1911. Preis geb. 7,50 Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 17(21):960, 1911. CO-DEN ZEAPAA. ISSN 0005-9021.

Juttner:1917:RDS

[Jüt17] F. Jüttner. [Rezensionen:] Dynamische und statistische Gesetzmäßigkeit. Rede, gehalten bei der Feier zum Gedächtnis des Stifters der Berliner Friedrich-Wilhelms-Universität am 3. Aug. 1914, von Max Planck. 32 S. Verlag von J. A. Barth, Leipzig. 1914. Preis geh. I Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 23(3–4):63, 1917. CODEN ZEA-PAA. ISSN 0005-9021.

Juttner:1918:RBN

[Jüt18a] F. Jüttner. [Rezensionen:] Die Naturwissenschaften. Herausgegeben von Arnold Berliner und August Pütter. 6. Jahrg., Heft 17, 26. April 1918. Max Planck zur Feier seines 60. Geburtstages. S. 195-264. Verlag von Julius Springer, Berlin. Preis geh. 4 Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 24(23–24):395, 1918. CODEN ZEAPAA. ISSN 0005-9021.

Juttner:1918:RBT

[Jüt18b] F. Jüttner. [Rezensionen:] Vorlesungen über Thermodynamik. Von Max Planck. 5. Auflage. X und 290 S. mit 5 Figuren im Text. Verlag von Veit & Co. Leipzig. 1917. Preis geb. 10 Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 24(23– 24):395, 1918. CODEN ZEAPAA. ISSN 0005-9021.

Juttner:1921:RBE

[Jüt21] F. Jüttner. [Rezensionen:] Die Entstehung und bisherige Entwicklung der Quantentheorie. Nobel-Vortrag, gehalten vor der Königlich Schwedischen Akademie der Wissenschaften zu Stockholm am 2. Juni 1920 von Max Planck. 32 Seiten. Verlag von Joh. Ambr. Barth, Leipzig 1920. Preis geh. 4 Mk. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 27(23–24):583, 1921. CODEN ZEAPAA. ISSN 0005-9021.

Juttner:1922:RBT

[Jüt22] F. Jüttner. [Rezensionen:] Vorlesungen über die Theorie der Wärmestrahlung. Von Max Planck. 4., abermals umgearbeitete Auflage. XII und 224 Seiten mit 6 Abbildungen. Verlag von Joh. Ambr. Barth, Leipzig 1921. Preis geh. 36 Mk., geb. 44 Mk., dazu Teuerungszuschläge. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 28(5–6):156–157, 1922. CODEN ZEAPAA. ISSN 0005-9021.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kangro:1969:PWS

[Kan69] H. Kangro. Das Paschen-Wiensche Strahlungsgesetz und seine Abänderung durch Max Planck. (German) [The Paschen–Wien Law and its amendment by Max Planck]. Physikalische Blätter, 25(5):216–220, 1969. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Kangro:1970:VPS

[Kan70] Hans Kangro. Vorgeschichte des Planckschen Strahlungsgesetzes; Messungen und Theorien der spektralen Energieverteilung bis zur Begründung der Quantenhypothese. (German) [Prehistory of Planck’s radiation law: Measurements and theories of spectral energy distribution up to the creation of quantum theory], volume 11 of Boethius; Texte und Abhandlungen zur Geschichte der exakten Wissenschaften. F. Steiner., Wiesbaden, West Germany, 1970. xv + 271 pp. LCCN QC484 .K35.

Planck:1972:POP

[Kan72a] Hans Kangro. Planck’s original papers in quantum physics, volume 1 of Classic papers in physics. Taylor and Francis, Boca Raton, FL, USA, 1972. ISBN 0-85066-060-2. viii + 60 pp. LCCN ????

Kangro:1972:POP

[Kan72b] Hans Kangro, editor. Planck’s original papers in quantum physics, volume 1 of Classic papers in physics. Taylor and Francis, Boca Raton, FL, USA, 1972. ISBN 0-85066-060-2. viii + 60 pp.

LCCN QC475 .P55 1972. URL https://archive.org/stream/
PlancksOriginalPapersInQuantumPhysics/Kangro-PlancksOriginalPapersInQuantumPhysics_

djvu.txt. German and English edition annotated by Hans Kangro, Translated by D. ter Haar and Stephen G. Brush.

Kangro:1976:EHP

[Kan76] Hans Kangro. Early history of Planck’s radiation law. Taylor and Francis, Boca Raton, FL, USA, 1976. ISBN 0-85066-063-7. xvii + 282 pp. LCCN QC484 .K3513 1976.

Kangro:2008:PMK

[Kan08] Hans Kangro. Planck, Max Karl Ernst Ludwig. In Complete Dictionary of Scientific Biography, page ?? Charles Scribner’s Sons, New York, NY, USA, 2008. ISBN 0-684-16962-2, 0-684-31559-9 (e-book). LCCN Q141 .C66 2008.

Kiefer:1989:RAU

[KAR89] C. Kiefer, H. Albrecht, and K. H. Rieder. [Rezensionen:] Audretsch u. Mainzer: Philosophie und Physik der Raum–Zeit/ Heilbron: Max Planck Ein Leben für die Wissenschaft 1858–1947/ Cottam u. Tilley: Introduction to Surface and Superlattice Excitations. Physikalische Blätter, 45(9):387, 1989. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Kast:1964:JFW

[Kas64] W. Kast. James Franck/Wilhelm Maier/Max-Planck-Medaille 1964. Physikalische Blätter, 20(6):276–277, 1964. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Kayser:1902:HS

[Kay02] H. (Heinrich) Kayser. Handbuch der Spectroscopie. S. Hirzel, Leipzig, Germany, 1902. xi + 696 pp. LCCN ???? This was the first textbook to incorporate Planck’s Law of blackbody radiation.

Kemble:1938:EBM

[Kem38] E. C. Kemble. The eightieth birthday of Max Planck. Science, 87(2260):360–361, April 22, 1938. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/view/10.2307/1663419.

Kendall:1946:PRF

[Ken46] David G. Kendall. Planck’s radiation formula. Nature, 157(3996): 737, June 1, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v157/n3996/pdf/157737a0.pdf.

Kruger:1955:VMP

[KGSO55] H. Krüger, H. Gartmann, H. Schardin, and W. Oppelt. Verleihung der Max-Planck-Medaille an Enrico Fermi in Heidelberg/Der 5. Internationale Astronautische Kongreß in Innsbruck/2. Internationales Symposium über Kurzzeitphotographie und Hochfrequenzkinematographie in Paris/FA Regelungsmathematik über ,,Anwendung der Laplace-Transformation” in Essen. (German) [Awarding the Max Planck Medal to Enrico Fermi in Heidelberg/ the 5th International Astronautical Congress in Innsbruck/2nd International Symposium on short-time photography and highfrequency cinematography in Paris/FA mathematical control application on “Use of Laplace transform” in Essen]. Physikalische Blätter, 11(1):35–39, 1955. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Khofman:1987:MPT

[Kho87] D. Khofman. Max Planck — teacher, organizer and popularizer of science. In Studies in the history of physics and mechanics, 1987 (Russian), pages 159–176, 245. “Nauka”, Moscow, USSR, 1987.

Kirchhoff:1891:VBM

[Kir94] Gustav Kirchhoff. Vorlesungen über mathematische Physik. (German) [Lectures on mathematical physics]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1891–1894. 466 + 272 + 228 + 210 pp. Four volumes, of which 3 and 4 are edited by Max Planck.

Kirchhoff:1898:ABE

[Kir98a] Gustav Kirchhoff. Abhandlungen über Emission und Absorption: I. Über die Fraunhoferschen Linien (1859). II. Über den Zusammenhang zwischen Emission und Absorption von Licht und Wärme (1859). III. Über das Verhältnis zwischen dem Emmissionsvermögen und dem Absorptionsvermögen der Korper für Wärme und Licht (1860–1862). (German) [Treatises on emission and absorption. I. On the Fraunhofer lines (1859). II. On the relationship between emission and absorption of light and heat (1859). III. On the relationship between emission strength and the absorption body for heat and light (1860–1862)], volume 100 of Ostwalds Klassiker. W. Engelmann, Leipzig, Germany, 1898. 41 pp. Edited by Max Planck.

Kirchhoff:1898:AIM

[Kir98b] Gustav Kirchhoff. Abhandlungen über mechanische Wärmetheorie. I. Über einen Satz der mechanischen Wärmetheorie und einige Anwendungen desselben (1858). II. Bemerkung über die Spannung des Wasserdampfes bei Temperaturen, die dem Eispunkt nahe sind (1858). III. Über die Spannung des Dampfes von Mischungen von Wasser und Schwefelsäure (1858). (German) [reatises on the mechanical theory of heat. I. On mechanical heat theory and some of its applications (1858). II. Comments on the vapor tension of water at temperatures close to the freezing point (1858). III. Above the vapor tension of mixtures of water and sulfuric acid (1858)], volume 101 of Ostwalds Klassiker. W. Engelmann, Leipzig, Germany, 1898. ???? pp. Edited by Max Planck.

Kirsten:1979:PPA

[Kir79] Christa Kirsten, editor. Physiker über Physiker II: Antrittsreden, Erwiderungen bei der Aufnahme von Physikern in die Berliner Akademie, Gedächtnisreden. (German) [Physicists about Physicists II: inaugural addresses, responses during the recording of physicists in the Berlin Academy, memorial speeches], volume 2. Akademie-Verlag, Berlin, Germany, 1979. ISBN ???? 298 pp. LCCN ????

Kirsten:1975:PPW

[KK75] Christa Kirsten and H.-G. Körber, editors. Physiker über Physiker: Wahlvorschläge zur Aufnahme von Physikern in die Berliner Akademie 1870 bis 1929 von Hermann von Helmholtz bis Erwin Schrödinger. (German) [Physicists about Physicists: Nominations for receiving physicists in the Berlin Academy 1870– 1929], volume 1 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie-Verlag, Berlin, Germany, 1975. ISBN ???? 299 pp. LCCN ???? With an introduction by Hans-Jürgen Treder.

Klein:1996:CPA

[KKRS96] Martin J. Klein, A. J. Kox, Jürgen Renn, and Robert Schulman. The collected papers of Albert Einstein. Vol. 4. The Swiss years: Writings, 1912–1914. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02610-6. xii + 314 pp. URL

http://einsteinpapers.press.princeton.edu/vol4-doc;
http://einsteinpapers.press.princeton.edu/vol4-trans;

 http://press.princeton.edu/titles/5603.html;   http:
//press.princeton.edu/titles/5837.html;   http://press.
princeton.edu/TOCs/c5603.html;   http://press.princeton.

edu/TOCs/c5837.html. The Swiss years: writings, 1912–1914, English translation by Anna Beck in consultation with Don Howard, with a preface by Beck and Howard.

Klein:1961:MPB

[Kle61] Martin J. Klein. Max Planck and the beginnings of the quantum theory. Archive for History of Exact Sciences, 1 (5):459–479, October 1961. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0003-
9519&volume=1&issue=5&spage=459.

Klein:1963:PEQ

[Kle63] Martin J. Klein. Planck, entropy, and quanta, 1901–1906. The Natural Philosopher, 1:83–108, 1963. CODEN NPHLAC. ISSN 0547-9592.

Klein:1966:TQP

[Kle66] Martin J. Klein. Thermodynamics and quanta in Planck’s work. Physics Today, 19(11):23–32, November 1966. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://scitation.aip.org/content/aip/magazine/physicstoday/
article/19/11/10.1063/1.3047812. Reprinted in [WP85, 294–

302].

Klein:1985:TQP

[Kle85] Martin J. Klein. Thermodynamics and quanta in Planck’s work. In Weart and Phillips [WP85], pages 294–302. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Kle66].

Klose:1963:QBR

[Klo63] W. Klose. Quantentheorie: [book review:] E. Schrödinger, M. Planck, A. Einstein, and H. A. Lorentz, Briefe zur Wellenmechanik, Herausgegeben von K. Przibram. Wien: Springer-Verlag 1963. 67 S. ????, ??(??):454, ???? 1963.

Kentosh:2012:GPS

[KM12] J. Kentosh and M. Mohageg. Global positioning system test of the local position invariance of Planck’s constant. Physical Review Letters, 108(11):110801:1–110801:4, March 16, 2012. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://journals.aps.org/prl/abstract/10.
1103/PhysRevLett.108.110801. See comment [BF12] and news

story [Ano12].

Kockel:1959:MPF

[KMP59] B. (Bernhard) Kockel, Wilhelm Macke, and Achilles Papapetrou. Max-Planck-Festschrift 1958. Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1959. 412 pp. LCCN QC3 .M45.

Koch:1991:BPC

[Koc91] Martin Koch. From Boltzmann to Planck: On continuity in scientific revolutions. In Woodward and Cohen [WC91], pages 141– 150. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-

0346. LCCN ???? URL https://link.springer.com/chapter/
10.1007/978-94-011-3164-3_13.

Koenig:2001:MPS

[Koe01] Robert Koenig. Max Planck [Society] offers historic apology. Science, 292(5524):1979–1982, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/292/5524/1979.2.short. From the

introduction: “For half a century, survivors of cruel experiments at Nazi death camps have been seeking a formal apology as well as more details about the research abuses they endured from Germany’s scientific societies. On 7 June, a few of those victims finally got an explicit apology from the head of the country’s premier basic research organization, the Max Planck Society, on behalf of its forerunner, the Kaiser Wilhelm Society (KWG), some of whose scientists were implicated in the nefarious research.”.

Kohl:2002:PKW

[Koh02] Ulrike Kohl. Die Präsidenten der Kaiser Wilhelm-Gesellschaft im Nationalsozialismus: Max Planck, Carl Bosch und Albert Vögler zwischen Wissenschaft und Macht. (German) [The Presidents of the Kaiser Wilhelm Society under National Socialism: Max Planck, Carl Bosch and Albert Vogler between science and power], volume 5 of Pallas Athene. Steiner, Stuttgart, Germany, 2002. ISBN 3-515-08049-X (hardcover). 281 pp. LCCN Q49 .K64 2002.

Konagaya:2002:MPR

[Kon02] Daisuke Konagaya. The methodology of Planck’s radiation theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 12(1):43–58, 2002. CODEN HISCDU. ISSN 0285-4821.

Korber:1961:WBW

[Kör61] Hans Günther Körber, editor. Aus dem Wissenschaftlichen Briefwechsel Wilhelm Ostwalds. Teil I: Briefwechsel mit Ludwig Boltzmann, Max Planck, Georg Helm und Josiah Willard Gibbs. (German) [From the scientific correspondence of Wilhelm Ostwald. Part I: correspondence with Ludwig Boltzmann, Max Planck, Georg Helm and Josiah Willard Gibbs]. Akademie-Verlag, Berlin, Germany, 1961. xxiv + 143 + 19 pp.

Koyre:1952:BRB

[Koy52] Alexandre Koyré. Book review: Max Planck in seinen Akademie-Ansprachen. Erinnerungsschrift der Deutschen Akademie der Wissenschaften zu Berlin. Revue philosophique de la France et de l’Étranger, 142(??):276–279, ???? 1952. ISSN 0035-3833 (print),

2104-385X (electronic). URL http://www.jstor.org/stable/
view/10.2307/41087384.

Kracklauer:1976:CCD

[Kra76] Aloysius F. Kracklauer. Comment on classical derivations of Planck’s spectrum. Physical Review D (Particles and Fields), 14(1):654, July 15, 1976. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://journals.aps.org/prd/abstract/10.
1103/PhysRevD.14.654. See [TP76].

Krafft:1997:RDU

[Kra97] Fritz Krafft. Rezensionen: Dirk Ullmann: Quelleninventar Max Planck. (Veröffentlichungen aus dem Archiv zur Geschichte der Max-Planck-Gesellschaft, 8) Berlin 1996. 176 Seiten. ISBN 3-927579-08-4 (Auslieferung: Archiv zur Geschichte der Max-Planck-Gesellschaft, Boltzmannstraße 14, D-14195 Berlin-Dahlem). Berichte zur Wissenschaftsgeschichte, 20(4):296, ???? 1997. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kragh:2000:MPR

[Kra00] Helge Kragh. Max Planck: the reluctant revolutionary. Physics World, 13(12):31–36, December 2000. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/

13/12/phwv13i12a34.pdf. This paper describes Max Planck’s steps in 1900 to the notion of quantization, which Albert Einstein exploited so well in 1905.

Krafft:2002:RBP

[Kra02a] Fritz Krafft. Rezension: Max Planck: Vorträge, Reden, Erinnerungen von Hans Ross und Armin Hermann. Berichte zur Wissenschaftsgeschichte, 25(2):80, June 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Krafft:2002:RMP

[Kra02b] Fritz Krafft. Rezension: Max Planck: Vorträge, Reden, Erinnerungen von Hans Ross und Armin Hermann. Berichte zur Wissenschaftsgeschichte, 25(2):80, June 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Krafft:2003:RPH

[Kra03] Fritz Krafft. Rezension: Petra Hauke (unter Mitarbeit von Maren Dißmann): Literatur über Max Planck. Berichte zur Wissenschaftsgeschichte, 26(2):156, July 2003. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kragh:2012:PDE

[Kra12] Helge Kragh. Preludes to dark energy: zero-point energy and vacuum speculations. Archive for History of Exact Sciences, 66(3):199–240, May 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0003-
9519&volume=66&issue=3&spage=199.

Kretzschmar:1967:MPP

[Kre67] Hermann Kretzschmar. Max Planck als Philosoph. (German) [Max Planck as Philosopher]. E. Reinhardt, München, West Germany, 1967. 115 pp. LCCN B3323.P544 K7.

Krips:1986:APP

[Kri86] H. Krips. Atomism, Poincaré and Planck. Studies in History and Philosophy of Science Part A, 17(1):43–63, March 1986. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
003936818690018X.

Kropp:1952:PGM

[Kro52] Gerhard Kropp. Die philosophischen Gedanken Max Plancks. (German) [The philosophical thoughts of Max Planck]. Zeitschrift für philosophische Forschung, 6(3):434–458, ???? 1952. ISSN

0044-3301. URL http://www.jstor.org/stable/view/10.
2307/20480574.

Kirchner:1963:HMP

[KSB63] F. Kirchner, H. Schreiber, and E. Brüche, E.Brüche. Heiligers Max Planck-Plastik/Einigkeit und Recht und Stärke/Sehr geehrter Herr Schreiber!/Fast zu ernst. (German) [Heiliger’s Max Planck plastic/unity and law and strength/Dear Mr. Schreiber!/ almost too serious]. Physikalische Blätter, 19(9):432–435, 1963. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Kuhn:1978:BBT

[Kuh78] Thomas S. Kuhn. Black-body theory and the quantum discontinuity, 1894–1912. Clarendon Press, New York, NY, USA, 1978. ISBN 0-19-502383-8. xvi + 356 pp. LCCN QC484 .K83.

Kuhn:1984:RP

[Kuh84] Thomas S. Kuhn. Revisiting Planck. Historical Studies in the Physical Sciences, 14(2):231–252, ???? 1984. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757534.

Kuhn:1987:BBT

[Kuh87] Thomas S. Kuhn. Black-body theory and the quantum discontinuity, 1894–1912. University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-45800-8 (paperback). xviii + 378 pp. LCCN QC484 .K83 1987. US$17.95 (est.). URL http://www.loc.

gov/catdir/description/uchi051/86030833.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/86030833-t.html;
 http://www.loc.gov/catdir/enhancements/fy0609/86030833-

b.html. Reprint of [Kuh78] with a new afterword.

Kutschera:2014:DFN

[Kut14a] Ulrich Kutschera. Design-Fehler in der Natur: Alfred Russel Wallace und die Gott-lose Evolution. (German) [Design flaw in nature: Alfred Russel Wallace and the God-less Evolution], volume 12 of Science and religion; Naturwissenschaft und Glaube; v.

12. Lit, Berlin, Germany, 2014. ISBN 3-643-12133-4 (paperback). 378 pp. LCCN QH31.W2 K887 2014.

Kutschera:2014:MPJ

[Kut14b] Ulrich Kutschera. Max Planck und Jehova, der Intelligente Designer : ein Schlusswort. (German) [Max Planck and Jehovah, the intelligent designer: a closing remark]. In Design-Fehler in der Natur: Alfred Russel Wallace und die Gott-lose Evolution. (German) [Design flaw in nature: Alfred Russel Wallace and the God-less Evolution] [Kut14a], pages 51–?? ISBN 3-643-12133-4 (paperback). LCCN QH31.W2 K887 2014.

Kranzer:1973:RBD

[KZR^{+}73] W. Kranzer, H. Zwicker, H. Rechenberg, F. Lenz, and W. A. P. Luck. [Rezensionen:] Ballif, Dibble: Physics: Fundamentals & Frontiers/Arzimowitsch: Elementare Plasmaphysik/Smith: The Calculation of Atomic Collision Processes/Arzeliès: Relativistic Point Dynamics/Lipson: Optical Transforms/Greenwood: Computing Methods in Quantum Organic Chemistry/Max-Planck-Medaille 1973. Physikalische Blätter, 29(5):240–242, 1973. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Lankenau:1957:MPP

[Lan57] Ehrfried Lankenau. Max Planck und die Philosophie. (German) [Max Planck and philosophy]. Verlagsdruckerei L. Leopold, Bonn, West Germany, 1957. 15 pp. LCCN ????

Lauje:1958:PMP

[Lau58] M. Lauje. Pamjati Maksa Planka. Uspekhi Fizicheskikh Nauk, 64(4):639–640, April 1958. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1958/4/b/.

Lawrence:2015:MPS

[Law15] Ernest Lawrence. Max Planck: 69 Success Facts — Everything You Need to Know About Max Planck. Emereo Publishing, ????, 2015. ISBN 1-4888-7911-7. 44 pp. LCCN ???? US$.

Langevin:1912:GWH

[LdB12] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911/sous les auspices de m. E. Solvay. Gauthier-Villars, Paris, France, 1912. URL https:

//archive.org/details/lathoriedurayo00inst.

Lemmerich:2008:MPR

[Lem08] Jost Lemmerich. Max Planck: Ein reiner Theoretiker: Der Beginn von Max Plancks akademischer Laufbahn im Spiegel der Quellen. (German) [Max Planck: A pure theorist: The beginning of Max Planck’s academic career in the mirror of sources]. Physik Journal, 7(3):35–37, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-physik.

de/details/articlePdf/1104315/issue.html.

Lenzen:1949:BRBa

[Len49] V. F. Lenzen. Book review: Erinnerungen I by Max Planck. Isis, 40(1):66–67, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/227430.

Leveugle:2004:RPE

[Lev04] Jules Leveugle. La relativité, Poincaré et Einstein, Planck, Hilbert: Histoire véridique de la théorie de la relativité. (French) [Relativity, Poincaré and Einstein, Planck, Hilbert: the true history of the Theory of Relativity]. Harmattan, Paris, France, 2004. ISBN 2-7475-6862-8. xi + 330 pp. LCCN QC173.52 .L48 2004.

Levin:2017:TCS

[Lev17] Frank S. Levin. On theory and consensus in science: a journey in time. Web blog., October 17, 2017. URL https://blog.oup.

com/2017/10/theory-consensus-in-science/.

Lewis:1930:QKP

[Lew30] Gilbert N. Lewis. Quantum kinetics and the Planck equation. Physical Review, 35(12):1533–1537, June 1930. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://journals.aps.org/pr/abstract/10.
1103/PhysRev.35.1533.

Lightman:2005:DGBa

[Lig05a] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. A. A. Knopf Canada, Toronto, ON, Canada, 2005. ISBN 0-676-97789-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55 D57 L53 2005.

Lightman:2005:DGBb

[Lig05b] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55.D57 .L53. URL http://www.loc.gov/

catdir/enhancements/fy0623/2005040854-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0623/2005040854-d.
html;   http://www.loc.gov/catdir/enhancements/fy0623/
2005040854-t.html.

Leibfried:2015:BWE

[LMOS15] Stephan Leibfried, Christoph Markschies, Ernst Osterkamp, and Günter Stock. Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy]. De Gruyter Akademie Forschung, Berlin, Germany, 2015. ISBN 3-11-037598-2 (hardcover). 517 pp. LCCN AS182.B38 B477 2015.

Lohr:2012:MPG

[Löh12] Gebhard Löhr. Max Planck — ein Gegner des Christentums? Die Debatte um Plancks Haltung zur Religion nach dem Zweiten Weltkrieg. (German) [Max Planck — an opponent of Christianity? The debate on Planck’s attitude to religion after the Second World War]. Berichte zur Wissenschaftsgeschichte, 35(1):42–61, March 2012. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

London:1945:PCL

[Lon45] F. London. Planck’s constant and low temperature transfer. Reviews of Modern Physics, 17(2–3):310–320, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://journals.aps.org/rmp/
abstract/10.1103/RevModPhys.17.310;   http://rmp.aps.
org/abstract/RMP/v17/i2-3/p310_1.

Longair:1984:TCP

[Lon84] Malcolm S. Longair. Theoretical concepts in physics: an alternative view of theoretical reasoning in physics for final-year undergraduates. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-25550-3, 0-521-27553-9 (paperback). xiii + 366 pp. LCCN QC20 .L66 1984. URL http://www.loc.

gov/catdir/description/cam022/83018928.html;   http:/
/www.loc.gov/catdir/toc/cam021/83018928.html.

Longair:2003:PSB

[Lon03a] Malcolm S. Longair. 1895–1900: Planck and the spectrum of black-body radiation. In Theoretical concepts in physics: an alternative view of theoretical reasoning in physics [Lon03c], chapter 12, pages 303–328. ISBN 0-521-52878-X (paperback), 0-521-82126-6. LCCN QC20 .L64 2003. URL http://www.loc.gov/

catdir/description/cam031/2002073612.html;   http://www.
loc.gov/catdir/samples/cam041/2002073612.html;   http:
//www.loc.gov/catdir/toc/cam031/2002073612.html.

Longair:2003:PTB

[Lon03b] Malcolm S. Longair. Planck’s theory of black-body radiation. In Theoretical concepts in physics: an alternative view of theoretical reasoning in physics [Lon03c], chapter 13, pages 303–328. ISBN 0-521-52878-X (paperback), 0-521-82126-6.

LCCN QC20 .L64 2003. URL http://www.loc.gov/catdir/
description/cam031/2002073612.html;   http://www.loc.
gov/catdir/samples/cam041/2002073612.html;   http://www.
loc.gov/catdir/toc/cam031/2002073612.html.

Longair:2003:TCP

[Lon03c] Malcolm S. Longair. Theoretical concepts in physics: an alternative view of theoretical reasoning in physics. Cambridge University Press, Cambridge, UK, second edition, 2003. ISBN 0-521-52878-X (paperback), 0-521-82126-6. xvii + 569 pp.

LCCN QC20 .L64 2003. URL http://www.loc.gov/catdir/
description/cam031/2002073612.html;   http://www.loc.
gov/catdir/samples/cam041/2002073612.html;   http://www.
loc.gov/catdir/toc/cam031/2002073612.html.

Lowood:1977:MPB

[Low77] Henry Lowood. Max Planck: a bibliography of his non-technical writings, volume 1 of Berkeley papers in history of science. Office for History of Science and Technology, University of California, Berkeley, CA, USA, 1977. ISBN 0-918102-00-6. ISSN 1045-0379. v + 69 pp. LCCN Z8695.7 .C34 1977 QC16.P6.

Levin:1995:PPR

[LSW95] Sharon G. Levin, Paula E. Stephan, and Mary Beth Walker. Planck’s principle revisited: a note. Social Studies of Science, 25(2):275–283, May 1, 1995. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL https://journals.

sagepub.com/doi/pdf/10.1177/030631295025002003.

Lucy:1930:MOM

[Luc30] Allen Lucy. A method offered as a means of computing Planck’s constant without involving the charge on the electron. Physical Review, 36(2):367–368, July 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://journals.

aps.org/pr/abstract/10.1103/PhysRev.36.367.2.

Luscher:2000:MPM

[Lüs00] Martin Lüscher. Max-Planck-Medaille: Von den Pionen zu den fundamentalen Parametern der Quantenchromodynamik: Ein Fall für die Gittereichtheorie. (German) [Max Planck Medal: From pions to fundamental parameters of quantum chromodynamics: a case for lattice gauge theory]. Physikalische Blätter, 56 (7–8):65–68, 2000. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Macrakis:1993:SSS

[Mac93] Kristie Macrakis. Surviving the swastika: scientific research in Nazi Germany. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-507010-0. xii + 280 + 16 pp. LCCN Q180.G4 M3 1993. URL

http://www.h-net.org/review/hrev-a0a0m1-aa;   http://
www.loc.gov/catdir/enhancements/fy0604/93019919-d.html;
 http://www.loc.gov/catdir/enhancements/fy0724/93019919-
b.html.

Meyer-Arendt:1963:MPZ

[MAS63] J. R. Meyer-Arendt and M. Sachs. Max Planck und die Züchtungsforschung/Was ist in den USA anders geworden seit 1920?. (German) [Max Planck and plant breeding research/what’s changed in the USA since 1920?]. Physikalische Blätter, 19(5): 234–239, 1963. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Masters:2012:MPF

[Mas12] Barry R. Masters. Max Planck: Founder of quantum theory. Optics and Photonics News, 23(4):30–35, April 2012. CODEN OP-PHEL. ISSN 1047-6938 (print), 1541-3721 (electronic). URL

http://www.osa-opn.org/home/articles/volume_23/issue_
4/features/max_planck_founder_of_quantum_theory.

MPG:1958:MPP

[Max58] Max-Planck-Gesellschaft zur Förderung der Wissenschaften e.V., editor. Max Planck, physikalische Abhandlungen und Vorträge: Aus Anlass seines 100. Geburtstages, 23. April 1958. (German) [Max Planck, physical treatises and lectures: On the occasion of his 100th birthday, 23rd April 1958]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1958. xv + 776 (vol. 1), xi + 716 (vol. 2), xii + 426 (vol. 3) pp. LCCN ???? 3 volumes.

McKeown:1927:BRB

[McK27] A. McKeown. [Book Review:] I. Treatise on thermodynamics. By Dr. Max Planck. Third English edition, translated from the seventh German edition by Dr. A. Ogg. Pp. xiv + 297. London: Longmans, Green & Co., Ltd., 1927. Price 15s. net. Journal of the Society of Chemical Industry, London, 46(18):408, 1927. CODEN JSCIAN. ISSN 1934-9971.

Mollenstedt:1963:LMA

[MdB63] G. Möllenstedt and W. de Beauclair. Ludwig Mayer/Alwin Walther, 65 JahreMax-Planck-Medaille für R. E. Peierls/ Friedenspreis für von Weizsäcker. Physikalische Blätter, 19(5): 223–224, 1963. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Meissner:1947:MMP

[Mei47] Walther Meißner. In Memoriam: Max Planck zum Gedächtnis: (Anhang: Verzeichnis der Veröfftentlichunger Max Plancks). (German) [In memoriam: Max Planck remembrance (Appendix: List of publications of Max Planck)]. Zeitschrift für Naturforschung, 2(10):587–595, October 1947. CODEN ZNTFA2. ISSN

0372-9516. URL http://www.degruyter.com/view/j/zna.
1947.2.issue-10/zna-1947-1013/zna-1947-1013.xml?format=
INT;   http://www.znaturforsch.com/aa/v02a/2a0587.pdf.

Meissner:1951:MPM

[Mei51] Walter Meissner. Max Planck, the man and his work. Science, 113(2926):75–81, January 26, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/www.jstor.org/stable/view/10.2307/1678762;   http://

www.sciencemag.org/content/113/2926/75.full.pdf;   http:
//www.sciencemag.org/content/113/2926/75.short.

Meitner:1958:MPM

[Mei58] Lise Meitner. Max Planck als Mensch. (German) [Max Planck as a human being]. Die Naturwissenschaften, 45(17):406–408, September 1, 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Meitner:2001:MPM

[Mei01] Lise Meitner. Max Planck als Mensch. (German) [Max Planck as a human]. In Vorträge Reden Erinnerungen. (German) [Lectures Speeches Memories], pages 77–83. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Meitner:2003:FWH

[Mei03] Lise Meitner. Die Frau in der Wissenschaft. (German) [The Woman in science]. Supposé, Köln, Germany, 2003. ISBN 3-932513-46-0. LCCN ???? Audio CD (57 minutes). Produced by Klaus Sander and Jost Lemmerich.

Mendenhall:1917:DPR

[Men17] C. E. Mendenhall. A determination of the Planck radiation constant C_{2}. Physical Review, 10(5):515–524, November 1917. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://journals.aps.org/pr/abstract/
10.1103/PhysRev.10.515.

Mendelssohn:1973:WWN

[Men73] Kurt Mendelssohn. The World of Walther Nernst: the rise and fall of German science, 1864–1941. University of Pittsburgh Press, Pittsburgh, PA, 1973. ISBN 0-8229-1109-4. 191 + 13

pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/12148/
cb35327549d.

Merton:1938:SSO

[Mer38] Robert K. Merton. Science and the social order. Philosophy of Science, 5(3):321–337, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/184838.

Meschini:2007:PSP

[Mes07] Diego Meschini. Planck-scale physics: Facts and beliefs. Foundations of Science, 12(4):277–294, December 2007. CODEN FOS-CFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http:

//link.springer.com/article/10.1007/s10699-006-9102-3.

Meyer:2008:MPU

[Mey08] Hinrich Meyer. Max Planck und die universellen Konstanten. (German) [Max Planck and the universal constants]. Physik in unserer Zeit, 39(2):55, 2008. CODEN PHUZAH. ISSN 1521-3943.

Markl:1997:MPG

[MH97] G. H. Markl and E. Henning. Max Planck (1858–1947): Zum Gedenken au seinen 50. Todestagam 4. Oktober 1997. (German) [Max Planck (1858–1947): Commemorating the 50th anniversary of his death on 4 October 1997]. Archiv der Max-Planck-Gesellschaft Exhibit at the Magnus-Haus, Berlin, 1997. URL

http://www.max-planck.mpg.de/.

Michels:1968:BRB

[Mic68] Walter C. Michels. Book review: Letters on Wave Mechanics; Albert Einstein, Erwin Schroedinger, Max Planck, H. A. Lorentz, translated by Martin J. Klein from a collection edited by K. Przibram. American Journal of Physics, 36(8):766–767, August 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://link.aip.org/link/?AJP/36/766/2;
http://scitation.aip.org/content/aapt/journal/ajp/36/
8/10.1119/1.1975134.

Millikan:1916:DPD

[Mil16] R. A. Millikan. A direct photoelectric determination of Planck’s “h”. Physical Review, 7(3):355, March 1916. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

journals.aps.org/pr/abstract/10.1103/PhysRev.7.355.

Mises:1933:RPD

[Mis33] Mises. [Rezensionen:] Prof. Dr. Max Planck, Wege zur physikalischen Erkenntnis. Reden und Vorträge. Leipzig 1933, S. Hirzel Verlag. IX + 280 S. Preis geb. 8 M. Zeitschrift für Angewandte Mathematik und Mechanik, 13(6):453, 1933. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Methfessel:1991:RMC

[MKD^{+}91] S. Methfessel, H. Klar, D. Dubbers, W. Sandhas, K. Kromphardt, H. F. Ebel, E. Filos, H. Rechenberg, Th. Görnitz, and K. Stierstadt. [Rezensionen:] Mott: Can Scientists Believe?/Friedrich:

Theoretische Atomphysik/Ignatovich: The Physics of Ultracold Neutrons/Belyaev: Lectures on the Theory of Few-Body Systems/ Elster: Humanökologie als Aufgabe für Natur- u. Geisteswissenschaften/Vischer: Plane deinen Ruhm!/Fritzsch, Sucher: Von und über Physiker/Kay: Managing Creativity in Science and Hi-Tech/Scribal: Max Planck: Wissenschaftliche Selbstbiographie/ Laurikainen: Beyond the Atom: The Philosophical Thought of W. Pauli/Falk: Physik Zahl und Realität. Physikalische Blätter, 47(7):650–653, 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Moore:1923:RMP

[Moo23] C. L. E. Moore. Review: Max Planck, Einführung in die Mechanik Deformierbarer Körper. Bulletin of the American Mathematical Society, 29(10):478, 1923. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183485751.

Mosch:1942:RPD

[Mos42] E. Mosch. [Rezensionen:] Prof. Dr. Hugo Dingler, Max Planck und die Begründung der sogenannten modernen theoretischen Physik. 32 S. Berlin 1939, Ahnenerbe-Stiftung-Verlag. Preis 0,80 M. Zeitschrift für Angewandte Mathematik und Mechanik, 22(1): 60, 1942. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Muller:1919:AGM

[MP19] G. Müller and Max Planck. Antrittsrede von G. Müller und Erwiderung von M. Planck. (German) [Inaugural speech by G. Müller and reply of M. Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 1(??):4–5, ???? 1919. CODEN SPWPAI. ISSN 0371-2435.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82d, MR82e, MR87, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The formulation of matrix mechanics and its modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44 vol. 3. URL http://www.loc.gov/

catdir/enhancements/fy1005/82003253-d.html;   http://
www.loc.gov/catdir/enhancements/fy1005/82003253-t.html.

See other volumes in this series [MR82a, MR82c, MR82d, MR82e, MR87, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4. See other volumes in this series [MR82a, MR82b, MR82d, MR82e, MR87, MR00].

Mehra:1982:QTPa

[MR82d] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR87, MR00].

Mehra:1982:QTPb

[MR82e] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982.

ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375– 878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR87, MR00].

Mehra:1987:ESR

[MR87] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol. 5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR00].

Morrow:1998:PCD

[MR98] G. O. Morrow and L. E. Reichl. Planck’s-constant dependence of the scaling of localization length in quantum dynamics. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 57(5):5266–5270, May 1998. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://journals.aps.org/pre/abstract/10.
1103/PhysRevE.57.5266.

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The completion of quantum mechanics, 1926–1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp. LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82d, MR82e, MR87].

Menge:1963:WMP

[MSQM63] R. Menge, W. Stein, H. Quenzel, and E. Möller. Wurde Max Planck ,,verheiligert” ? Zu Phys. Bl. 19, 234 (1963)/Ist Biophysik eine selbständige Wissenschaft?/Einigkeit und Recht und Stärke/Vom Ursprung der Dinge. (German) [Was Max Planck “verheiligert”? on Phys. Bl. 19, 234 (1963)/is biophysics an independent science?/Unity and justice and strength/The origin of things]. Physikalische Blätter, 19(6):284–287, 1963. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Mulligan:1994:MPB

[Mul94] Joseph F. Mulligan. Max Planck and the “black year” of German physics. American Journal of Physics, 62(12):1089–1097, December 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/62/12/10.1119/1.17666.

Mulligan:2000:BRB

[Mul00] Joseph F. Mulligan. Book review: Brieftagebuch zwischen Max Planck, Carl Runge, Bernhard Karsten, und Adolf Leopold by Max Planck; Carl Runge; Bernhard Karsten; Adolf Leopold; Klaus Hentschel; Renate Tobies. Isis, 91(2):371–372, June 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236983.

Muller:2008:MPL

[Mül08a] Ingo Müller. Max Planck: a life for thermodynamics. Annalen der Physik (Berlin), 17 [i.e. 520](2–3):73–87, 2008. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Muller:2008:MPLb

[Mül08b] Ingo Müller. Max Planck: Ein Leben für die Thermodynamik Vom Zweiten Hauptsatz der Thermodynamik zum Planckschen Wirkungsquantum. (German) [Max Planck: A life for thermodynamics: From the Second Law of Thermodynamics to Planck’s constant]. Physik Journal, 7(3):39–45, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1104319/issue.html.

Mundlak:1936:CPR

[Mun36] Max Mundlak. The consequences of philosophy: a reply to Planck and Einstein. J. Bale, Sons and Danielsson, Ltd., London, UK, 1936. 86 pp. LCCN Q175 .M98.

Nagel:1982:DCN

[Nag82] Bengt Nagel. The discussion concerning the Nobel Prize for Max Planck. In Bernhard et al. [BCS82], pages 352–376. ISBN 0-08-027939-2. LCCN Q124.6 .N63 1981. US$49.50 (est.).

URL http://www.sciencedirect.com/science/article/pii/
B9780080279398500281.

Nauenberg:2016:MPB

[Nau16] Michael Nauenberg. Max Planck and the birth of the quantum hypothesis. American Journal of Physics, 84(9):709–720, September 2016. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Needell:1980:IFC

[Nee80] Allan A. Needell. Irreversibility and the Failure of Classical Dynamics: Max Planck’s Work on the Quantum Theory, 1900–1915. Ph.D. dissertation, Yale University, New Haven, CT, USA, 1980.

viii + 285 pp. URL http://search.proquest.com/pqdtglobal/
docview/303093622/A26623FFE32F423APQ.

Nadarajah:2006:GPD

[NK06] Saralees Nadarajah and Samuel Kotz. A generalized Planck distribution. TEST, 15(2):361–374, September 2006. CODEN ???? ISSN 1133-0686 (print), 1863-8260 (electronic). URL http://

link.springer.com/article/10.1007/BF02607057.

Noerdlinger:1973:PRE

[Noe73] Peter D. Noerdlinger. Primordial 2.7^{∘} radiation as evidence against secular variation of Planck’s constant. Physical Review Letters, 30(16):761–762, April 16, 1973. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145.  URL http://journals.aps.org/prl/abstract/10.
1103/PhysRevLett.30.761.

Nye:2006:SBH

[Nye06] Mary Jo Nye. Scientific biography: History of science by another means? Isis, 97(2):322–329, June 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/504738.

OFlaherty:1956:MPA

[O’F56] James C. O’Flaherty. Max Planck and Adolf Hitler. AAUP Bulletin, 42(3):437–444, Autumn 1956. ISSN 0277-3945 (print), 1930-

322X (electronic). URL http://www.jstor.org/stable/view/
10.2307/40222051.

OFlaherty:1959:HLM

[O’F59] James C. O’Flaherty. A humanist looks at Max Planck. American Scientist, 47(1):68–79, March 1959. CODEN AMSCAC.

ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://

www.jstor.org/stable/view/10.2307/27827247.

Olesko:1981:BRB

[Ole81] Kathryn M. Olesko. Book review: Max Planck. A Bibliography of His Non-Technical Writings by Henry Lowood; William Henry Bragg and William Lawrence Bragg. A Bibliography of Their Non-Technical Writings by Henry Lowood; Ernest Rutherford. A Bibliography of His Non-Technical Writings by Henry Lowood. Isis, 72(2):289–290, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230982.

Oittner-Torkar:1997:DMP

[OT97] G. Oittner-Torkar. Die Dissertation von Max Planck ,Von meinen Universitätslehrern hatte keiner ein Verständnis für ihren Inhalt. (German) [The thesis of Max Planck: none of my university teachers had an understanding of its content]. MPG-Spiegel, 4(??): 40–47, ???? 1997.

Padmanabhan:1983:UBP

[Pad83] T. Padmanabhan. Universe before Planck time: A quantum gravity model. Physical Review D (Particles and Fields), 28(4):756– 760, August 1983. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

journals.aps.org/prd/abstract/10.1103/PhysRevD.28.756.

Partovi:1994:QCP

[Par94] M. Hossein Partovi. QED corrections to Planck’s radiation law and photon thermodynamics. Physical Review D (Particles and Fields), 50(2):1118–1124, July 1994. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500

(CD-ROM). URL http://journals.aps.org/prd/abstract/
10.1103/PhysRevD.50.1118.

Paradowski:2012:BRB

[Par12] Robert J. Paradowski. Book review: One Hundred Years at the Intersection of Chemistry and Physics: The Fritz Haber Institute of the Max Planck Society, 1911–2011. Isis, 103(4):770– 771, December 2012. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/670078.

Planck:1990:MPW

[PB90] Max Planck and Wieland Berg. Max Planck: wissenschaftliche Selbstbiographie, volume 19 of Acta historica Leopoldina. Deutsche Akademie der Naturforscher Leopoldina, Halle/Salle, Germany, 1990. ISSN 0001-5857. 42 pp. LCCN QC16.P6 A3 1990.

Pelogia:2017:AJI

[PB17] Karla Pelogia and Carlos Alexandre Brasil. Analysis of the Jun Ishiwara’s “The universal meaning of the quantum of action”. European Physical Journal H, 42(4–5):507–521, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL https://link.springer.com/article/10.1140/
epjh/e2017-80034-x. See [Ish17].

Planck:1993:IPF

[PD93] Max Planck and J. (Joachim) Du Plessis de Grenédan. Initiations á la physique. (French) [Introduction to physics], volume 204 of Champs. Flammarion, Paris, France, nouvelle éd., revue et augmentée. edition, 1993. ISBN 2-08-081204-1. 287

pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/12148/

cb35713473h. French translation by Joachim Du Plessis de Grenédan of Wege zur Physikalischen Erkenntnis.

Pedersen:1975:RBO

[Ped75] K. M. Pedersen. Review: Planck’s Original Papers in Quantum Physics — Classic Papers in Physics — by H. Kangro. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 19(1):81–82, ???? 1975. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Pegg:1977:FVP

[Peg77] D. T. Pegg. Future variation of Planck’s constant. Nature, 267 (5610):408–409, June 2, 1977. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Petruccioli:2014:CPV

[Pet14] Sandro Petruccioli. Correspondence principle versus Planck-type theory of the atom. Archive for History of Exact Sciences, 68(5): 599–639, September 2014. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/article/10.1007/s00407-014-0137-5.

Planck:2007:ASS

[PHR07] Max Planck, Dieter Hoffmann, and Fritz Reiche. Die Ableitung der Strahlungsgesetze: 1768–1783: [Sieben Abhandlungen aus dem Gebiete der elektromagnetischen Strahlungstheorie]. (German) [The derivation of the radiation law: 1768–1783: [Seven treatises on the topic of electromagnetic radiation theory]]. Verlag Harri Deutsch, Frankfurt am Main, Germany, fourth edition, 2007. ISBN 3-8171-3419-3. 120 pp. LCCN ????

Picolet:1925:SPC

[Pic25] Lucien E. Picolet. A survey of physics: a collection of lectures and essays. by Max Planck, translated by R. Jones and D. H. Williams. viii–184 pages, 7 case 12 × 5 inches, cloth. New York, E. P. Dutton and Company, no date. Price, $2.40. Journal of The Franklin Institute, 200(5):701–702, 1925. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0016003225907557.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN

Q175.32.R45 P53 2008. URL http://catdir.loc.gov/catdir/
toc/ecip087/2007051167.html;   http://www.gbv.de/dms/
goettingen/555633241.pdf;   http://www.oup.com/us/catalog/
general/subject/Physics/?view=usa&ci=9780195336115.

Planck:1879:BZH

[Pla79] Max Planck. Über den zweiten Hauptsatz der mechanischen Wärmetheorie. (German) [On the second law of thermodynamics]. Doctoral dissertation, Universität München, Munich, Germany, February 1879. 61 pp. URL http://edoc.hu-berlin.de/

dissertationen/historisch/planck-max/HTML/;   http://
genealogy.math.ndsu.nodak.edu/id.php?id=20750.

Planck:1880:GIK

[Pla80] Max Planck. Gleichgewichtszustände isotroper Körper in verschiedenen Temperaturen. (German) [Equilibrium states of isotropic bodies at different temperatures]. Habilitation thesis, Universität München, Munich, Germany, June 1880. 63 pp.

URL http://hdl.handle.net/2027/mdp.39015086650622;
http://hdl.handle.net/2027/uc1.b4252054.

Planck:1881:TSG

[Pla81] Max Planck. Die Theorie des Sättigungsgesetzes. (German) [The theory of the law of saturation]. Annalen der Physik (Berlin), 249(7):535–543, ???? 1881. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.18812490710/abstract.

Planck:1882:VSS

[Pla82] Max Planck. Verdampfen, Schmelzen und Sublimiren. (German) [Evaporation, melting and subliming]. Annalen der Physik (Berlin), 251(3):446–475, 1882. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1883:TGG

[Pla83] Max Planck. Ueber das thermodynamische Gleichgewicht von Gasgemengen. (German) [On the thermodynamic equilibrium of gas mixtures]. Annalen der Physik (Berlin), 255(6):358–378, ???? 1883. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1884:TFG

[Pla84] Max Planck. Zur Theorie der Flüssigkeitsstrahlen. (German) [On the theory of fluid jets]. Annalen der Physik (Berlin), 257(3): 499–509, ???? 1884. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1887:PEE

[Pla87a] Max Planck. Das Prinzip der Erhaltung der Energie. (German) [The principle of conservation of energy]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1887. ???? pp.

Planck:1887:MKV

[Pla87b] Max Planck. Über die molekulare Konstitution verdünnter Lösungen. (German) [On the molecular constitution of dilute solutions]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 1(??):577–582, ???? 1887. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1887:PVEa

[Pla87c] Max Planck. Ueber das Princip der Vermehrung der Entropie. 1. Abhandlung. (German) [On the principle of increase of entropy. First treatment]. Annalen der Physik (Berlin), 266(4):562–582, ???? 1887. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1887:PVEb

[Pla87d] Max Planck. Ueber das Princip der Vermehrung der Entropie. 2. Abhandlung. (German) [On the principle of increase of entropy]. Annalen der Physik (Berlin), 267(6):189–203, ???? 1887. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1887:PVEc

[Pla87e] Max Planck. Ueber das Princip der Vermehrung der Entropie. 3. Abhandlung. (German) [On the principle of increase of entropy: third treatment]. Annalen der Physik (Berlin), 268(11):462–503, ???? 1887. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1887:MVL

[Pla87f] Max Planck. Ueber die Molecularconstitution verdünnter Lösungen. (German) [On the molecular constitution of dilute solutions]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 1(11–12):577–582, ???? 1887. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1888:CGV

[Pla88a] Max Planck. Das chemische Gleichgewicht in verdünnten Lösungen. (German) [Chemical equilibrium in dilute solutions]. Annalen der Physik (Berlin), 270(5):139–154, ???? 1888. CO-DEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1888:RCM

[Pla88b] Max Planck. [Rezensionen]: Ch. M. van Deventen, Einfache Herleitung einiger für die Chemie wichtiger thermodynamischer Beziehungen. Beiblätter zu den Annalen der Physik, XII(11):763– 764, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:RJB

[Pla88c] Max Planck. [Rezensionen]: J. Buchanan: Über ein Gesetz der Vertheilung der Moleculargeschwindigkeiten unter den Molecülen einer Flüssigkeit. Beiblätter zu den Annalen der Physik, XII(12): 846, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:RJP

[Pla88d] Max Planck. [Rezensionen]: J. Parker. Über eine Erweiterung des Carnof sehen Satzes. Beiblätter zu den Annalen der Physik, XII (11):760, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:RMBa

[Pla88e] Max Planck. [Rezensionen]: M. Brillouin: Die specifische Wärme für eine beliebige Zustandsänderung, vom Standpunkt der Thermodynamik. Beiblätter zu den Annalen der Physik, XII(11):761, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:RMBb

[Pla88f] Max Planck. [Rezensionen]: M. Brillouin: Permanente Deformationen vom Standpunkt der Thermodynamik. Beiblätter zu den Annalen der Physik, XII(11):761–763, ???? 1888. CODEN BEA-PAI. ISSN 0375-846X.

Planck:1888:RMPa

[Pla88g] Max Planck. [Rezensionen]: M. Planck: Das Princip der Erhaltung der Energie. Beiblätter zu den Annalen der Physik, XII(2): 134, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:RMPb

[Pla88h] Max Planck. [Rezensionen]: M. Planck: Ueber die Dampfspannung von verdünnten Lösungen flüchtiger Stoffe. Beiblätter zu den Annalen der Physik, XII(11):776, ???? 1888. CODEN BEA-PAI. ISSN 0375-846X.

Planck:1888:DVLa

[Pla88i] Max Planck. Über die Dampfspannung von verdünnten Lösungen flüchtiger Stoffe. (German) [On the vapor pressure of dilute solutions of volatile matter]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 2(??):405–414, ???? 1888. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1888:HDS

[Pla88j] Max Planck. Über die hypothese der dissoziation der salze in sehr verdünnten lösungen. (German) [On the hypothesis of dissociation of salts in very dilute solutions]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 2(??): 343, ???? 1888. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1888:DVLb

[Pla88k] Max Planck. Ueber die Dampfspannung von verdünnten Lösungen flüchtiger Stoffe. (German) [On the vapor pressure of dilute solutions of volatile materials]. Beiblätter zu den Annalen der Physik, XII(11):776, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1888:MVL

[Pla88l] Max Planck. Ueber die Molecularconstitution verdünnter Lösungen. (German) [On the molecular constitution of dilute solutions]. Beiblätter zu den Annalen der Physik, XII(3):142–143, ???? 1888. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:CAA

[Pla89a] Max Planck. Ch. Antoine: Ausdehnung und Compression der atmosphärischen Luft. (German) [Ch. Antoine: Expansion and compression of atmospheric air]. Beiblätter zu den Annalen der Physik, XIII(7):473, ???? 1889. CODEN BEAPAI.

ISSN 0375-846X. URL http://echo.mpiwg-berlin.mpg.de/
ECHOdocuView?url=/permanent/library/DN7SCHS4/pageimg&
start=71&mode=imagepath&pn=79.

Planck:1889:CAV

[Pla89b] Max Planck. Ch. Antoine: Volumina gesättigter Dämpfe. (German) [Ch. Antoine: Volumes of saturated vapors]. Beiblätter zu den Annalen der Physik, XIII(7):483, ???? 1889. CODEN BEA-

PAI. ISSN 0375-846X. URL http://echo.mpiwg-berlin.mpg.
de/ECHOdocuView?url=/permanent/library/DN7SCHS4/pageimg&
start=71&mode=imagepath&pn=80.

Planck:1889:ROB

[Pla89c] Max Planck. [Rezensionen]: A. von Oettingen: Die thermodynamischen Beziehungen antithetisch entwickelt. Beiblätter zu den Annalen der Physik, XIII(7):466–469, ???? 1889. CODEN BEA-PAI. ISSN 0375-846X.

Planck:1889:RAV

[Pla89d] Max Planck. [Rezensionen]: Aroldo Violi: Die Isotherme der Gase. Beiblätter zu den Annalen der Physik, XIII(2):66–67, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RCP

[Pla89e] Max Planck. [Rezensionen]: C. Puschl: Über das Verhalten der Gase zum Mariotte’schen Gesetze bei sehr hohen Temperaturen. Beiblätter zu den Annalen der Physik, XIII(5):289–290, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RCAb

[Pla89f] Max Planck. [Rezensionen]: Ch. Antoine: Ausdehnung und Compression der atmosphärischen Luft. Beiblätter zu den Annalen der Physik, XIII(5):292–293, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RCAc

[Pla89g] Max Planck. [Rezensionen]: Ch. Antoine: Temperaturänderung eines Gases oder eines Dampfes, der bei constantem Wärmeinhalt comprimirt oder ausgedehnt wird und Über die Temperaturänderungen der Gase und Dämpfe, die eine constante Wärmemenge enthalten, unter verschiedenen Drucken. Beiblätter zu den Annalen der Physik, XIII(5):292–293, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RCAa

[Pla89h] Max Planck. [Rezensionen]: Ch. Antoine: Volumina gesättigter Dämpfe. Beiblätter zu den Annalen der Physik, XIII(5):292–293, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RDB

[Pla89i] Max Planck. [Rezensionen]: Duhem: Über die Verflüssigung der Kohlensäure bei Gegenwart von Luft. Beiblätter zu den Annalen der Physik, XIII(3):156, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RFW

[Pla89j] Max Planck. [Rezensionen]: F. Wald: Über den zweiten Hauptsatz der mechanischen Wärmetheorie. Zweite Abhandlung. Beiblätter zu den Annalen der Physik, XIII(3):139–140, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RGDa

[Pla89k] Max Planck. [Rezensionen]: G. D. Liveing: Über Lösung und Krystallisation. Beiblätter zu den Annalen der Physik, XIII(5): 281–284, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RGDb

[Pla89l] Max Planck. [Rezensionen]: G. D. Liveing: Ueber Lösung und Krystallisation. Beiblätter zu den Annalen der Physik, XIII(7): 453–455, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RGB

[Pla89m] Max Planck. [Rezensionen]: Gouy: Ueber ein allgemeines Gesetz bezüglich der Effecte reversibler Umwandlungen. Beiblätter zu den Annalen der Physik, XIII(7):463–465, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RJF

[Pla89n] Max Planck. [Rezensionen]: J. Farkas: Über die Allgemeinheit des zweiten Hauptsatzes der mechanischen Wärmetheorie. Beiblätter zu den Annalen der Physik, XIII(9):796, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RNB

[Pla89o] Max Planck. [Rezensionen]: L. Natanson: Über die Erklärung eines Experiments von Joule nach der kinetischen Theorie der Gase; G. Hirn: Bemerkungen dazu. Beiblätter zu den Annalen der Physik, XIII(3):138–139, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RCB

[Pla89p] Max Planck. [Rezensionen]: Le Chatelier: Über die Gesetze des chemischen Gleichgewichts; P. Duhem: Dasselbe; Le Chatelier: Dasselbe; P. Duhem: Dasselbe; Le Chatelier: Dasselbe. Beiblätter zu den Annalen der Physik, XIII(3):143, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RMB

[Pla89q] Max Planck. [Rezensionen]: M. Brillouin: Permanente Deformationen vom Standpunkt der Thermodynamik. Beiblätter zu den Annalen der Physik, XIII(6):352–353, ???? 1889. CODEN BEA-PAI. ISSN 0375-846X.

Planck:1889:RPDa

[Pla89r] Max Planck. [Rezensionen]: P. Duhem: Einige Bemerkungen über die Lösungs- und Verdünnungswärme. Beiblätter zu den Annalen der Physik, XIII(3):128–129, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RPDb

[Pla89s] Max Planck. [Rezensionen]: P. Duhem: Ueber den Einfluss der Schwere auf die Lösungen. Beiblätter zu den Annalen der Physik, XIII(3):129–131, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RPB

[Pla89t] Max Planck. [Rezensionen]: Pellat: Anwendung des Carnofsehen Princips auf endothermische Reactionen. Beiblätter zu den Annalen der Physik, XIII(3):141–142, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RVWc

[Pla89u] Max Planck. [Rezensionen]: Van der Waals: Moleculartheorie für eine Mischung von zwei Stoffen. Beiblätter zu den Annalen der Physik, XIII(7):465–466, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RVWb

[Pla89v] Max Planck. [Rezensionen]: Van der Waals: Über das thermodynamische Gleichgewicht. Beiblätter zu den Annalen der Physik, XIII(3):140–141, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RVWa

[Pla89w] Max Planck. [Rezensionen]: Van der Waals: Über die Dichte in der Übergangsschicht vom Dampf zur Flüssigkeit. Beiblätter zu den Annalen der Physik, XIII(3):134–135, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:RWN

[Pla89x] Max Planck. [Rezensionen]: W. Nernst: Zur Kinetik der in Lösung befindlichen Körper. Beiblätter zu den Annalen der Physik, XIII(3):131–133, ???? 1889. CODEN BEAPAI. ISSN 0375-846X.

Planck:1889:TTM

[Pla89y] Max Planck. Zur Theorie der Thermoelectricität in metallischen Leitern. (German) [On the theory of thermoelectriciy in metallic conductors]. Annalen der Physik (Berlin), 272(3):624–643, 936, 1889. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1890:ODG

[Pla90a] Max Planck. Über den osmotischen Druck. (German) [On osmotic pressure]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 6(??):187–189, ???? 1890. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1890:PZZa

[Pla90b]  Max Planck.  Über die Potentialdifferenz zwischen zwei

verdünnten Lösungen binärer Elektrolyte. (German) [On the potential difference between two dilute solutions of binary electrolytes]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 9(7):45–49, April 18, 1890. CODEN VDPEAZ. ISSN 0372-5448.

Planck:1890:EEW

[Pla90c] Max Planck. Ueber die Erregung von Electricität und Wärme in Electrolyten. (German) [On the excitement of electricity and heat in electrolytes]. Annalen der Physik (Berlin), 275(2):161– 186, 1890. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1890:EEE

[Pla90d] Max Planck. Ueber die Erregung von Elektrizität und Wärme in Elektrolyten. (German) [On the excitation of electricity and heat in electrolytes]. Annalen der Physik (Berlin), 275(2):161–181, ???? 1890. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.18902750202/abstract.

Planck:1890:PZZb

[Pla90e] Max Planck. Ueber die Potentialdifferenz zwischen zwei verdünnten Lösungen binärer Electrolyte. (German) [On the potential difference between two dilute solutions of binary electrolytes]. Annalen der Physik (Berlin), 276(8):561–576, ???? 1890. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1891:ANE

[Pla91a]  Max Planck.  Allgemeines zur neueren Entwicklung der

Wärmetheorie. (German) [General information on the recent development of the theory of heat]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 8(??):647– 656, ???? 1891. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1891:PVE

[Pla91b] Max Planck. Ueber das Princip der Vermehrung der Entropie. 4. Abhandlung. (German) [On the principle of increase of entropy. 4th paper]. Annalen der Physik (Berlin), 280(11):385–428, ???? 1891. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1892:BCC

[Pla92a] Max Planck. Bemerkungen über das Carnot–Clausius’sche Princip. (German) [Remarks on the Carnot–Clausius Principle]. Annalen der Physik (Berlin), 282(5):162–166, ???? 1892. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1892:EHA

[Pla92b] Max Planck. Erwiderung auf einen von Herrn Arrhenius erhobenen Einwand. (German) [Response to an objection raised by Mr Arrhenius]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 9(??):636–637, ???? 1892. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1892:NTD

[Pla92c] Max Planck. Notiz zur Theorie der Diffusion und Elektrolyse. (German) [Note on the theory of diffusion and electrolysis]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 9(??):347–348, ???? 1892. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1893:KZH

[Pla93a] Max Planck. Der Kern des zweiten Hauptsatzes der Wärmetheorie. (German) [The core of the Second Law of Thermodynamics]. Z. physik. chem. Unterricht, 6(??):217–221, ???? 1893.

Planck:1893:NHN

[Pla93b] Max Planck. Ein neues Harmonium in natürlicher Stimmung nach dem System C. Eitz. (German) [On new harmony in natural mood in the system of C. Eitz]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 12(??):8–9, ???? 1893. CODEN VD-PEAZ. ISSN 0372-5448.

Planck:1893:GAT

[Pla93c] Max Planck. Grundriß der allgemeinen Thermochemie. (German) [Outline of general thermochemistry]. Trewendt, Breslau, Germany, 1893. ???? pp.

Planck:1893:RCE

[Pla93d] Max Planck. [Rezensionen;] C. Eitz: Das mathematisch-reine Tonsystem. Leipzig 1891, Breitkopf u. Härtel. 36 S. und eine lithographische Tafel. Zeitschrift für Psychologie und Physiologie der Sinnesorgane, 5(??):112–114, ???? 1893. ISSN 0233-

2302. URL http://vlp.mpiwg-berlin.mpg.de/references?
id=lit15119.

Planck:1894:HRH

[Pla94a] Max Planck. Heinrich Rudolf Hertz. Rede zu seinem Gedächtnis in der Sitzung der Physik. Ges. zu Berlin am 16. Februar 1893. Leipzig 1894. (German) [Heinrich Rudolf Hertz. Speech in his memory at the meeting of the Physical Society of Berlin on 16 February 1893. Leipzig 1894]. Naturwissenschaftliche Rundschau, 9(11):140–143, March 17, 1894. CODEN NARSAC. ISSN

0028-1050. URL http://babel.hathitrust.org/cgi/pt?id=
wu.89119127140.

Planck:1894:RCH

[Pla94b] Max Planck. [Rezensionen]: C. Henry: Une transformation de l’orchestre. Paris 1892, A. Hermann. Zeitschrift für Psychologie und Physiologie der Sinnesorgane, 6(??):67–68, 1894. ISSN 0233-

2302. URL http://vlp.mpiwg-berlin.mpg.de/references?
id=lit15119.

Planck:1894:BMG

[Pla94c] Max Planck. Über den Beweis des Maxwellschen Geschwindigkeitsverteilungsgesetzes unter Gasmolekülen. (German) [On the evidence for Maxwell’s law of velocity distribution of gas molecules]. Sitzungsberichte der Mathematisch-Naturwissenschaftlichen Abteilung der Bayerischen Akademie der Wissenschaften zu München, 24(??): 391–394, ???? 1894. CODEN SMNBAC. ISSN 0376-1037.

Planck:1895:AEE

[Pla95a] Max Planck. Absorption und Emission elektrischer Wellen durch Resonanz. (German) [Absorption and emission of electric waves by resonance]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 16(??):289–301, March 21, 1895. CODEN SPWPAI. ISSN 0371-2435.

Planck:1895:BMG

[Pla95b] Max Planck. Ueber den Beweis des Maxwell’schen Geschwindigkeitsvertheilungsgesetzes unter Gasmolecülen. (German) [On the evidence for Maxwell’s law of velocity distribution of gas molecules]. Annalen der Physik (Berlin), 291(5):220–222, 1895. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1896:AEE

[Pla96a] Max Planck. Absorption und Emission electrischer Wellen durch Resonanz. (German) [Absorption and emission of electrical waves through resonance]. Annalen der Physik (Berlin), 293(1):1–14, ???? 1896. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.18962930102/abstract.

Planck:1896:GNE

[Pla96b] Max Planck. Gegen die neuere Energetik. (German) [Against recent energetics]. Annalen der Physik (Berlin), 293(1):72–78, ???? 1896. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1896:ESW

[Pla96c] Max Planck. Über elektrische Schwingungen, welche durch Resonanz erregt und durch Strahlung gedämpft werden.. (German) [On electrical oscillations, which through resonance are excited and damped by radiation]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 10(??): 151–170, February 20, 1896. CODEN SPWPAI. ISSN 0371-2435.

Planck:1897:NTD

[Pla97a] Max Planck. Notiz zur Theorie der Dämpfung electrischer Schwingungen. (German) [Note on the theory of damping of electrical vibrations]. Annalen der Physik (Berlin), 299(13):419–422, ???? 1897. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1897:ISGc

[Pla97b] Max Planck. Über irreversible Strahlungsvorgänge. (Dritte Mitteilung). (German) [On irreversible radiation processes. (Third communication)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 52(??):1122–1145, December 16, 1897. CODEN SPWPAI. ISSN 0371-2435.

Planck:1897:ISGa

[Pla97c] Max Planck. Über irreversible Strahlungsvorgänge. (Erste Mitteilung). (German) [On irreversible radiation processes. (First communication)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 6(??):57–68, February 4, 1897. CODEN SPWPAI. ISSN 0371-2435.

Planck:1897:ISGb

[Pla97d] Max Planck. Über irreversible Strahlungsvorgänge. (II. Mitt.). (German) [On irreversible radiation processes. (Second communication)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 34(??):715–717, July 8, 1897. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1897:ESW

[Pla97e] Max Planck. Ueber electrische Schwingungen, welche durch Resonanz erregt und durch Strahlung gedämpft werden. (German) [On electrical vibrations that are excited by resonance and dampened by radiation]. Annalen der Physik (Berlin), 296(4):577–599, ???? 1897. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1897:VTG

[Pla97f] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Verlag von Veit & Comp., Leipzig, Germany, 1897. vi + 1 + 248 pp. LCCN QC311 .P7.

Planck:1898:RCA

[Pla98a] Max Planck. [Rezensionen]: Ch. Antoine: Dampfspannung: Neue Relation zwischen den Spannungen und den Temperaturen; Berechnung der Spannung verschiedener Dämpfe. Beiblätter zu den Annalen der Physik, XXII(??):293–??, ???? 1898. CODEN BEAPAI. ISSN 0375-846X.

Planck:1898:MTE

[Pla98b] Max Planck. Über die Maxwellsche Theorie der Elektrizität. (German) [On Maxwell’s theory of electricity]. Naturwissenschaftliche Rundschau, 13(42):541, October 15, 1898. CODEN NARSAC. ISSN 0028-1050.

Planck:1898:ISI

[Pla98c] Max Planck. Über irreversible Strahlungsvorgänge. (IV. Mitt.). (German) [On irreversible radiation processes. (Fourth communication)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 34(??):449–476, July 7, 1898. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1899:MTE

[Pla99a] Max Planck. Die Maxwellsche Theorie der Elektrizität von der mathematischen Seite betrachtet. (German) [Maxwell’s theory of electricity from the mathematical side]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 7(1):77–89, ???? 1899. ISSN 0012-0456 (print), 1869-7135 (electronic).

Planck:1899:ISG

[Pla99b] Max Planck. Irreversible strahlungsvorgäne. (German) [Irreversible radiation processes]. Naturwissenschaftliche Rundschau, 14(44):564, November 4, 1899. CODEN NARSAC. ISSN 0028-1050.

Planck:1899:ISV

[Pla99c] Max Planck. Über irreversible Strahlungsvorgänge. (Fünfte Mitteilung). (German) [On irreversible radiation processes. (Fifth communication)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 23(??):440–480, May 18, 1899. CODEN SPWPAI. ISSN 0371-2435.

Planck:1900:BAT

[Pla00a] Max Planck. Bemerkungen zu einer Abhandlung über Thermodynamik des Hrn. K. Wesendonck. (German) [Comments on a discussion of thermodynamics by Mr. K. Wesendonck]. Annalen der Physik (Berlin), 306(3):621–624, 1900. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1900:VWM

[Pla00b] Max Planck. Ein vermeintlicher Widerspruch des magnetooptischen Faraday-Effektes mit der Thermodynamik. (German) [A supposed contradiction of the magneto-optical Faraday effect with thermodynamics]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(14):206–210, November 2, 1900. CODEN VD-

PEAZ. ISSN 0372-5448. URL https://hdl.handle.net/2027/
coo.31924056107224?urlappend=%3Bseq=516.

Planck:1900:ETS

[Pla00c] Max Planck. Entropie und Temperatur strahlender Wärme. (German) [Entropy and temperature of radiant heat]. Annalen der Physik (Berlin), 306(4):719–737, ???? 1900. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/andp.19003060410/
abstract.

Planck:1900:KZS

[Pla00d] Max Planck. Kritik zweier Sätze des Hrn. W. Wien. (German) [Criticism of two statements of Mr. W. Wien]. Annalen der Physik (Berlin), 308(12):764–766, ???? 1900. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/andp.19003081215/
abstract.

Planck:1900:LPT

[Pla00e] Max Planck. Lekcii po termodinamike Maksa Planka. Tipografiâ Ministerstva Putej Soobšeniâ (T-va I. N. Kušnerev i Ko), St-Peterburg, Russia, 1900. 2 + iv + 2 + 234 pp. LCCN ???? Russian translation by Ivan Ivanovič Borgman and V. A. Kašerininova.

Planck:1900:EES

[Pla00f] Max Planck. Über die von einem elliptisch schwingenden Ion emittierte und absorbierte Energie. (German) [On energy emitted and absorbed by an elliptically vibrating ion]. Archives Néerlandaises des Sciences Exactes et Naturelles, ??(??):164–174, ???? 1900. CODEN ANSNAI. ISSN 0365-5059. Jubelband für H. A. Lorentz, 1900.

Planck:1900:VWS

[Pla00g] Max Planck. Ueber eine Verbesserung der Wiens’chen Spektralgleichung. (German) [On an improvement of Wien’s spectral equation]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(13):202–204, October 19, 1900. CODEN VD-

PEAZ. ISSN 0372-5448. URL http://web.ihep.su/dbserv/
compas/src/planck00/eng.pdf;   http://web.ihep.su/owa/
dbserv/hw.part2?s_c=PLANCK+1900;   https://hdl.handle.
net/2027/coo.31924056107224?urlappend=%3Bseq=520. En-

glish translation in [Pla69a].

Planck:1900:ISG

[Pla00h] Max Planck. Ueber irreversible Strahlungsvorgänge. (German) [On irreversible radiation processes]. Annalen der Physik (Berlin), 306(1):69–122, ???? 1900. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19003060105/abstract.

Planck:1900:TGE

[Pla00i] Max Planck. Zur Theorie des Gesetzes der Energieverteilung im Normalspektrum. (German) [On the theory of distribution of energy in the normal spectrum]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(17):237–245, December 14, 1900. CODEN VDPEAZ. ISSN 0372-5448. URL http://web.ihep.

su/dbserv/compas/src/planck00b/eng.pdf;   http://web.
ihep.su/owa/dbserv/hw.part2?s_c=PLANCK+1900B;   https:
//hdl.handle.net/2027/coo.31924056107224?urlappend=%3Bseq=

551. English translation in [Pla69b]. In this paper, Planck introduces a formula for the blackbody radiation law, and also derives estimates of Loschmidt’s number L (proportional to Avogadro’s number), the Boltzmann–Drude constant α, and the electron charge, e.

Planck:1901:VEZ

[Pla01a] Max Planck. Über die Verteilung der Energie zwischen Äther und Materie. (German) [On the distribution of energy between ether and matter]. Archives Néerlandaises des Sciences Exactes et Naturelles, 6 (série 2)(??):55–66, ???? 1901. CO-DEN ANSNAI. ISSN 0365-5059. URL https://archive.org/

stream/archivesnerland261901laha.pdf. Jubelband für J.

Bosscha.

Planck:1901:ISNa

[Pla01b] Max Planck. Über irreversible Strahlungsvorgänge (Nachtrag). (German) [On irreversible radiation processes (addendum)]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 25(??):544–555, May 9, 1901. CODEN SPWPAI. ISSN 0371-2435.

Planck:1901:ISNb

[Pla01c] Max Planck. Über irreversible Strahlungsvorgänge. Nachtrag. (German) [On irreversible radiation processes. Addendum]. Annalen der Physik (Berlin), 311(12):818–831, ???? 1901. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1901:GEI

[Pla01d] Max Planck. Ueber das Gesetz der Energieverteilung im Normalspectrum. (German) [On the law of distribution of energy in the normal spectrum]. Annalen der Physik (Berlin), 309(3):553–563, ???? 1901. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19013090310/abstract;   http://web.ihep.su/
dbserv/compas/src/planck01/eng.pdf;   http://web.ihep.
su/owa/dbserv/hw.part2?s_c=PLANCK+1901.

Planck:1901:EME

[Pla01e] Max Planck. Ueber die Elementarquanta der Materie und der Elektricität. (German) [On elementary quanta of matter and electricity]. Annalen der Physik (Berlin), 309(3):564–566, ???? 1901. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1901:ISG

[Pla01f] Max Planck. Ueber irreversible Strahlungsvorgänge. (German) [On irreversible radiation processes]. Annalen der Physik (Berlin), 311(12):818–831, 1901. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1901:VAS

[Pla01g] Max Planck. Vereinfachte Ableitung der Schwingungsgesetze eines linearen Resonators im stationär durchstrahlten Felde. (German) [Simplified derivation of the laws of vibration of a linear resonator in a stationary irradiated field]. Physikalische Zeitschrift, 2(36):530–534, June 8, 1901. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://hdl.handle.net/2027/
mdp.39015068319659?urlappend=%3Bseq=550.

Planck:1902:NWL

[Pla02a] Max Planck. Ueber die Natur des weissen Lichtes. (German) [On the nature of white light]. Annalen der Physik (Berlin), 312 (2):390–400, ???? 1902. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19023120210/abstract.

Planck:1902:VEZ

[Pla02b] Max Planck. Ueber die Verteilung der Energie zwischen Aether und Materie. (German) [On the distribution of energy between ether and matter]. Annalen der Physik (Berlin), 314(11):629– 641, ???? 1902. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). Jubelband für J. Bosscha.

Planck:1902:EES

[Pla02c] Max Planck. Ueber die von einem elliptisch schwingenden Ion emittierte und absorbierte Energie. (German) [On energy emitted and absorbed by an elliptically vibrating ion]. Annalen der Physik (Berlin), 314(11):619–628, ???? 1902. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1902:ETD

[Pla02d] Max Planck. Zur elektromagnetischen Theorie der Dispersion in isotropen Nichtleitern. (German) [On the electromagnetic theory of dispersion in isotropic insulators]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 24(??):470–494, May 1, 1902. CODEN SPWPAI. ISSN 0371-2435.

Planck:1902:TDB

[Pla02e] Max Planck. Zur Thermodynamik und Dissoziationstheorie binärer Elektrolyte. (German) [On the thermodynamics and dissociation of binary electrolytes]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 41(2):212– 223, ???? 1902. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1903:MMT

[Pla03a] Max Planck. Metalloptik und Maxwell’sche Theorie. (German) [Metal optics and Maxwell’s theory]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 26(??): 558–559, May 14, 1903. CODEN SPWPAI. ISSN 0371-2435.

Planck:1903:MSE

[Pla03b] Max Planck. Mr. Swinburne and entropy. The Electrician: the Oldest Weekly Illustrated Journal of Electrical Engineering, Industry, Science and Finance, 50(??):694–695, 821, ???? 1903. CODEN ELETAU. ISSN 0367-0805.

Planck:1903:RJW

[Pla03c] Max Planck. [Rezensionen]: J. W. Gibbs: Elementary Principles in statistical Mechanics, developed with especial reference to the rational foundation of Thermodynamics. Beiblätter zu den Annalen der Physik, 27(8):748–753, ???? 1903. CODEN BEAPAI. ISSN 0375-846X.

Planck:1903:TT

[Pla03d] Max Planck. Treatise on Thermodynamics. Longmans, Green, London, UK, 1903. xii + 272 pp. LCCN QC311 .P72. Authorized English translation by Alexander Ogg of [Pla97f].

Planck:1903:ODL

[Pla03e] Max Planck. Über den osmotischen Druck einer Lösung von merklich variabler Dichte. (German) [On the osmotic pressure of a solution of markedly-variable density]. Zeitschrift für Physikalische Chemie — Stöchiometrie und Verwandtschaftslehre, 42(5):584– 590, ???? 1903. CODEN ZEPCAC. ISSN 0372-8501.

Planck:1903:GLE

[Pla03f] Max Planck. Über die Grundlage der Lösungstheorie; eine Erwiderung. (German) [On the foundation of solution theory; a reply]. Annalen der Physik (Berlin), 315(2):436–445, ???? 1903. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1903:OEM

[Pla03g] Max Planck. Über die optischen Eigenschaften der Metalle für lange Wellen. (German) [On the optical properties of metals for long waves]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 13(??):278–280, March 5, 1903. CODEN SPWPAI. ISSN 0371-2435.

Planck:1903:EMT

[Pla03h] Max Planck. Zur elektro-magnetischen Theorie der selektiven Absorption in isotropen Nichtleitern. (German) [On an electromagnetic theory of selective absorption in isotropic dielectrics]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 24(??):480–498, April 30, 1903. CODEN SPWPAI. ISSN 0371-2435.

Planck:1904:ELE

[Pla04a] Max Planck. Über die Extinktion des Lichtes in einem optisch homogenen Medium von normaler Dispersion. (German) [The extinction of light in an optical homogeneous media from normal dispersion]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 22(??):740–750, April 21, 1904. CODEN SPAWAV. ISSN 0371-196X.

Planck:1904:BMB

[Pla04b] Max Planck. Über die mechanische Bedeutung der Temperatur und der Entropie. (German) [On the mechanical importance of temperature and entropy]. In Boltzmann-Festschrift 1904, pages 113–122. ????, ????, 1904.

Planck:1905:NAD

[Pla05a] Max Planck. Normale und anormale Dispersion in nichtleitenden Medien von variabler Dichte. (German) [Normal and abnormal dispersion in non-conductive products of variable density]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 18(??):382–394, April 6, 1905. CODEN SPWPAI. ISSN 0371-2435.

Planck:1905:CTI

[Pla05b] Max Planck. On Clausius’ Theorem for irreversible cycles and on the increase of entropy. The London, Edinburgh and Dublin Philosophical Magazine and Journal of Science, 6(9):167–168, ???? 1905. CODEN PHMAA4. ISSN 0031-8086.

Planck:1905:VTG

[Pla05c] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Veit, Leipzig, Germany, second edition, 1905. vii + 256 pp. LCCN QC311 .P71 2d ed.

Planck:1906:JDJ

[Pla06a] Max Planck. Am 5. Juli dieses Jahres ist Paul Drude . . .. (German) [On July 5 of this year is Paul Drude …]. Annalen der Physik (Berlin), 325(9):i–iv, 1906. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1906:BKW

[Pla06b] Max Planck. Bemerkung über die Konstante des Wienschen Verschiebungsgesetzes. (German) [Remarks about the constant of the Wien’s displacement law]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 8(24):695–696, December 30, 1906. CODEN VDPEAZ. ISSN 0372-5448.

URL https://hdl.handle.net/2027/coo.31924056112125?
urlappend=%3Bseq=709.

Planck:1906:PRG

[Pla06c] Max Planck. Das Prinzip der Relativität und die Grundgleichungen der Mechanik. (German) [The Principle of Relativity and the basic equations of mechanics]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 8(7):136–141, April 15, 1906. CO-DEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.

net/2027/coo.31924056112125?urlappend=%3Bseq=150.

Planck:1906:KMAa

[Pla06d] Max Planck. Die Kaufmannschen Messungen der Ablenkbarkeit der α-Strahlen in ihrer Bedeutung für die Dynamik der Elektronen. (German) [Kaufmann’s measurements of deflection of β rays and their importance for the dynamics of electrons]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 8(20): 418–432, October 30, 1906. CODEN VDPEAZ. ISSN 0372-5448.

URL https://hdl.handle.net/2027/coo.31924056112125?
urlappend=%3Bseq=432.

Planck:1906:KMAb

[Pla06e] Max Planck. Die Kaufmannschen Messungen der Ablenkbarkeit der α-Strahlen in ihrer Bedeutung für die Dynamik der Elektronen. (German) [Kaufmann’s measurements of deflection of β rays and their importance for the dynamics of electrons]. Physikalische Zeitschrift, 7(21):753–761, October 25, 1906. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015068320962%3Bview=1up%3Bseq=797.

Planck:1906:PDG

[Pla06f] Max Planck. Paul Drude. Gedächtnisrede, gehalten in der Sitzg. d. Dtsch. Physik. Ges. am 30. November 1906. (German) [Paul Drude. Memorial speech, held in the meeting of the German Physical Society, 30 November 1906]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 8(23):599– 630, December 15, 1906. CODEN VDPEAZ. ISSN 0372-5448.

URL https://hdl.handle.net/2027/coo.31924056112125?
urlappend=%3Bseq=613.

Planck:1906:VTW

[Pla06g] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, 1906. xiii + 222 pp. LCCN QC331 .P7.

Planck:1907:KMA

[Pla07a] Max Planck. Die Kaufmannschen Messungen der Ablenkbarkeit der α-Strahlen in ihrer Bedeutung für die Dynamik der Elektronen. (German) [Kaufmann’s measurements of deflection of β rays and their importance for the dynamics of electrons]. Verhandlungen der Gesellschaft Deutscher Naturforscher und Ärzte, 2(1): 61–62, ???? 1907. CODEN VGDNAN. ISSN 0172-0651.

Planck:1907:DPC

[Pla07b] Max Planck. Dr. phil. Carl Liebenow. Verhandlungen der Deutschen Physikalischen Gesellschaft, 9(1):1, January 15, 1907. CODEN VDPEAZ. ISSN 0372-5448. URL https://hdl.

handle.net/2027/chi.21403652?urlappend=%3Bseq=15.

Planck:1907:NBK

[Pla07c] Max Planck. Nachtrag zu der Besprechung der Kaufmannschen Ablenkungsmessungen. (German) [Subsequent to the review of Kaufmann’s deflection measurements]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 9(14):301–305, July 30, 1907. CODEN VDPEAZ. ISSN 0372-5448. URL https://hdl.

handle.net/2027/chi.21403652?urlappend=%3Bseq=317.

Planck:1907:TSS

[Pla07d] Max Planck. Theorie der stationären Strahlungs in einem gleichförmig bewegten Hohlraum. Von Kurd von Mosengeil. (German) [Theory of stationary radiation in a uniformly moving cavity. From Kurd von Mosengeil]. Annalen der Physik (Berlin), 327(5):867–904, ???? 1907. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). Planck’s pupil Kurd von Mosengeil died prematurely in 1907.

Planck:1907:WB

[Pla07e] Max Planck. Wilhelm von Bezold. Verhandlungen der Deutschen Physikalischen Gesellschaft, 9(5–6):91–93, March 30, 1907. CO-DEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.

net/2027/chi.21403652?urlappend=%3Bseq=107.

Planck:1907:DBS

[Pla07f] Max Planck. Zur Dynamik bewegter Systeme. (German) [The dynamics of moving systems]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 29(??):542–570, June 13, 1907. CODEN SPWPAI. ISSN 0371-2435.

Planck:1907:TDG

[Pla07g] Max Planck. Zur Theorie der Dispersion. (German) [On the theory of dispersion]. Physikalische Zeitschrift, 8(24):

906, December 1, 1907. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015068320954%3Bview=1up%3Bseq=938.

Planck:1908:BPAa

[Pla08a] Max Planck. Bemerkungen zum Prinzip der Aktion und Reaction der allgemeinen Dynamik. (German) [Remarks on the principle of action and reaction of general dynamics]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 10(18– 20):728–732, October 30, 1908. CODEN VDPEAZ. ISSN 0372-

5448. URL https://hdl.handle.net/2027/chi.105146030?
urlappend=%3Bseq=750.

Planck:1908:BPAb

[Pla08b] Max Planck. Bemerkungen zum Prinzip der Aktion und Reaction der allgemeinen Dynamik. (German) [Remarks on the principle of action and reaction of general dynamics]. Physikalische Zeitschrift, 9(23):828–830, November 15, 1908. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://babel.hathitrust.org/
cgi/pt?id=mdp.39015068320947%3Bview=1up%3Bseq=860.

Planck:1908:PEE

[Pla08c] Max Planck. Das Prinzip der Erhaltung der Energie. (German) [The principle of conservation of energy]. Teubner, Stuttgart, Germany; Leipzig, Germany, second edition, 1908. ???? pp.

Planck:1908:AP

[Pla08d] Max Planck. Dr. Adolf Paalzow. Verhandlungen der Deutschen Physikalischen Gesellschaft, 10(2):17, January 30, 1908. CODEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.net/

2027/chi.105146030?urlappend=%3Bseq=35.

Planck:1908:KZE

[Pla08e] Max Planck. Über die kanonische Zustandsgleichung einatomiger Gase. (German) [On the canonical equation of state of monatomic gases]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 32(??):633–647, June 25, 1908. CODEN SPAWAV. ISSN 0371-196X.

Planck:1908:DBS

[Pla08f] Max Planck. Zur Dynamik bewegter Systeme. (German) [On the dynamics of moving systems]. Annalen der Physik (Berlin), 331(6):1–34, ???? 1908. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19083310602/abstract.

Planck:1908:TDG

[Pla08g] Max Planck. Zur Theorie der Dispersion. (German) [On the theory of dispersion]. Physikalische Zeitschrift, 9(11):354, June 1, 1908. CODEN PHZTAO. ISSN 0369-982X. URL http://babel.

hathitrust.org/cgi/pt?id=mdp.39015035421141%3Bview=1up%
3Bseq=380.

Planck:1909:BPA

[Pla09a] Max Planck. Bemerkungen zum Prinzip der Aktion und Reaktion der allgemeinen Dynamik. (German) [Remarks on the principle of action and reaction of general dynamics]. Verhandlungen der Gesellschaft Deutscher Naturforscher und Ärzte zu Köln, 2(??): 53–??, ???? 1909.

Planck:1909:EPWa

[Pla09b] Max Planck. Die Einheit des physikalischen Weltbildes. Leipzig 1909. (German) [The unity of physical pictures of the world. Leipzig 1909]. Physikalische Zeitschrift, 10(2): 62–75, January 15, 1909. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015023919049%3Bview=1up%3Bseq=90.

Planck:1909:EPWb

[Pla09c] Max Planck. Die einheit des physikalischen weltbildes; vortrag gehalten am 9. dezember 1908 in der naturwissenschaftlichen fakultät des studentenkorps an der Universität Leiden. (German) [The unity of the physical picture of the world; lecture held on 9 December 1908 in the Faculty of Science of the student corps at the Leiden University]. S. Hirzel, Leipzig, Germany, 1909. 38 pp. LCCN QC71 .P713.

Planck:1909:DEB

[Pla09d] Max Planck. Diskussion (zu Einstein, Konstitution der Strahlung). (German) [Discussion (on Einstein, Constitution of radiation)]. Physikalische Zeitschrift, 10(22):825–826, November 10, 1909. CODEN PHZTAO. ISSN 0369-982X. URL http://

babel.hathitrust.org/cgi/pt?id=mdp.39015023919049%3Bview=
1up%3Bseq=879.

Planck:1909:ELT

[Pla09e] Max Planck. Eight Lectures on Theoretical Physics: Delivered at Columbia University in 1909. Forgotten Books, ????, 1909. ISBN 1-330-06618-9. 150 pp. LCCN ????

Planck:1909:KGK

[Pla09f] Max Planck. Zur kinetischen Gastheorie. Eine kritische Untersuchung. (German) [On kinetic gas theory. a critical investigation]. Physikalische Zeitschrift, 10(6):195–196, March 15, 1909. CODEN PHZTAO. ISSN 0369-982X. URL http://babel.

hathitrust.org/cgi/pt?id=mdp.39015023919049%3Bview=1up%
3Bseq=229.

Planck:1910:AVT

[Pla10a] Max Planck. Acht Vorlesungen über theoretische Physik, gehalten an der Columbia university in the city of New York im frühjahr 1909. (German) [Eight Lectures on Theoretical Physics, held at Columbia University in the City of New York in the spring of 1909]. S. Hirzel, Leipzig, Germany, 1910. 4 + 127

pp. LCCN QC71 .P6. URL http://catalog.hathitrust.org/
api/volumes/oclc/982335.html.

Planck:1910:SNPd

[Pla10b] Max Planck. Die Stellung der neueren Physik zur mechanischen Naturanschauung. (German) [The position of modern physics in the mechanical view of nature]. F. C. W. Vogel, Leipzig, Germany, 1910. 20 pp. LCCN ????

Planck:1910:SNPa

[Pla10c] Max Planck. Die Stellung der neueren Physik zur mechanischen Naturanschauung. Vortrag gehalten am 29.9.1910 in Königsberg . (German) [The position of modern physics on the mechanical view of nature. Lecture held on 29 September 1910 in Königsberg]. Naturwissenschaftliche Rundschau, 25(41):521–525, October 13, 1910. CODEN NARSAC. ISSN 0028-1050.

Planck:1910:SNPb

[Pla10d] Max Planck. Die Stellung der neueren Physik zur mechanischen Naturanschauung. Vortrag gehalten am 29.9.1910 in Königsberg . (German) [The position of modern physics on the mechanical view of nature. Lecture held on 29 September 1910 in Königsberg]. Naturwissenschaftliche Rundschau, 25(42):533–536, October 20, 1910. CODEN NARSAC. ISSN 0028-1050.

Planck:1910:SNPc

[Pla10e] Max Planck. Die Stellung der neueren Physik zur mechanischen Naturanschauung. Vortrag gehalten am 29.9.1910 in Königsberg . (German) [The position of modern physics on the mechanical view of nature. Lecture held on 29 September 1910 in Königsberg]. Physikalische Zeitschrift, 11(21–22):922–932, November 1, 1910. CODEN PHZTAO. ISSN 0369-982X. URL http://hdl.handle.

net/2027/njp.32101054770845?urlappend=%3Bseq=472.

Planck:1910:GRL

[Pla10f] Max Planck. Gleichförmige Rotation und Lorentz-Kontraktion. (German) [Uniform rotation and Lorentz contraction]. Physikalische Zeitschrift, 11(7):294, April 1, 1910. CODEN PHZTAO. ISSN

0369-982X. URL http://babel.hathitrust.org/cgi/pt?id=
mdp.39015035421174%3Bview=1up%3Bseq=370.

Planck:1910:MTN

[Pla10g] Max Planck. The mechanical theory of nature. Scientific American, 70(??):387, ???? 1910. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v70/
n1824supp/full/scientificamerican12171910-387asupp.html.

Planck:1910:IBN

[Pla10h] Max Planck. Über den Inhalt und die Bedeutung des Nernstschen Wärmetheorems für die reine Thermodynamik. (German) [On the content and meaning of Nernst’s Heat Theorem for pure thermodynamics]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 40(??):865, October 20, 1910. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1910:MTP

[Pla10i] Max Planck. Zur Machschen Theorie der physikalischen Erkenntnis. Eine Erwiderung. (German) [On Mach’s theory of physical cognition. A reply]. Physikalische Zeitschrift, 11(25):1186–1190, December 8, 1910. CODEN PHZTAO.

ISSN 0369-982X. URL http://hdl.handle.net/2027/njp.
32101054770845?urlappend=%3Bseq=746.

Planck:1910:TWG

[Pla10j] Max Planck. Zur Theorie der Wärmestrahlung. (German) [On the theory of heat radiation]. Annalen der Physik (Berlin), 336(4):

758–768, ???? 1910. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1911:SNP

[Pla11a] Max Planck. Die Stellung der neueren Physik zur mechanischen Naturanschauung. Vortrag gehalten am 29.9.1910 in Königsberg . (German) [The position of modern physics on the mechanical view of nature. Lecture held on 29 September 1910 in Königsberg]. Verhandlungen der Gesellschaft Deutscher Naturforscher und Ärzte, 1(??):58–75, ???? 1911. CODEN VGDNAN. ISSN 0172-0651.

Planck:1911:NSG

[Pla11b] Max Planck. Eine neue Strahlungshypothese. (German) [A new radiation hypothesis]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 13(3):138–148, February 15, 1911. CODEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.net/

2027/coo.31924056113495?urlappend=%3Bseq=154.

Planck:1911:ETVa

[Pla11c] Max Planck. Énergie et température: Vortrag, gehalten am 21. April 1911 in Paris bei der Ostertagung der Französ. Physik. Ges.. (French) [Energy and temperature: Lecture held on 21 April 2011 in Paris at the Easter meeting of the French Physical Society]. Journal de Physique Th’eorique et Appliqu’ee, 1(1):345–359, 1911. CODEN JPTAAT. ISSN 0368-3893.

Planck:1911:ETVb

[Pla11d] Max Planck. Énergie et température: Vortrag, gehalten am 21. April 1911 in Paris bei der Ostertagung der Französ. Physik. Ges. (French) [Energy and temperature: Lecture held on 21 April 2011 in Paris at the Easter meeting of the French Physical Society]. Physikalische Zeitschrift, 12(16): 681–687, August 15, 1911. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015011417709%3Bview=1up%3Bseq=739.

Planck:1911:VTG

[Pla11e] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Verlag von Veit & Comp., Leipzig, Germany, third edition, 1911. vii + 286 pp. LCCN QC311 .P7

Planck:1911:HQG

[Pla11f] Max Planck. Zur Hypothese der Quantenemission. (German) [On the hypothesis of quanta emissions]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 35(??): 723–731, July 13, 1911. CODEN SPWPAI. ISSN 0371-2435.

Planck:1912:LDR

[Pla12a] Max Planck. La loi du rayonnement noir et l’hypothèse des quantités élémentaires d’action. (French) [The law of black body radiation and the hypothesis of the quantum of action]. In Langevin and de Broglie [LdB12], pages 93–114 + 115–132 (discussion).

URL https://archive.org/details/lathoriedurayo00inst.

Planck:1912:BGS

[Pla12b] Max Planck. Über die Begründung des Gesetzes der schwarzen Strahlung. (German) [On the justification of the law of black [body] radiation]. Annalen der Physik (Berlin), 342(4):642–656, ???? 1912. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1912:NTTa

[Pla12c] Max Planck. Über neuere thermodynamische Theorien. (Nernstsches Wärmetheorem und Quanten-Hypothese). (German) [On recent thermodynamic theories. (Nernst heat theorem and quantum hypothesis)]. Berichte der deutschen chemischen Gesellschaft, 45(1):5–23, 1912. CODEN BDCGAS. ISSN 0365-9496.

Planck:1912:NTTb

[Pla12d] Max Planck. Über neuere thermodynamische Theorien. (Nernstsches Wärmetheorem und Quanten-Hypothese). (German) [On recent thermodynamic theories. (Nernst heat theorem and quantum hypothesis)]. Physikalische Zeitschrift, 13(4):165– 175, February 15, 1912. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015023176806%3Bview=1up%3Bseq=223.

Planck:1913:AGS

[Pla13a] Max Planck. Ansprache. (German) [Speech]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 4(??):73–76, January 23, 1913. URL https://

biodiversitylibrary.org/page/40701758.

Planck:1913:PEE

[Pla13b] Max Planck. Das Prinzip der Erhaltung der Energie. (German) [The principle of conservation of energy]. Teubner, Stuttgart, Germany; Leipzig, Germany, third edition, 1913. ???? pp.

Planck:1913:GBQa

[Pla13c] Max Planck. Die gegenwärtige Bedeutung der Quantenhypothese für die kinetische Gastheorie. Wolfskehl-Vortrag, Göttingen, April 1913. (German) [The contemporary significance of the quantum hypothesis for the kinetic theory. Wolfskehl lecture, Göttingen, April 1913]. Physikalische Zeitschrift, 14(6):258– 259, March 15, 1913. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015021268936%3Bview=1up%3Bseq=304.

Planck:1913:GBQb

[Pla13d] Max Planck. Die gegenwärtige Bedeutung der Quantenhypothese für die kinetische Gastheorie. Wolfskehl-Vortrag, Göttingen, April 1913. (German) [The contemporary significance of the quantum hypothesis for the kinetic theory. Wolfskehl lecture, Göttingen, April 1913]. Nachrichten von der Gesellschaft der Wissenschaften zu Göttingen. Mathematisch-Physikalische Klasse, ??(??):137– 140, ???? 1913. CODEN NWGAAN. ISSN 0369-6650.

Planck:1913:GWH

[Pla13e] Max Planck. Die Gesetze der Wärmestrahlung und die Hypothese der elementaren Wirkungsquanten. Vortrag, gehalten auf dem Solvay-Kongreß, Brüssel, 1911. (German) [The laws of thermal radiation and the hypothesis of the elementary quantum effect. Paper presented at the Solvay Congress, Brussels, 1911]. In Arnold Eucken, editor, Die Theorie der Strahlung und der Quanten. Verhandlungen auf einer von E. Solvay einberufenen Zusammenkunft [30.Oktober bis 3.November 1911]; mit einem Anhange über die Entwicklung der Quantentheorie vom Herbst 1911 bis zum Sommer 1913. (German) [Theory of radiation and the quantum: Proceedings of a conference convened by E. Solvay [30 October to 3 November 1911]; with an appendix on the development of quantum theory from the fall of 1911 until the summer of 1913], pages 77–94 (??). S. Wilhelm Knapp, Halle, Germany, 1913.

Planck:1913:ESH

[Pla13f] Max Planck. Erwiderrung des Sekretärs Hrn. Planck. (German) [Reply of the secretary Mr. Planck]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften zu Berlin, 32(??):600–602, June 26, 1913. URL https://

biodiversitylibrary.org/page/40702346.

Planck:1913:LTT

[Pla13g] Max Planck. Leçons de thermodynamique: Le théorème de Nernst et l’hypothèse des quanta. (French) [Lectures on thermodynamics: Nernst’s theorem and the quantum hypothesis]. Librairie scientifique A. Hermann et Fils, Paris, France, 1913. 310 pp. LCCN ????

Planck:1913:NBP

[Pla13h] Max Planck. Neue Bahnen der physikalischen Erkenntnis: Rede, gehalten in der Berliner Universität anläßlich der Übernahme des Rektorats. (German) [New paths of physical knowledge: Speech held, in the University of Berlin in connection with the assumption of the Rector’s Office. Berlin 1913], volume 1913 of Rektorwechsel an der Friedrich-Wilhelms-Universität zu Berlin. Norddeutsche Buchdruckerei, Berlin, Germany, 1913. 23–45 pp. LCCN ????

Planck:1913:GZO

[Pla13i] Max Planck. Über das Gleichgewicht zwischen Oszillatoren, freien Elektronen und strahlender Wärme. (German) [On the balance between oscillators, free electrons, and radiant heat]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 18(??):350–363, April 3, 1913. CODEN SPWPAI. ISSN 0371-2435.

Planck:1913:VTW

[Pla13j] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, second edition, 1913. xii + 206 + 7 pp. LCCN QC331 .P7.

Planck:1913:VTG

[Pla13k] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Verlag von Veit & Comp., Leipzig, Germany, fourth edition, 1913. vii + 288 pp. LCCN QC311 .P7

Planck:1914:PKWa

[Pla14a] Max Planck. Das Prinzip der kleinsten Wirkung. (German) [The principle of least action]. Die Kultur der Gegenwart, 3(3):1–??, ???? 1914.

Planck:1914:PKWb

[Pla14b] Max Planck. Das Prinzip der kleinsten Wirkung. (German) [The principle of least action]. In Physik, volume 1 (Part 3) of Die Kultur der Gegenwart. Ihre Entwicklung und ihre Ziele, pages 692–702. Emil Warburg, Leipzig and Berlin, Germany, 1914.

Planck:1914:GBQ

[Pla14c] Max Planck. Die gegenwärtige Bedeutung der Quantenhypothese für die kinetische Gastheorie. Wolfskehl-Vortrag, Göttingen, April 1913. (German) [The contemporary significance of the quantum hypothesis for the kinetic theory. Wolfskehl lecture, Göttingen, April 1913]. In David Hilbert, editor, Vorträge über die kinetische Theorie der Materie und der Elektrizität: Gehalten in Göttingen auf Einladung der Kommission der Wolfskehlstiftung, volume 6 of Mathematische Vorlesungen an der Universität Göttingen, pages 2–16. Teubner, Stuttgart, Germany; Leipzig, Germany, 1914.

Planck:1914:DSG

[Pla14d] Max Planck. Dynamische und statistische Gesetzmässigkeit. Rede, gehalten bei der Feier zum Gedächtnis des Stifters der Berliner Friedrich-Wilhelms-Universität am 3. August 1914. (German) [Dynamic and statistical regularity. Speech held at the ceremony in memory of the founder of Berlin’s Friedrich Wilhelm University on 3 August 1914]. Rede zur Gedächtnisfeier des Stifters der Berliner Universität, König Friedrich Wilhelm III. J. A. Barth, Leipzig, Germany, 1914. 31 pp. LCCN Q171 .P65.

Planck:1914:VFQ

[Pla14e] Max Planck. Eine veränderte Formulierung der Quantenhypothese. (German) [A modified formulation of the quantum hypothesis]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 34(??):918–923, July 30, 1914. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1914:NBP

[Pla14f] Max Planck. Neue Bahnen der physikalischen Erkenntnis: rede Gehalten beim Antritt des Rektorats der Friedrich Wilhelm-Universitaet, Berlin, am 15. oktober, 1913. (German) [New paths of physical knowledge: speech held at the commencement of the Rector of the Friedrich Wilhelm University, Berlin, on 15 October 1913]. J. A. Barth, Leipzig, Germany, 1914. 27 + 1 pp. LCCN ????

Planck:1914:NPP

[Pla14g] Max Planck. [New paths of physical knowledge: Speech held, in the University of Berlin in connection with the assumption of the Rector’s Office. Berlin 1913]. The London, Edinburgh and Dublin Philosophical Magazine and Journal of Science, 6(??):60– 71, 1914. CODEN PHMAA4. ISSN 0031-8086.

Planck:1914:THR

[Pla14h] Max Planck. The Theory of Heat Radiation. P. Blakiston’s Son and Co., Philadelphia, PA, USA, 1914. xiv + 1 + 225 pp. LCCN QC331 .P73 1914. Authorized translation by Morton Masius.

Planck:1914:VTZa

[Pla14i] Max Planck. Verhältnis der Theorien zueinander. (German) [Ratio of the theories to each other]. Die Kultur der Gegenwart, 3 (3):1–??, ???? 1914.

Planck:1914:VTZb

[Pla14j] Max Planck. Verhältnis der Theorien zueinander. (German) [Ratio of the theories to each other]. In Physik, volume 1 (Part 3) of Die Kultur der Gegenwart. Ihre Entwicklung und ihre Ziele, pages 714–731. Emil Warburg, Leipzig and Berlin, Germany, 1914.

Planck:1915:AGO

[Pla15a] Max Planck. Ansprache, gehalten in der öffentl. Sitzung der Berliner Akademie zur Feier des Leibnizschen Jahrestages vom 1. Juli. (German) [Speech held at the public session of the Berlin Academy to celebrate the anniversary of Leibniz on 1 July]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 33(??):481–484, July 1, 1915. CODEN SPWPAI. ISSN 0371-2435.

Planck:1915:BES

[Pla15b] Max Planck. Bemerkung über die Emission von Spektrallinien. (German) [Comment on the emission of spectral lines]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 53(??):909–913, December 16, 1915. CODEN SPW-PAI. ISSN 0371-2435.

Planck:1915:BEZ

[Pla15c] Max Planck. Bemerkung über die Entropiekonstante zweiatomiger Gase. (German) [Remarks on the entropy constant of diatomic gases]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(22):418–419, November 30, 1915. CODEN VD-

PEAZ. ISSN 0372-5448. URL https://hdl.handle.net/2027/
coo.31924056113297?urlappend=%3Bseq=430.

Planck:1915:QMMa

[Pla15d] Max Planck. Die Quantenhypothese für Molekeln mit mehreren Freiheitsgraden. (I. Mitt.). (German) [The quantum hypothesis for molecules with several degrees of freedom. (First communication)]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(22):407–418, November 30, 1915. CODEN VD-

PEAZ. ISSN 0372-5448. URL https://hdl.handle.net/2027/
coo.31924056113297?urlappend=%3Bseq=419.

Planck:1915:QMMb

[Pla15e] Max Planck. Die Quantenhypothese für Molekeln mit mehreren Freiheitsgraden. (II. Mitt.). (German) [The quantum hypothesis for molecules with several degrees of freedom. (Second communication)]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(22):438–451, November 30, 1915. CODEN VD-

PEAZ. ISSN 0372-5448. URL https://hdl.handle.net/2027/
coo.31924056113297?urlappend=%3Bseq=450.

Planck:1915:ER

[Pla15f] Max Planck. Eduard Riecke. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(13):217–218, July 15, 1915. CO-DEN VDPEAZ. ISSN 0372-5448. URL https://hdl.handle.

net/2027/coo.31924056113297?urlappend=%3Bseq=227.

Planck:1915:ELT

[Pla15g] Max Planck. Eight lectures on theoretical physics: delivered at Columbia University in 1909, volume 3 of Publications of the Ernest Kempton Adams Fund for physical research. Columbia University Press, New York, NY, USA, 1915. ix + 130 pp. LCCN QC71 .P7. URL http://catalog.

hathitrust.org/api/volumes/oclc/3393318.html;   http:
//www.archive.org/details/cu31924005649185;   http://
www.archive.org/details/eightlecturesont00planuoft/. Trans-

lation from German to English by A.P. (Albert Potter) Wills.

Planck:1915:BEE

[Pla15h] Max Planck. Über die Energieverteilung in einem System rotierender Dipole. (German) [On the energy distribution in a system of rotating dipoles]. In Elster-Geitel-Festschrift 1915, pages 313–317. ????, ????, 1915.

Planck:1915:QEG

[Pla15i] Max Planck. Über Quantenwirkungen in der Elektrodynamik. (German) [On quantum effects in electrodynamics]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 35(??):512–519, July 8, 1915. CODEN SPWPAI. ISSN 0371-2435.

Planck:1916:BQD

[Pla16a] Max Planck. Bemerkung zur quantentheoretischen Deutung der Rubens–Hettnerschen Spektralmessung. (German) [Remarks on the quantum interpretation of Rubens–Hettner spectral measurements]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 18(5):168–172, March 15, 1916. CODEN VD-

PEAZ. ISSN 0372-5448. URL https://hdl.handle.net/2027/
coo.31924056112240?urlappend=%3Bseq=178.

Planck:1916:PSP

[Pla16b] Max Planck. Die physikalische Struktur des Phasenraumes. (German) [The physical structure of phase space]. Annalen der Physik (Berlin), 355(12):385–418, ???? 1916. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1916:EAM

[Pla16c] Max Planck. Einführung in die allgemeine Mechanik: zum Gebrauch bei Vorträgen sowie zum Selbstunterricht. (German) [Introduction to General Mechanics: For use in presentations, as well as self teaching], volume 1 of Einführung in die theoretische Physik. Hirzel, Leipzig, Germany, 1916. iv + 218 pp. LCCN ????

Planck:1916:EMD

[Pla16d] Max Planck. Einführung in die Mechanik deformierbarer Körper. (German) [Introduction to the Mechanics of Deformable Bodies], volume 2 of Einführung in die theoretische Physik. Hirzel, Leipzig, Germany, 1916. iv + 218 pp. LCCN ????

Planck:1916:NBP

[Pla16e] Max Planck. Neue Bahnen der physikalischen Erkenntnis: rede Gehalten beim antritt des Rektorats der Friedrich Wilhelm-Universitaet, Berlin, am 15. oktober, 1913. (German) [New paths of physical knowledge: speech held at the commencement of the Rector of the Friedrich Wilhelm University, Berlin, on 15 October 1913]. Norddeutsche Buchdruckerei, Leipzig, Germany, second edition, 1916. 27 pp. LCCN ????

Planck:1916:AEE

[Pla16f] Max Planck. Über die absolute Entropie einatomiger Körper. (German) [On the absolute entropy of single-atom bodies]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 31(??):653–667, June 8, 1916. CODEN SPWPAI. ISSN 0371-2435.

Planck:1917:AGB

[Pla17a] Max Planck. Ansprache, gehalten in der Berliner Akademie zur Feier des Geburtsfestes des Kaisers und des Jahrestages König Friedrichs II. (German) [Speech held at the Berlin Academy in celebration of the birth of the Emperor and the anniversary of King Frederick II]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 4(??):35–40, January 25, 1917. CODEN SPWPAI. ISSN 0371-2435.

Planck:1917:SSD

[Pla17b] Max Planck. Über einen Satz der statistischen Dynamik und seine Erweiterung in der Quantentheorie. (German) [On a theorem of statistical dynamics and its extension in quantum theory]. Sitzungsberichte der Königlich Preussischen Akademie der Wissenschaften, 24(??):324–341, May 10, 1917. CODEN SPW-

PAI. ISSN 0371-2435. URL https://en.wikipedia.org/wiki/
Fokker%E2%80%93Planck_equation. This paper introduces the

famous equation now known as the “Fokker–Planck equation”; see also [Fok14].

Planck:1917:VTG

[Pla17c] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Verlag von Veit & Comp., Leipzig, Germany, fifth edition, 1917. viii + 290 pp. LCCN QC311 .P7

Planck:1917:TRMa

[Pla17d] Max Planck. Zur Theorie des Rotationsspektrums. 1. Mitteilung. (German) [On the theory of the rotational spectrum. First communication]. Annalen der Physik (Berlin), 357(5):491–505, ???? 1917. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19173570505/abstract.

Planck:1917:TRMb

[Pla17e] Max Planck. Zur Theorie des Rotationsspektrums. 2. Mitteilung. (German) [On the theory of the rotational spectrum. Second communication]. Annalen der Physik (Berlin), 356(11):241–256, ???? 1917. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19173581107/abstract.

Planck:1917:TRGa

[Pla17f] Max Planck. Zur Theorie des Rotationsspektrums. (German) [On the theory of the rotational spectrum]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19(5–6):43– 47, March 30, 1917. CODEN VDPEAZ. ISSN 0372-5448.

URL https://hdl.handle.net/2027/coo.31924056113289?
urlappend=%3Bseq=53.

Planck:1917:TRGb

[Pla17g] Max Planck. Zur Theorie des Rotationsspektrums. (German) [On the theory of the rotational spectrum]. Annalen der Physik (Berlin), 357(5):491–505, 1917. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1917:TRGc

[Pla17h] Max Planck. Zur Theorie des Rotationsspektrums. (German) [On the theory of the rotational spectrum]. Annalen der Physik (Berlin), 358(11):241–256, 1917. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1918:EAA

[Pla18a] Max Planck. Erwiderung auf die Ansprachen vom 26. April 1918 zu Max Plancks 60. Geburtstag in der Deutschen Physikalischen Gesellschaft. (German) [Reply to the speeches of 26 April 1918 for Max Planck’s 60th birthday in the German Physical Society]. Unknown, 1918.

Planck:1918:QAK

[Pla18b] Max Planck. Zur Quantelung des asymmetrischen Kreisels. (German) [On quantization of the asymmetric top]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 50(??):1166–1174, December 5, 1918. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1919:AMP

[Pla19a] Max Planck. Ansprache von Max Planck. (German) [Speech by Max Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–8, ???? 1919. CODEN SPWPAI. ISSN 0371-2435.

Planck:1919:WLG

[Pla19b] Max Planck. Das Wesen des Lichts. (German) [The beings of light]. Die Naturwissenschaften, 7(48):903–909, November 28, 1919. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF01485818.

Planck:1919:DWN

[Pla19c] Max Planck. Die Dissoziationswärme des Wasserstoffs nach dem Bohr–Debyeschen Modell. (German) [The dissociation of hydrogen according to the Bohr–Debye model]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 48(??):914–931, November 27, 1919. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1919:MHD

[Pla19d] Max Planck. Eine Mitteilung von Hrn. Dr. A. Lande i Öberhambach bei Heppenheim: Elektronenbahnen im Polyederverband. (German) [A message from the Hon. Dr. A. Lande in Öberhambach in Heppenheim: electronic orbits in linked polyhedra]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 1(??):1, January 1, 1919. CODEN SPWPAI. ISSN 0371-2435.

Planck:1919:EMD

[Pla19e] Max Planck. Einführung in die Mechanik deformierbarer Körper: Zum Gebrauch bei Vorträgen, sowie zum Selbstunterricht. (German) [Introduction to the Mechanics of Deformable Bodies: For use in presentations, as well as self teaching], volume 2 of Planck: Einführung in die theoretische Physik. Hirzel, Leipzig, Germany, 1919. 3 + 193 pp. LCCN QA935 .P53. URL http://catalog.

hathitrust.org/api/volumes/oclc/8314942.html;   http:
//name.umdl.umich.edu/ABR6264.

Planck:1919:VTW

[Pla19f] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, third edition, 1919. x + 206 pp. LCCN QC331 .P7 1919.

Planck:1920:WLV

[Pla20a] Max Planck. Das Wesen des Lichts: Vortrag, gehalten in der Hauptversammlung der Kaiser Wilhelm-Gesellschaft am 28.10.1919. (German) [Particles of Light: Paper presented at the Annual Meeting of the Kaiser Wilhelm Society on 10.28.1919]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1920. 22 pp. LCCN ????

Planck:1920:EBEa

[Pla20b] Max Planck. Die Entstehung und bisherige Entwicklung der Quantentheorie. Nobel-Vortrag, (mit Lebenslauf). Gehalten vor der königl. Schwed. Akad. d. Wiss. zu Stockholm am 2.6.1920. (German) [The emergence and current development of quantum theory. Nobel Lecture, (with curriculum vitae). Held before the Royal Swedish Academy of Sciences in Stockholm on 2.6.1920]. J. A. Barth, Leipzig, Germany, 1920. 32 pp. LCCN ????

Planck:1920:EBEb

[Pla20c] Max Planck. Entstehung und bisherige Entwicklung der Quantentheorie: Nobelvortrag, gehalten vor der Königlich Schwedischen Akademie der Wissenschaften zu Stockholm am 2. Juni 1920. (German) [Emergence and current development of Quantum Theory: Nobel Lecture, held in front of the Royal Swedish Academy of Sciences in Stockholm on 2 June 1920]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1920. 14 pp. LCCN ????

Planck:1920:GPS

[Pla20d] Max Planck. The genesis and present state of development of the quantum theory. Nobel Lecture, June 2, 1920, June

2, 1920. URL http://www.nobelprize.org/nobel_prizes/
physics/laureates/1918/planck-lecture.html. The Nobel

Prize in Physics for 1918 was awarded to Max Karl Ernst Ludwig Planck “in recognition of the services he rendered to the advancement of Physics by his discovery of energy quanta”.

Planck:1921:AEC

[Pla21a] Max Planck. Absolute Entropie und chemische Konstante. (German) [Absolute entropy and chemical constant]. Annalen der Physik (Berlin), 371(21):365–372, ???? 1921. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1921:PEE

[Pla21b] Max Planck. Das Prinzip der Erhaltung der Energie. (German) [The principle of conservation of energy]. Teubner, Stuttgart, Germany; Leipzig, Germany, fourth edition, 1921. xvi + 278 pp. LCCN QC73 .P6.

Planck:1921:EAM

[Pla21c] Max Planck. Einführung in die allgemeine Mechanik: zum Gebrauch bei Vorträgen sowie zum Selbstunterricht. (German) [Introduction to General Mechanics: For use in presentations, as well as self teaching], volume 1 of Einführung in die theoretische Physik. Hirzel, Leipzig, Germany, third edition, 1921. 226 pp. LCCN ????

Planck:1921:HPQ

[Pla21d] Max Planck. Henri Poincaré und die Quantentheorie. (German) [Henri Poincaré and quantum theory]. Acta Mathematica, 38(1): 387–397, 1921. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Planck:1921:TT

[Pla21e] Max Planck. Treatise on Thermodynamics. Longmans, Green and Co., Ltd., London, UK, 1921. 272 pp. LCCN QC311 .P72 1921. Authorized English translation by Alexander Ogg of [Pla97f]. Reprint of [Pla03d].

Planck:1921:VTW

[Pla21f] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, fourth edition, 1921. x + 224 pp. LCCN QC331 .P7 1921.

Planck:1921:VTG

[Pla21g] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, sixth edition, 1921. viii + 292 pp. LCCN QC311 .P7 1921.

Planck:1922:AEC

[Pla22a] Max Planck. Absolute Entropie und chemische Konstante. (German) [Absolute entropy and chemical constant]. Annalen der Physik (Berlin), 371(21):365–372, 1922. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1922:AMP

[Pla22b] Max Planck. Ansprache von M. Planck zur öffentlichen Sitzung zur Feier des Leibnizischen Jahrestages. (German) [Speech by M. Planck at the public session in celebration of Leibniz’s anniversary]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–7, ???? 1922. CODEN SPWPAI. ISSN 0371-2435.

Planck:1922:EBE

[Pla22c] Max Planck. Die Entstehung und bisherige Entwicklung der Quantentheorie. (German) [The emergence and current development of quantum theory]. Nörstedt, Stockholm, Sweden, 1922. 1–14 pp. Nobel Prize in Physics lecture, 2 June 1920.

Planck:1922:EMD

[Pla22d] Max Planck. Einführung in die Mechanik deformierbarer Körper: Zum Gebrauch bei Vorträgen, sowie zum Selbstunterricht. (German) [Introduction to the Mechanics of Deformable Bodies: For use in presentations, as well as self teaching], volume 2 of Planck: Einführung in d. theoret. Physik. Hirzel, Leipzig, Germany, second edition, 1922. 3 + 193 + 1 pp. LCCN QA935 .P6

1922. URL http://catalog.hathitrust.org/api/volumes/
oclc/4703387.html.

Planck:1922:ETE

[Pla22e] Max Planck. Einführung in die Theorie der Elektrizität und des Magnetismus. (German) [Introduction to the theory of electricity and magnetism]. ????, Leipzig, Germany, 1922. ???? pp.

Planck:1922:EVM

[Pla22f] Max Planck. Eröffnungsrede des I. Vorsitzenden M. Planck. (German) [Opening speech of the First Chairman M. Planck]. Verhandlungen der Gesellschaft Deutscher Naturforscher und Ärzte, ??(??):1–9, ???? 1922. CODEN VGDNAN. ISSN 0172-0651.

Planck:1922:ODQ

[Pla22g] Max Planck. The Origin and Development of the Quantum Theory. Clarendon Press, New York, NY, USA, 1922. 23 + 1 pp. LCCN QC174 .P8. English translation by Hans Thacher Clarke and Ludwik Silberstein of the Nobel Prize address delivered before the Royal Swedish Academy of Sciences at Stockholm, 2 June, 1920.

Planck:1922:PRG

[Pla22h] Max Planck. Physikalische Rundblicke, gesammelte Reden und Aufsätze. (German) [A Survey of Physics; a Collection of Lectures and Essays]. S. Hirzel, Leipzig, Germany, 1922. 2 + 168 pp. LCCN 4QC 27.

Planck:1922:FEG

[Pla22i] Max Planck. Über die freie Energie von Gasmolekülen mit beliebiger Geschwindigkeitsverteilung. (German) [On the free energy of gas molecules with arbitrary velocity distribution]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 8(??):63–70, March 9, 1922. CODEN SPWPAI. ISSN 0371-2435.

Planck:1922:VTG

[Pla22j] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, seventh edition, 1922. viii + 292 pp. LCCN QC311 .P7 1922.

Planck:1923:GOQ

[Pla23a] Max Planck. ????. (German) [Observations on the quanta statistics of energy variations]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):355–358, ???? 1923. CODEN SPWPAI. ISSN 0371-2435.

Planck:1923:ASS

[Pla23b] Max Planck. Die Ableitung der Strahlungsgesetze: sieben Abhandlungen aus dem Gebiete der elektromagnetischen Strahlungstheorie. (German) [The derivation of the radiation law: seven treatises on the topic of electromagnetic radiation theory], volume 206 of Ostwalds Klassiker der exakten Wissenschaften. Akad. Verl.- Ges., Leipzig, Germany, 1923. 95 pp. LCCN ????

Planck:1923:BAG

[Pla23c] Max Planck. Die Bohrsche Atomtheorie. (German) [The Bohr atomic theory]. Die Naturwissenschaften, 11(27):535–537, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Planck:1923:ESP

[Pla23d] Max Planck. Die Energieschwankungen bei der Superposition periodischer Schwingungen. Bemerkung zur Quantenstatistik der Energieschwankungen. (German) [Energy fluctuations in the superposition of periodic vibrations. Remarks on the quantum statistics of power fluctuations]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):350–358, ???? 1923. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1923:EPR

[Pla23e] Max Planck. Erwiderung von Planck auf die Reden von Ludendorff, Guthenick, Johnsen. (German) [Reply by Planck to the speeches of Ludendorff, Guthenick, Johnsen]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–2, ???? 1923. CODEN SPWPAI. ISSN 0371-2435.

Planck:1923:GHR

[Pla23f] Max Planck. Gedächtnisrede auf Heinrich Rubens. (German) [Memorial speech for Heinrich Rubens]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–5, ???? 1923. CODEN SPWPAI. ISSN 0371-2435.

Planck:1923:KWV

[Pla23g] Max Planck. Kausalgesetz und Willensfreiheit. Vortrag i. d. preuß. Akad. Wiss. Berlin. (German) [Law of causality and free will. Lecture at the Prussian Academy of Sciences, Berlin]. ????, Berlin, Germany, 1923. 52 pp.

Planck:1923:NBP

[Pla23h] Max Planck. Neue Bahnen der physikalischen Erkenntnis: rede Gehalten beim antritt des Rektorats der Friedrich Wilhelm-Universitaet, Berlin, am 15. oktober, 1913. (German) [New paths of physical knowledge: speech held at the commencement of the Rector of the Friedrich Wilhelm University, Berlin, on 15 October 1913]. J. A. Barth, Leipzig, Germany, third edition, 1923. 27 pp. LCCN ????

Planck:1923:VTW

[Pla23i] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, fifth edition, 1923. ix + 221 pp. LCCN QC331 .P7 1923.

Planck:1924:APF

[Pla24a] Max Planck. Ansprache von Planck zur Feier des Jahrestages König Friedrichs II. (German) [Speech by Planck to celebrate the anniversary of King Frederick II]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–8, ???? 1924. CODEN SPWPAI. ISSN 0371-2435.

Planck:1924:NWG

[Pla24b] Max Planck. Über die Natur der Wärmestrahlung. (German) [On the nature of heat radiation]. Annalen der Physik (Berlin), 378 (3–4):272–288, ???? 1924. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1924:QBA

[Pla24c] Max Planck. Zur Quantenstatistik des Bohrschen Atommodells. (German) [On quantum statistics of the Bohr atomic model]. Annalen der Physik (Berlin), 380(23):673–684, ???? 1924. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1925:PKW

[Pla25a] Max Planck. Das Prinzip der kleinsten Wirkung. (German) [The principle of least action]. Die Kultur der Gegenwart, ??(??):772– 782, ???? 1925.

Planck:1925:NSD

[Pla25b] Max Planck. Eine neue statistische Definition der Entropie. (German) [A new statistical definition of entropy]. Zeitschrift für Physik, 35(3):155–169, March 1925. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/
10.1007/BF01379996.

Planck:1925:EPA

[Pla25c] Max Planck. Erwiderung von Planck auf die Antrittsrede von Hahn und Bodenstein. (German) [Reply by Planck to the inaugural address of Hahn and Bodenstein]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–2, ???? 1925. CODEN SPWPAI. ISSN 0371-2435.

Planck:1925:SPC

[Pla25d] Max Planck. A Survey of Physics; a Collection of Lectures and Essays. Methuen and Co. Ltd., London, UK, 1925. vii + 184 pp. LCCN QC71 .P72. English translation by R. Jones and D. H. Williams of [Pla22h].

Planck:1925:SEG

[Pla25e] Max Planck. Über die statistische Entropiedefinition. (German) [On statistical entropy]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):442–451, ???? 1925. CODEN SPWPAI. ISSN 0371-2435.

Planck:1925:VTZ

[Pla25f] Max Planck. Verhältnis der Theorien zueinander. (German) [Ratio of the theories to each other]. Die Kultur der Gegenwart, ?? (??):816–822, ???? 1925.

Planck:1925:RAG

[Pla25g] Max Planck. Vom Relativen zum Absoluten. (German) [From the relative to the absolute]. Die Naturwissenschaften, 13(3): 52–59, January 16, 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01559357.

Planck:1925:FQE

[Pla25h] Max Planck. Zur Frage der Quantelung einatomiger Gase. (German) [On the question of quantization of monatomic gases]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):49–57, ???? 1925. CODEN SPWPAI. ISSN 0371-2435.

Planck:1926:AMP

[Pla26a] Max Planck. Ansprache von M. Planck zum Leibnizischen Jahrestag. (German) [Speech by M. Planck on the Leibniz anniversary]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–6, ???? 1926. CODEN SPWPAI. ISSN 0371-2435.

Planck:1926:PGI

[Pla26b] Max Planck. Physikalische Gesetzlichkeit im Lichte neuerer Forschung. (German) [Physical laws in the light of new research]. Die Naturwissenschaften, 14(13):249–261, March 26, 1926. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01491761.

Planck:1926:TT

[Pla26c] Max Planck. Treatise on Thermodynamics. Dover Publications, Inc., New York, NY, USA, third edition, 1926. xiv + 1 + 297 pp. LCCN QC311 .P53 1926. Authorized English translation by Alexander Ogg of [Pla22j].

Planck:1926:BZH

[Pla26d] Max Planck. Über die Begründung des zweiten Hauptsatzes der Thermodynamik. (German) [On the justification of the Second Law of Thermodynamics]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):453–463, ???? 1926. CODEN SPWPAI. ISSN 0371-2435.

Plank:1926:FZV

[Pla26e] Max Plank. Fizicheskaja zakonomernost’ v svete novyh issledovanij. (Russian) [Physical laws in the light of new research]. Uspekhi Fizicheskikh Nauk, 6(3):177–199, March 1926. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1926/3/a/.

Planck:1927:DVL

[Pla27a] Max Planck. Dankworte bei Verleihung der Lorentz-Medaille am 28 Mai 1927. (German) [Words of thanks at the Lorentz Medal ceremony on 28 May 1927]. Königliche Akademie der Wissenschaften. Amsterdam, ??(??):532–538, ???? 1927.

Planck:1927:PRLa

[Pla27b] Max Planck. Die physikalische Realität der Lichtquanten: Mitteilung an das Franklin-Institut in Philadelphia, vorgelegt am 18.

Mai 1927. (German) [The physical reality of photons announcement from the Franklin Institute in Philadelphia presented on 18 May 1927]. Die Naturwissenschaften, 15(26):529–531, June 30, 1927. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF01504809.

Planck:1927:PRLb

[Pla27c] Max Planck. The physical reality of light-quanta. Journal of The Franklin Institute, 204(1):13–18, 1927. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0016003227913167. Translation from German by T. D. Cope.

Planck:1927:TT

[Pla27d] Max Planck. Treatise on Thermodynamics. Longmans, Green and Co., Ltd., London, UK, third edition, 1927. xiv + 297 + 1 pp. LCCN QC311 .P72 1927. Authorized English translation by Alexander Ogg of [Pla22j].

Planck:1927:PVL

[Pla27e] Max Planck. Über die Potentialdifferenz verdünnter Lösungen. (German) [On the potential difference in dilute solutions]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):285–288, ???? 1927. CODEN SPWPAI. ISSN 0371-2435.

Planck:1927:VTG

[Pla27f] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, eighth edition, 1927. ix + 287 pp. LCCN ????

Planck:1928:AFJ

[Pla28a] Max Planck. Ansprache zur Feier des Jahrestages König Friedrichs II. (German) [Speech to celebrate the anniversary of King Frederick II]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??): 1–8, ???? 1928. CODEN SPWPAI. ISSN 0371-2435.

Planck:1928:EAM

[Pla28b] Max Planck. Einführung in die allgemeine Mechanik: zum Gebrauch bei Vorträgen sowie zum Selbstunterricht. (German) [Introduction to General Mechanics: For use in presentations, as well as self teaching], volume 1 of Einführung in die theoretische Physik. Hirzel, Leipzig, Germany, 1928. iv + 218 pp. LCCN ????

Planck:1928:HAL

[Pla28c] Max Planck. Hendrik Antoon Lorentz: Gedächtnisrede, gehalten in der Physikalischen Gesellschaft zu Berlin am 29. Juni 1928. (German) [Hendrik Antoon Lorentz. commemoration speech, held in the Physical Society of Berlin on the 29th of June 1928]. Die Naturwissenschaften, 16(28):549–555, July 13, 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01506805.

Planck:1928:AWE

[Pla28d] Max Planck. Über die Abhandlungen zur Wellenmechanik von Erwin Schrödinger. (German) [On the essays on wave mechanics by Erwin Schrödinger]. Deutsche Literaturzeitung, ??(??):59–62, ???? 1928. ISSN 0012-043X.

Planck:1929:GPD

[Pla29a] Max Planck. ????. (German) [The potential difference of diluted solutions. (Second communication)]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):9–11, ???? 1929. CODEN SPW-PAI. ISSN 0371-2435.

Planck:1929:AMP

[Pla29b] Max Planck. Adresse an M. Planck zum fünfzigjährigen Doktorjubiläum am 28. Juni 1929. (German) [Address to M. Planck on the fifty-years anniversary of his doctorate on 28 June 1929]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):??, ???? 1929. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1929:AHS

[Pla29c] Max Planck. Antrittsrede des Herrn Schrödinger, Erwiderung der Sekretars Herrn Planck. Offentliche Sitzung vom. 4 Juli 1929. (German) [Inaugural speech of Mr. Schrödinger, reply from Secretary Mr. Planck. Public meeting on 4 July 1929]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):100–109, ???? 1929. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1929:NPGa

[Pla29d] Max Planck. Aus der neuen Physik. (German) [On the new physics]. Forschungen und Fortschritte: Nachrichtenblatt der deutschen Wissenschaft und Technik, 5(??):16–17, ???? 1929. CODEN FOFOAI. ISSN 0367-2794.

Planck:1929:NPVa

[Pla29e] Max Planck. Aus der neuen Physik: Vortrag gehalten in der öffentlichen Kundgebung der Notgemeinschaft der deutschen Wissenschaft in Dresden am 2 Dezember 1928. (German) [On the new physics: Lecture presented at the public rally of the Emergency Association of German Science in Dresden on 2 December 1928]. Deutsche Forschung [aus der Arbeit der Notgemeinschaft der deutschen Wissenschaft], 7(??):5–11, ???? 1929.

Planck:1929:NPVb

[Pla29f] Max Planck. Aus der neuen Physik: Vortrag gehalten in der öffentlichen Kundgebung der Notgemeinschaft der deutschen Wissenschaft in Dresden am 2 Dezember 1928. (German) [On the new physics: Lecture presented at the public rally of the Emergency Association of German Science in Dresden on 2 December 1928]. Zeitschrift für Angewandte Chemie, 42(??):84–??, ???? 1929. ISSN 0932-2132.

Planck:1929:NPGb

[Pla29g] Max Planck. Aus der neuen Physik. Vortrag über die Krisis der physikalische Weltanschauung. (German) [On the new physics. Lecture on the crisis of the physical world view]. Zeitschrift des Vereines deutscher Ingenieure, 73(??):353–355, March 16, 1929. CODEN VDIZAT. ISSN 0372-543X.

Planck:1929:WNPa

[Pla29h] Max Planck. Das Weltbild der neuen Physik. Gastvortrag in Wien. (German) [The worldview of the new physics. Guest lecture in Vienna]. Monatshefte für Mathematik und Physik, 36(1):387– 410, 1929. CODEN MMPHA8. ISSN 0026-9255 (print), 1436-5081 (electronic).

Planck:1929:WNPb

[Pla29i] Max Planck. Das Weltbild der neuen Physik. (German) [The world view of the new physics]. Forschungen und Fortschritte: Nachrichtenblatt der deutschen Wissenschaft und Technik, 5(??): 210–211, ???? 1929. CODEN FOFOAI. ISSN 0367-2794.

Planck:1929:WNPc

[Pla29j] Max Planck. Das Weltbild der neuen Physik. (German) [The worldview of the new physics]. J. A. Barth, Leipzig, Germany, 1929. 52 pp.

Planck:1929:MPFa

[Pla29k] Max Planck. Max Planck zur Feier seines goldenen Doktorjubiläums am 28 Juni 1929. (German) [Max Planck to celebrate his golden doctor jubilee on 28 June 1929]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):341–342, ???? 1929. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1929:MPFb

[Pla29l] Max Planck. Max Planck zur Feier seines goldenen Doktorjubiläums. (German) [Max Planck to celebrate his golden doctor jubilee]. Die Naturwissenschaften, 17(26):482–529, June 28, 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Planck:1929:MLO

[Pla29m] Max Planck. Max von Laue zum 9. Oktober 1929. Die Naturwissenschaften, 17(41):787–788, October 11, 1929. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01506652.

Planck:1929:PVL

[Pla29n] Max Planck. Über die Potentialdifferenz verdünnter Lösungen, II. Mitt.. (German) [On the potential difference of dilute solutions, second communication]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–5, ???? 1929. CODEN SPWPAI. ISSN 0371-2435.

Plank:1929:KMS

[Pla29o] Max Plank. Kartina mira sovremennoj fizik. (Russian) [The world view of modern physics]. Uspekhi Fizicheskikh Nauk, 9(4):407–436, April 1929. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1929/4/b/.

Planck:1922:ETP

[Pla30a] Max Planck. Einführung in die theoretische Physik. (German) [Introduction to Theoretical Physics]. S. Hirzel, Leipzig, Germany, 1922–1930. ???? pp. LCCN QC20 .P6. English translation by Henry L. (Henry Leopold) Brose.

Planck:1929:EAH

[Pla30b] Max Planck. Erinnerungen an Anna von Helmholtz. ,,Vom Schreibtisch und aus der Werkstatt”. (German) [Memories of Anna von Helmholtz. “from the desk and from the workshop”]. Velhagen und Klasings Monatshefte, 44(??):37–39, ???? 1929– 1930. ISSN 0176-9014.

Planck:1930:APO

[Pla30c] Max Planck. Ansprache von Planck in der öffentlichen Sitzung zum Leibnizischen Jahrestag. (German) [Speech by Planck at the meeting on the Leibniz anniversary]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–5, ???? 1930. CODEN SPWPAI. ISSN 0371-2435.

Planck:1930:WNPa

[Pla30d] Max Planck. Das Weltbild der neuen Physik. (German) [The worldview of the new physics]. ????, ????, third edition, 1930. ?? pp.

Planck:1930:WNPb

[Pla30e] Max Planck. Das Weltbild der neuen Physik. (German) [The worldview of the new physics]. ????, ????, fourth edition, 1930. ?? pp.

Planck:1930:TPF

[Pla30f] Max Planck. Theoretische Physik. Festschrift aus fünfzig Jahren deutscher Wissenschaft. Friedrich Schmidt-Ott dargebracht. (German) [Theoretical Physics. Festschrift of fifty years of German science. Offered by Friedrich Schmidt-Ott]. ????, ????, 1930. ?? pp.

Planck:1930:GVE

[Pla30g] Max Planck. Über die Grenzschicht verdünnter Elektrolyte. (German) [On the boundary layer of depleted electrolytes]. Sitzungsberichte der Preußischen Akademie der Wissenschaften:

Physikalisch-Mathematische Klasse, ??(??):1–9, ???? 1930. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1930:VTG

[Pla30h] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, ninth edition, 1930. x + 287 pp. LCCN QC311 .P58 1930.

Planck:1930:WGG

[Pla30i] Max Planck. Wissenschaft und Glaube. (German) [Science and faith]. Berliner Tageblatt, ??(607):??, December 25, 1930.

Planck:1931:GBL

[Pla31a] Max Planck. ????. (German) [The boundary layer of the depleted electrolyte]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):113–122, ???? 1931. CODEN SPWPAI. ISSN 0371-2435.

Planck:1931:EMD

[Pla31b] Max Planck. Einführung in die Mechanik deformierbarer Körper: Zum Gebrauch bei Vorträgen, sowie zum Selbstunterricht. (German) [Introduction to the Mechanics of Deformable Bodies: For use in presentations, as well as self teaching], volume 2 of Einführung in die Theoretische Physik / Max Planck. S. Hirzel, Leipzig, Germany, third edition, 1931. 193 pp. LCCN QC355 .P53x 1931.

Planck:1931:JCM

[Pla31c] Max Planck. James Clerk Maxwell in seiner Bedeutung für die theoretische Physik in Deutschland. (German) [James Clerk Maxwell and his importance to theoretical physics in Germany]. Die Naturwissenschaften, 19(44):889–894, October 30, 1931. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01516331.

Planck:1931:MIT

[Pla31d] Max Planck. Maxwell’s influence on theoretical physics in Germany. In Thomson [Tho31], page ?? LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Planck:1931:PRA

[Pla31e] Max Planck. Positivismus und reale Außenwelt: Vortrag, gehalten am 12. Nov. 1930 im Harnack-Haus der Kaiser Wilhelm-Gesellschaft zur Förderung der Wissenschaften. (German) [Positivism and real external world: lecture held, on 12 November 1930 Harnack House of the Kaiser Wilhelm Society for the Advancement of Sciences]. Akademische Verlagsgesellschaft M.B.H., Leipzig, Germany, 1931. 35 pp. LCCN ????

Planck:1931:GVE

[Pla31f] Max Planck. Über die Grenzschicht verdünnter Elektrolyte. (II. Mitt.). (German) [On the boundary layer of depleted electrolytes. (Second communication)]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–12, ???? 1931. CODEN SPWPAI. ISSN 0371-2435.

Planck:1931:ULM

[Pla31g] Max Planck. The Universe in the Light of Modern Physics. George Allen and Unwin, Ltd., London, UK, 1931. 110 pp. LCCN QC71 .P732.

Planck:1932:GCN

[Pla32a] Max Planck. ????. (German) [The causal notions in physics]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):18–19, ???? 1932. CODEN SPWPAI. ISSN 0371-2435.

Planck:1932:AGO

[Pla32b] Max Planck. Ansprache, gehalten in der öffentlichen Sitzung am 21. Januar 1932. (German) [Speech held at the public meeting on 21 January 1932]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??): 1–8, ???? 1932. CODEN SPWPAI. ISSN 0371-2435.

Planck:1932:CC

[Pla32c] Max Planck. The concept of causality. Proceedings of the Physical Society, London, 44(5):529–539, September 1, 1932. CODEN PPSOAU. ISSN 0370-1328. URL http://iopscience.iop.

org/article/10.1088/0959-5309/44/5/301/meta;   http:/
/stacks.iop.org/0959-5309/44/i=5/a=301. Guthrie lecture.

Planck:1932:KPGa

[Pla32d] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–2, ???? 1932. CODEN SPWPAI. ISSN 0371-2435.

Planck:1932:KPGb

[Pla32e] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, 1932. 26 pp. LCCN QC6 .P62.

Planck:1932:ITP

[Pla32f] Max Planck. Introduction to theoretical physics. ????, London, UK, 1932. ???? pp.

Planck:1932:MDB

[Pla32g] Max Planck. The Mechanics of Deformable Bodies, being Volume II of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1932. 4 + 234 + 1 pp. LCCN QC20 .P62 1932 vol. 2. English translation by Henry L. (Henry Leopold) Brose.

Planck:1932:TEM

[Pla32h] Max Planck. Theory of electricity and magnetism: being volume III of “Introduction to theoretical physics;”. MacMillan Publishing Company, New York, NY, USA, second edition, 1932. xii + 247 + 1 pp. LCCN ???? English translation by Henry L. (Henry Leopold) Brose.

Planck:1932:THB

[Pla32i] Max Planck. Theory of Heat, being Volume V of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1932. viii + 301 pp. LCCN QC20 .P62 1932 vol.5. English translation by Henry L. (Henry Leopold) Brose.

Planck:1932:TLB

[Pla32j] Max Planck. Theory of Light. being Volume IV of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1932. vi + 1 + 216 pp. LCCN QC20

.P62 1932 vol.4. URL http://catalog.hathitrust.org/api/

volumes/oclc/958604.html. English translation by Henry L. (Henry Leopold) Brose.

Planck:1932:WSG

[Pla32k] Max Planck. Where is Science Going? W. W. Norton & Co., New York, NY, USA, 1932. 221 pp. LCCN Q175 .P57. Prologue by Albert Einstein. Translation and biographical note by James Murphy.

Planck:1933:GOI

[Pla33a] Max Planck. ????. (German) [Origin and impact of scientific ideas]. Zeitschrift des Vereines deutscher Ingenieure, 77(??):185– 190, ???? 1933. CODEN VDIZAT. ISSN 0372-543X.

Planck:1933:GBL

[Pla33b] Max Planck. ????. (German) [The boundary layer of thinned electrolytes]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):362–368, ???? 1933. CODEN SPWPAI. ISSN 0371-2435.

Planck:1933:KIN

[Pla33c] Max Planck. Die Kausalität im Naturgeschehen. (German) [Causality in natural events]. Scientia (Milan), 53(??):153–164, ???? 1933. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic).

Planck:1933:GMB

[Pla33d] Max Planck. General mechanics, being volume I of “Introduction to Theoretical Physics”, volume 1. MacMillan Publishing Company, New York, NY, USA, 1933. ix + 272 pp. LCCN ???? English translation by Henry L. Brose.

Planck:1933:GVE

[Pla33e] Max Planck. Über die Grenzschicht verdünnter Elektrolyte. (III. Mitt.). (German) [On the boundary layer of depleted electrolytes. (Third communication)]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):3–9, ???? 1933. CODEN SPWPAI. ISSN 0371-2435.

Planck:1933:UWI

[Pla33f] Max Planck. Überproduktion von wissenschaftlichen Ideen. (German) [Overproduction of scientific ideas]. Berliner Tageblatt, ?? (281):??, June 18, 1933.

Planck:1933:UAW

[Pla33g] Max Planck. Ursprung und Auswirkung wissenschaftlicher Ideen: [Vorgetragen in der Sitzung des Vorstandes und des Wiss. Beirates des VDI 1933]. (German) [Origin and impact of scientific ideas: [Presented at the meeting of the VDI Management and Scientific Advisory Board 1933]]. VDI-Verlag, Berlin, Germany, 1933. 6 pp. Reprint from Zeitschrift des Vereines deutscher Ingenieure, 77(8) (1933).

Planck:1933:WPE

[Pla33h] Max Planck. Wege zur physikalischen Erkenntnis : Reden und Vorträge. (German) [Where is science going? Speeches and lectures]. S. Hirzel, Leipzig, Germany, fourth edition, 1933. ix + 280 pp. LCCN ????

Planck:1933:WSG

[Pla33i] Max Planck. Where is Science Going? George Allen and Unwin, London, UK, 1933. 224 pp. LCCN Q175 .P57. Prologue by Albert Einstein. Translation and biographical note by James Murphy.

Planck:1933:GHG

[Pla33j] Max Planck. Zum Geleit des Heftes ,,Der Geist meistert den Stoff”. (German) [Preface of the booklet “The spirit masters the material”], volume 6 of Forschung tut not, pages 1–2. ????, Berlin, Germany, 1933.

Planck:1934:PCBa

[Pla34a] Max Planck. Das Prinzip von Le Chatelier und Braun. (German) [The law of Le Chatelier and Braun]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):79–83, ???? 1934. CODEN SP-WPAI. ISSN 0371-2435.

Planck:1934:PCBb

[Pla34b] Max Planck. Das Prinzip von Le Chatelier und Braun. (German) [The law of Le Chatelier and Braun]. Annalen der Physik (Berlin), 411(7):759–768, ???? 1934. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19344110706/abstract.

Planck:1934:PCBc

[Pla34c] Max Planck. Das Prinzip von Le Chatelier und Braun (Nachtrag). (German) [The law of Le Chatelier and Braun (addendum)]. Annalen der Physik (Berlin), 412(2):196, ???? 1934. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19344120210/abstract.

Planck:1934:GGF

[Pla34d] Max Planck. Geleitwort. (German) [Foreword]. In Umsturz im Weltbild der Physik. (German) [Upheaval in the world of physics] [Zim34], page ?? LCCN QC7 .Z48 1934. Mit einem Geleitwort von Dr. Max Planck und 58 Abbildungen [With a foreword by Dr. Max Planck and 58 pictures].

Planck:1934:WPE

[Pla34e] Max Planck. Wege zur physikalischen Erkenntnis. (German) [Where is science going?]. S. Hirzel, Leipzig, Germany, second edition, 1934. ix + ?? pp. LCCN QC3 .P65 1934.

Planck:1935:AMP

[Pla35a] Max Planck. Ansprache von M. Planck, gehalten in der öffentlichen Sitzung am 27. Juni 1935. (German) [Speech by M. Planck, held at the public meeting on 27 June 1935]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–6, ???? 1935. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1935:BQI

[Pla35b] Max Planck. Bemerkungen über Quantitätsparameter, Intensitätsparameter und stabiles Gleichgewicht. (German) [Comments on quantity parameters, intensity parameters, and stable balance]. Physica, 2(??):1029–1032, ???? 1935. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Planck:1935:PIKa

[Pla35c] Max Planck. Die Physik im Kampf um die Weltanschauung; Vortrag gehalten am 6. März 1935 im Harnack-Haus Berlin-Dahlem. (German) [Physics in the battle for belief; lecture held on 6 March 1935 in Harnack House, Berlin-Dahlem]. J. A. Barth, Leipzig, Germany, 1935. 32 pp. LCCN QC6 .P63.

Planck:1935:PIKb

[Pla35d] Max Planck. Die Physik im Kampf um die Weltanschauung; Vortrag gehalten am 6. März 1935 im Harnack-Haus Berlin-Dahlem. (German) [Physics in the battle for belief; lecture held on 6 March 1935 in Harnack House, Berlin-Dahlem]. J. A. Barth, Leipzig, Germany, second edition, 1935. 32 pp. LCCN QC6 .P63.

Planck:1935:PEV

[Pla35e] Max Planck. Persönliche Erinnerungen. Vortrag gehalten in der Festsitzung zur Feier des 90 jährigen Bestehens der Physikalischen Gesellschaft zu Berlin. (German) [Personal memories. Lecture held at the ceremonial meeting in honor of the 90 anniversary of the Berlin Physical Society]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 13(3):11–16, ???? 1935. CODEN VDPEAZ. ISSN 0372-5448.

Planck:1935:TDE

[Pla35f] Max Planck. Zur Theorie der Diffusion von Elektrolyten. (German) [On the theory of diffusion of electrolytes]. Zeitschrift für Physik, 93(9–10):696, September 1935. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/
10.1007/BF01330544.

Planck:1935:TEEa

[Pla35g] Max Planck. Zur Theorie der Elektrizitätserregung in Elektrolyten. (German) [On the theory of excitation of electricity in electrolytes]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–3, ???? 1935. CODEN SPWPAI. ISSN 0371-2435.

Planck:1935:TEEb

[Pla35h] Max Planck. Zur Theorie der Elektrizitätserregung in Elektrolyten. (German) [On the theory of excitation of electricity in electrolytes]. Zeitschrift für Physik, 94(7–8):469, July 1935. CO-DEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.

com/article/10.1007/BF01330612.

Planck:1936:WWG

[Pla36a] Max Planck. Das Wesen der Willensfreiheit. (German) [The nature of free will]. Forschungen und Fortschritte: Nachrichtenblatt der deutschen Wissenschaft und Technik, 12(??):213–214, ???? 1936. CODEN FOFOAI. ISSN 0367-2794.

Planck:1936:PP

[Pla36b] Max Planck. The Philosophy of Physics. W. W. Norton & Co., New York, NY, USA, 1936. 128 pp. LCCN ???? English translation by H. W. Johnston of four essays: Physik im Kampfe um die Weltanschauung, Kausalität in der Natur, Ursprung und Auswirkung wissenschaftlicher Ideen, and Wissenschaft und Glaube.

Planck:1936:WWV

[Pla36c] Max Planck. Vom Wesen der Willensfreiheit, Vortrag gehalten zuerst in der Deutschen philosophischen Gesellschaft zu Berlin am 21. februar 1936. (German) [The nature of free will, lecture held first in the German Philosophical Society Berlin on 21 February 1936]. J. A. Barth, Leipzig, Germany, 1936. 30 pp. LCCN BJ1463 .P5 1936.

Planck:1937:JKW

[Pla37a] Max Planck. 25 Jahre Kaiser Wilhelm-Gesellschaft: Zur Förderung der Wissenschaften. Dritter Band. Die Geisteswissenschaften. (German) [25 years of the Kaiser Wilhelm Society for the Advancement of Science. Third volume. The humanities]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1937. ISBN 3-642-93843-4, 3-642-94243-1. viii

+ 398 pp. LCCN ???? URL http://link.springer.com/
openurl?genre=book&isbn=978-3-642-93843-6.

Planck:1937:AJK

[Pla37b] Max Planck. Ansprache zum Jahrestag König Friedrichs II. (German) [Speech on the anniversary of King Frederick II]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–7, ???? 1937. CO-DEN SPWPAI. ISSN 0371-2435.

Planck:1937:KPG

[Pla37c] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, second edition, 1937. 26 pp. LCCN QC6.4.C3 P56 1937.

Planck:1937:PIK

[Pla37d] Max Planck. Die Physik im Kampf um die Weltanschauung; Vortrag gehalten am 6. März 1935 im Harnack-Haus Berlin-Dahlem. (German) [Physics in the battle for belief; lecture held on 6 March 1935 in Harnack House, Berlin-Dahlem]. J. A. Barth, Leipzig, Germany, third edition, 1937. 32 pp. LCCN QC6 .P63.

Planck:1937:EAM

[Pla37e] Max Planck. Einführung in die allgemeine Mechanik. (German) [Introduction to general mechanics]. ????, Leipzig, Germany, fifth edition, 1937. vii + 226 pp. LCCN ????

Planck:1937:EPA

[Pla37f] Max Planck. Erwiderung von Planck auf die Antrittsreden von Debye und Geiger. (German) [Reply by Planck to the inaugural speeches of Debye and Geiger]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):1–2, ???? 1937. CODEN SPWPAI. ISSN 0371-2435.

Planck:1937:ULM

[Pla37g] Max Planck. The Universe in the Light of Modern Physics. George Allen and Unwin, Ltd., London, UK, second edition, 1937. 140 pp. LCCN QC71 .P732 1937.

Planck:1937:WWG

[Pla37h] Max Planck. Vom Wesen der Willensfreiheit. (German) [The nature of free will]. J. A. Barth, Leipzig, Germany, second edition, 1937. 34 pp. LCCN BJ1463 .P5 1937.

Planck:1937:JJE

[Pla37i] Max Planck. Zum 25-jährigen Jubiläum der Entdeckung von W. Friedrich, P. Knipping und M. v. Laue (Sonderdruck). (German) [On the 25th anniversary of the discovery of W. Friedrich, P. Knipping and M. v. Laue (reprint)]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 18(??):77–80, ???? 1937. CODEN VDPEAZ. ISSN 0372-5448.

Planck:1938:AMP

[Pla38a] Max Planck. Adresse an M. Planck zum- achtzigsten Geburtstage am 23. April 1938. (German) [Address to M. Planck on his eightieth birthday on 23 April 1938]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):2, ???? 1938. CODEN SPWPAI. ISSN 0371-2435.

Planck:1938:ASG

[Pla38b] Max Planck. Arnold Sommerfeld zum 70. Geburtstag. (German) [Arnold Sommerfeld’s 70th birthday]. In Max Planck, editor, Physikalische Abhandlungen und Vorträge. (German) [Physical Discourses and Lectures], volume 3, pages 368–371. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1938.

Planck:1938:ASS

[Pla38c] Max Planck. Arnold Sommerfeld zum siebzigsten Geburtstag. (German) [Arnold Sommerfeld on his seventieth birthday]. Die Naturwissenschaften, 26(48):777–779, December 2, 1938. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01774337.

Planck:1938:WNG

[Pla38d] Max Planck. Das Wunder der Naturgesetzlichkeit. (German) [The wonders of the natural law]. In Eberhard Dennert, editor, Die Natur, das Wunder Gottes. (German) [Nature, the miracle of God], page ?? M. Warneck, Leipzig, Germany, 1938.

Planck:1938:DIG

[Pla38e] Max Planck. Determinismus oder indeterminismus?. (German) [Determinism or indeterminism?]. J. A. Barth, Leipzig, Germany, 1938. 32 pp. LCCN Q175 .P565.

Planck:1938:EPA

[Pla38f] Max Planck. Erwiderung von Planck auf die Antrittsreden von Kraft, Nägel und Vahlen. (German) [Reply by Planck to the inaugural speeches of Kraft, Nagel and Vahlen]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):2, ???? 1938. CODEN SPWPAI. ISSN 0371-2435.

Planck:1938:FGE

[Pla38g] Max Planck. Feier des 80. Geburtstages des Ehrenmitglieds der Dtsch. physik. Ges. M. Planck. Sonnabend, d. 23. April 1938, im Harnack-Haus. (German) [Celebration of the 80th birthday of honorary member of the German Physical Society M. Planck. Saturday, 23 April, 1938, in Harnack House]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19 (reihe 3)(??):58– 76, ???? 1938. CODEN VDPEAZ. ISSN 0372-5448.

Planck:1938:RNV

[Pla38h] Max Planck. Religion und Naturwissenschaft. Vortrag, gehalten im Mai 1937. (German) [Religion and science. Lecture held in May 1937]. ????, Leipzig, Germany, 1938. 30 pp.

Planck:1939:E

[Pla39a] Max Planck. Électromagnétisme. F. Alcan, Paris, France, 1939. xiv + 448 pp. LCCN ???? Translation to French by Éduard Labin.

Planck:1939:MLO

[Pla39b] Max Planck. Max von Laue zum 9. Oktober 1939. Die Naturwissenschaften, 27(40):665–666, October 6, 1939. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01494989.

Planck:1939:WWN

[Pla39c] Max Planck. Vom Wesen der Willensfreiheit nach einem Vortrag in der Ortsgruppe Leipzig der Deutschen Philosophischen Gesellschaft am 27. November 1936. (German) [The nature of free will: from a lecture in the local Leipzig chapter of the German Philosophical Society on 27 November 1936]. J. A. Barth, Leipzig, Germany, third edition, 1939. 34 pp. LCCN BJ1463 .P5 1939.

Planck:1940:RAN

[Pla40a] Max Planck. RETRACTED ARTICLE: Naturwissenschaft und reale Außenwelt. (German) [Natural science and real outside world]. Die Naturwissenschaften, 28(50):778–779, December 13, 1940. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/

10.1007/BF01488952. This article has been withdrawn due to copyright violation.

Planck:1940:VSZa

[Pla40b] Max Planck. Versuch einer Synthese zwischen Wellenmechanik und Korpuskularmechanik. (German) [Attempt at a synthesis between wave mechanics and corpuscular mechanics]. Annalen der Physik (Berlin), 429(4):261–277, ???? 1940. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1940:VSZb

[Pla40c] Max Planck. Versuch einer Synthese zwischen Wellenmechanik und Korpuskularmechanik (Nachtrag). (German) [Attempt at a synthesis between wave mechanics and corpuscular mechanics (addendum)]. Annalen der Physik (Berlin), 430(4):272–273, ???? 1940. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/andp.19404300404/abstract.

Planck:1941:KPG

[Pla41a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, third edition, 1941. 26 pp. LCCN QC6.4.C3 P56 1941.

Planck:1941:IPF

[Pla41b] Max Planck. Initiations á la physique. (French) [Introduction to physics]. Flammarion, Paris, France, 1941. 278 + 2 pp. LCCN QC3 .P654. French translation by Joachim Du Plessis de Grenédan of Wege zur Physikalischen Erkenntnis.

Planck:1941:VSZ

[Pla41c] Max Planck. Versuch einer Synthese zwischen Wellenmechanik und Korpuskularmechanik (Zweite Mitteilung). (German) [Attempt at a synthesis between wave mechanics and corpuscular mechanics. Second communication]. Annalen der Physik (Berlin), 432(7):481–492, ???? 1941. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1942:SEW

[Pla42a] Max Planck. Der Sinn der exakten Wissenschaften. (German) [The meaning of exact sciences]. Deutsche Allgemeine Zeitung, 27(463):??, September 27, 1942.

Planck:1942:CDM

[Pla42b] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world], volume VIII of Biblioteca di cultura scientifica. bicilio Einaudi, Torino, Italia, second edition, 1942. 300 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1942:MPS

[Pla42c] Max Planck. Max Planck Selbstdarstellung im Filmportrait. (German) [Max Planck self-portrayal in the film]. Web site video, 1942.

Planck:1942:RAS

[Pla42d] Max Planck. RETRACTED ARTICLE: Sinn und Grenzen der exakten Wissenschaft. (German) [The signification and limits of exact science]. Die Naturwissenschaften, 30(9–10):125–133, February 27, 1942. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/

10.1007/BF01475382. This article has been withdrawn due to copyright violation.

Planck:1942:SGE

[Pla42e] Max Planck. Sinn und Grenzen der exakten Wissenschaft. Vortrag, gehalten in der Kaiser-Wilhelm-Gesellschaft. (German) [Meaning and limits of exact science. Lecture held in the Kaiser Wilhelm Society]. ????, Leipzig, 1942. 33 pp.

Planck:1942:WKW

[Pla42f] Max Planck. Warum kann Wissenschaft nicht populär sein?. (German) [Why can not science be popular?]. Das Reich: deutsche Wochenzeitung, ??(32):9, August 9, 1942. ISSN 0932-2868.

Planck:1943:CDM

[Pla43a] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world], volume 8 of Biblioteca di cultura scientifica / [Einaudi]. Einaudi, Torino, Italia, third edition, 1943. viii + 295 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1943:WPE

[Pla43b] Max Planck. Wege zur physikalischen Erkenntnis: Reden und Vorträge. (German) [Where is science going? Speeches and lectures]. S. Hirzel, Leipzig, Germany, third edition, 1943. vi + 238 pp. LCCN QC3 .P65 1943.

Planck:1943:GAP

[Pla43c] Max Planck. Zur Geschichte der Auffindung des physikalischen Wirkungsquantums. (German) [On the history of the discovery of the physical quantum of action]. Die Naturwissenschaften, 31(14– 15):153–159, April 2, 1943. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Planck:1944:WMG

[Pla44a] Max Planck. Das Wesen der Materie. (German) [The nature of matter]. Speech in Florence, Italy, recorded in Archiv zur Geschichte der Max-Planck-Gesellschaft, Abt. Va, Rep. 11 Planck, Nr. 1797, 1944.

Planck:1944:WPE

[Pla44b] Max Planck. Wege zur physikalischen Erkenntnis: Reden und Vorträge. (German) [Where is science going? Speeches and lectures]. S. Hirzel, Leipzig, Germany, fourth edition, 1944. vi + 1 + ?? pp. LCCN QC3 .P65 1944.

Planck:1945:PIKa

[Pla45a] Max Planck. Die Physik im Kampf um die Weltanschauung; Vortrag gehalten am 6. März 1935 im Harnack-Haus Berlin-Dahlem. (German) [Physics in the battle for belief; lecture held on 6 March 1935 in Harnack House, Berlin-Dahlem]. J. A. Barth, Leipzig, Germany, third edition, 1945. 32 pp. LCCN QC6 .P63.

Planck:1945:PIKb

[Pla45b] Max Planck. Die Physik im Kampf um die Weltanschauung; Vortrag gehalten am 6. März 1935 im Harnack-Haus Berlin-Dahlem. (German) [Physics in the battle for belief; lecture held on 6 March 1935 in Harnack House, Berlin-Dahlem]. J. A. Barth, Leipzig, Germany, fifth edition, 1945. 32 pp. LCCN QC6 .P63.

Planck:1945:TT

[Pla45c] Max Planck. Treatise on Thermodynamics. Dover Publications, Inc., New York, NY, USA, third edition, 1945. xiv + 1 + 297 pp. LCCN QC311 .P72 1945. English translation, with author’s permission, by Alexander Ogg of [Pla22j].

Planck:1945:WWG

[Pla45d] Max Planck. Vom Wesen der Willensfreiheit. (German) [The nature of free will]. J. A. Barth, Leipzig, Germany, fourth edition, 1945. 34 pp. LCCN BJ1463 .P5 1945.

Planck:1945:WPE

[Pla45e] Max Planck. Wege zur physikalischen Erkenntnis: Reden und Vorträge. (German) [Where is science going? Speeches and lectures]. S. Hirzel, Leipzig, Germany, fifth edition, 1945. ???? pp. LCCN QC3 .P65 1945.

Planck:1945:WSG

[Pla45f] Max Planck. Wissenschaftliche Streitfragen. (German) [Scientific issues]. Forschungen und Fortschritte: Nachrichtenblatt der deutschen Wissenschaft und Technik, 21(??):2, ???? 1945. CO-DEN FOFOAI. ISSN 0367-2794.

Planck:1946:MGW

[Pla46a] Max Planck. In memoriam: Gottfried Wilhelm Leibniz zur 300. Wiederkehr seines Geburtstags (1. Juli 1646). (German) [Gottfried Wilhelm Leibniz on the 300th anniversary of his birth (1 July 1646)]. Zeitschrift für Naturforschung, 1(5): 298–300, May 1946. CODEN ZNTFA2. ISSN 0372-9516.

URL http://www.degruyter.com/view/j/zna.1946.1.issue-
5/zna-1946-0511/zna-1946-0511.xml?format=INT.

Planck:1946:PEA

[Pla46b] Max Planck. Persönliche Erinnerungen aus alten Zeiten. (German) [Personal memories of old times]. Die Naturwissenschaften, 33(8):230–235, October 30, 1946. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.

springer.com/article/10.1007/BF01204476.

Planck:1946:SWG

[Pla46c] Max Planck. Scheinprobleme der Wissenschaft. (German) [Pseudo-problems of science]. Physikalische Blätter, 2(7):161–168, 1946. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1947:DIS

[Pla47a] Max Planck. ‘?Determinismo o indeterminismo?. (Spanish) [Determinism or indeterminism?]. Cuatro Pliegos, Madrid, Spain, 1947. 131 pp. LCCN ????

Planck:1947:MBA

[Pla47b] Max Planck. Mein Besuch bei Adolf Hitler. (German) [My visit to Adolf Hitler]. Physikalische Blätter, 3(5):143, 1947. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1948:KPG

[Pla48a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, fourth edition, 1948. 22 pp. LCCN QC6.4.C3 P56 1948.

Planck:1948:DIG

[Pla48b] Max Planck. Determinismus oder Indeterminismus?. (German) [Determinism or Indeterminism?]. J. A. Barth, Leipzig, East Germany, second edition, 1948. 29 pp. LCCN QC71 .P714 1948.

Planck:1948:MPW

[Pla48c] Max Planck. Max Planck, Wissenschaftliche Selbstbiographie mit e. Bildnis u. der von Max Laue gehaltenen Traueransprache. (German) [Max Planck, Scientific autobiography with a Portrait and a funeral address by Max [von] Laue], volume 5 of Lebensdarstellungen deutscher Naturforscher. J. A. Barth, Leipzig, East Germany, 1948. 33 pp. LCCN ????

Planck:1948:RNA

[Pla48d] Max Planck. Religion und Naturwissenschaft (Auszug). (German) [Religion and science (excerpt)]. Physikalische Blätter, 4(4):175– 178, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/phbl.19480040407/abstract. Special issue

on Max Planck.

Planck:1948:WWG

[Pla48e] Max Planck. Vom Wesen der Willensfreiheit. (German) [The nature of free will]. J. A. Barth, Leipzig, Germany, fifth edition, 1948. 30 pp. LCCN BJ1463 .P5 1948.

Planck:1948:MPS

[Pla48f] Max (Max Karl Ernst Ludwig) Planck. Max Planck in seinen Akademie-Ansprachen: Erinnerungsschrift der Deutschen Akademie der Wissenschaften zu Berlin. (German) [Max Planck’s Academy speeches: Memorial writings of the German Academy of Sciences in Berlin]. Akademie-Verlag, Berlin, Germany, 1948. 204 pp. LCCN ????

Planck:1949:GMB

[Pla49a] Max Planck. General mechanics, being volume I of “Introduction to Theoretical Physics”, volume 1. MacMillan Publishing Company, New York, NY, USA, 1949. ix + 272 pp. LCCN ???? English translation by Henry L. Brose.

Planck:1949:GVW

[Pla49b] Max Planck. Gibt es eine vernünftige Weltordnung?. (German) [Is there a rational world order?], volume 9 of Kleine Kostbarkeiten. Verlag “Die Fähre”, Bern, Switzerland, 1949.

13 pp. LCCN ???? URL http://permalink.snl.ch/bib/
sz000558836.

Planck:1949:CDM

[Pla49c] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world]. G. Einaudi, Torino, Italia, fourth edition, 1949. 295 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1949:MLE

[Pla49d] Max Planck. The meaning and limits of exact science. Science, 110(2857):319–327, September 30, 1949. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/110/2857/319.extract.

Planck:1949:MDB

[Pla49e] Max Planck. The Mechanics of Deformable Bodies, being Volume II of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1949. 234 pp. LCCN

QC20 .P625 vol.2. URL http://catalog.hathitrust.org/api/

volumes/oclc/996970.html. English translation by Henry L. (Henry Leopold) Brose.

Planck:1949:RNS

[Pla49f] Max Planck. Religion and natural science. In Scientific Autobiography and Other Papers. ????, New York, Ny, USA, 1949. URL

http://en.wikiquote.org/wiki/Max_Planck. Lecture given in

1937.

Planck:1949:SAO

[Pla49g] Max Planck. Scientific Autobiography, and Other Papers. Philosophical Library, New York, NY, USA, 1949. ISBN 1-4976-7588-X, 0-8065-3634-9. 192 pp. LCCN Q171 .P7.

Planck:1949:TEM

[Pla49h] Max Planck. Theory of electricity and magnetism, being volume III of Introduction to theoretical physics. MacMillan Publishing Company, New York, NY, USA, 1949. xii + 247 pp. LCCN QC20 .P625 vol.3. English translation by Henry L. Brose.

Planck:1949:THB

[Pla49i] Max Planck. Theory of Heat, being Volume V of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1949. 301 pp. LCCN QC20 .P625

v.5. URL http://catalog.hathitrust.org/api/volumes/

oclc/3705242.html. English translation by Henry L. (Henry Leopold) Brose.

Planck:1949:TLB

[Pla49j] Max Planck. Theory of Light, being Volume IV of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1949. vi + 216 pp. LCCN QC20 .P625 vol.4. English translation by Henry L. (Henry Leopold) Brose.

Planck:1949:VEG

[Pla49k] Max Planck. Vorträge und Erinnerungen. (German) [Lectures and memories]. S. Hirzel, Stuttgart, West Germany, fifth edition, 1949. vi + 380 pp. LCCN ????

Planck:1950:PGG

[Pla50a] Max Planck. Die Physikalische Gesellschaft um 1890. (German) [The Physical Society in 1890]. Physikalische Blätter, 6(10):433– 434, 1950. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1950:SAO

[Pla50b] Max Planck. Scientific Autobiography, and Other Papers. Williams and Norgate, London, UK, 1950. 192 pp. LCCN Q171 .P7 1950. With a memorial address on Max Planck by Max von Laue. Translated from German by Frank Gaynor.

Planck:1952:KPG

[Pla52a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, fifth edition, 1952. 22 pp. LCCN QC6.4.C3 P56 1952.

Planck:1952:DIG

[Pla52b] Max Planck. Determinismus oder Indeterminismus? (German) [Determinism or Indeterminism?]. J. A. Barth, Leipzig, East Germany, third edition, 1952. 29 pp. LCCN ????

Planck:1952:PIK

[Pla52c] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, sixth edition, 1952. 29 pp. LCCN QC6 .P63 1952.

Planck:1952:WWN

[Pla52d] Max Planck. Vom Wesen der Willensfreiheit nach einem Vortrag in der Ortsgruppe Leipzig der Deutschen Philosophischen Gesellschaft am 27. November 1936. (German) [The nature of free will: from a lecture in the local Leipzig chapter of the German Philosophical Society on 27 November 1936]. J. A. Barth, Leipzig, Germany, sixth edition, 1952. 31 pp. LCCN BJ1463 .P5 1952.

Planck:1953:KPG

[Pla53a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, sixth edition, 1953. 22 pp. LCCN QC6.4.C3 P56 1953.

Planck:1953:DIG

[Pla53b] Max Planck. Determinismus oder Indeterminismus?. (German) [Determinism or Indeterminism?]. Barth, Leipzig, East Germany, fourth edition, 1953. 29 pp. LCCN ????

Planck:1953:PIK

[Pla53c] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, seventh edition, 1953. 29 pp. LCCN QC6 .P63 1953.

Planck:1953:WWN

[Pla53d] Max Planck. Vom Wesen der Willensfreiheit nach einem Vortrag in der Ortsgruppe Leipzig der Deutschen Philosophischen Gesellschaft am 27. November 1936. (German) [The nature of free will: from a lecture in the local Leipzig chapter of the German Philosophical Society on 27 November 1936]. J. A. Barth, Leipzig, Germany, seventh edition, 1953. 30 pp. LCCN BJ1463 .P5 1953.

Planck:1954:CDMa

[Pla54a] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world]. G. Einaudi, Torino, Italia, fifth edition, 1954. 299 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1954:CDMb

[Pla54b] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world]. Boringhieri, Torino, Italia, ???? edition, 1954. 413 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1954:VTG

[Pla54c] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, tenth edition, 1954. xiii + 306 pp. LCCN QC311 .P7 1954. Revised and extended by Max von Laue.

Planck:1955:KPGa

[Pla55a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, sixth edition, 1955. 22 pp. LCCN QC6.4.C3 P56 1955.

Planck:1955:KPGb

[Pla55b] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, seventh edition, 1955. 22 pp. LCCN QC6.4.C3 P56 1955.

Planck:1955:DIG

[Pla55c] Max Planck. Determinismus oder Indeterminismus?. (German) [Determinism or Indeterminism?]. Barth, Leipzig, Germany, fifth edition, 1955. 29 pp. LCCN ????

Planck:1955:PIK

[Pla55d] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, eighth edition, 1955. 29 pp. LCCN QC6 .P63 1955.

Planck:1955:WWN

[Pla55e] Max Planck. Vom Wesen der Willensfreiheit nach einem Vortrag in der Ortsgruppe Leipzig der Deutschen Philosophischen Gesellschaft am 27. November 1936. (German) [The nature of free will: from a lecture in the local Leipzig chapter of the German Philosophical Society on 27 November 1936]. J. A. Barth, Leipzig, Germany, eighth edition, 1955. 30 pp. LCCN BJ1463 .P5 1955.

Planck:1957:MDB

[Pla57a] Max Planck. The Mechanics of Deformable Bodies: being Volume II of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1957. 234 pp. LCCN QC20 .P62 1932 vol.2. English translation by Henry L. (Henry Leopold) Brose.

Planck:1957:TEM

[Pla57b] Max Planck. Theory of electricity and magnetism: being volume III of “Introduction to theoretical physics”. MacMillan Publishing Company, New York, NY, USA, 1957. xii + 247 pp. LCCN QC20 .P62 1932 vol.3. English translation by Henry L. Brose.

Planck:1957:THB

[Pla57c] Max Planck. Theory of Heat. being Volume V of “Introduction to Theoretical Physics”. MacMillan Publishing Company, New York, NY, USA, 1957. 301 pp. LCCN QC20 .P625 v.

5, 1957. URL http://catalog.hathitrust.org/api/volumes/

oclc/11253083.html. English translation by Henry L. (Henry Leopold) Brose.

Planck:1957:TMP

[Pla57d] Max Planck. Zum 10. Todestag von Max Planck: Antrittsrede zur Aufnahme in die Akademie. (German) [On the 10th anniversary of the death of Max Planck: inaugural address to a reception in the Academy]. Physikalische Blätter, 13(10):433–435, 1957. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1958:WS

[Pla58a] Max Planck. Aus der wissenschaftlichen Selbstbiographie. (German) [From my scientific self-biography]. Physikalische Blätter, 14(4):145–152, 1958. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1958:KPG

[Pla58b] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, eighth edition, 1958. 22 pp. LCCN QC6.4.C3 P56 1958.

Planck:1958:DIG

[Pla58c] Max Planck. Determinismus oder Indeterminismus?. (German) [Determinism or Indeterminism?]. J. A. Barth, Leipzig, Germany, sixth edition, 1958. 29 pp. LCCN ????

Planck:1958:PIK

[Pla58d] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, ninth edition, 1958. 29 pp. LCCN QC6 .P63 1958.

Planck:1958:EBE

[Pla58e] Max Planck. Entstehung und bisherige Entwicklung der Quantentheorie. (German) [Emergence and current development of quantum theory]. Vieweg+Teubner Verlag, Wiesbaden, West Germany, 1958. ISBN 3-322-98792-2, 3-322-98129-0 (print). LCCN

???? URL http://link.springer.com/openurl?genre=book&
isbn=978-3-322-98129-5.

Planck:1958:PAV

[Pla58f] Max Planck. Physikalische Abhandlungen und Vorträge. (German) [Physical treatises and lectures]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1958. ???? + ???? + ???? pp. Three volumes.

Plank:1958:NAR

[Pla58g] Max Plank. Nauchnaja avtobiografija. (Russian) [Scientific autobiography]. Uspekhi Fizicheskikh Nauk, 64(4):625–637, 1958. CO-DEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

URL http://ufn.ru/ru/articles/1958/4/a/.

Planck:1959:NSC

[Pla59a] Max Planck. The new science. 3 complete works: Where is science going? The universe in the light of modern physics; The philosophy of physics. Greenwich editions. Meridian Books, New York, NY, USA, 1959. 328 pp. LCCN QC6 .P623. Translated from the German by James Murphy and W. H. Johnston.

Planck:1959:THR

[Pla59b] Max Planck. The Theory of Heat Radiation. Dover Publications, Inc., New York, NY, USA, 1959. xiv + 224 pp. LCCN QC331 .P73 1959. Authorised translation by Morton Masius.

Planck:1960:ASD

[Pla60a] Max Planck. Autobiographie scientifique et derniers écrits. (French) [Scientific autobiography and last writings], volume 247 of Champs/Flammarion. Flammarion, Paris, France, 1960. ISBN 2-08-081247-5. 220 pp. LCCN ???? Introduction, translation, and notes by André George.

Planck:1960:SPT

[Pla60b] Max Planck. A Survey of Physical Theory (formerly titled: A survey of physics). (Formerly titled: A survey of physics) Translated by R. Jones and D. H. Williams. Dover Publications, Inc., New York, NY, USA, 1960. vii + 117 pp. LCCN QC71 .P72 1960. English translation by R. Jones and D. H. Williams of [Pla22h].

Planck:1962:FJI

[Pla62] Max Planck. Fünfzig Jahre Interferenz der Röntgenstrahlen. (German) [Fifty years of X-ray interference]. Physikalische Blätter, 18(6):241–244, 1962. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Planck:1963:LDM

[Pla63a] Max Planck. L’image du monde dans la physique modern. (French) [The images of the world in modern physics]. Gronthier, Paris, France, 1963. 158 pp.

Planck:1963:PP

[Pla63b] Max Planck. The philosophy of physics. W. W. Norton & Co., New York, NY, USA, 1963. 128 pp. LCCN ???? English translation by H. W. Johnston.

Planck:1964:CDM

[Pla64a] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world], volume 4 of Universale scientifica / [Boringhieri]. Bollati Boringhieri, Torino, Italia, 1964. 413 pp. LCCN ???? Italian translation and editing by Enrico Persico.

Planck:1964:MDB

[Pla64b] Max Planck. The Mechanics of Deformable Bodies: being Volume II of “Introduction to Theoretical Physics”. Academic Press, New York, USA, 1964. 234 pp. LCCN QA935 .P5.

Planck:1965:KPG

[Pla65a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, ninth edition, 1965. 22 pp. LCCN QC6.4.C3 P56 1965.

Planck:1965:DIG

[Pla65b] Max Planck. Determinismus oder Indeterminismus?. (German) [Determinism or Indeterminism?]. J. A. Barth, Leipzig, East Germany, seventh edition, 1965. 29 pp. LCCN ????

Planck:1965:PIK

[Pla65c] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, tenth edition, 1965. 29 pp. LCCN QC6 .P63 1965.

Planck:1966:VTW

[Pla66] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. J. A. Barth, Leipzig, Germany, sixth edition, 1966. x + 221 pp. LCCN QC331 .P7 1966.

Planck:1967:KPG

[Pla67a] Max Planck. Der Kausalbegriff in der Physik. (German) [The concept of causality in physics]. J. A. Barth, Leipzig, Germany, tenth edition, 1967. 22 pp. LCCN QC6.4.C3 P56 1967.

Planck:1967:DIG

[Pla67b] Max Planck. Determinismus oder indeterminismus?. (German) [Determinism or indeterminism?]. J. A. Barth, Leipzig, East Germany, eighth edition, 1967. 29 pp. LCCN Q175 .P565 1967. 1.50.

Planck:1967:PIK

[Pla67c] Max Planck. Die Physik im kampf um die Weltanschauung. (German) [Physics in the battle for belief]. J. A. Barth, Leipzig, Germany, eleventh edition, 1967. 29 pp. LCCN QC6 .P63 1967.

Planck:1968:SAO

[Pla68] Max Planck. Scientific Autobiography, and Other Papers. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1968. 192 pp. LCCN Q171 .P7 1968. Reprint of [Pla49g]. With a memorial address on Max Planck by Max von Laue. Translated from German by Frank Gaynor.

Planck:1969:IWE

[Pla69a] Max Planck. On an improvement of Wien’s equation for the spectrum. In The Old Quantum Theory [tH67], pages 79–81. LCCN QC174 .H33 1967. URL http://www.ffn.

ub.es/luisnavarro/nuevo_maletin/Planck%20%281900%29,%
20Improvement%20of%20Wien%27s.pdf. English translation of

[Pla00g].

Planck:1969:TED

[Pla69b] Max Planck. On the theory of the energy distribution law of the normal spectrum. In The Old Quantum Theory [tH67], pages 82–90. LCCN QC174 .H33 1967. URL http://www.ffn.

ub.es/luisnavarro/nuevo_maletin/Planck%20%281900%29,%

20Distribution%20Law.pdf. English translation of [Pla00i].

Planck:1969:TT

[Pla69c] Max Planck. Treatise on Thermodynamics. Dover Publications, Inc., New York, NY, USA, third edition, 1969. ISBN 0-486-66371-X. xiv + 297 pp. LCCN QC311 .P53 1969. English translation, with author’s permission, by Alexander Ogg of [Pla22j].

Planck:1970:JFO

[Pla70] Max Planck. Jedność fizycznego obrazu świata: Wybór pism filozoficznych. (Polish) [The unity of the physical picture of the world: the choice of philosophical writings]. KiW, Warsawa, Poland, 1970. 273 pp. LCCN ???? Polish translation by Ryszard Kerner and Samuel Kerner.

Planck:1971:SAO

[Pla71] Max Planck. Scientific Autobiography, and Other Papers. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1971. ISBN 0-8371-0194-8. 192 pp. LCCN Q171 .P7 1971. Second Greenwood reprinting of [Pla49g]. With a memorial address on Max Planck by Max von Laue. Translated from German by Frank Gaynor.

Planck:1977:WSG

[Pla77] Max Planck. Where is Science Going? American Mathematical Society, Providence, RI, USA, 1977. ISBN 0-404-14696-1. 221 pp. LCCN Q175 .P57 1977. Prologue by Albert Einstein. English translation and biographical note by James Murphy.

Planck:1978:GEI

[Pla78a] Max Planck. Über das Gesetz der Energieverteilung im Normalspektrum. (German) [On the law of distribution of energy in the normal spectrum]. In Schöpf [Sch78c], pages 178–190. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Planck:1978:EME

[Pla78b] Max Planck. Über die Elementarquanta der Materie und der Elektrizität. (German) [On elementary quanta of matter and electricity]. In Schöpf [Sch78c], pages 191–194. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Planck:1978:VWS

[Pla78c] Max Planck. Über eine Verbesserung der Wienschen Spektralgleichung. (German) [On an improvement of Wien’s spectral equation]. In Schöpf [Sch78c], pages 175–177. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Planck:1978:IS

[Pla78d] Max Planck. Über irreversible Strahlungsvorgänge. (German) [On irreversible radiation processes]. In Schöpf [Sch78c], pages 166– 171. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Plank:1978:HL

[Pla78e] Maks Plank. H. A. Lorentz (1853–1928). Fiz.-Mat. Spis. B′′lgar. Akad. Nauk., 21(54)(1):75–83, 1978. CODEN FMBMAC. ISSN 0015-3265. Translated from the German by D. Vačov.

Planck:1981:WSG

[Pla81] Max Planck, editor. Where is science going? Ox Bow Press, Woodbridge, CT, USA, 1981. ISBN 0-918024-21-8, 0-918024-22-6 (paperback). 224 pp. LCCN Q175 .P436x 1981. Edited by James Murphy, with a preface by Albert Einstein.

Planck:1982:SGO

[Pla82] Max Planck. Selbstdarstellung (1942). (German) [Overview (1942)]. Akademie-Verlag, Berlin, Germany, 1982. 14 pp. LCCN ????

Planck:1988:THR

[Pla88a] Max Planck. The Theory of Heat Radiation, volume 11 of The History of modern physics. Tomash, Los Angeles, CA, USA, 1988. ISBN 0-88318-597-0. ISSN 1800-1950. xlv + 470 pp.

LCCN QC331 .P73 1988. URL http://www.loc.gov/catdir/
enhancements/fy1306/88034240-d.html. Introduction by Al-

lan A. Needell. English translation by Morton Masius.

Planck:1988:WGG

[Pla88b] Max Planck. Wissenschaft und Glaube. (German) [Science and Faith]. Union Verlag, Berlin, Germany, 1988. ISBN 3-372-00268-7. 62 pp. LCCN ????

Planck:1989:IPF

[Pla89] Max Planck. Initiations á la physique. (French) [Introduction to physics], volume 204 of Champs. Flammarion, Paris, France, 1989. ISBN 2-08-081204-1. 320 pp. LCCN ???? French translation by Joachim Du Plessis de Grenédan of Wege zur Physikalischen Erkenntnis.

Planck:1990:TT

[Pla90a] Max Planck. Treatise on Thermodynamics. Dover Publications, Inc., New York, NY, USA, 1990. ISBN 0-486-66371-X. xiv + 297 pp. Translated from the German by Alexander Ogg, Reprint of the 1945 translation.

Planck:1990:WWA

[Pla90b] Max Planck. Vom Wesen der Willensfreiheit: und andere Vorträge. (German) [The nature of free will, and other lectures]. Philosophie Fischer. Fischer Taschenbuch, Germany, Frankfurt am Main, Germany, 1990. ISBN 3-596-10472-6. 251 pp. LCCN BJ1463 .P55 1990. With an introduction by and Armin Hermann.

Planck:1990:WSG

[Pla90c] Max Planck. Wissenschaftliche Selbstbiographie. (German) [Scientific autobiography]. Acta Historica Leopoldina, 19(??):7–42, ???? 1990. ISSN 0001-8557. With a collection of documents on the genesis of Planck’s autobiography compiled by W. Berg.

Planck:1991:ASD

[Pla91a] Max Planck. Autobiographie scientifique et derniers écrits. (French) [Scientific autobiography and last writings], volume 247 of Champs. Flammarion, Paris, France, 1991. ISBN 2-08-081247-5. 220 pp. LCCN ???? Introduction, translation, and notes by André George.

Planck:1991:THR

[Pla91b] Max Planck. The Theory of Heat Radiation. Dover Publications, Inc., New York, NY, USA, 1991. ISBN 0-486-66811-8. xiv + 224 pp. LCCN QC331 .P73 1991. US$7.95.

URL http://www.loc.gov/catdir/description/dover032/
91014983.html;   http://www.loc.gov/catdir/enhancements/

fy1318/91014983-t.html. Authorized translation by Morton Masius.

Planck:1992:MPP

[Pla92] Max Planck. The Mach–Planck polemics [1908–1909]. In Ernst Mach — A Deeper Look: Documents and New Perspectives [Bla92], chapter 6, pages 127–150. ISBN 94-010-5230-1, 94-011-2771-9 (e-book). ISSN 0068-0346. LCCN C5.53; QC6.9.

URL https://link.springer.com/chapter/10.1007/978-94-
011-2771-4_6.

Planck:1993:CDM

[Pla93a] Max Planck. La conoscenza del mondo fisico. (Italian) [Knowledge of the physical world], volume 261 of Universale scientifica. Bollati Boringhieri, Torino, Italia, 1993. ISBN 88-339-0737-6. 413 pp. LCCN ???? Italian translation and editing by Enrico Persico and Augusto Gamba.

Planck:1993:SPT

[Pla93b] Max Planck. A Survey of Physical Theory. Dover Publications, Inc., New York, NY, USA, 1993. ISBN 0-486-67867-9. 117 pp. LCCN QC71 .P7213 1993. URL http://www.loc.gov/

catdir/enhancements/fy1106/93006110-d.html;   http://
www.loc.gov/catdir/enhancements/fy1318/93006110-t.html;
 http://www.loc.gov/catdir/enhancements/fy1413/93006110-

b.html. English translation by R. Jones and D. H. Williams of [Pla22h].

Planck:1993:TEM

[Pla93c] Max Planck. Theory of electricity and magnetism. UMI Books on Demand, Ann Arbor, MI, USA, 1993. xii + 247 pp. LCCN ???? English translation by Henry L. Brose of [Pla32h].

Planck:1995:ASS

[Pla95] Max Planck. Die Ableitung der Strahlungsgesetze: (1895– 1900): sieben Abhandlungen aus dem Gebiete der elektromagnetischen Strahlungstheorie. (German) [The derivation of the radiation law: (1895–1900): seven treatises on the topic of electromagnetic radiation theory], volume 206 of Ostwalds Klassiker der exakten Wissenschaften. Verlag Harri Deutsch, Thun, Switzerland and Frankfurt am Main, Germany, second edition, 1995. ISBN 3-8171-3206-9 (hardcover). v + 95 pp. LCCN

???? URL http://d-nb.info/945207220/04;   http://www.
gbv.de/dms/ilmenau/toc/187252386.PDF. With a foreword by

W. Trageser and notes by F. Reiche.

Planck:1996:ULM

[Pla96] Max Planck. The universe in the light of modern physics. In Schirmacher [Sch96], pages 38–57. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996. Contains works by Albert Einstein, Werner Heisenberg, Max Planck, and others.

Planck:1997:ASS

[Pla97] Max Planck. Die Ableitung der Strahlungsgesetze: (1895–1900): sieben Abhandlungen aus dem Gebiete der elektromagnetischen Strahlungstheorie. (German) [The derivation of the radiation law: (1895–1900): seven treatises on the topic of electromagnetic radiation theory], volume 206 of Ostwalds Klassiker der exakten Wissenschaften. Verlag Harri Deutsch, Thun, Switzerland and Frankfurt am Main, Germany, third edition, 1997. ISBN 3-8171-3206-9 (paperback). v + 95 pp. LCCN ????

Planck:1998:ELT

[Pla98] Max Planck. Eight Lectures on Theoretical Physics. Dover Publications, Inc., New York, NY, USA, 1998. ISBN 0-486-69730-4 (paperback). xiii + 145 pp. LCCN QC71

.P613 1998. URL http://www.loc.gov/catdir/description/

dover031/98010716.html. English translation by A. P. Wills. With a new introduction and notes by Peter Pesic.

Planck:2008:LTA

[Pla08a] Max Planck. Leçons de thermodynamique: avec une conférence du même auteur à la société chimique de Berlin [sur] le théorème de Nernst et l’hypothèse des quanta. (French) [Lectures on thermodynamics: with a talk by the same author at the Berlin Chemical Society on Nernst’s theorem and the quantum hypothesis]. J. Gabay, Paris, France, 2008. ISBN 2-87647-303-8 (paperback). 310 pp. LCCN QC311.

Planck:2008:TGG

[Pla08b] Max Planck. Über thermodynamische Gleichgewichte. (German) [On thermodynamic equilibrium], volume 299 of Ostwalds Klassiker der exakten Wissenschaften. Verlag Harri Deutsch, Frankfurt am Main, Germany, 2008. ISBN 3-8171-3299-9 (paperback). xxiv + 236 pp. LCCN QC318.T47 P53 2008. Includes Planck’s 1880 habilitation thesis [Pla80], edited and introduced by Werner Ebeling and Dieter Hoffmann.

Planck:2010:ODQ

[Pla10a] Max Planck. The Origin and Development of the Quantum Theory. Project Gutenberg, ????, 2010. 23 pp. URL https://

www.gutenberg.org/ebooks/33663. English translation by Hans

Thacher Clarke and Ludwik Silberstein of the Nobel Prize address delivered before the Royal Swedish Academy of Sciences at Stockholm, 2 June, 1920.

Planck:2010:VTG

[Pla10b] Max Planck. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Project Gutenberg, ????, seventh

edition, 2010. LCCN ???? URL http://www.gutenberg.org/
etext/31564;   http://www.gutenberg.org/license. Reprint

of [Pla22j].

Planck:2012:ELT

[Pla12a] Max Planck. Eight Lectures on Theoretical Physics: Delivered at Columbia University in 1909. Project Gutenberg, ????, 2012.

LCCN ???? URL http://www.gutenberg.org/ebooks/39017;
 http://www.gutenberg.org/license. Translation from Ger-

man to English by A. P. (Albert Potter) Wills.

Planck:2012:THR

[Pla12b] Max Planck. The Theory of Heat Radiation. Project Gutenberg,

????, 2012. 289 pp. URL https://www.gutenberg.org/ebooks/

40030. Authorized translation by Morton Masius of [Pla14h].

Planck:2014:SAO

[Pla14] Max Planck. Scientific Autobiography, and Other Papers. Open Road Media, ????, 2014. 196 pp. LCCN Q171 .P7 2014. Reprint of [Pla49g]. With a memorial address on Max Planck by Max von Laue. Translated from German by Frank Gaynor.

Planck:2015:AMW

[Pla15a] Max Planck. Abhandlungen über Mechanische Wärmetheorie. (German) [Treatises on the Mechanical Theory of Heat]. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-297-35098-7. 52 pp. LCCN ????

Planck:2015:PEE

[Pla15b] Max Planck. Das Princip der Erhaltung der Energie. (German) [The principle of the conservation of energy]. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-297-31095-0. 264 pp. LCCN ????

Planck:2015:MWK

[Pla15c] Max Planck. Die Mechanische Wärmetheorie…: Die Kinetische Theorie Der Gase … Hrsg. Von Dr. Max Planck Und Dr. Carl Pulfrich. 1889/91, Dritter Band. (German) [The Mechanical Theory of Heat …: The Kinetic Theory of gases …, Ed. By Dr. Max Planck and Dr. Carl Pulfrich. 1889/91, volume 3]. War College Series, ????, 2015. ISBN 1-296-48005-4. 294 pp. LCCN ????

Planck:2015:ELTa

[Pla15d] Max Planck. Eight Lectures on Theoretical Physics. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-298-20372-4. 144 pp.

Planck:2015:ELTb

[Pla15e] Max Planck. Eight lectures on theoretical physics, delivered at Columbia University in 1909. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-295-99499-2. 146 pp. LCCN ????

Planck:2015:EAM

[Pla15f] Max Planck. Einführung in die Allgemeine Mechanik: Zum Gebrauch bei Vorträgen, Sowie zum Selbstunterricht. (German) [Introduction to General Mechanics: For use in lectures as well as for self-instruction]. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-297-29118-2. 236 pp. LCCN ????

Planck:2015:MDA

[Pla15g] Max Planck. Modern Doga Anlayisi ve Kuantum Teorisine Giris (Turkish) [Introduction to Modern Perceptions of Nature and Quantum Theory]. Belge Yayınları, Istanbul, Turkey, 2015. ???? pp. LCCN ????

Planck:2015:ULM

[Pla15h] Max Planck. The Universe in the Light of Modern Physics. Scholar’s Choice, Rochester, NY, USA, 2015. ISBN 1-296-03286-8. 120 pp. LCCN ????

Planck:2015:VTW

[Pla15i] Max Planck. Vorlesungen über die Theorie der Wärmestrahlung. (German) [Lectures on the Theory of Heat Radiation]. War College Series, ????, 2015. ISBN 1-297-48500-9. 216 pp.

Planck:2015:WSG

[Pla15j] Max Planck. Where is Science Going? Facsimile Publisher, Delhi, India, 2015. ISBN 93-333-5776-9. 226 pp. LCCN ????

Planck:1947:ELG

[Pla7 ] Max Planck. Eighth lecture: General dynamics. principle of relativity. Web site and 44m49s audio recording.,

1947 (??). URL http://ia801900.us.archive.org/5/items/
8thanniversary_1308_librivox/16_planck8_pjk.mp3;   https:
//librivox.org/author/9626.

Paschen:1925:APE

[PP25] L. K. H. Friedrich Paschen and Max Planck. Antrittsrede von Paschen. Erwiderung von Planck. (German) [Inaugural speech by Paschen. Reply of Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):4–5, ???? 1925. CODEN SPWPAI. ISSN 0371-2435.

Planck:1964:VTG

[PP64] Max Planck and Max Päsler. Vorlesungen über Thermodynamik. (German) [Lectures on thermodynamics]. Walter de Gruyter, Berlin, Germany, eleventh edition, 1964. ISBN 3-11-082543-0, 3-11-177422-8. xxxiv + 343 pp. LCCN QC311 .P7 1964. Expanded to include a biography of Max Planck and a chapter by Max Päsler on some basic concepts from thermodynamics of irreversible processes.

Paschen:1979:AFP

[PP79] Friedrich Paschen and Max Planck. Antrittsrede von Friedrich Paschen (1865–1947), Erwiderung von Max Planck, gehalten in der Akademie der Wissenschaften am 2. Juli 1925. (German) [Inaugural speech by Friedrich Paschen (1865–1947), Reply by Max Planck, held in the Academy of Sciences on 2 July 1925]. In Kirsten [Kir79], pages 261–263. ISBN ???? LCCN ????

Przibram:1963:SPE

[Prz63] K. Przibram, editor. Schrödinger–Planck–Einstein–Lorentz: Briefe zur Wellenmechanik. (German) [Schrödinger–Planck– Einstein–Lorentz: Letters on Wave Mechanics]. Herausgegeben im Auftrage der österreichischen Akademie der Wissenschaften von K. Przibram. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1963. vi + 68 pp.

Przibram:1986:LWM

[Prz86] Karl Przibram, editor. Letters on wave mechanics: Albert Einstein, Erwin Schrödinger, Max Planck and Hendrik A. Lorentz. Philosophical Library, New York, NY, USA, 1986. ISBN ???? ???? pp. LCCN ????

Przibram:2011:AEL

[Prz11] Karl Przibram, editor. Albert Einstein: Letters on wave mechanics: correspondence with H. A. Lorentz, Max Planck, and Erwin Schrödinger. Open Road Integrated Media, New York, NY, USA, 2011. ISBN 1-4532-0468-7. xv + 101 pp. LCCN QC174.2.

Planck:1933:ZWN

[PS33] Max Planck and Erwin Schrödinger. Zagadnienia wspólczesnej nauki: indeterminizm, wplyw srodowiska na nauki przyrodnicze. (Polish) [Issues of modern science: indeterminism, environment and science natural]. Nakladem “Mathesis Polskiej”, Warszawa, Poland, 1933. ???? pp. LCCN ????

Plaskowski:1958:MPL

[PS58] Z. Plaskowski and H. Sprenger. Max Planck: Lebenslauf bis 1920/ Jakob Ackeret 60 Jahre. (German) [Max Planck: life up to 1920/ Jakob Ackeret’s 60th birthday]. Physikalische Blätter, 14(5):224– 226, 1958. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Pungs:1967:HGM

[Pun67] L. Pungs. Hans-Georg Möller zum 85. Geburtstag/J. R. Oppenheimer/Unruhe um Mössbauer/Max-Planck-Medaille 1967. Physikalische Blätter, 23(5):223–227, 1967. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Pyenson:1983:PSR

[Pye83] L. Pyenson. Physical sense in relativity: Max Planck edits the Annalen der Physik, 1906–1918. In Ernst Schmutzer, editor, Proceedings of the Ninth International Conference on General Relativity and Gravitation, Jena, 14-19 July 1980, pages 285–302. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24669-5. LCCN QC173.6 .I57 1980.

Pyenson:1985:PSR

[Pye85a] Lewis Pyenson. Physical sense in relativity: Max Planck edits the Annalen der Physik, 1906–1918. In The young Einstein: the advent of relativity [Pye85b], pages 176–189. ISBN 0-85274-779-9.

LCCN QC16.E5 P93 1985. URL http://www.loc.gov/catdir/
enhancements/fy0745/85120280-d.html.

Pyenson:1985:YEA

[Pye85b] Lewis Pyenson. The young Einstein: the advent of relativity. Adam Hilger Ltd., Bristol, UK, 1985. ISBN 0-85274-779-9. xiv + 255 + 9 pp. LCCN QC16.E5 P93 1985. URL http://www.loc.

gov/catdir/enhancements/fy0745/85120280-d.html.

Pyenson:1988:BRBa

[Pye88] Lewis Pyenson. Book review: The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. Isis, 79(1):122– 126, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234453.

Pyenson:2008:PSR

[Pye08] Lewis Pyenson. Physical sense in relativity: Max Planck edits the Annalen der Physik, 1906–1918. Annalen der Physik (Berlin), 17 [i.e. 520](2–3):176–189, 2008. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). Reprinted (with minor changes) from ıt The young Einstein. The advent of relativity (1985), pp. 194–214 [Adam Hilger, Ltd., Bristol].

Quinn:2013:PMP

[Qui13] Dan Quinn. Photograph of Max Planck monument in Kiel. Web

site, May 15, 2013. URL http://capturethelight.zenfolio.

com/kiel/h69564526. The monument reads “Max Planck: h = 6.62×10^{−27} erg sec: er wurde am 23. April 1858 in Kiel im Hause Küterstrasse 17 geboren und stärb am 4. Okt. 1947 in Göttingen” [Max Planck: h = 6.62 × 10^{−27} erg sec: He was born on 23 April 1858 in Kiel in the house at Küterstrasse 17 and died on 4 October 1947 in Göttingen].

Quinn:2017:AAN

[Qui17] Terry Quinn. From artefacts to atoms — a new SI for 2018 to be based on fundamental constants. Studies in History and Philosophy of Science Part A, 65–66(??):8–20, October/December 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117301851.

Robotti:2001:MPC

[RB01] Nadia Robotti and Massimiliano Badino. Max Planck and the ‘constants of nature’. Annals of Science: A Quarterly Review of the History of Science Since the Renaissance, 58(2):137–162, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Regener:1934:RWP

[Reg34] E. Regener. [Rezensionen:] Wege zur physikalischen Erkenntnis. Von Max Planck. 280 S. Verlag von S. Hirzel, Leipzig 1933. Preis geh. RM 6. geb. RM 8. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 40(2):107–108, 1934. CODEN ZEAPAA. ISSN 0005-9021.

Rechenberg:1997:TMP

[RH97] Helmut Rechenberg and Dieter Hoffmann. Zum 50. Todestag von Max Planck. (German) [On the 50th anniversary of the death of Max Planck]. Physikalische Blätter, 53(10):999–1002, 1997. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Roos:2001:VRE

[RH01] Hans Roos and Armin Hermann, editors. Vorträge Reden Erinnerungen. (German) [Lectures Speeches Memories]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-642-56594-8, 3-642-62520-7 (print). LCCN

QC1-75. URL http://link.springer.com/book/10.1007/978-
3-642-56594-6.

Risken:1996:FPE

[Ris96] H. (Hannes) Risken. The Fokker–Planck equation: methods of solution and applications, volume 18 of Springer series in synergetics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 3-540-61530-X. ISSN 0172-7389. xiv + 472 pp. LCCN QC20.7.D5 R57 1996. URL http://www.loc.gov/

catdir/enhancements/fy0817/96033182-d.html;   http://
www.loc.gov/catdir/enhancements/fy0817/96033182-t.html.

Rosenfeld:1959:MPDa

[Ros59a] Léon Rosenfeld. Max Planck et la definition statistique de l’entropie. (French) [Max Planck and the statistical definition of entropy]. VEB Deutscher Verlag der Wissenschaften, Berlin, Germany, 1959. ???? pp.

Rosenfeld:1959:MPDb

[Ros59b] Léon Rosenfeld. Max Planck et la definition statistique de l’entropie. (French) [Max Planck and the statistical definition of entropy]. In Max-Planck-Festschrift 1958 [KMP59], pages 203– 211. LCCN QC3 .M45.

Rosenfeld:1979:MPS

[Ros79] Léon Rosenfeld. Max Planck and the statistical definition of entropy. In Cohen and Stachel [CS79a], pages 235–246. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21

QC7. URL http://link.springer.com/chapter/10.1007/
978-94-009-9349-5_17.

Rossi:2008:MPC

[Ros08] Arcangelo Rossi. Max Planck: la continuità fisica e il quanto d’azione. (Italian) [Max Planck: physical continuity and the quantum of action]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 45(1–2):349–363, 2008. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Rossi:2012:BRB

[Ros12] Arcangelo Rossi. Book review: Il professore e il suo demone: La lunga lotta di Max Planck contro la statistica (1896–1906). Isis, 103(4):790–791, December 2012. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/670100.

Rosenthal-Schneider:1949:BRB

[RS49] Ilse Rosenthal-Schneider. Book review: Wissenschaftliche Selbstbiographie by Max Planck. Isis, 40(1):67–69, February 1949. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227431.

Rosenthal-Schneider:1980:RST

[RS80] Ilse Rosenthal-Schneider. Reality and scientific truth: discussions with Einstein, von Laue, and Planck. Wayne State University Press, Detroit, MI, USA, 1980. ISBN 0-8143-1650-6. 148 pp. LCCN QC6 .R5696. Edited by Thomas Braun. Foreword by Arthur I. Miller.

Rompe:1975:PPG

[RTKK75] Robert Rompe, Hans-Jürgen Treder, Hans-Günther Körber, and Christa Kirsten, editors. Physiker über Physiker. (German) [Physicists about Physicists], volume 1 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie-Verlag, Berlin, Germany, 1975. ISBN ???? ???? pp. LCCN ????

Rubens:1912:VFD

[Rub12] Heinrich Rubens. Vérification de la formule du rayonnement de Planck dans le domaine des grandes longueurs d’onde. (French) [Verification of Planck’s formula in the case of long wavelengths]. In Langevin and de Broglie [LdB12], pages 87–

91 + 92 (discussion). URL https://archive.org/details/
lathoriedurayo00inst.

Rusk:1950:BRB

[Rus50] Rogers D. Rusk. Book reviews: Scientific Autobiography and Other Papers, by Max Planck. American Journal of Physics, 18 (2):117, 1950. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://scitation.aip.org/content/
aapt/journal/ajp/18/2/10.1119/1.1932511.

Rutherford:1929:NPS

[Rut29] Ernest Rutherford. Note by Professor Sir Ernest Rutherford [Max Planck]. Die Naturwissenschaften, 17(26):483, June 28, 1929. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Ruthardt:1969:RJW

[Rut69] K. Ruthardt. A. Rubinowicz 80 Jahre/Winfried Schumann 81 Jahre/Josef Hengstenberg 65 Jahre/Max-Planck-Medaille 1969. Physikalische Blätter, 25(5):221–222, 1969. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3

S36 1994. URL http://www.loc.gov/catdir/description/
cam025/92041633.html;   http://www.loc.gov/catdir/samples/
cam031/92041633.html;   http://www.loc.gov/catdir/toc/
cam028/92041633.html.

Sachsse:1969:RMP

[Sac69] H. Sachsse. [Rezensionen:] Max Planck als Philosoph. Von H. Kretzschmar. Ernst Reinhard Verlag, München-Basel 1967. 1. Aufl., 115 S., kart. DM 11,-. Chemie in unserer Zeit, 3(1):29, February 1969. CODEN CUNZAW. ISSN 0009-2851 (print), 1521-3781 (electronic).

Sachidanandam:1984:CBD

[Sac84] S. Sachidanandam. Comment on Boyer’s derivation of the Planck spectrum. Physical Review D (Particles and Fields), 29(10):2417, May 1984. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://journals.

aps.org/prd/abstract/10.1103/PhysRevD.29.2417. See [Boy83,

Boy84].

Sangster:1987:BRJ

[San87] James Sangster. Book review: J. L. Heilbron. The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. Berkeley: University of California Press, 1986. Pp. xiii + 238. ISBN 0-520-05710-4. £15.25. British Journal for the History of Science, 20(1):123–124, January 1987. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026475.

Sommerfeld:1951:FYQ

[SB51] Arnold Sommerfeld and F. Bopp. Fifty years of quantum theory. Science, 113(2926):85–92, January 26, 1951. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//science.sciencemag.org/content/113/2926/85. Part A by

Sommerfeld, Part B by Bopp.

Scerri:2016:TSS

[Sce16] Eric R. Scerri. A tale of seven scientists and a new philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-023299-4 (hardcover), 0-19-023300-1 (e-book). xxxiv + 228 pp. LCCN Q175 .S3044 2016.

Schrodinger:1929:AHM

[Sch29] Erwin Schrödinger. Adresse an Hrn. Max Planck zum fünfzigjährigen Doktorjubiläum am 28. Juni 1929 [unsigned]. (German) [Address to Mr. Max Planck on the fiftieth anniversary of his doctorate on 28th June 1929]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):341–342, ???? 1929. CODEN SP-WPAI. ISSN 0371-2435.

Schafer:1948:MP

[Sch48] Klaus Schäfer. Max Planck. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 52(1):3–5, 1948. CODEN ZEAPAA. ISSN 0005-9021.

Schaefer:1949:RMP

[Sch49] Clemens Schaefer. [Rezensionen:] Max Planck in seinen Akademie-Ansprachen. Erinnerungsschrift der Deutschen Akademie der Wissenschaften zu Berlin. Berlin 1948. Akademie-Verlag, 212 Seiten, 1 Bild, DM 8.75. Angewandte Chemie, 61(8):346, 1949. CODEN ANCEAD. ISSN 1521-3757.

Schulze:1955:RZU

[Sch55] Gustav E. R. Schulze. [Rezensionen:] E. Zimmer, Umsturz im Weltbild der Physik. Mit einem Geleitwort von Dr. Max Planck. Zehnte, völlig überarb. Aufl. 300 S. m. 103 Abb. München 1954. Carl Hanser-Verlag. Preis geb. 12,80 DM. Zeitschrift für Angewandte Mathematik und Mechanik, 35(5):195, 1955. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Schrodinger:1961:AES

[Sch61] Erwin Schrödinger. Antrittsrede von Erwin Schrödinger (1887– 1961). Erwiderung von Max Planck. (German) [Inaugural speech by Erwin Schrödinger (1887–1961). Reply from Max Planck]. In Ch. Kirsten and H.-G. Körber, editors, Physiker über Physiker 2. Antrittsreden, Erwiderungen bei der Aufnahme von Physikern in die Berliner Akademie. Gedächtnisreden. 1870–1929. (German) [Physicist on physicist: second inaugural speeches, responses during the meeting of physicists in the Berlin Academy. Memorial speeches. 1870–1929], volume 8 of Studien zur Geschichte der Akademie der Wissenschaften der DDR. Akademie der Wissenschaften der DDR, Berlin, DDR, 1961.

Schlogl:1965:RMP

[Sch65] R. Schlögl. [Rezensionen:] Max Planck: Vorlesungen über Thermodynamik. 11. Auflage, erweitert um eine Biographie von Max Planck und ein Kapitel über einige Grundbegriffe aus der Thermodynamik irreversibler Prozesse von Max Päsler. Walter de Gruyter & Co., Berlin 1964. 343 Seiten. Preis: DM 30,-. Berichte der Bunsen-Gesellschaft f ”ur Physikalische Chemie, 1963–1990, 69(7):652, 1965. CODEN BBPCAX. ISSN 0005-9021.

Schopf:1978:MPE

[Sch78a] Hans-Georg Schöpf. Max Planck und das endgültige Strahlungsgesetz. (German) [Max Planck and the final radiation law]. In Von Kirchhoff bis Planck: Theorie der Wärmestrahlung in historischkritischer Darstellung. (German) [From Kirchhoff to Planck: the theory of thermal radiation in historical-critical representation] [Sch78c], pages 99–104. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Schopf:1978:MPQ

[Sch78b] Hans-Georg Schöpf. Max Planck und die Quantisierung des harmonischen Oszillators. (German) [Max Planck and the quantization of the harmonic oscillator]. In Von Kirchhoff bis Planck: Theorie der Wärmestrahlung in historisch-kritischer Darstellung. (German) [From Kirchhoff to Planck: the theory of thermal radiation in historical-critical representation] [Sch78c], pages 105–130. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Schopf:1978:KBP

[Sch78c] Hans-Georg Schöpf, editor. Von Kirchhoff bis Planck: Theorie der Wärmestrahlung in historisch-kritischer Darstellung. (German) [From Kirchhoff to Planck: the theory of thermal radiation in historical-critical representation]. Reihe Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1978. ISBN 3-528-06840-X. 199 pp. LCCN QC338 .S34. DM24.80.

Schopf:1978:WWM

[Sch78d] Hans-Georg Schöpf. Willy Wien, Max Planck und das Energieverteilungsgesetz. (German) [Willy Wien, Max Planck and the energy distribution law]. In Von Kirchhoff bis Planck: Theorie der Wärmestrahlung in historisch-kritischer Darstellung. (German) [From Kirchhoff to Planck: the theory of thermal radiation in historical-critical representation] [Sch78c], pages 45–67. ISBN 3-528-06840-X. LCCN QC338 .S34. DM24.80.

Schuler:1991:IMP

[Sch91] Jürgen Schüler. Zur Indeterminismuskritik von Max Planck. (German) [On the indeterminism criticism of Max Planck]. Wiss. Z. Martin-Luther-Univ. Halle-Wittenberg Math.-Natur. Reihe, 40 (2):81–86, 1991. ISSN 0138-1504.

Schirmacher:1996:GES

[Sch96] Wolfgang Schirmacher, editor. German Essays on Science in the 20th Century, volume 82 of The German library. Continuum, New York, NY, USA, 1996. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). xviii + 314 pp. LCCN QC9.G3 G47 1996. Contains works by Albert Einstein, Werner Heisenberg, Max Planck, and others.

Scheibe:2000:OSR

[Sch00a] Erhard Scheibe. The origin of scientific realism: Boltzmann, Planck, Einstein. In The Reality of the Unobservable: Observability, Unobservability and Their Impact on the Issue of Scientific Realism [AP00], chapter 2, pages 31–44. ISBN 90-481-5458-8, 94-015-9391-4 (e-book). ISSN 0068-0346. LCCN ????

URL https://link.springer.com/chapter/10.1007/978-94-
015-9391-5_2.

Schirrmacher:2000:RKH

[Sch00b] Arne Schirrmacher. Rezension: Klaus Hentschel/Renate Tobies (Hrsgg.): Brieftagebuch zwischen Max Planck, Carl Runge, Bernhard Karsten und Adolf Leopold. Mit den Promotions- und Habilitationsakten Max Plancks und Carl Runges im Anhang. (Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik, Bd 24) Berlin: ERS-Verlag 1999. 273 Seiten, broschiert DM 55,50. Berichte zur Wissenschaftsgeschichte, 23(1):62–63, ???? 2000. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Schirrmacher:2003:ETP

[Sch03] Arne Schirrmacher. Experimenting theory: The proofs of Kirchhoff’s radiation law before and after Planck. Historical Studies in the Physical and Biological Sciences, 33(2):299–335, March 2003. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Schmidt:2004:BRB

[Sch04] W. Schmidt. Book review: Brieftagebuch zwischen Max Planck, Carl Runge, Bernhard Karsten und Adolf Leopold. by Klaus Henschel und Renate Tobies (Hrsg.). Zeitschrift für Angewandte Mathematik und Mechanik, 84(8):575, 2004. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Schiesser:2017:SCM

[Sch17] William E. Schiesser. Spline Collocation Methods for Partial Differential Equations: with Applications in R. Wiley, New York, NY, USA, 2017. ISBN 1-119-30103-3 (hardcover), 1-119-30105-X (PDF), 1-119-30104-1 (ePub), 1-119-30106-8 (online). xv + 549 pp. LCCN QA377 .S355 2017.

Schneider:2022:SCP

[Sch22] Mike D. Schneider. A (strictly) contemporary perspective on trans-Planckian censorship. Foundations of Physics, 52(4):??, August 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00593-7.

Schmidt:1919:ASC

[SCP19] ?. Schmidt, Constantine Carathéodory, and Max Planck. Antrittsreden von Schmidt und Carathéodory und Erwiderung von M. Planck. (German) [Inaugural addresses of Schmidt and Carathéodory and reply by M. Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):6–7, ???? 1919. CODEN SPWPAI. ISSN 0371-2435.

Shen:2014:BPE

[SDF14] Wen Shen, T. P. Devereaux, and J. K. Freericks. Beyond Planck– Einstein quanta: Amplitude-driven quantum excitation. Physical Review B: Condensed Matter and Materials Physics, 90(19): 195104:1–195104:6, November 2014. CODEN PRBMDO. ISSN

1098-0121. URL http://journals.aps.org/prb/abstract/10.
1103/PhysRevB.90.195104.

Segre:2007:APH

[Seg07a] Emilio Segrè. Appendix 2: Planck’s hunt for the blackbody radiation formula. In From X-rays to quarks: modern physicists and their discoveries [Seg07b], pages 61–77. ISBN 0-486-45783-4.

LCCN QC7 .S4413 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006102450-d.html.

Segre:2007:XRQ

[Seg07b] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. Dover classics of science and mathematics. Dover Publications, Inc., New York, NY, USA, 2007. ISBN 0-486-45783-4. ix + 339 pp. LCCN QC7 .S4413 2007. URL http://www.loc.

gov/catdir/enhancements/fy0702/2006102450-d.html.

Segre:2007:PUR

[Seg07c] Emilio Segrè. Planck, unwilling revolutionary: the idea of quantization. In From X-rays to quarks: modern physicists and their discoveries [Seg07b], chapter 4, pages 61–77. ISBN 0-486-45783-

4. LCCN QC7 .S4413 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006102450-d.html.

Seidler:2008:NUG

[Sei08] Christoph Seidler. Namens-Überraschung: Gestatten, Marx Planck. (German) [Name surprise: Allow, Marx Planck]. Spiegel Online Wissenschaft, ??(??):??, April 22, 2008. ISSN 0038-7452.

URL http://www.spiegel.de/wissenschaft/mensch/namens-
ueberraschung-gestatten-marx-planck-a-549404.html.

Seth:2003:PPC

[Set03] Suman Seth. Principles and problems: Constructions of theoretical physics in Germany, 1890–1918. Ph.D. dissertation, Department of History, Princeton University, Princeton, NJ, USA, November 2003. xiii + 315 pp. URL https://search.proquest.

com/pqdtglobal/docview/288241277/.

Sewig:1953:GGJ

[SH53] R. Sewig and W. Hanle. Günther Güntherschulze 75 Jahre/Paul Cermak 70 Jahre alt/25 Jahre Max-Planck-Medaille. Physikalische Blätter, 9(8):363–365, 1953. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Shpolskii:1958:CPM

[Shp58a] E. V. Shpol’skii. The creative path of M. Planck. Physics-Uspekhi, 1(2):326–332, February 1958. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://ufn.ru/en/

articles/1958/2/k/.

Shpolskij:1958:TPM

[Shp58b] Je. V. Shpol’skij. Tvorcheskij put’ M. Planka. Uspekhi Fizicheskikh Nauk, 66(11):535–542, November 1958. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1958/11/g/.

Simon:2006:AEA

[Sim06] Dieter Simon. Albert Einstein: Akademie-Vorträge: Sitzungsberichte der Preußischen Akademie der Wissenschaften, 1914– 1932. (German) [Albert Einstein: Academy lectures: Proceedings of the Prussian Academy of Sciences, 1914–1932]. Wiley-VCH, Weinheim, Germany, 2006. ISBN 3-527-40609-3 (hardcover), 3-527-60895-8 (e-book), 3-527-62244-6 (e-book). xiv + 431 pp. LCCN QC71 .E36 2006.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-011-0065-6;
 http://www.springerlink.com/content/k12202vg92147h68/

.

Singh:2008:MPG

[Sin08] Rajinder Singh. Max Planck and the genesis of the energy quanta in historical context. Current Science, 95(6):788–792, 2008. CO-DEN CUSCAM. ISSN 0011-3891.

Schuss:2001:DPN

[SNE01] Z. Schuss, B. Nadler, and R. S. Eisenberg. Derivation of Poisson and Nernst–Planck equations in a bath and channel from a molecular model. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 64(3):036116:1– 036116:14, August 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://journals.aps.

org/pre/abstract/10.1103/PhysRevE.64.036116.

Sommerfeld:1911:PWS

[Som11] Arnold Sommerfeld. Das Plancksche Wirkungsquantum und seine allgemeine Bedeutung für die Molekularphysik. (German) [Planck’s quantum of action and its overall importance for molecular physics]. Physikalische Zeitschrift, 12(24):1057– 1069, December 1, 1911. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015011417709%3Bview=1up%3Bseq=1115.

Sommerfeld:1912:ATE

[Som12] Arnold Sommerfeld. Application de la théorie de l’élément d’action aux phénomènes moléculaires non périodiques. (French) [Application of the theory of the quantum of action to nonperiodic molecular phenomena]. In Langevin and de Broglie [LdB12], pages 313–373 + 374–392 (discussion). URL https://archive.

org/details/lathoriedurayo00inst.

Sommerfeld:1918:MPS

[Som18] Geheimrat Prof. Dr. Arnold Sommerfeld. Max Planck zum sechzigsten Geburtstage. (German) [Max Planck on his sixtieth birthday]. Die Naturwissenschaften, 6(17):195–199, April 26, 1918. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF01491326.

Sommerfeld:1933:PMW

[Som33] Arnold Sommerfeld. Planck, M.: Wege zur physikalischen Erkenntnis. Reden und Vorträge von Max Planck — Besprechung. (German) [Planck, M.: Paths to physical knowledge. Speeches and Lectures by Max Planck — Review]. ????, ??(??):??, ???? 1933.

Sommerfeld:1936:RMP

[Som36] Arnold Sommerfeld. Rezension: Max Planck, Die Physik im Kampf um die Weltanschauung. (German) [Review: Max Planck, Physics in the Battle for World Views]. Physikalische Zeitschrift, 37(??):144–??, ???? 1936. CODEN PHZTAO. ISSN 0369-982X.

Sommerfeld:1938:MPG

[Som38a] Arnold Sommerfeld. Max Planck zum 80. Geburtstag. (German) [Max Planck on his 80th birthday]. Münchener Medizinische Wochenschrift, 85(??):829–830, ???? 1938.

Sommerfeld:1938:TBF

[Som38b] Arnold Sommerfeld. Tischrede beim Festessen im Goethe-Saal des Hamack-Hauses (zu Ehren von Max Planck). (German) [Table talk at the feast in the Goethe Hall of the Hamack House (in honor of Max Planck)]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19(??):69–72, ???? 1938. CODEN VDPEAZ. ISSN 0372-5448.

Sommerfeld:1947:MPG

[Som47] Arnold Sommerfeld. Max Planck zum Gedächtnis. (German) [In memory of Max Planck]. Neue Zeitung, ??(??):??, October 6, 1947.

Sommerfeld:1948:GPG

[Som48] Arnold Sommerfeld. Gedächtnisfeier der Physikalischen Gesellschaft in Württemberg-Baden zu Heidenheim am 15. November 1947 [für Max Planck]. (German) [Memorial of the Physical Society in Württemberg-Baden to Heidenheim, 15 November 1947 [for Max Planck]]. Annalen der Physik (1900), 438(1):3–6, 1948. ISSN 0003-3804 (print), 1521-3889 (electronic).

Sommerfeld:1949:RBP

[Som49] A. Sommerfeld. [Rezensionen:] Max Planck in seinen Akademie-Ansprachen Erinnerungsschrift der Deutschen Akademie der Wissenschaften in Berlin. DIN A 5 broschiert, 212 Seiten, 1 Bild, Ladenpreis DM 8,75. Akademie-Verlag Berlin. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 53(4): 257, 1949. CODEN ZEAPAA. ISSN 0005-9021.

Schrodinger:1929:ASE

[SP29] Erwin Schrödinger and Max Planck. Antrittsrede von Schrödinger. Erwiderung von Planck [Sonderdruck]. (German) [Inaugural speech by Schrödinger. Reply of Planck [reprint]]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):??, ???? 1929. CODEN SPWPAI. ISSN 0371-2435.

Spretke:1989:BPT

[Spr89] Ulf Spretke. Bemerkungen zur Physik- und Theorieauffassung Max Plancks. (German) [Comments on Max Planck’s physics and theory]. Wiss. Z. Martin-Luther-Univ. Halle-Wittenberg Math.- Natur. Reihe, 38(1):115–119, 1989. ISSN 0138-1504.

Sommerfeld:1937:NRG

[SR37] Arnold Sommerfeld and E. Rüchardt. Nachruf Rutherford, Gratulation Millikan, Planck Gedächtnisrede auf Mach. Rede zu Plancks 80. Geburtstag. (German) [Obituary of Rutherford, congratulations to Millikan, Planck’s memorial speech on Mach. Speech for Planck’s 80th birthday]. München. med. Wochenschr., ??(??):829–??, ???? 1937.

Szekely:1953:HBD

[SS53] A. Székely and E. Spiro. Hans Benndorf/Max Planck. Physikalische Blätter, 9(4):176–177, 1953. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Staley:2010:BRB

[Sta10] Richard Staley. Book review: Max Planck und die Max-Planck-Gesellschaft: Zum 150, by Lorenz Friedrich Beck. Geburtstag am 23. April 2008 aus den Quellen zusammengestellt. Isis, 101 (4):891–892, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/659696.

Stichweh:1984:EMS

[Sti84] Rudolf Stichweh. Zur Entstehung des modernen Systems wissenschaftlicher Disziplinen: Physik in Deutschland, 1740–1890. (German) [On the origin of the modern system scientific disciplines: physics in Germany, 1740–1890]. Suhrkamp, Frankfurt am Main, West Germany, 1984. ISBN 3-518-57688-7. 558 pp. LCCN QC9.G3 S75 1984.

Stillinger:1991:PCE

[Sti91] Frank H. Stillinger. Planck’s-constant expansions for bound states. Physical Review A (Atomic, Molecular, and Optical Physics), 43(7):3317–3324, April 1991. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http:/

/journals.aps.org/pra/abstract/10.1103/PhysRevA.43.3317.

Stone:2013:EQQ

[Sto13] A. Douglas Stone. Einstein and the quantum: the quest of the valiant Swabian. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-13968-7 (hardcover), 1-4008-4834-2 (ebook), 1-299-87423-1 (e-book). x + 332 pp. LCCN QC16.E5

S76 2013. URL http://www.npr.org/2013/11/01/242356997/
einsteins-real-breakthrough-quantum-theory.

Straumann:2011:FSC

[Str11a] Norbert Straumann. On the first Solvay Congress in 1911. European Physical Journal H, 36(3):379–399, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/
epjh/e2011-20043-9.

Strickland:2011:WSC

[Str11b] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Swenson:1989:BRB

[Swe89] Loyd S. Swenson. Book review: The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. The American Historical Review, 94(1):171–172, February 1989. ISSN 0002-8762 (print), 1937-5239 (electronic). URL http://www.jstor.

org/stable/view/10.2307/1862175.

Teichmann:2002:PPM

[TEW02] Jürgen Teichmann, Michael Eckert, and Stefan Wolff. Physicists and physics in Munich. Physics in Perspective (PIP), 4 (3):333–359, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8372-6.

terHaar:1967:OQT

[tH67] D. ter Haar. The Old Quantum Theory. The Commonwealth and international library. Selected readings in physics. Pergamon Press, Oxford, UK, 1967. x + 206 pp. LCCN QC174 .H33 1967.

Thiele:1969:ZUK

[Thi69] Joachim Thiele. Ein zeitgenössisches Urteil über die Kontroverse zwischen Max Planck und Ernst Mach. (German) [A contemporary assessment of the controversy between Max Planck and Ernst Mach]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 13(1):85–90, March 1969. CO-DEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thiele:1971:BRV

[Thi71] Joachim Thiele. Book review: Vorgeschichte des Planckschen Strahlungsgesetzes. Messungen und Theorien der spektralen Energieverteilung bis zur Begründung der Quantenhypothese (Boethius, Bd. XI) by Hans Kangro. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 55(2):221–223, ???? 1971. CO-DEN SUARAH. ISSN 0039-4564. URL http://www.jstor.org/

stable/20775907.

Thiessen:1988:MPP

[Thi88] Peter Adolf Thiessen. Max Planck prüft und verbessert. (German) [Max Planck reviewed and improved]. Annalen der Physik (Berlin), 500(3):161–162, 1988. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Thomson:1931:JCMb

[Tho31] Sir Joseph John Thomson, editor. James Clerk Maxwell; a commemoration volume, 1831–1931. Cambridge University Press, Cambridge, UK, 1931. 3 + 146 pp. LCCN QC16.M4 J3. Essays by Sir J. J. Thomson, Max Planck, Albert Einstein, Joseph Larmor, James Jeans, William Garnett, Ambrose Fleming, Oliver Lodge, R. T. Glazebrook, and Horace Lamb.

Tibbetts:2013:MPF

[Tib13] Gary G. Tibbetts. 7 — Max Planck, the first superhero of quantum theory, saves the universe from the ultraviolet catastrophe: Assemble two flawed hypotheses about a key phenomenon into a model that fits experiment exactly and people will listen to you even if you must revolutionize physics. In Gary G. Tibbetts, editor, How the Great Scientists Reasoned, pages 89–99. Elsevier, Amsterdam, The Netherlands, 2013. ISBN 0-12-398498-X.

URL http://www.sciencedirect.com/science/article/pii/
B9780123984982000072.

Tobies:2008:MPT

[Tob08] Renate Tobies. Max Planck: ,,Treulichst Euer Max”: Max Planck und das Brieftagebuch mit seinen Freunden aus der Münchner Studienzeit. (German) [Max Planck: “faithfully your Max”: Max Planck and the letter diary with his friends from his Munich study time]. Physik Journal, 7(3):47–50, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1104321/issue.html.

Theimer:1976:RCC

[TP76] O. Theimer and P. R. Peterson. Reply to “Comment on classical derivations of Planck’s spectrum”. Physical Review D (Particles and Fields), 14(2):656, July 15, 1976. CODEN PRV-DAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500

(CD-ROM). URL http://journals.aps.org/prd/abstract/
10.1103/PhysRevD.14.656. See [Kra76].

Treder:2001:BRN

[Tre01] Hans-Jürgen Treder. Book review: Naturwissenschaft und Religion (Versuch einer Verhältnisbestimmung) Science Edition by Wilfried Schröder, Max Planck, Werner Heisenberg. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 85(1):119–120, ???? 2001. CODEN SUARAH. ISSN 0039-4564. URL http:

//www.jstor.org/stable/20777800.

Ulimann:1997:MPW

[Uli97] Dirk Ulimann. Max Planck als Wissenschaftsorganisator im Spiegel der archivalischen Überlieferung. (German) [Max Planck as a scientific organizer in the mirror of archival tradition]. Physikalische Blätter, 53(10):1017–1018, 1997. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Ullmann:1996:QMP

[Ull96] Dirk Ullmann. Quelleninventar Max Planck. (German) [Max Planck Source Inventory], volume 8 of Veröffentlichungen aus dem Archiv zur Geschichte der Max-Planck-Gesellschaft. Archiv zur Geschichte der Max-Planck-Gesellschaft, Berlin, Germany, 1996. ISBN 3-927579-08-4. 174 pp. LCCN ????

URL http://www.gbv.de/dms/goettingen/229671136.pdf;
http://www.gbv.de/dms/ohb-opac/229671136.pdf.

Ullmann:2013:MPW

[Ull13] Dirk Ullmann. Max Planck als Wissenschaftsorganisator im Spiegel der archivalischen Überlieferung. (German) [Max Planck as a science organizer in the mirror of archival tradition]. Physikalische Blätter, 53(10):1017–1018, October 2013. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19970531018/full.

Unsold:1958:MPD

[Uns58] A. Unsòld. Max-Planck-Denkmal in Kiel. (German) [Max Planck monument in Kiel]. Physikalische Blätter, 14(10):470, 1958. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Unsold:1967:MPS

[Uns67] A. Unsöld. Max Planck, seine Zeit und unsere Zeit: Gedanken zur Max-Planck-Ausstellung in Hamburg und Kiel. (German) [Max Planck, his time and our time: Rethinking the Max Planck Exhibition in Hamburg and Kiel]. Physikalische Blätter, 23(9):405– 416, 1967. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

vanderWaals:1947:LVM

[van48] J. D. van der Waals, Jr. Levensbericht van M. Planck (23 April 1858–4 October 1947). (Dutch) [Life report of M. Planck (23 April 1858–4 October 1947)]. In Huygens Institute — Royal Netherlands Academy of Arts and Sciences (KNAW) Jaarbook, pages 177–182. Royal Netherlands Academy of Arts and Sciences, Amsterdam, The Netherlands, 1947–1948. URL http://www.dwc.

knaw.nl/DL/levensberichten/PE00002335.pdf.

vanderWaals:2010:EMK

[van10] J. D. van der Waals, Jr. Eulogy: Max Karl Ernst Ludwig Planck (1858–1947). Web site, September 24, 2010. URL http://www.

dwc.knaw.nl/biografie/pmknaw/?pagetype=authorDetail&aId=
PE00002335;   http://www.dwc.knaw.nl/DL/levensberichten/

PE00002335.pdf. See also [van48].

vBichowsky:1918:NPA

[v.B18] F. Russell v.Bichowsky. The necessary physical assumptions underlying a proof of the Planck radiation law. Physical Review, 11(1):58, January 1918. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://journals.aps.

org/pr/abstract/10.1103/PhysRev.11.58.

vomBrocke:1996:KWS

[vBL96] Bernhard vom Brocke and Hubert Laitko, editors. Die Kaiser-Wilhelm-/Max-Planck-Gesellschaft und ihre Institute: Studien zu ihrer Geschichte. (German) [The Kaiser Wilhelm/Max Planck Society and their institutions: studies on their history]. Walter de Gruyter, Berlin, Germany, 1996. ISBN 3-11-015483-8 (vol. 1). xiii + 672 pp. LCCN Q180.G42 K35 1996. URL http://

www.degruyter.com/viewbooktoc/product/146831;   http:
//www.reference-global.com/doi/book/10.1515/9783110802443.

VanDusen:1947:FPC

[VD47] M. S. Van Dusen and A. I. Dahl. Freezing points of cobalt and nickel and a new determination of Planck’s constant C_{2}. Science, 106(2757):428–429, 1947. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/106/2757/428.full.pdf;   http:
//www.sciencemag.org/content/106/2757/428.short.

vanKampen:1997:FPG

[vK97] N. G. van Kampen. Die Fokker–Planck-Gleichung. (German) [The Fokker–Planck equation]. Physikalische Blätter, 53(10): 1012–1013, October 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19970531016/abstract.

vonLaue:1918:PTA

[vL18] Max von Laue. Plancks Thermodynamische Arbeiten. (German) [Planck’s thermodynamical work]. In Zu Max Plancks sechzigstem Geburtstag: Ansprachen, gehalten am 26. April 1918 in der Deutschen Physikalischen Gesellschaft, nebst einer Erwiderung von M. Planck [WvLuAE18], pages 6–15. ISSN 0005-9021.

vonLaue:1938:GMP

[vL38] M. von Laue. Zum 80. Geburtstag von Max Planck. (German) [On the 80th birthday of Max Planck]. Annalen der Physik (Berlin), 424(1–2):2–4, 1938. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

vonLaue:1947:MP

[vL47] Max von Laue. Max Planck. Physikalische Blätter, 3(8):249–252, 1947. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

vonLaue:1948:MP

[vL48] Max von Laue. Max Planck. Die Naturwissenschaften, 35(1):1–7, January 1948. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://link.springer.com/article/
10.1007/BF00626622.

vonLaue:1950:NMP

[vL50] Max von Laue. Nachruf auf Max Planck. (German) [Obituary of Max Planck]. In Jahrbuch der Deutschen Akademie der Wissenschaften zu Berlin, 1946–1949, pages 217–220. Deutschen Akademie der Wissenschaften, Berlin, West Germany, 1950.

vonLaue:1958:MPG

[vL58] Max von Laue. Zu Max Planck’s 100. Geburtstag. Vortrag vor der Deutsche Akademie er Wissenschaften zu Berlin am 24.4.1958. (German) [On Max Planck’s 100th birthday. Lecture before the German Academy of Sciences, Berlin, 24 April 1958]. Die Naturwissenschaften, 45(10):221–226, May 2, 1958. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

vonLaue:1921:AMV

[vLP21] Max von Laue and Max Planck. Antrittsrede M. v. Laues und Erwiderung von M. Planck. (German) [Inaugural speech by M. v. Laue and reply of M. Planck]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, ??(??):3–4, ???? 1921. CODEN SPWPAI. ISSN 0371-2435.

Vogel:1961:PWM

[Vog61] Heinrich Vogel. Zum philosophischen Wirken Max Plancks. Seine Kritik am Positivismus. (German) [On the philosophical work of Max Planck. His criticism of positivism]. Akademie-Verlag, Berlin, Germany, 1961. 255 pp.

vonPufendorf:2006:PFZ

[vP06] Astrid von Pufendorf. Die Plancks: eine Familie zwischen Patriotismus und Widerstand. (German) [The Plancks: a family between patriotism and resistance]. Propyläen, Berlin, Germany, 2006. ISBN 3-549-07277-5. 512 + 16 pp. LCCN DD256.3

.P84 2006. URL http://www.loc.gov/catdir/toc/fy0609/
2006388371.html.

Wilber:1987:CCE

[W^{+}87] Ken Wilber et al., editors. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, 1987. ISBN 84-7245-172-0. 298 pp. LCCN ????

Walker:1994:SNS

[Wal94] Mark Walker. Science, National Socialism, and the longue durée. Historical Studies in the Physical and Biological Sciences, 24(2): 395–402, ???? 1994. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Warburg:1912:VEF

[War12] Emil Warburg. Vérification expérimentale de la formule de Planck pour le rayonnement du corps noir. (French) [Experimental verification of Planck’s formula for blackbody radiation]. In Langevin and de Broglie [LdB12], pages 78–85 + 86 (discussion). URL

https://archive.org/details/lathoriedurayo00inst.

Wasserman:1978:TIP

[WB78] Ira Wasserman and Kenneth Brecher. Time invariance of Planck’s constant. Physical Review Letters, 41(14):920–923, October 1978. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://journals.aps.org/prl/
abstract/10.1103/PhysRevLett.41.920.

Woodward:1991:WVSb

[WC91] William R. Woodward and Robert S. Cohen, editors. World Views and Scientific Discipline Formation: Science Studies in the German Democratic Republic: Papers from a German–American Summer Institute, 1988, volume 134 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-5395-2, 94-011-3164-3 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-011-3164-3.

Wilber:2002:CCE

[WdC02] Ken Wilber and Pedro de Casso. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, seventh edition, 2002. ISBN 84-7245-172-0. 296 pp. LCCN ????

Wegener:2010:AEE

[Weg10] Daan Wegener. De-anthropomorphizing energy and energy conservation: The case of Max Planck and Ernst Mach. Studies in History and Philosophy of Modern Physics, 41(2):146–159, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000158.

Weisskopf:1958:MPO

[Wei58] V. F. Weisskopf. Max Planck: One hundredth birthday celebration. Physics Today, 11(8):16–19, August 1958. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weil:1968:RMP

[Wei68] K. G. Weil. [Rezensionen:] Max Planck: Theorie der Wärmestrahlung. 6. Auflage. Johann Ambrosius Barth Verlag, Leipzig 1965. 221 Seiten. Preis: DM 16,50. Berichte der Bunsen-Gesellschaft f ”ur Physikalische Chemie, 1963–1990, 72(3):490A–490, 1968. CODEN BBPCAX. ISSN 0005-9021.

Weindling:1989:HRG

[Wei89] Paul Weindling. Health, race, and German politics between national unification and Nazism, 1870–1945. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36381-0. x + 641 pp. LCCN RA418.3.G3

W45 1989. URL http://www.loc.gov/catdir/description/
cam023/88037009.html;   http://www.loc.gov/catdir/toc/
cam028/88037009.html.

Weir:2007:MPU

[Wei07] Jane Weir. Max Planck: Uncovering the World of Matter. Physical Science (Science Readers). Shell Education, Huntington Beach, CA, USA, 2007. ISBN 0-7439-0568-7, 1-4333-9097-3 (ebook). 18 pp. LCCN QC16.P6 .W452 2007.

Weir:2009:MPR

[Wei09] Jane Weir. Max Planck: revolutionary physicist. Compass Point Books, Minneapolis, MN, USA, 2009. ISBN 0-7565-4073-9 (hardcover). 40 pp. LCCN QC16.P6 W45 2009. URL http://catdir.

loc.gov/catdir/toc/ecip0827/2008037622.html.

Wessel:1940:SWK

[Wes40] W. Wessel. Zur Synthese von Wellen- und Korpuskularmechanik. Anmerkungen zu der Abhandlung Plancks. (German) [On the synthesis of wave and particle mechanics. Comments on the paper by Planck]. Annalen der Physik (Berlin), 430(4):261–271, ???? 1940. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Westphal:1957:CSJ

[Wes57] Wilhelm H. Westphal. Clara von Simson 60 Jahre/Zur zehnten Wiederkehr des Todes von Max Planck/Der Pour le Mérite auch für Erwin Schrödinger. (German) [Clara von Simson 60th birthday/the tenth anniversary of the death of Max Planck/the Pour le Mérite for Erwin Schrödinger]. Physikalische Blätter, 13(10):470– 471, 1957. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Westphal:1958:MPM

[Wes58] Wilhelm H. Westphal. Max Planck als Mensch. (German) [Max Planck as a human being]. Die Naturwissenschaften, 45(10):234– 236, January 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Westphal:1963:HMP

[WF63] Wilhelm Westphal and V. Fragstein. Heiligers Max Planck-Plastik/Bergmann-Schäfer Lehrbuch der Experimentalphysik III. Physikalische Blätter, 19(10):481–483, 1963. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Wadlinger:1988:MPN

[WH88] Robert L. Wadlinger and Geoffrey Hunter. Max Planck’s natural units. The Physics Teacher, 26(8):528–529, November 1988. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928

(electronic). URL http://scitation.aip.org/content/aapt/
journal/tpt/26/8/10.1119/1.2342611.

Whyte:1950:PCF

[Why50] L. L. Whyte. Planck’s constant and the fine-structure constant [α]. Nature, 166(4228):824–825, November 11, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v166/
n4228/pdf/166824b0.pdf.

Willers:1937:REZ

[Wil37] Willers. [Rezensionen:] Ernst Zimmer, Umsturz im Weltbild der Physik, mit einem Geleitwort von Dr. Max Planck. 3. Aufl. 270 S. m. 59 Abb. München 1936, Verlag Knorr & Hirth, G. m. b. H. Preis geb. 5,70 M. Zeitschrift für Angewandte Mathematik und Mechanik, 17(6):375, 1937. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Wolff:2008:MPK

[Wol08] Stefan L. Wolff. Max Planck: Kontrovers, aber kooperativ Max Planck und Wilhelm Wien — eine Zusammenarbeit über Gegensätze hinweg. (German) [Max Planck: controversial, but cooperative: Max Planck and Wilhelm Wien — cooperation on opposite ways]. Physik Journal, 7(3):51–55, March 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1104323/issue.html.

Worrell:1963:BRB

[Wor63] Francis T. Worrell. Bok review: A Survey of Physical Theory, by Max Planck. American Journal of Physics, 31(7):555–556, 1963. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://scitation.aip.org/content/aapt/
journal/ajp/31/7/10.1119/1.1969663.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Williams:1998:AMP

[WSNO98] Edwin R. Williams, Richard L. Steiner, David B. Newell, and Paul T. Olsen. Accurate measurement of the Planck constant. Physical Review Letters, 81(12):2404–2407, September 1998. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://journals.aps.org/prl/abstract/10.
1103/PhysRevLett.81.2404.

Warburg:1918:MPS

[WvLuAE18] Emil Warburg, Max von Laue, and Arnold Sommerfeld und Albert Einstein. Zu Max Plancks sechzigstem Geburtstag: Ansprachen, gehalten am 26. April 1918 in der Deutschen Physikalischen Gesellschaft, nebst einer Erwiderung von M. Planck. C. F. Müllerschen Hofbuchhandlung m.b.H., Karlsruhe, Germany, 1918. ISSN 0005-9021. 36 pp.

Yost:1995:BPK

[Yos95a] Res Yost. Boltzmann und Planck: Die Krise des Atomismus um die Jahrhundertwende und ihre Überwindung durch Einstein. (German) [Boltzmann and Planck: The crisis of atomism at the turn of the century and its surmounting by Einstein]. In Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers] [Jos95], pages 35–51. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. LCCN QC16.J67 J67 1995. URL http://

www.springerlink.com/openurl.asp?genre=issue\%26issn=

0940-7677\%26volume=34. Edited posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Yost:1995:EMM

[Yos95b] Res Yost. Ernst Mach und Max Planck. In Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers] [Jos95], pages 53–66. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. LCCN QC16.J67 J67 1995. URL http://

www.springerlink.com/openurl.asp?genre=issue\%26issn=

0940-7677\%26volume=34. Edited posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Yost:1995:PKK

[Yos95c] Res Yost. Planck — Kritik des T. Kuhn. [Planck — criticism by T. Kuhn]. In Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers] [Jos95], pages 67–78. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. LCCN QC16.J67 J67 1995. URL http://

www.springerlink.com/openurl.asp?genre=issue\%26issn=

0940-7677\%26volume=34. Edited posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Yost:1995:VPS

[Yos95d] Res Yost. Zur Vorgeschichte des Planckschen Strahlungsgesetzes. (German) [On the prehistory of Planck’s radiation law]. In Das Märchen vom Elfenbeinernen Turm: Reden und Aufsätze. (German) [The tale of the ivory tower: Speeches and Papers] [Jos95], pages 23–34. ISBN 3-540-59476-0, 3-540-49276-3. ISSN 0940-7677. LCCN QC16.J67 J67 1995. URL http://

www.springerlink.com/openurl.asp?genre=issue\%26issn=

0940-7677\%26volume=34. Edited posthumously by Abraham Pais, K. (Klaus) Hepp, W. Hunziker, and Walter Kohn.

Yourgrau:1968:BRB

[You68] W. Yourgrau. Book review: Max Planck als Philosoph. Hermann Kretzschmar. Reinhardt, Munich, 1967. 115 pp. Paper, DM 11. Science, 162(3859):1264, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/162/3859/1264.1.short.

Zacher:1992:MPG

[Zac92] Hans F. Zacher. The Max-Planck-Gesellschaft and German unification. Science and Public Policy, 19(2):71–74, April 1992. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zajac:1983:MPR

[Zaj83] Rudolf Zajac. Max Planck and the rise of modern physics. Pokroky Mat. Fyz. Astronom., 28(6):311–319, 1983. CODEN PM-FAA4. ISSN 0032-2423.

Zenneck:1938:MPA

[Zen38] J. Zenneck. Max Planck zum achtzigsten Geburtstag. (German) [Max Planck on his 80th birthday]. Die Naturwissenschaften, 26 (16):241–242, April 1938. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://link.springer.

com/article/10.1007/BF01774026.

Zilsel:1933:RWP

[Zil33] E. Zilsel. [Rezensionen:] Wege zur physikalischen Erkenntnis. Reden und Vorträge. Von Max Planck. IX, 279 Seiten. Verlag S. Hirzel, Leipzig 1933. Preis geh. RM. 6, geb. RM. 8. Angewandte Chemie, 46(24):355, 1933. CODEN ANCEAD. ISSN 1521-3757.

Zimmer:1934:UIW

[Zim34] Ernst Zimmer. Umsturz im Weltbild der Physik. (German) [Upheaval in the world of physics]. Knorr and Hirth GmbH, München, Germany, 1934. 263 pp. LCCN QC7 .Z48 1934. Mit einem Geleitwort von Dr. Max Planck und 58 Abbildungen [With a foreword by Dr. Max Planck and 58 pictures].